A Complete Bibliography of Publications in Computer Physics Communications: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.21

Title word cross-reference

(2 + 1) [Sin94]. (θ, δ) [OF92]. +6 [Str91b]. +γ [vOFB94]. −1 [Str91b]. 1 [AB99, BT92a, Fij99b, HKG90, IIM92, KB98, RCL98]. 1 + 1 [App99]. 1.10 [CO92]. 1.5 [RI97]. 2 [BL99b, BMPS99, ÇYG99, Fij99c, FMIM93, FV99, HHH99, KW90, KKMS99, MD96, PD90, PS98a, Rea95, RTLM92, SPA99, SS94a, SVV99, TAMSW90, VBS96]. 2.0 [CMNP97, MNPP96, PP94]. 2.00 [Odo96]. 2.2 [IRS97, KM98]. 3 [ADS92, CHMG92, DJUM99, Eas91a, Fer97, HKA^{+}94, JBB99, JV99, JMR^{+}93, KR95b, Kud99, LC90a, MMCH93, MOSM96, NS90, Ott90, PC99, PCSD99, PG94, RMMUJ99, Scr95, SA99b, Sou99, SW91b, TBS98, WLD95, XZ91]. 3 + 1 [MRC98, Sal94]. 4 [Dik97, Lön92, Pas96]. 5 [FT95]. 5′′ × 5′′ [NCC^{+}96]. 6 [AC92, LC90a, RWL^{+}98, TBS98]. 6.5 [IER97]. 64 [DH97b]. 8 [ÇYG99, CBB99]. 9 [LC92]. ^{+} [PI95, Zie98]. _{2}^{+} [Hea91]. ^{0} [JY97b, MPN^{+}93b, MNPP96]. ^{1} [WW93]. ^{14+} [MSP^{+}98]. ^{1}Σ_{u}^{+} [PLC91]. ^{2} [MK96]. ^{3} [PG96]. ^{4} [CPCF99, MBC99a]. ^{89} [Car94]. ^{90} [Car94]. ^{N} [Koo93, Mon91]. _{1} [MCCC99]. _{2} [GE99, MCCC99, PLC91, PRRL99]. _{3} [MCCC99]. _{f} [Gra98]. _{t} [MSW92b]. _{τ} [Shu94, BTC99a]. α [Glü97, Joó97]. α_{S} [PB93]. B [BDL92, HWS92, HS93b, JY97b, SPP^{+}98, BB93a]. B^{00} [JY97a]. β [CMC^{+}99, Glü97, Joó97]. D [CHM91, HPV99, Koo93]. Δ_{T} q [HKM98a]. E [AO97, AS91]. e^{+}e^{−} [AB97, CMNP97, FR96, FIK^{+}97, FHI^{+}98, JPS^{+}99, KM91, KM98, Odo92, Pap99]. e^{+}e^{−} → 4 [Sig97, vOFB94]. e^{+}e^{−} → 4f [BBL^{+}97]. e^{+}e^{−} → f^{+} + f^{−} [BBR^{+}90]. e^{+}e^{−} → ff [KS92a]. e^{+}e^{−} → v_{e}vν_{e}W^{+}W^{−} [NK96]. e^{+}e^{−} → vvνγ(γ) [MNP96]. e^{+}e^{−} → W^{+}W^{−} [FJKvO95]. e^{+}e^{−} → W^{+}W^{−}γ [TKS91]. e^{+}e^{−} → W^{+}W^{−} → 4f [MNP95]. e^{+}e^{−} → Z → τ^{+}τ^{−} [Sti94]. e^{±}p → e^{±}X [JP92]. ep [ABB^{+}96, Jun95, KSM92]. f [GF96]. γ [GLF97, Gra92, JL93]. GF(2^{n}) [BTC99b]. H [NM97]. I [SMH98, WBS95]. j [LC90a, LC92, TBS98, Wei11]. J_{n}(z) [dT93]. K [HB90, STI99]. L [MCCC99]. l^{±}N [ABB^{+}96]. M [MCCC99]. M_{W} [BPP98]. μ(x, β, α) [Aya92]. N [BADP96, CW92, IMES90, IOM95, Kal93, Ryc90a, RJ90, Ryc90b, RJ91a, RJ91b, SGF99b, MCV98, RP91, Tau92]. O(α) [JWJK92]. O(h̄8) [Sha92]. O(N) [BG99, CGM^{+}96, Goe99]. p [AS91, FT99a, Odo90, Odo92]. Φ [MPN93a]. P m¯3m [DPD96b]. P m¯3n [DPD96d]. Q [AO97, CK99, MK96, CHM91]. q = exp[−πK(k′)K(k)] [Lee92]. Q^{2} [HKM98b, HKM98a]. R [CGHB98, HSSB98, NC96, SMB^{+}93, SHNS98, BEN95, BN95, Str93b]. R^{4} [BGT95]. S [FHI^{+}98]. S^{2} [Len98]. SO(3) [SS90b]. SU(3) [BD94, Kae95, MSTK93]. SU(3) ⊃ SU(2) ⊗ U(1) [KW96]. SUSY 2 [Pop97]. T [Sch93, FMH97]. T = 0 [FV99]. τ [JW91, JKW91, JWDK93, JWJK92]. τ → μ(e)vv [JWJK92]. → [BKMM93, WW93]. U [Tau92, RBB93]. V [BS99a]. W [ABH^{+}94, SJPW96, Wag92]. x [Tau92]. Y_{n}(z) [dT93]. Z [FR96, KR95a, MPN^{+}93b, MNPP96, Wag92]. Z^{0} [BBR^{+}90, KS92a].

-1 [Ano98a]. -algebra [JL93]. -bit [DH97b]. -body [BADP96, IMES90]. -brass [CMC^{+}99]. -butane [RP91]. -chloranil [FT99a]. -cubic [MRC98]. -D [BT92a, Eas91a, IIM92, JMR^{+}93, NS90, PD90, PS98a, Rea95, TAMSW90]. -dimensional [AC92, CHM91, KW90, RWL^{+}98, RTLM92, Scr95, Sin94, CW92]. -dimensions [HKA^{+}94]. -fermion [Pas96]. -fold [MCV98]. ’-Hubbard [FMH97]. -level [MOSM96]. -matrices [Gra92, HSSB98]. -matrix [SMB^{+}93, SHNS98, BEN95, BN95, HB90, Sch93]. -matrix-Floquet [CGHB98]. -pair [SJPW96]. -plane [DPD96b, DPD96d]. -point [Dik97, FT95]. -potentials [Mon91]. -quanta [GLF97]. -range [Str93b]. -splines [BDL92]. -state [ÇYG99, CBB99, CHM91, CK99]. -subshells [GF96]. -transform [RBB93].

0 [Art93, Art95a, CT96a, Cra98, Cyb90, Eas93, Fin91, Fly90, Hoc92, Par99, Rai91, Ryb98, Tay99, Top96, Tru93a]. 0-19-853443-4 [Eas93]. 0-19-856256-X [Art93]. 0-387-51906-8 [Fin91]. 0-387-94512-1 [CT96a]. 0-387-94630-6 [Ryb98]. 0-471-54884-7 [Top96]. 0-750-30435-9 [Par99].

0-7503-0277-1 [Art95a]. 0-7923-0549-3 [Rai91]. 0-7923-9798-3 [Cra98]. 0-8493-2016-X [Tay99]. 0-9518236-0-4 [Tru93a].

1 [Art95a, Bro92, CT96a, Deb92, IMES90, MNPP99]. 1.0 [AB97, GW94, KL99a]. 1.0.1 [Zha98b]. 1.00 [ABB^{+}96]. 1.02 [SJPW96]. 1.1 [JL93]. 1.2 [Kal93]. 1.42 [JPS^{+}99]. 100 [J̈K97a]. 101 [JY97a]. 109 [Zha98a]. 10Q [WCL90]. 116 [Fij00]. 120 [Wei11]. 13-17 [Cyb90]. 148-309 [RBG99]. 16-bit [MES90b]. 1990 [Whi93]. 1992 [Reb93, Tru93a]. 1PH [CC97a]. 1r [Mon91].

2 [BLGT94, MOSM93a, Mon99, MNPP99, Ric91, TRSM98, WENG95]. 2.0 [AIS^{+}97, AMM^{+}94, BBL^{+}97, BFK97, KS92b, Pan93, ST95]. 2.01 [JRWWW92]. 2.1 [AA98a, JW91, PL90]. 2.2 [AAKP98]. 230 [Wei11]. 2nd [Fle92, Fly93].

3 [Art96, Bis91, Bon93, Bro92, Cra98, Cyb90, Deb92, Fin90, Fle92, Fly90, Fly93, Gie90, Hoc92, Hop97, Mon99, Oln98, Rai91, Reb93, Tru93b, WW93, Weg91, Whi93]. 3-540-50339-0 [Cyb90]. 3-540-50449-4 [Fin90]. 3-540-51109-1 [Deb92]. 3-540-51541-0 [Fly90]. 3-540-51906-8 [Fin91]. 3-540-51933-5 [Gie90]. 3-540-52380-4 [Bis91]. 3-540-53058-4 [Fle92]. 3-540-53137-8 [Weg91]. 3-540-53418-0 [Hoc92]. 3-540-53571-3 [Bon93]. 3-540-53601-9 [Fle92]. 3-540-53782-1 [Bro92]. 3-540-54369-4 [Tru93b]. 3-540-54960-9 [Whi93]. 3-540-55722-9 [Fly93]. 3-540-55997-3 [Reb93]. 3-540-60214-3 [Art96]. 3-540-60529-0 [Hop97]. 3-540-60530-4 [Hop97]. 3-540-62649-2 [Mon99]. 3-540-62743-X [Oln98]. 3-7 [Whi93]. 3.0 [ABH99, PDL98a, PRL94, TMR94]. 3d [Bai91]. 3L [CA92, CA92]. 3nj [FLV99]. 3nj-coefficients [FLV99].

4 [Bis91, Eas93, Fin90, Hop97, hLTM^{+}92, MSW92b, Tru93a, Tru93b]. 4.0 [JWW94, MNPP99]. 4.1 [Gra95]. 4fan [BBL^{+}97].

5 [Gie90]. 5.1 [MWA^{+}92].

6 [Ryb98]. 6.23 [Odo92]. 6.5 [HLL^{+}95, SHL^{+}95]. 64 [RLPR94].

7 [Lön99, Top96]. 7.0 [Pi92]. 79 [Jam96b].

8 [Fin91, Weg91]. 83 [BCOY95]. 85 [Ham95a, Stu95b]. 87 [Sul96].

9 [Fle92, Fly93, Par99, Whi93]. 90 [CSC^{+}97, DLLR96, DNS98, Hen95, KK97, MKS^{+}96, Sat97a, SKM94]. 92 [VRS^{+}99]. 93 [ABE^{+}94, DBME96].

A-particle [NC96]. AA [ST95]. Ab-initio [DKT96, PvBMN99, OE99].

ABCRATE [GLAT98]. ability [dSTC99, dMP99]. abrupt [Lam96a, Lam96b]. ABSFAC [CR98]. Absolute [Cap93]. absorption [BKMM93, BM92b, CR98, GDRR94, PGK90b]. abstract [MKS^{+}96, SKM94]. Academic [Cra98]. accelerated [KS96, WCJ92]. Acceleration [HJ97, Ryc90a, Swe91, JMSW99, Ole96, Wen91]. Accelerator [Mes90a, Boy95, Gib92, MTH^{+}99]. acceptability [DL98]. according [ABE^{+}94]. accretion [GA90, RHFR95, SRR95]. accumulation [LHL95, RdMP99]. accuracy [AAKP95b, AAKP95a, AAKP98, CZ92, IIM92, Meg94, MR95, PSV99, UKA97]. Accurate [KSA^{+}91, SM91a, SFVW95, TBS98, BPG97, BM95, Fuj94a, GH91b, MN97a, Mon91, Moo94, Sim97, YB99]. ACE [SMB^{+}93]. achieving [WF90]. acidophila [HB99a]. acids [LZS95, LHSD95]. acoustic [PGJ91, SW91a, Tar99]. acquisition [Lan98]. action [Luo96]. actions [BEF^{+}99]. activation [BM99a]. activation-relaxation [BM99a]. active [Lin90, PWG93]. Ada [LN93]. Adaptation [GPSM95, CP99, RG92]. adapted [Ská90]. adaption [NRG90]. Adaptive [Bal95, Dik94, DF95, Ohl99, DD95, FCS99, JK99, KSZ97, Kre97, Wuc95, YM92, Zie98, Zie99]. adaptive-grid [YM92]. Addendum [DBME96]. address [Gen98]. Adhara [SAJ^{+}97b]. adiabatic [AAS95, BvG97, Lim96, MTM99]. adiabaticity [LM91b]. admittance [HM91]. adsorbate [BWS94, SSPD92]. adsorption [FB99, FBGV99, PSR99]. advance [LK96]. Advanced [ZK99, BMW91, SRB94]. advances [AD91, Whi93]. advection [Bal95, XYNI96, XYI96, XYE99]. Advisory [Ano99o]. aerodynamic [WJ91]. aerodynamics [AD91, GL91]. Ag [BHS^{+}99]. aggregates [BWNR99]. aggregation [PFBdlN99]. aging [MBZM99b]. agreement [GGSR94]. aid [Fan90]. aided [Bra90, DNSP99, SK90, vHK99]. aimed [Kud99]. air [DATL99, KT91, Kas91b]. airfoils [CC91a]. Al [BBB99, Met96, BBB99]. Al/Ga [BBB99]. algebra [AP98, BLS96, BS99a, Can94a, Can94b, CTDdM97, CTDdM98, CDD^{+}96, CSS94b, CC98, FGR98, JL93, JOP98, Rob97, RF95, SS93a, SMH98, VV99, Wyb94]. Algebraic [Lio97, SGT90, Stu95a, CY95, CdG96, KBD90, MBD91, OMC93, SSW94, Tak92, WW93, Eas93, Stu95b]. algebras [Abl98, BTC99b, BdM93]. Algorithm [Fer99, Sin90, Ale93, Alf99, AL94, AAS91, AAD92a, AK91, AK99, AL93, BP90a, BM99b, BMMMMR99, BTC99b, Ber96, BSSH99, BC91a, BCOY94, BCOY95, BMM97, BE90, Cal96, CNPT91, Cat95, CNO98, CD91, CD92, Cle93, DL90, DB93, EK94, FGM98, Fra92, FJ95a, GH91a, GC95, Gra98, GHE99, HR91, Iri96, JBH99, Jor95, KS96, Kar91, KK99, Kno90, Kob94, KS98, KNK95, KK96b, Krö96, Krö99, Kru96, LS93, Luč95, LBDL99, MP95, Min91, ME91, MP90b, MB94, NKV94, Ohl93, OM93, PJ98, PS98b, Pit98, Por98, Rho92, SFM^{+}96, SP93b, SVBD92, SF94b, SL97, Sto98, SZ99, SKST91, TG96, TD99, TT99, VD94, WCL90, Wei93, WS95b, WCJ92, XYIT97, XZ91, Zei97, Zha98c, ZX91]. Algorithmic [HGCM98]. Algorithms [BP90b, CTJ^{+}95, PT92, SS93a, AAD92b, BSKO96, BK92, BI90, CGM99, CS99b, Ess95, Far98, FKMR96, GH91c, Hoc92, Hop97, KSZ97, KMz91, KBL98, LK96, Mid92, MDSF^{+}99, MTH^{+}99, MPB91, NVK94, Pai93, PJ99, SMR99, Sci94, Sey95, Smi92a, TF98, ZL99]. Allen [Rai91]. allowed [Glü97, Joó97]. alloy [Smi92a, VBS96]. Alloys [Cra98, BdT97, PFVC99]. alpha [DB93, GDGR97]. alpha-particle [DB93]. Altarelli [KKK95, KL92]. alternative [RGLV^{+}99]. aluminum [FBB99]. AMBER [PCC^{+}95]. ambiguities [FS99a]. American [Tru93a]. amorphous [Bin99]. amplitude [BSKO96, NK96, Wan93]. amplitudes [BA94, Kah91, KS93, KBD90, KL93, MBD91, Nor94, SL94b, Tan90, TKS91, Wes93]. AMYR [TRSM98]. ANACAL [LMS^{+}93]. analyses [KI99]. Analysing [QBP98, BBB^{+}98]. Analysis [Bha91, Jan99, KDH^{+}98, LBB91, PTMUJdFM99, Rho97, VG99, VdN92, YN91, Zlo90, AI90a, AI90b, AB90, Aya92, BR97b, BHL92, CNO98, Chi92, Chr98, CHC91, CI98, CT96b, CGR93, CB99, Deg99, DSD91, DGKS99, DdMdO^{+}99, Eas91a, FCS99, GPSW95, GW98, Hea93, Hea96, Hel99, HIK90, Hof99a, HHC90, HB99b, HPS92, IKH^{+}99, JWM^{+}96, KKK91, Kas91b, Ker90a, KKKS95, KLS91, Leh99, LDOO95, LMS94, LB99, LMS^{+}93, LEBB99, Luc94b, LMW94, MRPC94, Mal97, MM98, Mer90, POA92, PCSD99, PCC^{+}95, Pop96, QTR99, Ren92, RCVJ99, RTLM92, SM93, Sch91a, SSH94, Siv91, SAY93, TCE99b, Tay99, WENG95, You99, ZA99, dAASZ97, vHK99]. Analytic [BS99a, BM92a, JBSZ97, AFKT97, BBL^{+}97, CY95]. Analytical [Kal94, Bes98, CTvB95, HD92, HE91, Iri96, MF95a, Str91b, Var97, VD94]. analyzing [UP90]. Anderson [JBB99, SMR99]. Angle [VN98, AMM^{+}94, CMNP95, CHUS98, FR96, HE91, LC90b, MEA^{+}90, WENG95]. angle-resolved [CHUS98]. angles [JRWWW92, JPRW^{+}97]. angular [Bur98, FS92, Fri97, FVGF98, FGH91, HF91, HGF91, Koi92, KK95, KK97, Mar91, Pot98, SS90b, Tak92, Wei99]. angular-gridded [Mar91]. angular-momentum [FS92]. anharmonic [Kal94, RBB98, Zlo90]. animals [TT99]. animations [Coh92]. anion [RFMM99]. anisotropic [CHMG92]. annealing [AK91, ST99, YBS91, FH97]. annihilation [DCM99, Odo92]. Announcement [Ano94c, Ano94a, Ano94b]. Announcement-processing [Ano94c]. ansatz [BK93]. antenna [HB99a]. antennas [Bha91, DP91, Nak91]. Anthony [Hoc92]. anti [BV95, WA94]. anti-bound [BV95]. anti-quark [WA94]. antiphase [PFVC99]. antiproton [ABE^{+}94]. antiproton-proton [ABE^{+}94]. Antisymmetry [BBCM93]. any [Kan95]. anytime [Chr99]. anywhere [Chr99]. APE [LMM^{+}97]. APE100 [BST95]. APEmille [Ano98-27]. aperture [MC99]. apertures [HM91]. Appleseed [DDK99]. Application [AAD92a, BN92, Ein90, IKH^{+}99, Jaq90, Lee90a, MMS93, PS98a, PVR96, PVRR97, RDF^{+}97, SZS95, Ver99, ZL99, BJL^{+}96, BH92, CSS94b, DBGW99, GHM93, Hea91, JWM^{+}96, KN95a, Lin90, MTH^{+}99, NBJK91, Ple96, RBB98, SS90a, SLGB96, SVBD92, Sim93, TKS91, Zha91, Sul96]. Application-driven [MMS93]. Applications [Cab90, CS98a, GPM95, PS97, d’I98, BJL^{+}96, Bar98, BG99, DH97a, FLV99, Hew94, HAC^{+}99, Kli98, MTR94, Nad90a, OPP96, Scr95, UTI94, VS95]. Applied [Tru93b, Art96, BGG91, FBGV99, KTT98, KM95, KH95, KV90, MDMR99, NL91, PHN99, SZG99, VW95, VVGI99]. Applying [Bes98, PCC^{+}95]. approach [Abl98, AAS95, Bra98, BVVJA99, Bru95, CMM92, CGM^{+}96, CTvB95, CD92, CPCF99, CI98, CKW98, DVH90, DATL99, FB99, FBdR^{+}98, Flo97, Frü95, Gil99, HM91, HKA^{+}94, IDW99, Jac91, LM95, MAB^{+}94, MBC^{+}99b, Nak91, Nak97a, Nak97b, OPY92, OMC93, Pas96, PI95, RF94, SV94, SW99, TEJJ93, USW99, WZKH98a, WZKH98b, Wei99, Win91, Won90, ZS91]. approaches [HE99, Kut99, Lag99, RVCF99]. Approx [CT96a]. approximants [HZQ93, Moo93, RBB93, RBB96, RBB98, Sty90, VBO97]. approximate [PS98a]. Approximating [Gro90, HKA^{+}94, WZKH98a, WKH98, WZA^{+}98, ZKHG99]. approximation [BK92, CY99, Dem97, DS92, FZZ92, HvSA98b, HKSV90, Kup99, MLB93, MMC95, Sal91, SS90a, SLGB96, TON95, YB99]. approximations [AAKP95b, AAKP95a, AAKP98, KNN91, MD97, NE93, PT92, Riz99, WC90]. arbitrary [BMM97, Cat95, CC91b, DP91, DGK97, IWWY92, JL93, LVL90, MP95, Wol98b]. architecture [Ano98-27, DHK^{+}98, Len98, MM98]. architectures [GG96, Hum90a, Ker90a, Ols96, RFH^{+}95, SG91]. arcs [Kar92]. argon [VS95]. argument [GS97, KRVZ98, YM97, dT93]. arguments [SG98b]. Ariadne [Lön92]. arising [Abd90, BKMM93, FJP90]. arithmetic [Gri97]. arithmitic [Nie94]. arms [OPY92, Ohl93]. AROMA [IRS97]. array [BL99b, CKS91, Fan90, Goe93, HP97b, MPRS97, SY95]. arrays [BS91b, BF90b, CP91, DP91, RBP^{+}90, RBS92, RBS94, Yeu91]. articles [Ano96-31]. artifacts [Eva90, RF96]. Artificial [LLF97, LPR92, PRL94, VV99]. arts [KBD90]. ASAD [CBW97]. ASAP [MRPC94]. ASPECT [SRB94]. aspects [BJW90, CGP94, Gup92, HSSB98, KV90, MKWH91, RHFR95]. Assembler [Ham97]. assembly [BBE^{+}98, RJS99, WW90]. assess [MR95, Sor98]. assessment [DPP^{+}99, Hel99, MC95]. assisted [JWM^{+}96]. associated [vdB93]. associations [ISS91, TOSM91, TRSM98]. ASTRID [Mer90]. Astronomico [MRPC94]. astrophysical [Cal96, CDM99, HS95, MRPC94]. asymmetric [LL99b, MB94]. asymmetries [MPN^{+}93b]. asymmetry [AHH94, LL99b]. Asymptotic [Fuk94, BN95, GGH91, GPSM95, Hea91, Szm95, YN94, vHKH97]. ASYROT [Zau94]. ATENSOR [IK96]. ATF [JMR^{+}93]. ATLAS [Ano98a, BJM^{+}98, Kno98]. atmosphere [KMT^{+}99]. atmospheric [AR95, CBW97]. Atom [NOC98, AZR96, AM94, BS91a, BM99c, BN94, Bur98, CM92b, DM90, FS91, FB94, GLAT98, Hor91, HKA^{+}94, LGZL98, Lem96, MF95a, MF95b, Man94, MD97, MÖ93, Pop98, Szm95]. atom-diatom [BS91a, GLAT98]. atom-surface [AM94, HKA^{+}94, Man94]. atom/molecule [Lem96]. atom/molecule-surface [Lem96]. atomic [ACA99, BSKO96, BEN95, BS99a, BKNS97, BBS91, CSS94b, DKT96, Fro91c, Fro91d, FSMVM93, FJ94, FS99b, Gal99, Gra94, GM92, Han90, JPF96, KWS^{+}96, Lim91a, MHN99, MSP^{+}98, PWG93, PFG96, Pot98, Pul97, Sal91, SSPD92, SMV90, SAP^{+}95, SMB^{+}93, SF94a, SAZ90, Ste97, SS94b, TCE99b, WM99, YAD99, Zat96, Zei97]. atomic-structure [Fro91c, Fro91d, KWS^{+}96]. atomic-subshell [Ste97]. atomistic [Lil93]. atoms [BMSW91, BWS94, CB99, GFS99, KR91, LMW94, Mit96, MOSM96, MSW92b, MSW92a, Pop96, SM93, SP93a, SGB99, SR98, ZFZ92]. attenuated [KCK99]. attraction [HS93b]. attractor [dMP99]. Auger [CS98b, CHUS98, HUCS98]. Auger-electron [CS98b, CHUS98, HUCS98]. augmented [BSST90, KWS^{+}96]. Austria [Reb93]. Auteur [Ano98b]. Author [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91a, Ano91b, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c, Ano93d, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97i, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano94m]. authors [Ano91h, Ano94j]. autocorrelation [NVFNP93]. autoindexing [Zlo95]. autoionization [FB93]. automata [BB92, BC99c, CPM99, Gie90, JMK99, dMP99]. Automated [HPV99, BFK98, CD91, CD92, Hum90c, PD96, Pop96]. Automatic [LMW94, MF95a, SL94b, Wan93, AFKT97, NE93, PJSFM92, PJSF93, Sem98]. automatical [NK96]. automaton [KW99, SC99b]. AUTOX [Zlo95]. available [PB93]. Average [Kle92, MLB93, HHT95]. Average-case [Kle92]. averaged [MGMR97, WP96, Zyb99]. averages [dO99]. Averaging [Kot92, Sch93, NSE95]. avoiding [BS90]. axial [MR96]. axially [GCP95, RGL97]. AXIOM [Sei94]. axis [MCV98, SLGB96, YB99]. Axisymmetric [LBR92, Boy95, Cam95, DMM^{+}97, KK91, KCH92]. Axodraw [Ver94]. azimuthal [LM91b].

B [Bis91, Fin90, Fin91, Jon90, PLC91, Reb93, Weg91, JN94, JWM^{+}96, Lan98, Pös98, SKG91]. B-spline [Pös98, SKG91]. BaBar [Ged98]. back [Frü93]. backbones [SFM^{+}96]. background [WA94]. backpropagation [GC95]. backward [MPN^{+}93b]. bad [Lan99b]. balance [BAAD^{+}97, MSJ99, ZYQ94]. balancer [BF90c]. balancing [AL94, FH97, MSW92a]. Ballistic [PSR99, DCM99, GP95]. Band [RHGDM99, BPMW95, CY93, Lam96a, MNS99, WP98, Yeu91]. band-structure [Yeu91]. BANDAS [MNS99]. bands [SYM98]. Bank [MMPT94]. BARRIER [GRM^{+}99]. barriers [BBB99, FJ95b, GRM^{+}99, JBB99]. bars [GC95]. baryon [GW95]. based [AK91, Boc92, CHC91, CT96b, DB93, Far98, GLV99, GP99, HP97b, HK96a, HK96b, HK97b, HHR98, JD91b, MLM96, MB91, MTH^{+}99, MKPGS97, NVK94, PBDZ94, RdC93, Ren92, SdCR97, SCS^{+}96, SL97, VES99, Wee92, YY99, You99, FH97]. BASES [Kaw95]. BASES/SPRING [Kaw95]. Basic [Krö90]. Basis [Bil91, All93, BDL92, CL96, DD97a, DD97b, EHK^{+}92, FMSdC99, GJ99b, HP97a, HT93, HHR98, LPC94, Lem96, LM91b, LL96, MR96, Pac91, Sor98, SF93, Tau95, TZC96, Wil90]. Bavaria [Cyb90]. Bayesian [DGKS99, Frü95, JY97b, SW99, JY97a]. Baym [KKY99]. BBBREM [KB94]. Bead [VVGI99]. beam [Aut99, FWH95, JTAW97a, LB91, Ohl97, Pes93, SAP^{+}95, SRFN92, SJ91]. beam-plasma [LB91]. beams [GP95, JTAW97b]. beat [Gib92]. beat-wave [Gib92]. bed [CMSK99]. beginning [Ano94r]. Behavior [BR97a, Bai91, KS93, Poz91]. Behaviorly [Sol99]. Behaviour [LVC^{+}94, Alb99, BMeS99, ITW97]. Belfast [BEN95]. Belle [KIMS98]. benchmark [MPSvG94]. benchmarking [SMR99]. benefits [Krö90]. bent [SvNF92]. Berandungen [Weg91]. Berlin [Art96, Bis91, Bon93, Bro92, CT96a, Cyb90, Deb92, Fin90, Fin91, Fly90, Fly93, Gie90, Hoc92, Hop97, Jon90, Reb93, Ryb98, Tru93b, Weg91, Whi93]. Bernd [Bro92]. Bertero [Par99]. Bessel [SL95, VRS^{+}99, CRS90, KRVZ98, Lem97, RdC93, SdCR97, SG99, VRS^{+}95, WC90, YM97, dT93]. best [BPP98]. BeTe [RHGDM99]. Bethe [BK93, WA94]. between [BCD^{+}93, FLV99, GGSR94, JAT97, MEA^{+}90, Pi92]. Beyond [RVCF99, TG96, BPK95, BBG94, BBJ91, Ohl99, Pap97]. BGK [HCO99]. Bhabha [AMM^{+}94, CMNP95, CC97a, FR96, JRWWW92, JPRW^{+}97, KB94, MPN93a]. BHAGEN [CC97a]. BHAGEN-1PH [CC97a]. BHAGENE3 [FR96]. BHLUMI [JRWWW92, JPRW^{+}97]. bias [Ber92a]. bias-corrected [Ber92a]. bicubic [LBR92]. Bidaux [Gie90]. bidisperse [HSW99]. Biedenharn [LC90b]. Bifurcation [Win91]. bilayers [LDOO95]. billiards [Has99]. Binary [MSD93, FLV99, LH92, Smi92a]. Binder [Tru93b]. binding [CGM^{+}96, IOM95, DCR99]. bioelectromagnetics [Isk91]. biological [GLV99, MP99]. biologically [dSTC99]. biology [NHK^{+}95]. biomolecules [TSJ91]. biorthogonal [FMSdC99]. biorthogonal-basis [FMSdC99]. Birkhoff [BGT95, Kal93, Rut98]. Bistable [GM99]. bit [DH97b, MES90b]. black [SRR95]. bleaching [MOSM96]. blende [SMV90]. Blind [NK99]. Block [GH91b, DHL93, MCV98, MR95, SUYH99]. board [Ano93m, Ano90n, Ano90o, Ano90p, Ano91p, Ano91q, Ano91r, Ano91s, Ano91t, Ano91u, Ano91v, Ano92l, Ano92m, Ano92n, Ano92o, Ano92p, Ano93i, Ano93j, Ano93k, Ano93l, Ano94s, Ano94t, Ano94u, Ano94v, Ano94w, Ano94x, Ano95q, Ano95r, Ano95s, Ano95t, Ano95u, Ano95v, Ano95w, Ano95x, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano96x, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s]. Boccacci [Par99]. Boccara [Gie90]. bodies [Gri97, Kud99]. Body [EABH95, AAS91, ACA99, BMSW91, BADP96, BAAD^{+}97, BS95, CPCF99, Deg99, IMES90, Lia95, LBDL99, LMW94, MSW92b, MSW92a, NVK93, PHN99, Rit92, RAD99b, SGB99, SY95, Sta96, VS97, Zha94]. Body-fitted [EABH95]. Bogoliubov [PVR97]. Boltzmann [Bas99a, IBR98, KKH^{+}98, LLW98, MP90a, PW99]. bombardment [BBG94]. bond [ÇYG99, SR91, WK90]. Bonding [TSL^{+}99]. Book [Art93, Art95a, Art96, Bis91, Bon93, Bro92, CT96a, Cra98, Cyb90, Deb92, Eas93, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Hop97, Jon90, Mon99, Oln98, Par99, Rai91, Reb93, Ryb98, Tay99, Top96, Tru93a, Tru93b, Weg91, Whi93, Lan99a]. bookkeeping [AM90]. bootstrap [MR95, WP96]. Boris [MV97]. Boris/DADI [MV97]. Born [BS99a, TON95]. Bose [Gau93, KR99]. boson [DKS98, Kni90, xLzHjW95, HHCS95]. bosonic [BS96]. bosons [AAH98, BM99b, LSZ92]. bottom [CMSK99]. Bounce [MGMR97, WP96]. Bounce-averaged [MGMR97]. bound [BKWM91, BKMM93, BV95, Du93b, HHR98, Hut94, IL91, Jan90, Lem97, LHL98, Pap92a, WA94]. bound-free [Jan90]. bound-state [BKWM91, HHR98, Hut94]. boundaries [DdMdO^{+}99, MS91a]. Boundary [RZE99, AB99, AE99, CZ92, DWL91, Eva90, Fuj94a, LMF97, Lem97, NSE95, PD90, RF96, ŠČ91]. bounded [Mac91, Ris96, SV94]. bounds [FR90]. box [Kru96, WCL90]. box-counting [Kru96]. brackets [Hol96]. Brandt [Fly90, Fly93]. brass [CMC^{+}99]. Breakdown [Riz99, KK91]. breaking [KTT98]. Breit [HGF91]. Bremsstrahlung [GJ97, JP92, MJ90]. bridge [KL99b, RL98]. Bristol [Tay99]. brittle [BA95]. broad [MH99]. broadening [MKWH91]. broken [Has99, KK96a]. Broschiert [Weg91]. Brown [EHK^{+}92]. Brownian [DMLM91, MBW^{+}95, FR90, PFBdlN99]. brushes [Krö99]. brute [MK96]. brute-force [MK96]. bubble [Avd96, MKL99]. building [For91]. bulk [MCV^{+}90, Mar97, MDMR99]. bulk- [MDMR99]. bulk-synchronous [Mar97]. bundling [FKMR96]. Burgers’ [BP93, WZKH98a, Fog99]. Burke [Ano94-33]. Burkitt [Hoc92]. bursting [TVH^{+}99]. butane [RP91].

C [Can94a, Fle92, Rai91, Reb93, Tay99, Wyb94, BDKM91, BM91, Can94b, CSC^{+}97, GLF97, Hop97, Lön94, MPRS97, NCC^{+}96, Oln98, SB97, Ryb98]. C-code [GLF97]. C-language [BDKM91, BM91]. C90 [MSTK93]. CAD [DP91]. Calc [MBD91]. calculate [AAS95, BKMM93, Bar93, CY99, FPV94, FPV95, Kag92, KS92a, Kud99, LH92, LPSM92, Mar91, Nor94, RdC93, SP93a, STR91a, SAZ90]. calculated [JF92, JF99]. Calculating [Kah91, NvdHPG97, Pap92a, Pap92b, Pöt99, TR93, AAS91, BPMW95, BFK97, Bur98, CW92, CC97b, EHKK93, FWH95, HFD90, IRV95, Kno90, LP93, Lim91a, xLzSlLzH94, LC96a, Mal90, MS95a, MIO96, Rho92, TSK^{+}97, WP98]. Calculation [BLS96, Bra98, FPV97, GRM^{+}99, GDGR97, Gra92, GD93, JGP93, KL93, Lee92, LEBB99, MMC95, NOC98, OC98, PRH90, RLPR91, RLER97, YYK^{+}92, vVF96, AIS^{+}97, ABB^{+}96, AMS96, Bai91, BS91a, BKDMM91, BBL^{+}97, Ber92b, BdM93, BWS94, BFK98, CHW91, CS98b, CR98, CGHB98, Cop91, FS92, FT92, Fra92, GLAT98, GH91b, GGH91, Gra94, Han90, HD92, HUCS98, HLT93, HT93, HHC90, HB90, Iri96, KWS^{+}96, KA95, Lam96a, Lam96b, Lin90, hLTM^{+}92, LC93b, MR96, MBD91, MF94, NP96, Neu99, NSE95, OM93, PZH^{+}96, PI95, PvBMN99, PMM98, RT98, RKV91, RGL^{+}98, Rou95, RSW92, Sag91a, SGB99, Sci94, SHL^{+}95, SS94b, TMR94, Tak92, Tan90, TKS91, TML93, THF95, TAMSW90, The94, TK91, TOSM91, TRSM98, Var97, VRS^{+}95, VRS^{+}99, Wan93, Wei99, RLPR94]. Calculational [MKWH91]. calculations [Abg91, AG99, AA98b, BPL97, BKNS97, BG99, BF90a, BDL92, CSP92, CvN99, CGP94, CSS94b, DKT96, DS96, Eng91, ERB90, FKFE99, Fer97, FF97b, FJ94, FS99b, GRRG95, GCP95, GSS97, GJ96, Glü97, GLSW95, GHGB97, HRK99, HPV99, HK97a, HP97a, HLL^{+}95, Jac94, Jaq90, JKSC99, JPF96, JF97, Joó97, KC98, KKH^{+}92, KV90, KR91, Lag92, LN93, LPR^{+}93, Luč95, MCMD91, MR99a, Mic94, MR99b, MTNA98, NL91, NND98, Pac91, PWG93, PFG96, PCC^{+}95, Pit97, Pit98, PB93, PVR96, Pot98, RM90, RDB^{+}90, SAP^{+}95, Sei91, Ská90, SP95b, Sta99, SH97, Ten97, TYJ92, ThLL^{+}93, VST94, WP96, WDB^{+}91, WF90, WF93, dCdJDvL99]. calculus [AP98, Kad96]. called [AHZ90]. calorimetry [SCS^{+}96]. Cambridge [Mon99]. Can [RFMM99, ACG^{+}99, Bin99]. canonical [HCVM99, MB94]. CAPMULT [CCM95]. capture [CMC94, CCM95, Gal99, MCCC99]. capture-gamma [CMC94]. capturing [OY99]. Car [CC99, WJ94, WJ95]. Carlo [Jun95, Tru93b, ALR99, Alb99, AL94, AAD92b, APN^{+}91, ADM^{+}92, AMM^{+}94, ABH^{+}94, ALR97, BN92, BvEW91, BW94, BB93b, BM99b, BS91c, BSW97, BC99a, BPK95, Ber92b, BMT94, BS90, BHM99, BBCM93, Blo92, BC99b, BKW97, BKS99, BRM99, BVVJA99, BLLM94, BG94, CMNP95, CC97a, CDG^{+}90, CSS94a, CMNP97, CHM91, CF95, CvN99, Cop91, DB93, DS90, ES92a, ES92b, Fan94, FR95, FF97a, FKKJ91, FR96, FBGV99, FCS99, GP95, GJ96, GE99, GGMS95, GW94, HP93, HMH99, HSP93, HCVM99, Heu90, HKL93, HN99, HK96a, HK96b, HK97b, IRS97, IER97, IMG99, JW90, JWW91, JKW91, JRWWW92, JP92, JWW94, JPRW^{+}97, JPS^{+}99, Jan95, JLFS91, KM91, KM98, KTT98, KB94, KP94, KF99, KR99, LMF97, LPS99, LM95, LR92a, MMH99, MSV92, MWA^{+}92, MBC99a, MC97]. Carlo [MPB96, MEA^{+}90, Mit96, MH96, MNP95, MNP96, MSJ99, NP96, Odo90, Odo92, Ohl99, Pai93, PI99, Pan93, PP99, PK99, Pes93, Poz91, RB99, RP95, RI97, RKV91, RH99, SGB99, SMV90, Sci91, Sel96, Sik99, Sin90, SWT91, SJPW96, SL99, SF98, Sti94, Swe91, TLK94, TFY95, VS95, VVGI99, VVIBA99, Wag92, WM99, WW90, Wan99b, Wan99a, WCK99, WS95b, WBB99, Zhe99, dCRSF99]. CARRE [MD96]. carrier [Poz91]. carry [LMS^{+}93]. CARS [KD95]. Cartesian [DD97a, DD97b, SS90b, Zie99]. cartographic [Ker90b]. cascade [BCSS99, Kas91a, WW93, Zha98a, Zha98b]. cascade-shower [Kas91a]. cascades [Gei97, Lön92]. case [BM99c, Goe93, JWM^{+}96, Kle92, RG92, Sou99, YKO^{+}94, ZN93, MMSM93]. cases [JWM^{+}96]. Cassinian [GCP95]. Cataloguing [FBdC98]. catalysts [AI90b]. catalyze [RFMM99]. catalyzed [RM90]. Catlow [Rai91]. cautionary [Art91]. CAV [NN93]. cavities [BCP91, JTAW97b, NN93]. cavity [KV90]. ccMBPT [MSW92b]. ccMBPT-4 [MSW92b]. CDF [Lam98]. Ce [PI95]. CELEST1D [VB92b]. Cell [FBdR^{+}98, BF90a, CC99, KN95a, KN95b, Lan92, LB95, MR99a, MP99, MN90, MN92, MT92b, MTH^{+}99, SZS95, Sul96, VS95, RTLM92]. cell-linked [MR99a]. cell-neighbor [MT92b]. Cellular [BB92, BC99c, CPM99, JMK99, KW99, RTLM92, Gie90]. center [AB99, HL92, HN97b, HN97c, RLPR91, RLPR94]. centers [PB96]. Central [VS97, HK97b, Lio97, SM91a]. centre [Dub91b, Han90, HD92, MAH98, RLER97]. centrifugal [Kor99]. centroid [MTM99]. century [RGLV^{+}99]. CERN [Rob98a]. certain [DH97a]. CESD97 [FG97]. CFD [FMH93]. CFP [LSZ92]. CFPs [HHCS95, jWzHxL95]. CFPSIB [xLzHjW95]. Chaban [LC93b]. chain [CP99]. chains [BHS^{+}99, CBG99, FR90, Krö99, MPS99, NRG90, RG92]. challenge [Bin99, dGT99]. challenges [Lan98]. chamber [OPT99]. chambers [TG99]. change [CP92, CMSK99]. changes [SS90a, Tar99]. changing [KTM96]. channel [AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, BMOF94, CMSK99, HRK99, HKC98, Hut94, LLM91, PZ91, Szm95]. channel-like [HKC98]. chaos [PGJ91, SWG99]. Chaotic [BMeS99, ITW97, dMP99, GM99, GP99, LEBB99, TVH^{+}99]. character [IMG99]. Characterisation [HNS99, LDOO95]. characteristic [Wan99b, Yat99]. characteristics [BR95, BZO92, MY92]. characterization [MBC99a]. Charge [MV97, Alf99, FT99a, FBVdlN99, LL99b, Nak97b, Poz91, VB92a]. charge-carrier [Poz91]. charge-transfer [FT99a]. charged [ABB^{+}96, ALR97, BE90, BG94, KKK91, KN95a, KN95b, LH92, MEA^{+}90, Mor92, RVB99, SA91, VB96]. charged-particle [RVB99]. charges [Hni94, Hoe90]. CHEASE [LBS96]. Chebyshev [CG99, CGK^{+}93, Dun96, Mac92a]. checking [xLzSlLzH94]. Chemical [Tox99, Tru93a, VB91, Alb99, Aro96, BVVJA99, Cab90, CBW97, HLL^{+}95, Jan95, hLTM^{+}92, Pag94, PBL94, PC94c, SHL^{+}95, Tru94, ThLL^{+}93]. chemistry [Ano96k, CBW97, DD94, FT99b, LP99, Tru93a, VVA^{+}94]. CHEMSODE [Aro96]. CHEP97 [GUW98, Lan98]. CHEREN [CM98]. Cherenkov [CM98]. Cherwell [Tru93a]. Chichester [Top96]. chloranil [FT99a]. chlorine [RFMM99]. Choice [KH97, FCS99]. Cholesky [Wil90]. chromaticity [OF92]. chromodynamics [QAd^{+}95, Sug91]. chute [DW99]. CIP [Don99, IY95, YY99]. circe [Ohl97]. circle [MK91a]. circles [Nie94]. circuit [AG91]. circuits [CHC91]. circulant [MCV98]. circular [Bha91, Kar92]. CKM [JY97a, JY97b]. class [Lön94, LL96, RW97, RFH^{+}95, SPP^{+}98]. Classical [PGJ91, KBL98, Luč95, WRF91, CT96a]. classically [LEBB99]. classification [Ano96-31, CS91, KK96b, Krö96, RCVJ99]. clean [SSPD92]. Clebsch [Kae95]. CLEO [WGH^{+}98]. CLHEP [Lön94]. Clifford [Abl98]. Climate [MAB^{+}94]. CLOF [JN94]. close [BM91, BS95, Bra98, GPSM95, MK91b]. close-coupling [BS95, Bra98, GPSM95, MK91b]. closed [Kre99, RDB^{+}90]. closed-shell [RDB^{+}90]. closures [HRK95]. clouds [HFD90]. Cluster [SC99a, TT99, Bab92, BPL97, BHL92, CHM91, CS98b, DDK99, FBVdlN99, GDGR97, HUCS98, LMF97, LPR^{+}93, Min91, PCNO94, PCC94, PC94b, PC94c, PC94a, Pul97, Sto98, Var97]. clustering [DBGW99, Nak97a]. clusters [BL99a, CTJ^{+}95, RBG99, TVH^{+}99, TSL^{+}99, Zei97]. CM [Ric91]. CM-2 [Ric91]. CMPT [Str91b]. CMS [BBE^{+}98, CI98, Kru98, OPT99]. Coarse [HTK99]. coarsening [GST99a, GST99b]. coated [Kis91]. coaxial [FG95b]. coaxial-multiple-coil [FG95b]. Code [GRRG95, Sun97, AIS^{+}97, AG91, AS95, AZR96, AO97, AHH94, BBL^{+}99, BADP96, BAAD^{+}97, CDM99, CC99, CM92a, CHM91, CZ92, DMM^{+}97, DER90, DGK97, DS96, Eis98, Fij99b, Fij99c, FG95b, GRM^{+}99, GS98a, GS98b, GLF97, HIK90, JTAW97b, KB98, Kar92, KM91, KMT^{+}99, KCH92, KWS^{+}96, KKY99, xLzSlLzH94, LB99, LPSM92, LBS96, MCV^{+}90, MGMR97, MLM96, NVC96, PG94, PL99, PMJ93, RP95, RdC93, Rea95, RCL98, SP93a, SSPD92, SBK98, SdCR97, SS93b, SRFN92, SRB94, Str93a, SH97, TBA^{+}98, TYJ92, TGG97, VLG95, YM92, YK90, YB92, Zie98, Zie99]. codes [APaI99, ADS92, BEN95, Dec95, Gri97, JKSC99, JMR^{+}93, KRT95, Lan92, MV97, MTG98, Ols96, RBS92, SG99, WP96]. coefficients [BLS96, BPMW95, FVR92, FPV94, FPV95, FPV97, FLV99, FS92, Kae95, KW96, Kag92, Koi92, KK95, KK97, Lee92, Lim91b, Mar91, NRG90, RG92, SS90a, SYM98, TSK^{+}97, Wei99]. coevolution [Jac99]. coherent [BHY92, BDH99, Izv99]. cohesive [CMSK99]. coil [FG95b]. Cojets [Odo90, Odo92]. cold [GDGR97]. Collaborative [Bun98, GN98]. collaboratories [ASJ98]. collapse [BM95, DPRS90, Man93]. collection [GRS90]. Collective [SEZ99b, BGG91, HTI99, Ran93]. college [WY99]. collider [FKMR96, Ohl97]. colliders [AB97, CMNP97, FHI^{+}98, KM98, Pap99]. collinear [BS91a]. collision [BKDMM91, DS90, GM92, HL92, MSD93, Nun90, SMB^{+}93, SVBD92, Ten97, TK91, VS95]. collision-free [Nun90]. collisional [PC90, PdA99, SAP^{+}95]. collisionless [Tan95]. collisions [AS91, ABH99, Ask94, ABE^{+}94, BBS91, CM92b, DM90, Dub91b, FR96, FIK^{+}97, Gei97, GGMS95, GW94, Hor91, IRS97, JPS^{+}99, Jun95, Kno90, MF94, MÖ93, NOC98, Odo90, Odo92, Pan93, PLC91, RT98, RI97, ST95, Sal91, TP99, Zha98b, Zha98a]. collocation [HHR98]. colloidal [Hin95, PSR99]. colloids [MPS99]. color [CJ91, LHL98]. colour [HK97a, Lön92, MPB91, HK97a]. coloured [BRWN96, GV99]. combination [BTC99a]. combinations [BMT94, SS90a]. Combinatorial [BBMD99, MDSF^{+}99]. combine [PJ99]. combine-and-conquer [PJ99]. Combined [CSS94a, GRRG95, JMSW99, SZG99, VGK90, Var97]. Combining [SW98, GLF97]. Combustion [Jon90]. comet [GBWES97]. Comment [HS96, Met96, Nar92, Wyb94, Can94a]. Comments [JD91a, PG96]. committee [BAS99b]. Committees [Ano99o]. COMMON [CT90]. Commun [BCOY95, DBME96, Ham95a, Jam96b, J̈K97a, KK97, RLPR94, Stu95b, Sul96, VRS^{+}99, Zha98a, Nar92]. Communication [WY99]. Communications [Fij00, JY97a, Ano90g, Ano91h, Ano94j, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano97j, Ano97k, Ano96k]. community [CLM98]. commutator [Sei91]. Compact [Sch91b]. Comparative [HZQ93, Sci94, Hum90a, KMz91, VKSAN95]. comparing [CvN99].

Comparison [CSC^{+}97, FR90, JMR^{+}93, KRT95, MN92, OPP96, BCD^{+}93, BHP97, Ess95, FBGV99, FMH93, Goe99, Iqb95a, MV97, MDMR99, STJ90, Siv91, SKPP95, WA97, ZFZ92, dFZLS95]. compatible [MM98]. competing [Szo99]. competitive [RNGR99]. Complete [NK96, FG95a, Lee90a, PWG93]. completely [NRG90]. completeness [STI99]. Complex [SM91b, BPMW95, Chi90, FT99a, Gie90, GH91c, HAR91, HHGCM99, IRV95, JK99, KRVZ98, MS91a, ME91, Neu99, RNR^{+}99, SvdPS99, TH95, VFO94, YM97, ZAW94, dT93, vVF96]. Complex-time [SM91b]. complexes [PvBMN99]. complexity [BGCRRO99, Kle92]. compliance [Ask94]. complicated [CTE95, ECT93, Gri97]. component [ABE^{+}94]. composite [ZS91]. Compositional [BGCRRO99]. compound [Cap93]. compressibility [CC91a]. compressible [DSB91, GHM93, HKG90, KK91, STJ90, TBCH90, YY99, Zie99]. Compton [TLK94]. Comput [BCOY95, DBME96, Ham95a, Jam96b, J̈K97a, KK97, Nar92, RLPR94, Stu95b, Sul96, VRS^{+}99, Zha98a]. Computation [AS92, AAH98, BMOF92, BZO92, CS98b, CHUS98, Gra98, GL91, HUCS98, HEM95, IBR98, KK91, LVL90, Lob99, LHL98, MSH91, Str93b, TK91, AZM92, Bru91, CdG96, FVR92, FG95b, Gau93, Gro90, KCK99, Kli98, Kor99, KSA^{+}91, LC90a, LC92, Lob96, Mac92b, MOSM93b, MOSM96, Mon91, OA95, OA96, Sch92, Sha92, Shu94, SF94a, TBS98]. Computational [BJW90, Deu99, Di 95, Duc96, Fle92, HSSB98, IKTY99, Isk91, Lan99a, MC99, RNR^{+}99, Ryb98, Ver99, YAD99, AD91, Ano96k, AL93, CP92, CTvB95, DLS90, Hoc92, ITW97, JH91b, Kas91a, KH95, Lan99b, MMW91, MB91, NVK94, RGLV^{+}99, Tru94, Win91, ZX91, dGT99, Bon93, Fle92, Reb93]. computationally [Hof99a]. computations [BTC99a, Bra91, Fri97, KKS90, LL96, Ott90, Pag94, PM91, Sim97, SL99]. compute [BGT95, FSMVM93, KS96, KK95, KK97, Sat97a, Sat97b, Wei11]. computed [JMR^{+}93]. Computer [AP98, Ano90g, Ano91h, Ano94j, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano97j, Ano97k, AS99, BRWN96, BHY92, BVR95, Bra90, CTDdM97, CTDdM98, Coh92, DNSP99, FKFE99, Fau91, FGR98, FMSdC99, Fij00, FV99, HMR99, JY97a, JOP98, JWM^{+}96, LL99b, MMW96, MBD91, MP99, Nar96, Oln98, RGL97, RNGR99, RS99, Sal94, SK90, SMH98, VV99, WT99, vGHM^{+}95, vHK99, ABH99, AHH94, BKWM91, BKMM93, BLS96, Ber99, BdM93, BBJ95, Bin99, BS99a, CHW91, Chr98, CT96b, CSS94b, DG99, DBGW99, FS92, Fra92, GG96, GW98, Gri97, HK97a, HFD90, HE99, HLL^{+}95, HPS92, IYOY94, IEY^{+}96, Jan90, KBD90, Kud99, LC96a, hLTM^{+}92, MMS93, MSW92a, MNPP99, ME91, Nad90a, Ols96, OMC93, PMUV90, PCC^{+}95, PTU97, QAd^{+}95, Rit92, Rob97, Rou95, SSPD92]. computer [SS93a, SHL^{+}95, SH91, Swe91, TMR94, TSJ91, TR93, Tob99, TRSM98, Uka98, WLD95, Wie96, YM92, Ano96k, Fly90, Fly93, Fin90, Fin91, Rai91]. Computer-aided [Bra90, vHK99]. Computer-algebraic [MBD91, KBD90]. Computer-assisted [JWM^{+}96]. computer-based [CT96b].

Computer-simulation [AS99]. computers [AS95, BNSW92, CGM^{+}96, CDD^{+}96, CSP92, Dec95, EABH95, Fan90, GPSM95, GHGB97, Ham93a, Ham93b, Ham95b, IOM95, LMM^{+}97, MN92, MTNA98, NKV94, PRdM99, Sek92, SSHT91, Smi91, SHNS98, Tar99, Cyb90, Ham95a]. Computing [AN97, Din91, Du93a, Ern98, Eva92, Gal99, Gup92, Lam98, Mor98, Nie94, RU92, Rob98a, dM94, AHZ90, Bow99, BFK95, Bun98, BCP91, DDK99, DHK^{+}98, Eas93, Far98, FG91, FB93, Ged98, HS93a, HF91, HGF91, HMG92, JWF93, KIMS98, Kno98, KDH^{+}98, KRVZ98, Lag92, Lag99, Mar96, MPN93a, MPN^{+}93b, MNPP96, NHK^{+}95, RBS94, Van96, VD94, WW93, Wol98a, Zat96]. concentration [HMH99]. concept [Tar99]. concerning [MKS^{+}96]. Concurrent [BF90b, WW90, SSHT91]. condensation [JMSW99, KR99]. Condensed [Tru93b, AH99, Fin90, Fin91, RJ90, Ryc90b, RJ91a, RJ91b, Tox99, WRF91]. condensed-matter [Ryc90b]. condition [Fuj94a, HÓ99, NSE95]. conditional [Fan94]. Conditioning [DR90, BG98]. conditions [AE99, CZ92, Eva90, LMF97, Lem97, PD90, RF96, RZE99]. conductance [HP97b, Ver99]. conducting [BG95, PD90]. conduction [GFAALS98, Rea95, ŠČ91]. conductive [Gri97]. conductivity [GFAALS98]. confidence [Ber92b, BR97b]. Config [MSK98]. configurable [GLSW95]. Configuration [IST99, AZR96, AMS96, Fro91a, Fro91b, FL91, MSK98, PI95, SF93]. configuration-state [FL91, SF93]. configurational [LMS^{+}93]. configurations [CC91b, Koo93]. confined [BVVJA99, MBMM99, Pai93, WBB99]. conflict [MMPT94]. conformal [Fuj94b, LM91a, MS91b]. conformational [FCS99, WENG95]. conjugate [BI90, KS96, Nak97b, SG91]. conjugate-gradient [Nak97b]. connection [BBJ91, TR93]. connectivity [BMM97, JMK99]. conquer [PJ99]. conservation [KPD97, MV97, VPE94, VP96, VB92a]. conservative [KKMS99]. conserved [CZ99, Ito94]. conserving [Kro92]. Considerations [HDF97, MMS93, Rap91a, Rap91b, Rap93, SGB99]. Consistent [Kre99, AW94, BWS94, KNN91, PAC^{+}98, PVR96, SLGB96, TAMSW90, TK91, WF93, ZAW94]. consortium [Top96]. constant [BVVJA99, BC91a, FF97a, RARF96]. constant-pressure [BC91a, RARF96]. Constants [FG90, HLL^{+}95, Jan95, JWF93, Kan95, KH97, Kor99, LM91a]. constrained [CMV99]. constraint [ME91]. constraints [Kut99, MPB91]. construct [BP90a]. Constructing [XYNI96, XYI96, MSK98]. construction [FGM98, HZQ93, Koo93, Lil93]. constructive [Ren92]. contact [GY99]. contained [AI90a]. containing [CMC^{+}99]. Content [Ano97l]. Contents [Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano91m, Ano91n, Ano91o, Ano91i, Ano91j, Ano91k, Ano91l, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano93e, Ano93f, Ano93g, Ano93h, Ano94k, Ano94m, Ano94l, Ano94n, Ano94o, Ano94p, Ano94q, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p, Ano96q, Ano97q, Ano97m, Ano97n, Ano97o, Ano97p, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m]. contest [Don99]. context [Hof99a]. continued [RdC93, SdCR97]. continuity [ZYQ94]. continuous [BC99c, MH99, NCC^{+}96, SVBD92]. Continuum [BDL92, IBR98, PHT93, Ask94, BSKO96, GD93, MMC95, MKL99, PK91, TSJ91]. contours [MS91a]. contractions [BdM93]. contribute [ACG^{+}99]. contributors [Ano90r]. Control [LB95, GF98, IKTY99, Lan98, Mes90a, MG90, PDL98b, Rho97, RHGDM99, SR91, VW95]. controlled [CGdlR99, GRS90]. CONTUR [CT96b]. convection [HKG90, KM95, MY92, NS90, NN93, TBCH90, Wuc95]. convection-diffusion-reaction [KM95]. Convergence [JMSW99, DHL93, HR91, LB99, Ole96, Tau95, Wen91]. convergent [BS95]. converging [AK99, PTMUJdFM99]. conversion [Gei97]. convex [AN97, Dem97, Kud99]. convoluted [MNPP96]. convolution [LP93, VD94]. convolution-type [LP93]. CONVUS [LP93]. Cooley [Du90, JD91a]. cooperative [MXR98]. Coordinate [HLPT93, Bal95, DGK97, LM91b, MS91a, PVR97]. coordinates [DEAM91, Eas91a, FBdR^{+}98, KH97]. coordination [RSW92, TMR94]. copolymer [BHM99]. copolymers [AS99]. core [AZM92, BT92a, IMA94, Pac91]. Coriolis [GG96]. Coriolis-coupled [GG96]. corners [MS91b]. coronal [BPG97, DVH90, Goe90, PGK90a, PGK90b, RBP^{+}90, SMBV90]. correct [PC90]. corrected [Ber92a, DS92]. Correction [LHL95, Du90, GJ96, Glü97, Joó97, Luc94a, PC94a, SVBD92]. corrections [AIS^{+}97, ADM^{+}92, ABH^{+}94, ABB^{+}96, BvEW91, BW94, BBR^{+}90, FJKvO95, JWJK92, Kni90, KS92a, Sey95, Zha96]. corrector [AH91, Du93b, Kro92, YY99]. Correlated [HÓ99]. Correlation [HMH99, BZO92, CGdlR99, LC90b, MCMD91, NSE95, TB91]. correlation-angle [LC90b]. Correlations [RP91, HDF97, Kno90, MEA^{+}90]. correlators [PTMUJdFM99, RMMUJ99]. corresponding [PB93]. corrugated [MK91b, PD90]. cosine [BT91]. cosmic [ML92, Str93a]. COSMO [ML92]. cosmological [BNS^{+}95]. cosmology [Jak98]. Coulomb [Pap92a, Pap92b, AS92, BMOF94, Ber99, BS95, CT94b, FZZ92, HN97b, HN97c, Ran93, TON95, USW99, Zha94, ZFZ92]. Coulomb-like [Pap92a, Pap92b]. Coulombic [HD92, KdLNV93, PG96]. counting [CMC94, CM98, CCM95, Kru96, MCCC99, OA96, Sto98]. Coupled [BMOF94, DWL91, Hut94, AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, AH91, BDF^{+}97, CT94b, DHL93, FJP90, FJ91, FG95a, FG97, GG96, HRK99, HvSA98a, Her91, SWG99, SF93]. coupled-channel [AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, HRK99]. coupling [AB90, BS95, Bra98, BC91a, FLV99, FS92, Fri97, FVGF98, GPSM95, HM91, JWM^{+}96, KW96, Kag92, Kan95, Koi92, Kre99, MK91b, PCC94, Pur90, SS90b, Tak92, Wei99]. couplings [HRK99]. course [Lan99a]. courseware [WY99]. cover [Cyb90]. covered [SSPD92]. CP [Uka98, dAASZ97]. CP-PACS [Uka98]. CPC [Ano94r, Ano96r, Wei11]. crack [TMS^{+}99].

CRAY [BAAD^{+}97, MSTK93]. CRC [Tay99]. CRCWFN [CT94b]. creation [Krö99]. Critical [CVZ99, IDW99, vdB93, Bai91, Rie99, SC99a, Zhe99]. cross [BKDMM91, BKMM93, BM92a, BM92b, Cap93, CM92b, DM90, Gra95, HD96, KA95, LN93, LC96a, Mal90, MMC95, MJ90, MPN93a, MPN^{+}93b, NOC98, Pöt99, RT98, SG98a]. cross-section [BKMM93, LN93]. cross-sections [BKDMM91, BM92a, BM92b, CM92b, HD96]. crossing [ZN93]. Crossover [CZ99]. crosswind [ZmSjFyJ97, ZYQ94]. Crowley [Whi93]. crystal [BWS94, CGRV99, HMR99, Kot96, RP91, RNGR99, TEJJ93, WBS95, dM94]. crystalline [BSST90, Lil93, MPB96, TF98]. Crystallisation [RLF96]. crystallite [AI90b]. crystallites [Mic94]. crystallographic [Rei93]. crystals [All99, CR98, CDR91, DBGY99, GP95, HHH99, KF99, MSR96, OBRR92, PCNO94, PCC94, PC94b, PC94c, PC94a, PB96, RLP^{+}99, SYM98, TBA^{+}98, Zyb99]. Cubic [IY95, NY99, DER90, Eva90, GP95, MRC98, YA91, YIW^{+}91, UKA97]. Cubic-Interpolated [IY95]. cubic-polynomial [YA91, YIW^{+}91]. Cullum [Kal96]. curl [TSK^{+}97]. curl-free [TSK^{+}97]. current [ABB^{+}96, Ern98, LBB91, PD90, SS93b, WP96, ZYQ94]. currents [Eva92]. curvature [IJV92]. curve [HP97b, ZN93]. curves [AAS95, RdMP99]. curvilinear [Eas91a]. cusp [DJUM99]. Cutoff [TC97]. CXFTV2 [Dem97]. cycles [Kar91]. cyclic [BWNR99]. cycling [MDSF^{+}99]. cyclotron [CM92a, DJUM99, JTAW97a, JTAW97b]. cylinder [LL90, SG98b, Tau92, YB92]. cylinders [YK90]. cylindrical [JTAW97b, KB98, Lem97, QV95]. cylindrically [TDV96].

D [Art93, Fin90, Fin91, Fly90, Fly93, Top96, WW93, AB99, ADS92, BL99b, BT92a, BMPS99, ÇYG99, CHMG92, DJUM99, Eas91a, Fer97, Fij99b, Fij99c, FMIM93, FV99, HHH99, HKG90, IIM92, JBB99, JV99, JMR^{+}93, KB98, KW90, KKMS99, KR95b, Kud99, MD96, MMCH93, NS90, Ott90, PC99, PD90, PCSD99, PS98a, PG94, Rea95, RTLM92, RMMUJ99, SA99b, SPA99, SS94a, Sou99, SW91b, SVV99, TAMSW90, VBS96, WLD95, XZ91]. D-CELL [RTLM92]. D-FROTH [KW90]. D. [Mon99]. DADI [DHL93, MV97]. Dahmen [Fly90]. Damage [VS99]. Daniel [Top96]. DAQ [Ano98a, BJM^{+}98, WGH^{+}98]. Darwin [Hew94]. Data [BT92a, BT92b, Fri97, Pai93, You99, AA99, AI90a, All93, AA98b, BADP96, BAAD^{+}97, BR97b, BHL92, CS98b, CT96b, Coo99, CGR93, DBGW99, GKP93a, GKP93b, GC95, GRS90, Gut92, HKSV90, Kas91b, Lam96b, Lan98, LS93, MKS^{+}96, MM98, MPT98, MG93, PT92, PP94, Pod99, Rap93, RBS92, SAP^{+}95, SKM94, Smi92b, SF94a, SF94b, SG91, SA95, YJ92, ZFZ92]. Data-based [You99]. data-collection [GRS90]. data-domain [RBS92]. Data-parallel [BT92a, BT92b, Pai93, Rap93, SG91]. data-sharing [BADP96]. Database [Gli90, CBW97, PBDZ94]. Davidson [SF94c]. Davydov [LC93b]. Davydov-Chaban [LC93b]. dc [SSF92, HHV99]. DCS [Mal90]. de-convoluted [MNPP96]. Dealing [AB99]. dealloying [CMC^{+}99]. decay [GDGR97, JWDK93, JWJK92, Ste97, Wag92]. decays [BvEW91, Deg99, DKS98, GJ96, GJ97, Glü97, JW91, JKW91, Joó97, SA91]. decide [SW98]. decomposition [BCOY93, BCOY94, BCOY95, BMM97, CGM^{+}96, DBVS99, MAB^{+}94, PK98, SAJ^{+}97a, UP90, VN98, WJ94, WJ95, Wil90]. deconvoluted [MNPP99]. deconvolution [LBC^{+}93]. deconvolver [VD94]. decrease [ES92b, Tur99]. dedicated [PTU97, QAd^{+}95]. Deep [BKW97, AIS^{+}97, ADM^{+}92, ABB^{+}96, CSS94a, IER97, JP92, Ker90a]. deeply [Gra95]. deexcitation [Ste97]. Defect [HHGCM99, PB96]. Defect-freezing [HHGCM99]. defect-unbinding [HHGCM99]. defects [PCNO94, PCC94, PC94b, PC94c, PC94a]. Deformable [Bla94a, CC99, OPY92]. deformation [GCP95]. deformed [DD97a, DD97b, GRRG95, GCP95, RGL97]. degeneracies [dO99]. degrado [Sik99]. degrees [Kal93, SVV99]. delay [ABHR99, PMB99, ST98]. delay-differential [ABHR99]. delayed [MD97, Ohi99]. Delbrück [Kah91]. delivering [Chr99]. Delta [PP94, PZH^{+}96]. DENCOM [GRRG95]. DENISIS [SBK98]. denominator [RDMS95]. dense [FBH90, PR93, PG94]. densities [ACA99, AA98b, HHT95, Ito94, Sun97]. Density [BKNS97, BG99, TZC96, dCdJDvL99, BHP97, CP92, FS99b, GRRG95, GJ98, HLMM99, HKLM99, JBSZ97, J̈K97a, J̈K97b, KKM^{+}99, Leh99, MCMD91, MD97, PvBMN99, PB93, RF94, SS94b, TAMSW90, vGSB99]. Density-functional [BKNS97, FS99b, J̈K97a, J̈K97b, SS94b]. density-of-states [TAMSW90]. depacking [LC96b]. Dependence [SSK97]. dependencies [PP94]. dependent [AW94, BS91a, BKDMM91, BKMM93, BGG91, BK91, Bil91, BBS91, CK91a, CKW98, DEAM91, Eng91, Eva92, GH91a, GHE99, HKA^{+}94, ICS95, Jac91, Jan95, JH91a, KC98, KNN91, KR91, Lee90b, LM91b, Lin91, LPSM92, LD91, Mic91, MP90a, NBJK91, PLC91, PCC94, PSV99, RdMP99, Sch91a, SPT98, SM91b, SKST91, TK91, Wuc95, ZZGS91, ZAW94, vGSB99]. deposition [VG99]. derivation [HE94, Kre99, Wen91]. derivative [AP97, FJ95a, JD91b, JF99]. derivatives [CT94a, Deu99, JBSZ97, J̈K97a, J̈K97b]. derive [LK96]. deriving [Lim91b]. Dervieux [Jon90]. descendant [LM91a]. description [GKP93a, KR95a, MGMR97, RGL97, Kre99]. Design [Rap91a, Rap91b, Rap93, CDG^{+}90, CDD^{+}96, KDH^{+}98, Len98, MMS93, MES90b, MMSM93, Mor92, Siv91, KDH^{+}98]. designed [Sik99]. designing [GF98, Lön94]. designs [Jan99]. DESY [BF98, Ern98]. detailed [Lag99]. detected [BO98]. detection [ST98]. detector [Asa98, BCD^{+}93, CDG^{+}90, DL90, MS96, Rei93]. detectors [CNPT91, CGK^{+}93, KT91]. determinant [AW94, ZAW94]. determinants [FG95a, FG97]. Determination [HCVM99, HP97b, dlVM98, FFB99, GJ98, JOP98, Kar91, NE93, Sin90, VRS^{+}95, VRS^{+}99, Wie96]. determine [AO97, SRFN92, Ste97]. determining [KS98]. Deterministic [Lob96, CBB99, Pas96]. Developing [Van96, VV99]. Development [Lön99, MTR94, SKG91, BJL^{+}96, MXR98]. developmental [Jac99]. developments [BN94, Pag94, ST92, Fin90]. device [FKKJ91, Gus91, SSHT91, ZmSjFyJ97]. devices [BMW91, JLFS91, TSK^{+}97]. Dfl [Rai91]. DFT [BG98, GHGB97]. DFT-pseudopotential [GHGB97]. Dhmen [Fly93]. di-block [SUYH99]. diagonal [Kre99]. diagonalization [FMH97, JKSC99, Lia95, VES99, ZK99]. diagram [BL99b, BFK98, Lai98, OF92]. Diagrammatic [BMSW91, LMW94, MSW92b, MSW92a]. diagrams [Avd96, AFKT97, EHK^{+}92, FT92, KV99, LF92, Lev90, LMW94, Neu99, Ohl95, SLGB96, SGT90, Stu95a, Stu95b, Wan93, vO91]. dialectric [MT91]. diamond [DER90, KSSB99, SMV90]. diatom [BS91a, GLAT98]. diatomic [BKMM93, EHKK93, KLS96, Kor99, Mal90, Nor94, Sch91b]. diblock [BHM99]. dichotomic [BRWN96]. Dielectric [SP95a, DBGY99, IMG99, Kis91, LHL98, Nak91]. Dieter [Hoc92]. DIFCARS [KD95]. difference [AA94b, AA94a, AA98a, BS91b, CK91a, FT95, KKMS99, KSG91, Mic91, PT95, PS98a]. differences [BC99b]. differencing [AR95]. different [FR90, GGJG97, Iri96, Kre99, LMF97, NK99, STR92, STJ90]. Differential [SG98a, SVBD92, ABHR99, Aok97, AH97, Bil92, CHW91, CTvB95, DHL93, FJP90, FJ91, Har97, Hea93, Hea96, HGCM98, LK96, Mal90, Rod92, RW95, RSW96, SP93b, SP93c, Sei94, Sty90, YKM^{+}96]. differentiation [Dun96]. diffraction [CT97, CS98b, CHUS98, Fan90, HUCS98, KSSB99, MC99, SJ91]. diffractive [Jun95]. diffractometers [PTMUJdFM99]. diffuse [MS95a]. diffusing [FB99]. Diffusion [BC91b, BM99c, BKKS99, DMLM91, KM95, Les99, LBDL99, MBW^{+}95, MBC99a, MMGS93, MLM96, Pae91, SMV90, YK95, Zei94, ZmSjFyJ97]. diffusive [PFBdlN99]. Digital [Bro94, GFAALS98, MES90b, Bro92]. dihedral [WENG95]. dike [CvN99]. Dill [Luč95]. dilute [BRM99, dQRdS99]. dimension [Cat95, DdMdO^{+}99, HK96b, IY95]. dimensional [AN97, AC92, AG91, App99, Art95b, BS91b, BHS^{+}99, BPG97, BK92, CS99a, CRS90, CW92, CHM91, DSB91, DER90, DGK97, DCM99, DPRS90, FBH90, FG95b, Fuj94a, Goe93, Gra92, Gri97, GRS90, HSSB98, Heu90, IMA94, Jaq90, JN94, JTAW97b, Kaw95, KW90, KH97, KR95b, Kup99, LP93, LPC94, LM91a, Lee90b, LC93a, LP90, MC93, MOC91, MMTK94, NY99, NE93, NSE95, PC90, PI99, PL99, PFVC99, QV95, RLWW97, RWL^{+}98, RBP^{+}90, RTLM92, Rob97, SBK98, Scr95, SSHT91, Sim93, Sin94, SJ91, TBCH90, UKK98, Ver99, Wei93, XYNI96, YA91, YIW^{+}91, YB99, Zie99]. Dimensionality [Chi92, CTS99, Mac91]. dimensionally [BN92]. dimensions [BF90a, HPV99, HKA^{+}94, IYOY94, JL93, Kal94, Nar96, SKPP95, WZKH98a]. dimers [Sik99]. dioxide [OE99]. dipolar [HMR99, KN95b, MPS99, dGT99]. Dipole [Sal97, BL99a, GLF97, JWF93, KN95a, Lön92, STR91a]. DIPSI [ALR97]. Dirac [AZR96, CY99, Gau93, KS96, MN97a, NK93, PZH^{+}96, PI95, Sag91a, SM91a, SM93, SFVW95]. Direct [BRM99, Wei11, DH97a, HLL^{+}95, KA95, MK91a, STR91a, TP99, TGAW96, TDV96, ThLL^{+}93]. directed [HÓ99]. direction [KB94]. directional [Kor99]. directions [Ern98, Wol98a, Fin91]. Directory [Tru93a]. discharge [HHV99]. discharges [LC93a, VS95]. discontinuous [BSW97]. discovery [HN97a]. discrepancies [JHK97, vHKH97]. Discrepancy [HK96a, HK96b, HK97b]. Discrepancy-based [HK96a, HK96b, HK97b]. Discrete [AE99, GLV99, SGF99a, CGK^{+}93, Dem97, HLPT93, LM91b, Neu99, POA92, PW99, SL95, Zag99]. discrete-coordinate [LM91b]. discrete-lattice [Neu99]. discrete-velocity [PW99]. discretisation [Moo94]. discretization [Abd90, MSH91, RH99]. discriminant [Mal97]. disk [BF98]. diskette [Bis91, Top96]. disks [FB99, GA90, RHFR95, SRR95]. disorder [BWNR99, IMG99, SPA99]. Disordered [Cra98, Rie99, BN92, BM99a, BBMD99, BDH99, HHT95, HM99, MSTH92, SL94a, TCE99a, TCE99b, VES99, Wha94, ZK99, dVdRL93]. displacements [BDF^{+}97]. displacive [IMG99]. Display [CDdG^{+}98]. displaying [BBB^{+}98]. displays [dM94]. disposal [DPP^{+}99]. dissipative [BK91, RZE99, SEZ99b, SEZ99a, SM97]. dissociation [Lim96, TYJ92]. dissociative [BM99c]. distance [FLV99]. distorted [Eis98, MMC95]. distorted-wave [MMC95]. distortion [Kor99, PC90]. Distributed [HKA^{+}94, WZKH98a, ZKHG99, BF98, BCOY93, BF90c, CDD^{+}96, DHK^{+}98, KdLNV93, LBC91, LR92a, MTM92, MPT98, Pet92, Rap91b, RBS92, RFH^{+}95, SHNS98, TC93, WKH98, WZA^{+}98]. distributed-memory [KdLNV93, RBS92]. Distribution [Str93b, App99, BAAD^{+}97, BG95, Cha92, HKM98a, LL99c, POA92, SRFN92, YJ92, dM94, vVF96]. distributions [BKDMM91, BH90, GGSR94, HM92, HS99, HK97b, IWWY92, KKK91, Kle92, OC98, Pot98]. diurnal [AR95]. divergent [BRB92, Wen91]. divertor [AMS96]. divide [PJ99]. divide- [PJ99]. dividing [CK91b]. DIZET [BBR^{+}90]. DJANGO6 [CSS94a]. DM [Art96, Bon93, CT96a, Cyb90, Fin91, Fle92, Gie90, Hoc92, Reb93, Tru93b, Whi93]. DM64.00 [Weg91]. DM78.00 [Deb92]. DM88 [Jon90]. DM98.00 [Bro92, Fin90, Fly90, Fly93]. DNA [BGCRRO99]. DØ [Lam98]. document [Krö90, Won90]. documents [Bra90, Eic90]. does [dSB99]. doing [CLM98, Luč95]. Domain [GST99a, GST99b, McD90, SA99b, BCOY93, BCOY94, BCOY95, BMM97, GHE99, MAB^{+}94, MSH91, PFVC99, PK98, RBS92, SAJ^{+}97a]. Domany [Deb92]. dominant [You99]. Dominated [Mon99]. Dordrecht [Rai91]. dosemeters [RVB99]. dots [Ver99]. Double [Ber92a, DJUM99, Fuj94a, GRS90, MSFR96, Pae91, PL90]. down [IDW99]. Downs [Tru93a]. dragonfly [GL91]. Drawing [Ohl95, Lev90]. Drchal [Cra98]. Drift [PMB99, Art94, Art95b, BO98, Yat99]. drift-wave [Art94, Art95b]. drive [SS93b]. Driven [dlSG99, LL99a, MC95, Mar97, MMS93, Ple96, PGK90a, SPT98]. drop [MDMR99]. droplet [LMW96]. drops [DS99]. drum [KW99]. dry [MHHL99]. DSP [Tar99]. DTUJET [ABE^{+}94]. DTUJET-93 [ABE^{+}94]. Dual [BM99b, ABE^{+}94]. due [BDF^{+}97, Eva90, FWH95, RF96, Ste97, Tau95]. duration [TD99]. during [CR98, Pes93, TMG99]. dusty [SBK98]. DVR1D [HT93]. DVR3D [HLT93, THF95]. DYEFIC [OA95]. Dynamic [AH99, SAJ^{+}97a, AL94, BWNR99, BAAD^{+}97, BF90c, CC91a, CPCF99, FH97, HTI99, JH91b, Krö99, MDMR99, MSW92a, OGG91, SKG91, SA92, WENG95]. Dynamic-domain-decomposition [SAJ^{+}97a]. Dynamical [KKKK99, OBRR92, SSW94, BKMM93, DR90, DdMdO^{+}99, HEM95, Izv99, JL97, KNN91, RNR^{+}99, Sat97a, WZKH98b]. dynamically [BJW90, BJK92, Cat95, SWT91]. Dynamics [CR99, DMLM91, MBW^{+}95, Alf99, AF90, AM90, BJL^{+}96, BDKM91, BST95, BdT97, BSB99, Bas99a, BvdSvD95, BGG91, BMeS99, BBG94, BCSS99, BvG97, Bil91, BKNS97, BT92a, BC91a, BCOY93, BCOY94, BCOY95, BMM97, BF90b, BF90c, Bru95, BJK92, CGM^{+}96, CD90, CF95, CJ91, CPCF99, CP91, DW99, DNSP99, EK94, FHSO99, FKKJ91, FR90, GG96, GLF97, GY99, GJ99b, Gup92, HNR90, Hof99b, HLL^{+}95, IOM95, IK99, JAT97, JH91a, KdLNV93, KB99, KBL98, KLS91, Kud99, Kut99, LMM^{+}97, LDOO95, LHSD95, LBC91, Lil94, LMW96, Luo96, Mak91, MMW96, MC95, Mar97, MTM99, MPRS97, MKL99, MS95a, MHHL99, Min91, MT92b, MP90b, MPSvG94, NVK93, NKV94, NVK94, Nak97a, Nak97b, NND98, NKR99, Ohi99, OD99, OPP96, PCC^{+}95, PS98b, PFBdlN99, PK98, Rad99a, Rap91a, Rap91b, Rap93, RJS99]. dynamics [RF94, RZE99, Rob97, RS99, RLF96, Ryc90a, RJ90, Ryc90b, RJ91a, RJ91b, SWG99, SGB^{+}93, SEZ99b, SEZ99a, SM97, SSH94, SZG99, Smi91, Smi92b, SAJ^{+}97a, SAJ^{+}97b, SA99c, Szo99, TBA^{+}98, TC97, TD99, TvG91, Tox99, TMS^{+}99, Tru94, ThLL^{+}93, US91, VW95, WR99, WF90, YBS91, Zha94, Zha96, Zhe99, Zie98, Zyb99, vVF96, DCR99, KKKS95, Fle92]. dynamics/quantum [BvG97]. dynamo [Cam95]. dynodic [OA95, OA96]. Dyson [HvSA98a, HvSA98b].

earth [ML92]. ECAL [BBE^{+}98]. ECCO [Pan93]. ECMC [FG95b]. ecology [CPM99]. economical [BTC99b]. economics [ACG^{+}99]. Econophysics [ACG^{+}99]. ECPSSR [LC96a]. ECRCYL [CM92a]. ECRH [MGMR97]. ed [Fle92]. eddy [Eva92, PZ91]. edge [BKKS99, MD96]. edges [FBB99, MS91b]. Edited [Reb93, Tru93a, Bon93]. edition [Fly93]. editor [Cyb90, Jam96c, Lai98, Tru93b]. Editorial [Ano93m, Ano90n, Ano90o, Ano90p, Ano91p, Ano91q, Ano91r, Ano91s, Ano91t, Ano91u, Ano91v, Ano92l, Ano92m, Ano92n, Ano92o, Ano92p, Ano93i, Ano93j, Ano93k, Ano93l, Ano94s, Ano94t, Ano94u, Ano94v, Ano94w, Ano94x, Ano95q, Ano95r, Ano95s, Ano95t, Ano95u, Ano95v, Ano95w, Ano95x, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano96x, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano97w, Ano98o, Ano98p, Ano98q, Ano98r, Ano98s]. editors [Ano90g, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano97j, Ano97k]. eds [Art96, Deb92, Fin90, Fin91, Gie90, Jon90, Rai91, Whi93, Har97]. education [BH99, Lan99b]. Edwards [JBB99]. EEWW [FJKvO95]. Effect [FBVdlN99, GV99, PFVC99, HTI99, JY97a, JY97b, SVV99, TC97, TMG99, ZYQ94]. Effective [KR91, FB99, IMA94, Pac91]. Effects [JTAW97a, BHY92, CC91a, CSS94a, Cop91, GJ97, GE99, MOSM96, MBMM99, SEZ99b, SS93b, SS94a, VW95, VGM90]. Efficiency [TSJ91, CMC94, CM98, CCM95, MCCC99, MB91, OA95, Pös98]. Efficient [BSSH99, CK91a, Krö99, KN95a, LF92, Neu99, Ols96, RASS99, SP95b, APaI99, AM90, BLGT92, BLGT94, BE90, DLLR96, DB93, FPV94, FPV95, Fan94, FCS99, GS98b, LP93, LPSM92, LS93, Mon91, Moo94, PJSFM92, PJSF93, PW99, SFM^{+}96, SAZ90, VES99, WF93]. efficiently [MSTK93]. eigenenergies [JF92]. eigenfunction [YYK^{+}92]. eigenfunctions [BKWM91, LM91b, MOSM93b]. eigenmode [SL97]. eigenpairs [SF94c]. eigensolutions [IRV95]. eigenstates [Kal94, LGZL98]. Eigenvalue [Du93b, AK91, AHH94, FJP90, Jaq90, KRT95, xLzSlLzH94, SMR99, WF93, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90]. eigenvalues [BKWM91, EHKK93, JD91a, KS96, MIO96, MOSM93b, OM93, PZH^{+}96, Sci93a, Sci94]. eigenvectors [DM98, PZH^{+}96, Sci93a]. eight [LL99a]. eight-state [LL99a]. Eighth [Sim97, Sim99]. eikonal [DS92, MF95b, MMC95]. Einstein [DL98, Gau93, KR99]. Elastic [GH91c, KK96a, SM93, SR98, ALR97, BKNS97, FS99a, OPY92, Ohl93, WT99, ZL99, SR98]. electric [Boy95, GFS99, JWF93, PR93, Poz91, SSF92]. electrically [Bha91]. electro [PC94c]. electro-chemical [PC94c]. electrolytes [LL99b]. Electromagnetic [HM91, ADM^{+}92, ABH^{+}94, ADS92, Eas91b, EABH95, FMSdC99, GLSW95, Gri97, Hew94, IKH^{+}99, JTAW97b, KSA^{+}91, Lan92, LBB91, Lin91, MV97, MSH91, MAH98, SV94, SS91a, VLG95, VB92a, VB92b, WLD95, ZS91]. electromagnetics [DWL91, MMW91, MB91]. Electron [BA94, Fan90, MSP^{+}98, AAS95, Bar93, Bar98, Bra98, BN94, Bur98, CMC94, CCM95, CM92a, CS98b, CHUS98, CB99, DJUM99, DM90, Fuk94, Gal99, Gra95, GHE99, HUCS98, Hor91, JTAW97a, JTAW97b, Kas91a, KK95, KK97, LVL90, Leh99, LLW98, MCCC99, MF95a, MF95b, MD97, MF94, MÖ93, Nor94, PMB99, Pot98, RT98, RLPR91, RLPR94, RLER97, RGL^{+}98, SG98a, Sat97b, Sin94, ST92, Szm95, Ten97, Wil90, YN94, ZFZ92]. electron- [Bra98]. electron-atom [BN94, Bur98, MF95a, MF95b, Szm95]. electron-cyclotron [JTAW97a, JTAW97b]. electron-energy-loss [LVL90]. Electron-impact [MSP^{+}98]. electron-ion [Szm95, YN94]. electron-molecular [RT98]. electron-molecule [Ten97]. electron-photon [Kas91a]. electron/positron [SG98a]. Electronic [HBH91, JKSC99, MCV98, MOPR99, Mit96, Pag94, SMR99, AK99, BKNS97, BG99, DKT96, FS99b, MCV^{+}90, MCMD91, Mal90, MSTH92, Mic94, MR99b, PBL94, RKV91, SS94b, VW95, Cra98]. electronics [RGLV^{+}99, SWT91]. electrons [HM92, HSP93, SM93, SR98, USW99, VES99]. Electrostatic [Boy95, BB99, CC97b, IBR98, SFFDR94, SP95b, UOMO94, YB99]. Electrostatics [DMLM91, MBW^{+}95]. Electroweak [BBR^{+}90, ABB^{+}96, HKL93, KS92a, TKS91]. elegant [Iri96]. element [AAKP95b, AAKP95a, AAKP98, BD94, BF91, DWL91, FBdR^{+}98, Fuj94a, IKH^{+}99, Jaq90, KV90, Man93, MT92a, NL91, Ple96, PVR96, PVRR97, PVR97, Sey95, SZG99, SG91, VGK90, dFZLS95]. element/Fourier [VGK90]. elementary [Lag99]. Elements [Mon99, AAS95, AH92, Du93a, HEM95, JY97b, Jaq90, Lim91a, LPSM92, LC93b, LBR92, MHN99, MF95a, MR96, Pös98, SP93a, Sch93, Scr95, TS92, VSS91, Var97, Zat96, JY97a]. ELENDIF [MP90a]. elevation [CMSK99]. ELF [SG99]. Eliminating [MS91a]. ellipsometric [VdN92]. elliptic [DHL93, LL90, Lee90a, PS98a, VV99]. elliptical [SS91a]. Elmau [Cyb90]. elongational [EH91, TD99]. EMBED [PCNO94]. embedded [MD97]. embedded-atom [MD97]. embryos [IST99]. Emden [Moo94]. Emden-type [Moo94]. Emergence [JK99]. Emery [HI93]. EMI [CMC94]. EMI2 [MCCC99]. emission [MWA^{+}92, WBB99]. empirical [GGSR94, MC95]. employed [TK91]. employing [LGZL98, LM91b]. emulsion [TG99]. Encryption [Ohi99]. End [Luc94a]. End-correction [Luc94a]. ended [BCP91]. energetic [PCC^{+}95]. energetics [IST99]. energies [AS92, Chi92, JBSZ97, JWW91, JWW94, JF99, Kni90, KSM92, MCMD91, Pap97, RBB98, SJPW96, Sun97, TP99]. Energy [CMV99, SGRZ99, Smi92a, Zha96, dO99, ATP98, All93, Ask94, AO98, Aut99, BMSW91, BKMM93, CSP92, CS91, CGM99, DL90, Den99, Du90, FJ95b, FFS98, Fuk94, Gei97, GW94, HE94, HP93, HP97a, HCVM99, HBH91, Hum90b, Jac94, JBB99, JBH99, Jun95, KML98, KS93, Kro92, Lam96b, LVL90, LM95, LHSD95, LPR92, Lön94, LC93b, LMW94, OE99, Ohl92, Pan93, Pap99, PCC^{+}95, PCC94, RSW^{+}99, ST95, Sat97b, Sch91a, Sch92, SvNF92, SSK97, Sin90, Sjö94, SRFN92, SRF94, WENG95, Zei97, ZYQ94, dSVTM99]. energy-conserving [Kro92]. energy-weighted [JBH99]. enforcing [Lan92]. Engeln [Hop97]. Engineering [Tay99]. enhanced [DH97a, Lan99a]. Ensemble [Zyb99, BVVJA99, Han99, VVIBA99, FH97]. Ensemble-averaged [Zyb99]. Ensemble-Based [FH97]. entire [Wei11]. entries [BMT94]. entropic [RVCF99]. entropies [YAD99]. entropy [ACA99, RVCF99, VVIBA99]. environment [GLV99, Gli90, LN93, LMS94, MXR98, MMSM93, NHK^{+}95, PDL98a, SMB^{+}93, TEJJ93, Ves98]. environmental [ZA99]. environments [Bun98, GN98]. ep [IRS97, RI97]. epitaxial [BKKS99]. epitaxy [Pes93]. EPR [BLGT92, BLGT94]. EQSHELL [Boc92]. equally [Mar91]. Equation [MSTH92, Abd90, AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, BMOF94, BKWM91, Bal95, BBL^{+}99, BK91, BV95, BP93, CRS90, CK91a, CTE95, CP95, CMC^{+}99, DEAM91, Fij99a, Fij99b, Fij99c, Fij00, Fog99, FJ95a, HvSA98b, HK95, HAR91, HKM98a, HRK95, HKC98, HHGCM99, Hut94, ICS95, IBR98, IIM92, ITMI95, JN94, KKMS99, KH97, KKK95, Kul97, LPC94, LLW98, Li95, LM91b, Mac92a, MOSM93b, MOSM93a, NY99, PT95, PW99, PSV99, SSF92, Sim92, Sim93, Sim97, SPT98, TDV96, USW99, UKK98, Var93, VD94, WZKH98a, WZA^{+}98, WA94, WF93, XYNI96, XYI96, XYE99, Yat99, ZYQ94]. Equations [Mon99, ABHR99, AH91, Aok97, AE99, Aro96, AZM92, AK99, AH97, AAH98, BDD^{+}92, BDF^{+}97, Bes98, Boc92, BSSH99, BK93, BD99, Cam95, CHW91, CTvB95, CP95, CS98a, CZ93, CS99b, CZ92, DL98, DHL93, DD97a, DD97b, Du93a, Du93b, ECT93, FG90, FJP90, FJ91, HvSA98a, Hea93, Hea96, Her91, HGCM98, HCO99, HKM98b, Ito94, Jor95, KCC96, KM95, KKY99, Kre99, KL92, KSG91, Les99, LK96, Lin91, LP99, MMGS93, MLM96, MOC91, Mic91, MK96, Ome98, OGG91, OPP96, PD96, PS98a, Pös98, QV95, Ren92, Rod92, RW95, RSW96, SM91a, SFVW95, SB97, SP93b, SP93c, SSHT91, Szm95, TG96, Wee92, Wol98b, YA91, YIW^{+}91, YKM^{+}96, YN94, ZZGS91, vGSB99]. equilibria [CHMG92, JMR^{+}93, LBS96]. equilibrium [AZM92, AMS96, BPG97, Boc92, BS99b, FG95b, IDW99, LBR92, SA99c, SS94b, VB91, WF90, dSVTM99]. equivalence [BI90]. equivalent [Kag92]. ERAFITTER [BPP98]. ERATO [Pap97]. Ergodic [MT91]. ergodicity [Has99]. Erratum [Ano90q, Ano91w, Ano91x, Ano91y, Ano92q, Ano92r, Ano92s, Ano92t, Ano93n, Ano94y, BCOY95, Fij00, Ham95a, JY97a, Jam96a, Jam96b, J̈K97a, KK97, RLPR94, Stu95b, Sul96, VRS^{+}99, Wei11, Zha98a]. Error [JF92, JF99, DSD91, GC95, HHC90, HK96a, HK96b, HK97b, SSK97, HK97b]. errors [CBTL97, ES92b, ITW97, RH99, vdB93]. eruption [HSW91]. estimate [GGJG97, HN97a, JY97a, JY97b, ML92, NVFNP93]. estimates [HW95, HK96a, HK96b, HK97b, Mar91, MR95]. Estimating [HHT95, BMT94]. estimation [ACA99, JF92, JF99, MB91, RB99, SP93b]. estimators [Ber92a]. ETA [WCL90]. ETA-10Q [WCL90]. etc [Fly93]. Euler [OY99]. Eulerian [DATL99, HE91]. eutrophication [PCSD99]. evaluate [AI90b, BCD^{+}93, BPK95, CM92b, GS98a, MC99, Sci93a, SdCR97, SG99, Wie95, vO91]. evaluated [DDDL98, ZFZ92]. Evaluating [MMU99, Nie95, ZA99]. Evaluation [AH92, AF90, ERB90, FPV95, GS97, LL90, RDMS95, AO98, AFKT97, Blo92, DM90, Dub91b, FZZ92, Flo97, FVGF98, Hin95, HHK92, HWS92, HS93b, HS96, LF92, Lee90a, LG93b, MHN99, MN97a, MN97b, MNPP99, NK96, NK93, NM97, NN93, PTMUJdFM99, RW97, RLWW97, SSS91, Sec99, Sin94, SH91, SRB94, TGG97, TS92, CHUS98]. evaluator [PBN93]. evaporation [GLF97, LMW96]. even [Lio97, RGL97, VST94]. even-even [RGL97, VST94]. even-power-series [Lio97]. Event [CDdG^{+}98, Mar97, Pap97, ADM^{+}92, AMM^{+}94, ABH^{+}94, Ano98a, CC97a, CMNP97, CS91, FR96, FIK^{+}97, FHI^{+}98, KM98, Kaw95, Kru98, KSM92, LN92, Lön99, MWA^{+}92, MC95, MNP95, MNP96, MKPGS97, Pan93, Pi92, PP92, ST95, SA92, SY95, Sjö94, TG99]. Event-driven [Mar97, MC95]. event-list [SA92, SY95]. events [IRS97, MH96, MNP96, Ohl92, Wag92]. Evolution [Jac99, Sal97, ALR99, Aly90, BVR95, BK91, CS99a, CS99b, DeV91, Gei97, Her91, HKM98b, HKM98a, Ito94, IL91, KSG91, MK96, PD96, PGK90b, SDB96, TG96, VGM90, WW90]. evolutions [Ren92]. evolutive [BMMMMR99]. evolving [BS99b]. Ewald [DDDL98, Ome97, PG96, Smi92b, SP95b, SKPP95, TB96, WA97]. Exact [Her91, IL91, LC90a, LC92, Eng91, FMH97, JWJK92, Lia95, The94, Wei99, Wei11]. EXAFS [AI90a, KSSB99]. Example [ES92a, Mer90]. Excalc [SMH98]. EXCALIBUR [BPK95]. except [Cor99]. exchange [BS91a, BBS92, DER90, MCMD91, PRRL99]. exchange-correlation [MCMD91]. Excitation [CKW98, Eng91, Ber99, LM91b, Mal90, RT98]. excitations [HM99, TCE99b]. excited [BMSW91, CY99]. exciton [HB99a]. excitonic [BRWN96, BWNR99, DBGY99]. exclusion [HN97a]. exhaustive [SvdPS99]. expand [FG97]. Expanded [VVIBA99, BVVJA99]. Expansion [vMF94, AL93, BMSW91, CS99b, FG95a, Goz91, Lee92, LMW94, Mar91, MSW92b, MSW92a, PZH^{+}96, Ran93, Sha92, VGM90]. expansions [BMOF92, DGKS99, Fuj94b, HN97b, Jak98]. expectation [dlVM98]. experience [BS91c, BSW97, Krö90]. experiment [MG90]. experimental [BA94, FBGV99, Hen94, Lam96b, MG93]. experiments [BMSW91, Cop91, GF98, Hum90b, MSR96, MTH^{+}99, VFO94]. expert [CB99, Hum90c, Ker90a, McD90]. EXPFIT4 [IDVV97]. Explicit [Koi92, Sim99, LM91b, Mic91, Pop98, VDVV99, VB91]. explicitly [Zha98c]. exploration [SvdPS99]. explorations [Ohl93]. explore [GPM95]. Exploring [BM99a]. exponent [MP95]. Exponential [Sim92, CRS90, Fer99, Iqb95a, IVDV97, MG93]. exponential-fitted [IVDV97]. Exponentially [VDVV99, AH91, Sim98]. Exponentially-fitted [VDVV99, Sim98]. exponentiation [JRWWW92, SJPW96]. exponents [CVZ99, Sat97a]. exposed [MFG^{+}98]. exposure [ML92]. expressing [Wes93]. expressions [Por98]. Extended [VW95, Bla94a, Hea91, Hni94, HS99, MBZM99b, SM91b, Zau94]. Extending [MH99]. extensible [RGL^{+}98]. Extension [GF96, DKS98, Mre97, dBPB^{+}93]. Extensions [Ohl93, CA92, FKKJ91]. extensivity [dMCT99]. exterior [Har97]. extinction [MBZM99a]. extracting [LPPR91]. extraction [BMMMMR99, RTLM92]. extrapolation [AA94b, AA94a, AA98a, Alf99, Gro90, HRK95, Mid92]. Extreme [SWG99, DS96, OM93, Sci94, SF94c]. extrinsic [IJV92].

F [Eas93, Hop97, Top96, GUW98, HHCS95, JP92]. faces [HI93]. facet [FBB99]. Facility [MMSM93]. factor [HK97a, HMH99]. factorization [FT95, Ohl99, PS98a]. factors [BEFJ99, Jan90, LHL95, Ten97]. family [DH97a]. far [Gri97]. Farm [BN95]. farms [CGL^{+}98]. Fast [BT91, Dub91b, Kar92, KBL98, PDZ92, PJ99, RBS92, SAG91b, SH91, BTC99b, BHM99, CC97b, DM90, EK94, GH91a, Goe93, Gut95, Heu90, KT91, Kru96, LG93b, PJ98, PS98a, Rho92, SS94a, SKPP95, Sto98, Wee92, vMC91]. fast-electron [DM90]. fcc [CDUT99, HMH99, KF99]. FCT [Rea95]. FDEXTR [AA94a, AA98a]. FDTD [Gri97, WP98]. FE [Cam95, MSP^{+}98]. FE-time [Cam95]. feasible [SH97]. Feature [DD95]. features [KI99, LDOO95, LPPR91, PCNO94, TH95]. February [Reb93]. FEDER [WENG95]. FEDER/2 [WENG95]. feed [Frü93, LVC^{+}94]. feed-back [Frü93]. feed-forward [LVC^{+}94]. feedback [AGIS99]. Feedforward [Pod99]. Fermi [AA99, Gau93, Mac92a, MN97a, Moo93, NK93, PR93, Sag91a, ZZGS91]. Fermilab [CGL^{+}98, Wol98a]. Fermion [PK99, AB97, AW94, BPK95, BEF^{+}99, BI90, CMNP97, FIK^{+}97, HKL93, JW90, JPS^{+}99, KKY99, MNP95, Mon98, Pap97, Pas96, jWzHxL95, ZAW94]. fermions [DR90, JL97, Kal96, LBDL99, RH99, Tan90, vOFB94]. ferroelectric [dM94]. ferromagnetic [Fer97, HN99]. FESSDE [AAKP95a, AAKP98]. FET [Fau91]. few [SF94c, VS97, Zha94]. few-body [VS97]. Feyn [KBD90, MBD91]. Feynman [Kan95, Stu95b, BFK98, EHK^{+}92, FT92, KV99, KBD90, Lai98, LF92, Lev90, Mak91, MBD91, Ohl95, RDF^{+}97, Sem98, Sin94, SGT90, Stu95a, TKS91, Wan93, Wes93, vO91]. Feynman-graph [Kan95]. FeynmanParameter [Wes93]. FF [vO91]. FFT [Fer97, GH91a]. fibers [FMSdC99]. Field [Kre99, Aly90, BVR95, BG95, Boy95, CS98a, Coh92, CKW98, DB93, FV99, Fuj94b, Hea91, Hol96, Jac91, KNN91, Lin90, Lüs94, Mar91, MGMR97, MS91b, PGJ91, PVR96, PVRR97, Pös98, Reb93, RGL97, SA99b, SPA99, SW91a, SSF92, Sou99, US91, VB92a, VS99, ZS91]. fields [AW94, CJ91, DeV91, DG99, GFS99, Gri97, HE94, LM91a, PR93, Poz91, SM91a, SMBV90, SSW94, TSK^{+}97, YB99, ZAW94, vMF94]. fifth [LMW94]. fifth-order [LMW94]. figures [CT96a, Deb92, Fly90, Fly93, Tru93b, Weg91]. filament [KCH92]. filled [GF96]. Filling [KKM^{+}99]. film [HNS99, LPS99, WCK99]. films [HB99b, Wie96]. filter [Ano98a, Kru98, MES90b]. Filtered [SM97]. filtering [AAD92a, AAD92b, LG93a]. filters [Dik94, FS99a, SZ99]. final [APN^{+}91, JPS^{+}99]. financial [Deu99, Man99]. Finding [BO98, BRB92, DM98, Cle93, Dik94, Ito94, JD91a, LVC^{+}94, OPY92, PD96, PP94, RDF^{+}97, SF94c, SAG91b]. Finite [AA94b, AAKP95b, BS91b, FBdR^{+}98, PC99, PVR97, dFZLS95, AA94a, AAKP95a, AA98a, AAKP98, BB93b, BM99c, BF91, CK91a, DWL91, Fer97, Fuj94a, IKH^{+}99, Jac91, Jaq90, KCC96, KKMS99, KV90, KSG91, Lem96, Man93, MSFR96, Mic91, MT92a, MSH91, MN97b, NL91, PT95, Ple96, PVR96, PVRR97, Pös98, SS90a, Scr95, SZG99, SG91, VGK90]. Finite-difference [AA94b, AA94a, AA98a, CK91a, KSG91, Mic91, PT95]. Finite-element [AAKP95b, AAKP95a, AAKP98, MT92a, NL91]. finite-volume [MSH91]. Finsler [Rut98]. Firecracker [ABH99, ST95]. First [FT99a, Abd90, BC99a, BS99a, CTDdM97, CTR98, CY93, DS92, Ker90a, Lee90a, SP93b, SG99, Wee92, YM97]. first-generation [Ker90a]. first-order [BC99a, DS92]. First-principles [FT99a]. Fisher [RAD99b]. Fission [PP99, BGG91, GRM^{+}99, GDGR97, GLF97, VST94]. Fit [MG93, ATP98, BPP98, Kar92]. fitted [AH91, BM92a, EABH95, IVDV97, OE99, Sim92, Sim98, VDVV99]. Fitting [BH90, BB93b, BNSW92, Ber92b, CRS90, CGR93, Frü95, Frü97, FFS98, FS99a, LH97, MPN^{+}93b, MK91a, OPT99, Rob98b, SvNF92, SRF94, Swa91].

FitzHugh [IDW99]. five [Gal99]. five-shell [Gal99]. fixed [HK94, Mal90, Rei93]. flares [Lin90]. flavour [AS91, ABF^{+}99, IRS97]. FLEX [SLGB96]. flexible [BN95, FR90, KH95, KH96]. flight [GRS90]. Floquet [CGHB98]. flow [ADK^{+}99, BRM99, CS99a, DSB91, DW99, EH91, FMH93, GGH91, JAT97, KK91, LLM91, MCO99, MK95, NvdHPG97, PD90, PZ91, Sug99, TD99, VB91, WNS99, XYIT97, YY99]. flows [Bru91, Cal96, CK91b, Dub91a, DJMT91, FMIM93, Mac91, MT92a, MS95b, WJ91, ZY97]. fluctuation [WK90]. fluctuations [BRWN96, BG94, CBB99, STJ90, TMG99]. Fluid [BB99, Art93, BT92a, Cal96, CS99a, DL98, Fle92, GBWES97, GHM93, KL99b, Kud99, LB95, LB91]. fluids [MC93, Ome97, Rai91, Ris96, SC99b, Tox99, dGT99, vVF96]. flux [APaI99, BPG97, KCK99, RB99, SRFN92, Str93a, YK95]. flux-limited [YK95]. FMM [PG96]. foams [SLB99]. Fock [CY99, Kre99, AZR96, BBS91, DD97a, DD97b, Fro91b, KLS96, PI95, PCNO94, PCC94, PC94b, PC94c, PC94a, PB96, TK91, VES99]. FOCUS [LDOO95]. focused [Chr99]. focussing [GP95]. foils [SRFN92]. Fokker [KH97, MGMR97, MOC91, MMTK94, PS98a, SSF92, SS93b, SS94a, WP96]. fold [MCV98]. Folding [SUYH99, DMvG99, Han99]. Force [KWS^{+}96, Aly90, HE94, HS99, Lin90, MK96, Rou95]. force-free [Aly90, Lin90]. forces [CKS91, IBR98, MPB91, Ran93, SF94a, TR93, The94, VS97, WCL90]. forcing [BC91b, NKR99]. forecasts [Ker90b]. foreign [BM99c]. forest [CPM99, KKKK99]. Foreword [Ano91z, DT95, Tru95]. form [AH91, Gri97, Kal93, Ker90b, Ren92, Wan93, BS98, CvN99, FT92, SW98]. formal [MKS^{+}96, SKM94]. formalism [HK96a, MCMD91, VW95, dCRSF99]. formalisms [CC98]. formation [CGRV99, HSW91, LMF97, Min91, MBMM99, RJS99, RLF96]. formatted [Top96]. forms [BGT95, RU92]. formula [Du90, FPV94, FPV95, Ver99]. formulae [APaI99, JD91b, Koi92, Lim91b, Luc94a, Luc94b, MIO96]. formulation [JH91a, MT92a]. Formulations [Kis91]. forone [Sat97b]. forone-electron [Sat97b]. Fortnet [AHZ90, CA92]. FORTRAN [Tay99, Jam96b, AIS^{+}97, AI90a, AHZ90, CHM91, CJP^{+}94, GS98b, HHCS95, IDVV97, KKK95, KL92, LSZ92, NVC96, NVFNP93, RBS92, TYJ92, YB92, CSC^{+}97, CRS90, CA92, DLLR96, DNS98, Dem97, DCR99, Hen95, Heu90, HIK90, Jam94, KKY99, MKS^{+}96, RM90, Sat97a, SKM94, SH97, Var97, jWzHxL95, YK90, Hop97]. Forward [Ano91-27, KB94, LVC^{+}94, MPN^{+}93b, WF93]. forward-backward [MPN^{+}93b]. Four [IVDV97, AB97, BR97a, BPK95, CCM95, CMNP97, FIK^{+}97, HPV99, HKL93, JPS^{+}99, LGZL98, MNP95, Pap97, WDB^{+}91]. four-atom [LGZL98]. four-fermion [AB97, BPK95, CMNP97, FIK^{+}97, HKL93, JPS^{+}99, MNP95, Pap97]. four-index [WDB^{+}91]. four-shell [CCM95]. Four-step [IVDV97]. Fourier [CS98a, SL95, AI90a, AI90b, AL93, BKWM91, BH92, BH99, DEAM91, DVH90, FS91, GH91a, Goe93, Luc94b, Ole96, PJ98, SL95, VGK90, VVGI99].

Fourier-Riccati [SL95]. Fourier-spectral [CS98a]. fourth [BMSW91]. fourth-order [BMSW91]. FPPAC94 [MMTK94]. fractal [DdMdO^{+}99, FBVdlN99, dVdRL93]. fractally [SL97]. fraction [SdCR97]. fractional [Kag92, xLzHjW95]. fractions [RdC93]. fragment [KKK91]. Fragmentation [DS99]. frame [LC93b]. framework [GF98, HAC^{+}99, KH96]. Frautschi [JRWWW92, SJPW96]. Fredhol [Wee92]. Fredholm [Abd90]. Free [Whi93, Aly90, Fuj94b, HCVM99, IKH^{+}99, Jan90, JBB99, Lin90, Mac91, MBC99a, Nun90, PCC^{+}95, RH99, Sin90, SSW94, TSK^{+}97, TSL^{+}99, VSS91, VVIBA99]. free-energy [JBB99, Sin90]. free-free [Jan90, VSS91]. Free-Lagrange [Whi93]. free-mesh [IKH^{+}99]. freedom [Kal93, SVV99]. freezing [HHGCM99]. frequencies [Nie95]. frequency [Hew94, LEBB99, SLGB96, SYM98, TCE99a, VB92b, ZX91]. friction [KKM^{+}99, MHHL99]. FRITIOF [Pi92]. Fritts [Whi93]. frog [Li95]. frontal [Duf96]. FROTH [KW90]. froths [KW90]. frustrated [BL99b, HM99, SA99c]. FTN77 [Ham94]. Full [BSST90, KWS^{+}96, Luo96]. Full-potential [BSST90, KWS^{+}96]. fully [HLT93, THF95, VB92b]. Funaro [Mon99]. Function [Dik97, Str93b, vMC91, AO97, Aya92, BWS94, BHP97, CPCF99, DL91, FBB99, FZZ92, Fuk94, HKA^{+}94, KL99b, LHSD95, Lia95, LMS^{+}93, LG93b, NM97, NVFNP93, OPT99, PZH^{+}96, PD96, Sat97b, SVBD92, SRR95, WC90, WBS95, XYNI96, XYI96, YYOM90, YM97]. functional [BRB92, BKNS97, CT94a, FS99b, JBSZ97, J̈K97a, J̈K97b, Krö90, Lob96, Lob99, SM97, SS94b, TZC96, WZKH98a, WZA^{+}98, ZKHG99, vGSB99]. functionals [Cle93, Les99, WKH98]. functions [AS92, All93, Ask94, AL93, BMOF92, BT91, CL96, Cha92, CY99, ES92a, ES92b, FG95a, FG97, FG91, GJ98, GS97, GP99, Gra98, GD93, GHE99, Gro90, Gro91, HS93a, HKM98b, HN97b, HN97c, HWS92, HS93b, Ixa97, KRVZ98, LC90b, Lee90a, LL96, LL99c, MP95, MG93, NSE95, NCC^{+}96, PT92, PB93, RdC93, Sch91a, Sch91b, SdCR97, SG98b, SG99, Ská90, SAZ90, SF93, Tau92, TB91, VV99, VRS^{+}95, Wan99b, WP98, dT93, VRS^{+}99]. Further [ST92]. fused [Iri96]. fusion [DB93, HRK99, MDMR99, RM90, WBB99]. fusion-reactor [DB93]. Fuzzy [Nak97a].

G [Gie90, Hop97, Oln98, SMB^{+}93]. G2DEM [JTAW97b]. Ga [BBB99]. Gaemers [Nar92]. gain [VFO94]. Galerkin [HMG92]. GALUGA [Sch98]. gamma [CMC94, HIK90, NCC^{+}96]. gamma-spectrum [HIK90]. gap [Boy95]. gas [Ask94, BB92, DCM99, Dub91a, Jac94, KK99, LC93a, Pae91, SC99b, Sin94, Zha91, Zie98]. gas-surface [Ask94, Jac94, Zha91]. gaseous [SWT91]. gases [CR99, GSS97, HD96, MCO99, SLB99, dlSG99]. gauge [AUG92, AB90, Hoe90, MSTK93, RF95]. Gaunt [Jan90]. Gauss’ [Lan92]. Gaussian [Frü97, HN97b, HS96, Sag91a, SSS91, TC93, TZC96, vHKH97]. Gausterer [Reb93]. GCPM [DLS90]. GEANT [MS96, dBPB^{+}93, LR92a].

General [GKP93a, Han90, AE99, CW92, CC98, FPV94, FPV95, FPV97, FMIM93, Fro91b, FBdC98, Gro90, Hal90, HF91, HGF91, HB90, JOP98, KV99, MP95, PCNO94, Pit98, PS97, RG92, RH99, RDMS95, Sta99, Zat96, HK96a, Sal94, d’I98]. general-purpose [Hal90, Sta99]. generalised [NK93, You99]. generalization [Kle92]. Generalized [BH99, Jak98, SS91a, SKWH99, SL95, ZYQ94, BH92, BC99b, DP91, Gau93, GM99, Han99, HM91, HHR98, Lud91, PBN93, Sag91a, Sci91]. generalized-ensemble [Han99]. generate [BTC99a, BTC99b, FL91, MH96, SF93]. generated [RBB93]. generating [APN^{+}91, GS98b, HM92, Kae95, KW96, MSFR96, PWG93, Wag92]. Generation [AB90, FPV94, HK94, J̈K97b, MM91, MTM92, Shu94, TC93, Wan93, AMM^{+}94, Boc92, CGL^{+}98, Hen95, IKTY99, IWWY92, KD95, KI99, Kaw95, KTM96, Ker90a, KBD90, LMW94, MF95a, McD90, POA92, Rod92, Sem98, SS90b, Sjö94, Ská90, SL94b, J̈K97a]. Generator [Jun95, AS91, ADM^{+}92, ABH^{+}94, ALR97, BLLM94, CC97a, CSS94a, CMNP97, FR96, FJKvO95, FIK^{+}97, FHI^{+}98, GP99, Gut95, Ham95a, Ham95b, Hen94, IRS97, IER97, Jam94, Jam96b, Kan95, KM98, KSM92, LN92, Lön99, Lüs94, MSV92, MD96, MWA^{+}92, MNZ90, MNP95, MNP96, MKPGS97, NK96, Pan93, Pap97, Pi92, PP92, RI97, ST95, Sti94, vOFB94]. generators [CBTL97, DH97a, DH97b, Ham93a, Ham93b, Ham94, HDF97, Jam90, Shc99, VKSAN95]. Genetic [CGM99, Pul97, BS99b, Smi92a, Zei97]. GENTLE [BBL^{+}97]. GENTLE/4fan [BBL^{+}97]. Geoff [Tru93a]. geometric [Bru95]. geometric-systolic [Bru95]. geometrics [RSW92]. geometries [DP91, HKC98, MCV^{+}90, TMR94]. geometry [CTE95, Cop91, DS90, ECT93, HSP93, IMA94, LVL90, RBG99, STR92, TDV96]. geophysical [Cal96, Coo99]. George [Ano94-33]. germanium [OE99]. GESAMAC [DPP^{+}99]. Getting [PRdM99]. GFSR [Ham95a, Ham95b]. GGPS [MKPGS97]. giant [GLF97]. Ginzburg [HHGCM99]. GITA [BCM^{+}95]. given [GGJG97]. giving [Str91b]. glass [Bin99, JBB99, PTU97, RLF96]. glasses [Hof99b, Krö99, PC99, RARF96]. GLie [AH97]. Global [ATP98, JAVV95, PCSD99, FBdR^{+}98, SL99]. GLOMAC [LC93a]. glow [HHV99]. gluon [HvSA98b, KKK91]. gluons [MWA^{+}92]. GMRES [BSSH99]. GNOME [SG99]. good [Lan99b]. Gordan [Kae95]. GRACE [BBB^{+}98, MSK98, SJB^{+}94]. graded [BdM93]. gradient [AG99, BI90, KS96, Nak97b, SG91]. grain [Nar96]. grained [HTK99]. grains [KW99]. grand [HCVM99, MB94]. Granular [KK99, BRM99, DW99, Rad99a]. GRAPE [IMES90]. GRAPE-1 [IMES90]. graph [Kan95]. graphic [LP90]. Graphical [SMB^{+}93, FPV97]. Graphics [Len98, Lev90]. graphite [WCK99]. graphs [FBdC98, KBD90, PRH90, Sin94]. GRASP [PHT93]. GRASP92 [PFG96]. Grassmann [AH97]. Grassmann-valued [AH97]. gravitating [CDM99, MS95b]. gravitational [Kli98]. Gravity [MBMM99, BBJ95, Cat95, SMH98]. grc4f [FIK^{+}97]. greater [GS97]. Green [BWS94, AL93, HS93a, MCV98, WP98, WBS95]. Green-function [BWS94].

Greene [CT96a]. Grid [GJ99b, KTM96, BKWM91, BK93, EK94, MS91a, SSW94, Wei93, YM92, ZY97]. gridded [Mar91]. grids [LHL95, Wuc95, YK95]. GROMACS [BvdSvD95]. ground [BS96, CY99, RDF^{+}97, RGL97, RBB98, SS94b, dCdJDvL99]. group [Ber95, Flo97, Fra92, Gra98, MBC^{+}99b]. groups [CP99, DPD96a, DPD96c, MXR98, MSFR96]. Growth [MPS99, App99, BKKS99, CGRV99, GST99a, GST99b, JK99, Kot96, LPS99, MMW96, MBC^{+}99b, Pes93, PFVC99, RNGR99, SDB96]. GSAP [HIK90]. guiding [HL92, MAH98]. guiding-center [HL92]. Gummel [ZYQ94]. Gustavson [Kal93]. GW [RSW^{+}99]. Gyrokinetic [SV94, FBdR^{+}98, MSD93]. gyrotron [JTAW97a]. gyrotropic [GBWES97].

H [Fin90, Fin91, Reb93, Whi93, DM90, Hea91, Hoe90, PLC91]. H-atom [DM90]. H-gauge [Hoe90]. H.-B [Fin90, Fin91]. HABIT [CDR91]. Hadamard [MN97b]. hadron [Gei97, MWA^{+}92, Pan93, PP92]. hadron-hadron [Pan93]. hadronic [APN^{+}91, GW94, Kno90]. hadrons [Pi92]. HAGIS [PAC^{+}98]. Haken [Cyb90]. half [GDGR97, SG98b, SJ91]. half-integer [SG98b]. half-life [GDGR97]. Hamilton [LK96]. Hamiltonian [BKWM91, BC91b, CTdO96, Far98, HF91, HGF91, KMz91, Ren92, SS93a, Zha96]. Hamiltonians [BCM^{+}95, RU92, SGRZ99, SM91b]. handle [Kre99]. Handling [Rit92, KH96]. Hankel [Sec99]. Hans [Fly90, Fly93]. Hard [GJ97, Jun95, Str93b, BM91, BT92a, BT92b, Cyb90, JW90, LH92, Mar97]. hard-core [BT92a]. hard-particle [BT92b, Mar97]. Hard-Sphere [Str93b]. hardcore [BM99b]. Hardcover [Art96, Bis91, Bon93, CT96a, Deb92, Fin91, Fly90, Fly93, Gie90, Hoc92, Rai91, Reb93, Tru93b, Whi93, Fin90]. harmonic [BdT97, DD97a, DD97b, LPC94, LBDL99, MR96, Mar91, TSK^{+}97]. harmonics [AC92, GS98a, vMF94]. harness [AHZ90, CA92]. Hartree [DD97a, DD97b, Kre99, BBS91, CY99, Fro91b, JD91a, JF92, KLS96, PCNO94, PCC94, PC94b, PC94c, PC94a, PB96, PVR97, TK91, VES99]. Hashing [Lia95]. HCI [RFMM99]. HCP [KF99]. HDECAY [DKS98]. He-atom [FB94]. heart [AGIS99]. Heat [HHC90, BLS96, GFAALS98, NN93, ŠČ91]. heat-kernel [BLS96]. heating [PdA99, SS93b, UOMO94]. heavy [AS91, ABH99, BZO92, BG94, Dub91b, HRK99, IRS97, RI97, RVB99, SH97, Zha98a, Zha98b]. heavy-ion [ABH99, HRK99]. heavy-particle [Dub91b]. HECTOR [ABB^{+}96]. Heermann [Hoc92]. HEG [AH92]. Heidelberg [Fle92, Fly90, Fly93]. Heisenberg [Bes98, Flo97, HN99]. helical [Sik99, TSK^{+}97]. helicity [SL94b, Tan90]. helicon [KB98]. Helium [SL94a, OC98, SPT98, WW93, WCK99]. helpful [JH91b]. Hemmen [Deb92]. hemoglobin [GPM95]. HEP [Bun98, CLM98, GF98, Lön99]. hepawk [Ohl92]. HERA [ADM^{+}92, JP92, KSM92]. HERACLES [KSM92]. HERBVI [GW95]. Hermann [Cyb90]. Hermite [LBR92]. Hermitian [MIO96, OM93]. HERWIG [MWA^{+}92]. heterogeneous [Ves98]. heterojunctions [RHGDM99]. heteropolymers [RS99].

Heterostructures [SYM98, CL96]. hexagonal [IMA94]. HF [GF96]. HFODD [DD97b]. HFS92 [JPF96]. hidden [Zlo90]. hierarchical [Man99, Nak97a, PG94]. Higgs [AB97, DKS98, SAY93]. High [KS93, Mar96, Sim93, Sjö94, YB99, AAKP95b, AAKP95a, AAKP98, All93, Art91, AO98, Aut99, AB90, Bow99, Chi92, CS91, DLLR96, DG99, DL90, Den99, Gei97, GW94, HP93, HBH91, HMG92, Hum90b, IWWY92, Jam94, Jam96b, JMK99, Jun95, KMT^{+}99, KSSB99, LPR92, Lön94, Lüs94, Ohl92, Pan93, Pap99, PSV99, RdC93, ST95, Sek92, SL97, Sun97, Tar99, TP99, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90, DCR99]. High-energy [KS93, Gei97, Hum90b, Pap99]. High-energy-physics [Sjö94]. high-level [DLLR96]. High-order [Sim93, AAKP95b, AAKP95a, AAKP98, PSV99, SL97]. High-performance [Mar96]. high-quality [Jam94, Jam96b, Lüs94]. high-speed [IWWY92, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90]. higher [ADM^{+}92, ABH^{+}94]. highly [KH95, KH96, LL99b, MAH98, QAd^{+}95]. HIJING [GW94]. Hilbert [GJ99b, LM91a]. Hiller [Top96]. hindered [HSW99]. histogram [MH99]. histograms [LC96b]. HNCR [LH92]. HOC1 [RFMM99]. hole [FZZ92, SRR95]. holes [TAMSW90]. holographic [OBRR92]. holography [IEY^{+}96]. Holstein [MOSM93b, MOSM93a]. homoclinic [HAR91]. homogeneous [KKY99]. homothetic [KS98]. Hong [Fly90]. HOPS [AS95]. HORN [IYOY94, IEY^{+}96]. HORN-1 [IYOY94]. HORN-2 [IEY^{+}96]. hot [DG93, DS99, GLF97, JAVV95]. Houston [DMLM91, MBW^{+}95]. HPCN [BJL^{+}96]. HSTERM [AAS95]. Hubbard [FMH97, BS96, NSE95, Sci91]. HUBCOM [Sci91]. Hugenholtz [LMW94]. hull [AN97]. Hybrid [FCS99, HNR90, MMW91, RF94, ST99, Krö99, LB91, MC93, Nun90, WJ95, ZmSjFyJ97, ZFZ92, AS95]. HYDMATEL [SP93a]. hydrocodes [PC90]. hydrodynamic [Hin95, MCO99, MS95b, RBP^{+}90, Rea95, Rho92, SS91b]. hydrodynamical [APaI99]. hydrodynamically [HTI99]. hydrodynamics [BA95, BM95, BNS^{+}95, DF95, HAC^{+}99, KKH^{+}98, Man91, RHFR95, SEZ99a, YK95]. Hydrogen [DG99, DeV91, PMM98, SP93a]. hydrogen-like [SP93a]. hydrogenic [SH91]. HYDROLIB [Hin95]. hydromagnetic [VGM90]. hyper [NY99]. hyper-dimensional [NY99]. hyperbolic [YA91, YIW^{+}91]. hypercube [Smi91]. hypercubic [CHM91]. hyperfine [JWF93, JPF96, KK95, KK97]. hypergeometric [PBN93]. hyperon [GJ96, GJ97]. hyperplane [DH97a]. Hypersonic [Abg91, Bru91]. hyperspherical [AAS95]. hypervirial [Lio97]. Hypervolume [FF97a]. Hypervolumes [FR95].

I-NoLLS [LH97]. I. [Cra98]. Ian [Top96]. ice [IK99, PP99]. IDA [HKSV90, Mal97]. ideal [CNO98, CZ92, DVH90, DMM^{+}97, KCC96, LL96]. identical [Lia95, LSZ92, xLzHjW95, VVGI99, jWzHxL95]. Identification [BB93a, Bla94b, TT99]. IDO [Aok97]. II [Fin91, DD97b, FPV95, Fan90, FVGF98, GKP93b, Ham93b, HUCS98, HK96b, JMR^{+}93, Lam98, PCC94, Rap91b, RDB^{+}90, RSW96, Ryc90b, RJ91b, SJB^{+}94, SF94b, SGT90, YIW^{+}91, YYK^{+}92]. iii [Tru93a, CHUS98, Fau91, HK97b, HZQ93, Lam96a, Lam96b, Pes93, PC94b, Rap93, WDB^{+}91, WGH^{+}98, YKO^{+}94]. III-V [Fau91, Lam96a, Lam96b, Pes93]. ill [Abd90, BG98, Wee93]. ill-posed [Abd90, Wee93]. illustrated [CvN99]. ILU [FT95]. image [Bro92, Bro94, IYOY94, RTLM92]. Imaging [Par99, CNPT91]. IMAS [Rho97]. immersed [MC97]. immiscible [SC99b]. immune [BCSS99, dSB99]. Impact [MBZM99a, Bar93, Bra98, Lag92, MSP^{+}98, OC98, Sal91]. impedance [Kis91]. implantation [RGL^{+}98]. Implementation [BV99, CA92, CZ92, FH97, FHSO99, JL97, Kru96, MLM96, MPRS97, Sta99, vGSB99, APaI99, BDF^{+}97, BvdSvD95, Bla94a, BF90c, BMPS99, DLLR96, DCR99, GKP93b, JBSZ97, Jam94, Jam96b, KDH^{+}98, LR92a, MKS^{+}96, MC97, MES90b, RBS92, SKM94, SBGM91, Stu95a, Stu95b, Ver99, WRF91, WJ94, WJ95]. Implementing [KN95b, Lön92]. Implicit [IY95, Cam95, CS98a, PT95, VB92b, Li95]. implicitly [LGZL98]. Importance [Krö90, SGB99]. imposed [Lil94]. improve [DHL93]. Improved [HP93, KSZ97, LB99, TF98, BEF^{+}99, KSSB99, Luo96, OY99, SP93b]. Improvement [AMS96, Tau95, JL97, Krö96]. improving [MN90, dMP99]. impulsive [Lim96]. impurity [HMH99, MLB93]. impurity-vacancy [HMH99]. In-place [PJ98]. incidence [YK90, YB92]. incident [SRFN92]. include [Pac91]. including [ABH^{+}94, FJKvO95, GLF97, JP92, KM91, KR95b, KSM92, MOSM96, PMB99, Str91b, UTI94, USW99, ZYQ94]. inclusion [NVC96]. inclusive [BKW97, KL93]. incompressible [CS99a, DSB91, MT92a, YY99]. increase [MB91]. increasing [WW90]. independent [BTC99a, Bil91, HEM95, Hor91, KR95a, Lag99, Pot98, Rob98b, Wha94]. Index [Ano99n, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90s, Ano90t, Ano90u, Ano90v, Ano90w, Ano90y, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91a, Ano91b, Ano91-29, Ano91-30, Ano91-31, Ano91-32, Ano91-33, Ano91-34, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano93a, Ano93b, Ano93c, Ano93d, Ano93o, Ano93p, Ano93q, Ano93r, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94m, Ano94-27, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95z, Ano95y, Ano95-27, Ano95-28, Ano95-29, Ano95-30, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano96-29, Ano96-30, Ano97a]. index [Ano97i, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97-30, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, SL99, WDB^{+}91]. indices [KKKT99]. indicial [Kli98]. individual [GLV99, TBS98]. induced [BZO92, BG95, CP92, Di 95, IST99, KKM^{+}99, SA99a, SZG99, TC97, Zei94]. induction [VPE94]. industry [RGLV^{+}99]. Inelastic [BKW97, FB94, MK91b, AIS^{+}97, ADM^{+}92, ABB^{+}96, ABE^{+}94, CSS94a, CB99, Gra95, IER97, JP92, SKST91]. infinite [BS91b, BWS94, Fer97, HS96, SSS91]. infinitely [YB92]. Influence [CBB99, BWNR99, BC99a, BKKS99]. Information [BMMMMR99, Hoc92, Man99, KSA^{+}91, RAD99b, SA99a, VFO94, Whi98, YAD99]. inherent [ZmSjFyJ97]. inheritance [DNS98]. inhibition [GST99a, GST99b]. Inhomogeneous [BV95, Cal96, Du93a]. initial [FS91, HÓ99, IDVV97, JW90, KRT95, Krö90, MMC95, Sim93, Sim98, Sim99]. initial-state [MMC95]. initial-value [FS91, IDVV97, Sim98, Sim99]. initio [Alf99, Der99, DKT96, FS99b, OE99, PBL94, PvBMN99, ATP98, BPL97, BKNS97, CSP92, ERB90, ISS91, LPR^{+}93, MTM99, RHGDM99, RDB^{+}90, WDB^{+}91]. injection [FWH95]. inkompressible [Weg91]. Innovative [Lag99]. input [CS98b, Pod99, RASS99, SA95, Tur99]. input-output [RASS99]. Inst [Par99]. instabilities [HAR91, JTAW97a, JTAW97b, KRT95, PL99]. instant [SWG99]. Institute [Art95a]. instruction [DNSP99]. instructions [Ano91h, Ano94j]. Integer [Gri97, SG98b, dT93]. Integer-arithmetic [Gri97]. integrability [HGCM98, SW98]. integrable [SGF99a, Sei94]. Integral [AC92, AZM92, VBO97, VD94, WDB^{+}91, Abd90, BRWN96, BSSH99, Bra91, CP95, MN97a, MN97b, RDF^{+}97, RDMS95, Str91b, VVGI99, Wee92, Wil90, Yat99, dCRSF99]. integrals [BS98, BSKO96, BFK95, BFK97, Bur98, CW92, Dub91b, ES92a, ES92b, ERB90, EHK^{+}92, FZZ92, Gau93, GM92, Han90, HD92, HF91, HGF91, HWS92, HS93b, HS96, LP93, LL90, Lee90a, Lob96, Lob99, Mac92b, MC99, MS98, NK93, RW97, RLWW97, RWL^{+}98, RLPR91, RLPR94, RLER97, Sag91a, SSS91, SVBD92, SGT90, Stu95a, Stu95b, TON95, Wes93, ZFZ92]. integrands [BSW97]. integrate [Pop98]. Integrated [BBP^{+}90, Rho97, CHC91, Mer90]. Integration [VSS91, BS91c, BSW97, Bil92, BF91, BP93, CBW97, CRS90, JHK97, JD91b, Kaw95, Kle92, Lil94, LP99, Luc94a, Luc94b, Mak91, MP95, Ohl99, Ome98, PJSFM92, PJSF93, SP93c, Sim92, Sim93, Sim98, Sim99, SPT98, SM91b, Sta96, Ves98, Wan99b]. integrations [FGH91]. integrator [CMNP95, DG99]. integrators [KPD97, Kre97]. integrodifferential [Jor95]. Intel [Gut95]. intense [CJ91, DeV91, DG99, Hea91]. intensities [CHUS98, MS95a]. intensive [DDK99, Hof99a]. interacting [BR97a, HTI99, LBDL99, VES99, WCL90]. Interaction [Sal97, BR95, DJMT91, FS91, Fro91a, GBWES97, HMH99, HE99, LL99c, MHN99, NVK93, Ome97, PAC^{+}98, PP92, Riz99, SKG91, Sun97, TC97, USW99, Zha91]. interactions [AAS91, Bra91, Ess95, EK94, GJ99a, Hin95, KSM92, LB91, Nun90, Pi92, PFVC99, Rho92, RJ90, Ryc90b, RJ91a, RJ91b, SC99b, dMCT99].

Interactive [CDdG^{+}98, Tay99, BNSW92, BH92, HKSV90, LH97, Mor92, Rob98b]. Interatomic [OE99, CKS91, Pop98, PFVC99]. interchange [Krö90]. interest [GM92]. Interface [Lam96a, Lam96b, CMC^{+}99, Fit90, GPSW95, Hal90, Hen95, MCV^{+}90, MG90, Pag94, Zan90]. Interfaces [Cra98, Rei90, AS99, BBB99, Gli90]. interfering [MWA^{+}92]. intermediate [BMSW91]. Intermittency [JMK99]. internal [Gri97, MPB91]. International [Cyb90, Ano96r, Ano96k, Par90]. Internationale [Reb93]. Internet [CLM98]. internucleotide [BSB99]. interplanetary [VGM90]. Interpolated [Aok97, IY95, UKA97, NY99]. interpolating [SKWH99]. interpolation [BRB97, UTI94, YA91, YIW^{+}91, YJ92]. interpretability [MB91]. interpreting [DATL99, JH91b]. intersection [Kud99, LL90]. interstellar [HFD90]. interstitial [HMH99]. interstitial-substitutional [HMH99]. interval [HS96, SSS91]. intrinsic [FR90]. Introduction [Pet92, Cab90, NK91, Par99, Bis91, Ryb98]. Invariant [dAASZ97, LVC^{+}94]. Invariants [SPP^{+}98, FFB99, Fra92]. Inverse [Par99, AP97, AP98, BMT94, HHR98, XZ91, ZX91, vMC91]. Inversion [BI90, Iqb95a, Iqb95b, TC93, WCJ92]. invert [CNO98]. investigate [Ber96]. Investigation [Mac91, RF96, BN92, BBB99, BC99a, KW90]. inviscid [VB91]. involving [BMSW91, HS96, MÖ93, SSS91, Tan90]. Ion [AG99, NOC98, ABH99, CM92b, GBWES97, HRK99, Hor91, IST99, KI99, LPR^{+}93, MLB93, MF94, MÖ93, OC98, PMM98, RT98, SKG91, Szm95, YN94, Zha98a, Zha98b, Bar93, CM92b, BPL97, YN91]. ion-atom [CM92b, Hor91, MÖ93, NOC98]. Ion-Atom-Wave [NOC98]. ion-induced [IST99]. ion-metal [MF94, SKG91]. ION2 [MMC95]. ionic [BPL97, FT99a, LPR^{+}93, MC93, TBA^{+}98]. ionisation [CM92b, MMC95, Ste97]. ionization [AO97, Bar93, BM92a, Bra98, BG94, CKW98, MSP^{+}98, NOC98, OC98, PMM98]. ionized [LLW98, MP90a, SBK98]. ions [BZO92, MKWH91, MSP^{+}98, SR98, SRFN92]. iron [AZM92, Men99]. iron-core [AZM92]. irreducible [Ber95]. irregular [BM91]. Irreversible [Alb99]. ISBN [Art93, Art95a, Art96, Bis91, Bon93, Bro92, CT96a, Cyb90, Deb92, Eas93, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Hop97, Rai91, Reb93, Ryb98, Top96, Tru93a, Tru93b, Weg91, Whi93]. ISICS [LC96a]. Ising [PTU97, Bai91, DER90, FV99, Heu90, JBB99, JV99, LHA95, MR95, PC99, Sel96, SA99b, SPA99, Sou99, SVV99, VBS96, dQRdS99]. Ising-like [PTU97]. ISO [Krö90]. isobaric [CF95]. isoelectronic [MSP^{+}98]. isolated [BWS94, DL98]. Isolation [Hel99, Eva90]. isomeric [CMC94, VST94]. Isothermal [CF95, ŠČ91]. Isothermal-isobaric [CF95]. isotope [FSMVM93, JF97]. isotropic [BLGT92, CP95]. isovalue [SW91b]. issues [CDD^{+}96, MKS^{+}96]. ITAP [YN91]. iterated [SZ99, Wen91]. iteration [BRB92, KN95a, SL97]. iterations [AP97, RDB^{+}90]. Iterative [HRK95, MC97, AP97, Mal97, MDSF^{+}99, PZH^{+}96, SM97]. IV [PC94c, Sal91, YKM^{+}96]. ix [Fin90, Gie90, MSW92a].

J [Bis91, Bon93, Cra98, Deb92, Fle92, Whi93]. jackknife [Ber92a]. Jackson [Whi93]. Jahne [Bro92]. JASP [HN97a]. Java [Tob99]. JavaFit [Rob98b]. jellium [Ber90]. Jet [Gra95, FMH93, LPPR91, MH96, Pöt99]. JETNET [LPR92, PRL94]. jets [BB93a, KKK91, KM91, Sig97]. JETSET [Sjö94]. JetViP [Pöt99]. jj [FG95a, FG97, Kag92]. jj-coupled [FG95a, FG97]. jj-coupling [Kag92]. John [Top96]. Johnson [FJ95a, JF99]. Johnston [Top96]. Jones [BDKM91, RBG99, Zyb99]. Joseph [Ryb98]. Josephson [GV99]. journal [Ano96k]. JUMNA [LZS95]. jump [HI93]. jumps [AA99]. junction [GV99, HP97b, LZS95]. June [Cyb90, Whi93]. justification [VP96].

K- [LC96a]. K.J.F [Nar92]. Kac [RDF^{+}97]. Kadanoff [KKY99]. Kalman [LG93a, SZ99]. kaon [Di 95]. KAPSIES [KA95]. kB [AO97]. KdV [FG90, Li95]. KEK [Mor98]. KENO [RP95]. KENO-Va [RP95]. Kern [Reb93]. Kern- [Reb93]. kernel [BLS96]. kernels [Bra91]. Kerr [SRR95]. keV [BBG94]. keywording [Ano96-31]. killing [Wol98b]. kinase [GPM95]. kind [Abd90, SG99, Wee92, WC90, YM97]. kinds [Lee90a]. kinematical [BPP98]. Kinetic [ALR99, BBL^{+}99, DER90, KB98, KCC96, KRT95, LP99, RCL98, USW99, VB92b]. kinetic-type [KCC96]. Kinetics [DCM99, Pur99, AR95, Aro96, Pag94]. KINX [DMM^{+}97]. Kinzel [Oln98]. Kirkwood [Sha92]. kit [Whi98]. KL [MCCC99]. KLMN [CCM95]. Kluwer [Cra98, Rai91]. knock [Di 95]. knock-on [Di 95]. Kohn [J̈K97a, MCV^{+}90, AK99, BDF^{+}97, J̈K97b]. komplexen [Weg91]. Kong [Fly90]. Koonin [Bon93]. KORALB [JW91, JW95]. KORALW [SJPW96, JPS^{+}99]. KORALZ [JWW91, JWW94]. Korringa [MCV^{+}90]. Korteweg [HAR91, Var93]. KPZ [HCO99, SDB96]. Kr [PI95]. Kratzer [TS92]. Kree [Art96]. KRONOS [ADM^{+}92]. Krylov [KM95]. Kuba [MSTK93]. Kubo [Ver99]. Kudrnovský [Cra98]. KUIP [Zan90]. Kummer [BMOF92]. Kutta [Sim98, VDVV99].

L [CT96a, Deb92, Ryb98, LC96a]. L- [LC96a]. Lab [TG99]. laboratory [BBP^{+}90, Hal90, MG90]. lag [Sim93]. Lagrange [Whi93]. Lagrangian [Kre99, PC90, Sem98, VW95, XYE99]. Laguerre [CS99b, KL92, POA92]. Lake [Whi93]. lamellar [MBMM99]. laminar [FMH93, NN93]. Lanczos [BI90, CNO98, CGP94, JBH99, Kal96, Kal94, LGZL98, Zha98c, ZK99]. Landau [Fin90, Fin91, BBL^{+}99, HHGCM99, STR91a]. Lang [Reb93]. Langanke [Bon93]. Langdon [MV97]. Langevin [GLF97]. Langmuir [DPRS90]. language [BDKM91, BM91, Ein90, Hor90, Ohl92, PDL98b]. LanHEP [Sem98]. Laplace [Iqb95a]. Large [Chr98, CSP92, Fer97, GUW98, MY92, MHHL99, PZ91, AMM^{+}94, BAAD^{+}97, BCP91, CTS99, CHW91, CD91, CD92, DJMT91, FGR98, Gra98, IL91, JGP93, KH95, LHA95, MIO96, OM93, PZH^{+}96, PFG96, PG96, RSW^{+}99, Sci93a, Sci93b, Sci94, SW91a, SP95b, SF94c]. Large-eddy [PZ91]. Large-scale [CSP92, MY92, PFG96, SW91a]. Larrouturou [Jon90]. laser [CP92, CJ91, CKW98, DeV91, DG99, DNSP99, Eng91, FS91, GDRR94, MHN99, SZG99, SPT98]. laser-atom [FS91]. laser-driven [SPT98]. laser-induced [CP92]. lasers [Kot92]. LASPE [HM92]. Lattice [HCO99, Hoe90, KKH^{+}98, AUG92, AB90, BdT97, BMT94, BHM99, BB92, DL91, DCM99, FFG^{+}96, HMH99, HGCM98, IK99, JL97, KS96, KK99, Lil93, Lüs94, MCO99, Mon91, Neu99, Pae91, SC99b, TBA^{+}98, TT99, VVIBA99, Wie96, WK90, dlSG99]. Lattice-Boltzmann [KKH^{+}98]. lattices [CHM91, DER90, JV99, LF92, Sto98, TFY95]. Laue [Rei93]. law [Lan92, MP95]. laws [VPE94]. Layer [MCV^{+}90, BF98, WCK99]. layout [Eic90]. LCAO [MNS99]. LDCMC [KL99a]. lead [RLF96, RARF96]. lead-silicate [RARF96]. leading [FJKvO95, KM91, Sig97]. Leap [Li95]. Leap-frog [Li95]. learn [dSB99]. Learning [BCSS99, AFP99, CS91, FT99b, SW99]. Least [Bru91, Ber92b, BDL92, CGR93, Dem97, HHK92, LH97, Mar91, MK91a, Nie94, Rob98b, Swa91]. Least-squares [Bru91, BDL92, CGR93, HHK92, LH97, Mar91, MK91a, Rob98b, Swa91]. Lecture [Art96, Jon90, Whi93]. LEDs [Kot92]. leg [Pit97, Pit98]. Legendre [GS97]. Lehmer [DH97a, DH97b]. lemma [Kle92]. Length [BG98]. Length-scale [BG98]. Lennard [BDKM91, RBG99, Zyb99]. Lennard-Jones [BDKM91, RBG99, Zyb99]. LEP [AMM^{+}94, BPK95, FKMR96, JW90, JWW91, JWW94, MNP96, MNPP99]. LEP/SLC [JW90, JWW94, JWW91]. LEP2 [BPP98, Pap97]. LEP2NLC [SJPW96]. LEPTO [IER97]. lepton [ALR97, CSS94a, FR96, GW95, IER97, JWW91, JP92, JWW94]. lepton-nucleon [IER97]. lepton-proton [ALR97, CSS94a]. leptons [JKW91]. leptoproduction [MSV92]. leptoquark [Pap99]. LERG [Stu95a, Stu95b]. LERG-I [Stu95a, Stu95b]. LESKO [JP92]. LESKO-F [JP92]. level [AA98b, DLLR96, GRRG95, Lam96b, MOSM96, NK96, SCS^{+}96, SL94b, Sun97, Top96]. levels [STR91a, SGRZ99, Sch92]. Levenberg [Rob98b]. Levin [HW95, BRB97, RBB96, RBB98]. Levin-type [HW95]. Lewis [RDMS95]. lexicography [SK90]. LGA [BMPS99]. LHC [Asa98]. libraries [Fro91d]. library [Ano96r, Ano96k, BLGT94, CDD^{+}96, GH91b, Hin95, JW91, JKW91, JWDK93, JWJK92, Lön94, PB93, RFH^{+}95, SB97, Sty90]. lid [SvdPS99]. Lie [Can94a, Wyb94, Hea93, Hea96, AH97, AAH98, BdM93, Can94b, CHW91, CdG96, Rod92]. Lie-series [Rod92]. Liège [MG90]. Liesegang [CGRV99]. life [GDGR97]. light [GLF97, GFS99, HTI99, MHN99, MEA^{+}90, SRFN92, TC97]. like [BB99, CZ92, Fau91, HKC98, Pap92a, Pap92b, PTU97, RAD99b, RS99, SP93a]. likelihood [BCD^{+}93, GKP93a, GKP93b]. limit [KL99b]. limitations [Ker90a]. limited [YK95]. Limits [KKKT99, Ber92b, BR97b, HN97a, HK97b, Jor95, vHKH97]. line [BRWN96, BWNR99, BS96, DVH90, DHK^{+}98, FBdC98, GY99, Gut92, JP92, KTM96, KDH^{+}98, MKWH91, Nak91, SSBA91, VGK90]. line-spacing [KTM96]. line-tied [DVH90, VGK90]. Linear [AAD92b, GHGB97, Abd90, Ale93, AAD92a, BMT94, BG98, CY95, CDD^{+}96, CMC^{+}99, Duf96, FJ91, HD96, JD91b, JBH99, Jor95, Kno90, Moo93, Ohl97, OPP96, PD96, Rod92, SS90a, SFFDR94, Tau95, ZN93]. linear-analytic [CY95]. Linear-scaling [GHGB97, BG98]. linearized [BSST90, KWS^{+}96]. Linearly [HKG90]. link [MN90, MN92]. link-cell [MN90, MN92]. linkage [BSB99]. linked [MR99a]. Linux [AO98]. Liouville [YKM^{+}96, BK91, HK95, IDV99, PM91]. Liouville-type [YKM^{+}96]. LIPAOL [PP92]. lipid [FHSO99, LDOO95]. liquid [All99, BSSH99, DBGY99, HMR99, MDMR99, OA96, RLF96, dM94]. liquid-drop [MDMR99]. liquids [CKS91, ME91]. LISP [BBJ91, CT90]. List [Ano90r, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano97j, Ano97k, Ano90g, CD90, CD91, CD92, MR99a, Ryc90a, Ryc90b, SA92, SY95]. lists [FL91, SF93]. lithium [SAP^{+}95]. LMD [OC98]. LMTO [Yeu91]. load [AL94, BAAD^{+}97, BF90c, FH97, MSW92a]. Local [CvN99, FBB99, MCMD91, Pap92a, Pap92b, PJ98, PvBMN99, PCNO94, PCC94, PC94b, PC94c, Siv91, TZC96, VB92a]. Localised [HP97a, SGF99b]. localization [CGdlR99, CBG99, SMR99, dVdRL93]. LOCFES [JN94, NE93]. LOCFES-B [JN94]. Lodge [Whi93]. Log [JD91b, FJ95a, JF99]. Log-derivative [JD91b, FJ95a, JF99]. logarithmic [GJ99a]. Logical [Eic90, Wei93]. London [Fly90]. Long [SC99b, SEZ99a, Tru94, dMCT99, Ess95, Frü95, Gra92, WCL90, YB92]. Long-range [SC99b, Ess95]. long-ranged [WCL90]. Long-standing [Tru94]. long-tailed [Frü95]. longitudinal [HLMM99, HKLM99, OC98]. longwave [MSJ99]. look [DDDL98]. look-up [DDDL98]. loop [AUG92, Avd96, BS98, BBR^{+}90, BFK95, BFK97, EHK^{+}92, FT92, FJKvO95, HPV99, MS98, Pit97, Pit98, SGT90, Wan93, vO91]. loop-diagrams [Wan93]. loops [AUG92, DVH90, Goe90, PGK90a, PGK90b]. loose [BC91a]. loose-coupling [BC91a]. Lorentz [HL92, Sti94]. Lorentzians [HS96, SSS91]. loss [FFS98, LVL90, SRF94]. losses [BG94]. Low [Hew94, HM99, Rob97, TCE99a, AS92, Ask94, FKFE99, JRWWW92, JPRW^{+}97, KS96, LC93a, RF94, TBCH90, VB92b, ZX91]. Low-dimensional [Rob97]. Low-frequency [Hew94, TCE99a, VB92b]. low-lying [KS96]. Lower [FR90]. lowest [Zei97]. LP [BTC99a, Shu94]. LS [FG91, SF93]. LS-coupled [SF93]. LSC [OA95]. LSGEN [SF93]. LSJ [FG91]. LUCIAE [ABH99, ST95]. Lüscher [Jam96b, Jam94]. LVQNET [CO92]. Lyap [Sat97a]. Lyapunov [Sat97a]. lying [HM99, KS96].

M [Cra98, Eas93, Rai91, Whi93, LC96a, PG96]. M-shell [LC96a]. M. [Par99]. MacCallum [Eas93]. MacDonald [Art95a]. MACH [AHH94]. machine [BBJ91, IMES90, MMGS93, MLM96, TR93]. machines [BAS99b, BCOY93, ERB90, GS98b, KdLNV93, LBC91, RDB^{+}90, WDB^{+}91]. Macintosh [DDK99]. Macramé [HTZ^{+}98]. macro [Tan95, WENG95]. macro-EM [Tan95]. macro-molecules [WENG95]. macromolecular [SF94a, SF94b]. macromolecules [ERS^{+}95, GPM95]. macroscopic [BL99a, Wie95]. MACSYMA [Her91]. MAG [DGK97]. MAGIC [GLSW95]. magnetic [Aly90, BG95, Cam95, DJUM99, Dub91a, FBdR^{+}98, GFS99, HKG90, JWF93, Lee90b, Lin90, MGMR97, PSR99, PB96, SMBV90, Tan95, TSK^{+}97, VB96, VGK90, dMCT99, vMF94]. magnetically [Pai93, WBB99]. magnetism [BHS^{+}99]. magnetization [HN99]. magnetized [MAH98]. magnetoconvection [NS90]. magnetofluid [GHM93]. Magnetohydrodynamic [SMBV90, CHMG92, LL96, NvdHPG97, PL99, Sto99]. magnetohydrodynamical [HFD90, ZY97]. magnetohydrodynamics [KCC96, Meg94, YM92, Zie99]. magnetoplasma [SBK98]. magnetospheric [KV90, Ott90]. magnetostatic [BPG97, YB99]. magnetovolume [GE99]. magnetron [IKTY99]. magnitudes [NK99]. major [Goe99, RI97]. Majorana [RI97]. make [Par90, YB99]. Maley [Met96]. management [BBE^{+}98, Deu99, MM98]. Mandelstam [HvSA98b]. manipulation [AUG92, Lio97, RU92]. manipulations [CT90]. Manneville [Gie90]. manuscripts [Ano94c, Ano94a, Ano94b]. Many [LBDL99, PHN99, BMSW91, BK92, BKS99, CPCF99, Lia95, xLzSlLzH94, LMW94, MSW92b, MSW92a, Ran90, RLER97, RAD99b, SLGB96, SY95]. Many-body [LBDL99, PHN99, BMSW91, CPCF99, Lia95, LMW94, MSW92b, MSW92a, RAD99b, SY95]. many-centre [RLER97]. many-dimensional [BK92]. many-particle [Ran90, SLGB96]. map [Bor99, SS91b]. Maple [Ryb98, AH97, BS99a, CT94a, CSS94b, CC98, FFB99, Fri97, FVGF98, GS98b, Gro91, Hol96, PS97, Rod92, SW98, CT96a]. Mapping [GG96, DD95, MS91b, vMC91]. mappings [TGG97]. maps [BGT95, Leh99]. Marder [MV97]. market [KKKT99, Sol99]. markets [Man99]. Marquardt [Rob98b]. MARS [LB99]. Maruhn [Bon93]. MasPar [PS98b]. mass [Avd96, BF98, Cor99, PI95, TAMSW90]. masses [AFKT97]. massive [Tan90]. Massively [Mic94, NVK94, AS95, CGM^{+}96, GPSM95, LMM^{+}97, MTR94, MTNA98, PS98b, SG91, TR93]. master [SB97]. materials [BM99a, BPL97, BC91a, Kis91, LPR^{+}93, Nak91, NKV94, Nak97a, RLP^{+}99]. Mathematica [Oln98, Stu95b, Tay99, Bes98, DM98, JL93, MS98, Stu95a, Var97]. mathematical [BH99, KRVZ98]. MATLAB [Tay99]. matrices [BG99, Gra92, HSSB98, HE91, Koo93, MCV98, MIO96, OM93, PZH^{+}96, PCC94, Sci93a, Sci93b, Sci94, VN98]. Matrix [PZH^{+}96, Sey95, AAS95, AH92, AM94, AH99, BD94, Bar93, Bar98, BMT94, BEN95, BHP97, BBS92, BN94, BN95, BI90, CP95, CGHB98, CK99, Du93a, HEM95, HB90, JY97a, JY97b, KNK95, Lim91a, LPSM92, LC93b, MHN99, MF95a, MR96, MTG98, Ome98, PDZ92, QBP98, SP93a, Sch93, SGRZ99, SMB^{+}93, SSBA91, SF94c, ST98, SHNS98, TGAW96, TS92, VSS91, Var97, Wan99a, Wil90, WCJ92, Zat96, dCdJDvL99]. Matrix-element [Sey95]. Matsubara [Nie95]. Matter [Tru93b, AH99, Bin99, Fin90, Fin91, RJ90, Ryc90b, RJ91a, RJ91b, WRF91]. Maximum [ACA99, GKP93a, GKP93b, RVCF99, VFO94]. Maximum-entropy [ACA99]. Maxwell [BDD^{+}92, ECT93]. MC [Gei97, Kre97, LN92, TP99]. MCBETH [OA96]. McComb [Art93]. MCHF [Fro91c, Fro91d, FG91, FJ94, JWF93]. MCRG [Bai91]. McTrap [MOSM96]. MD [BMPS99, RARF96, WCL90]. MDPYRS1 [RJ90]. MDPYRV1 [RJ91a]. MDScope [NHK^{+}95]. MDSLAB1 [RJ91b]. MDSPNL [Ryc90b]. mean [BVR95, Jac91, PVR96, PVRR97, Pös98, RGL97, Rou95, SP93b, US91]. mean-field [BVR95, Jac91, PVRR97, US91]. means [BHM99, Gro90, HMH99, Lio97, LG93a, PMUV90, SS93a]. measurable [BPP98]. measured [NCC^{+}96]. measurement [Kot92, Tar99]. measurements [BG95, HLMM99, HKLM99, JY97a, JY97b, UP90, VdN92]. measures [CvN99, JHK97, MT91]. Measuring [IJV92]. mechanical [BPL97, CDUT99, Gil99, JGP93, LPR^{+}93, NL91, Scr95, TSJ91]. mechanics [CT96a, Fly90, Fly93, LLF97, MKL99, PCC^{+}95, SZS95, Top96, Sul96]. mechanism [BM99c, HMH99, KCK99]. mechanisms [BM99a, PMB99]. mechanistic [You99]. media [ADK^{+}99, Chr99, GJ99a]. media-focused [Chr99]. medium [FB99, Goz91, KK99, LVL90, VGM90, ZL99]. Mellin [Iqb95b]. Mellor [KDH^{+}98]. melting [CP92]. melts [BG95, Krö99, WS95b]. Members [Ano99o]. membrane [FHSO99, KS93]. membranes [TF98]. memory [BTC99b, BCOY93, CDD^{+}96, HTI99, KdLNV93, LBC91, MTM92, PJ98, Pet92, RBS92, SHNS98]. memory-economical [BTC99b]. MENLO PARC [Sig97]. mercury [LC93a]. mercury-noble [LC93a]. Merlin [PDL98b, PL90, PDL98a]. MERLIN-2.1 [PL90]. Merlin-3.0 [PDL98a]. MESFET [Fau91]. mesh [Art94, Eas91b, IKH^{+}99, KZS95, KH96, MD96, PC90, Zie98, Zie99]. mesh-handling [KH96]. meshes [ADS92]. meson [ALR97]. mesons [WA94]. Mesoscale [GY99]. Mesoscopic [BL99a, BL99b, FBVdlN99, PFBdlN99, vVF96]. message [BvdSvD95, CA92, MAB^{+}94]. message-passing [BvdSvD95, CA92, MAB^{+}94]. metal [AI90b, BHS^{+}99, CB99, FB94, HP97b, HB99b, MF94, MSTH92, Mic94, SKG91, TSL^{+}99, ZFZ92]. metals [BdT97, CDUT99, RGL^{+}98]. metastable [SvdPS99]. Method [Tru93b, Whi93, AH92, AH91, AM90, AH99, ADS92, AZM92, AMS96, BKWM91, Bal95, Bar93, Bar98, BCD^{+}93, BF91, BWS94, BC99b, BK93, BS95, BHP97, BFK95, BNS^{+}95, Bur98, CW92, CL96, CY95, CS98a, CC97b, CGHB98, DSD91, DWL91, DD97a, Dun96, Eas91b, EHKK93, FS91, Fan94, FF97a, FMSdC99, Fuj94a, GRRG95, GH91a, GH91b, Gro90, GPM95, HMH99, HS93a, HHR98, IY95, IKH^{+}99, IIM92, ITMI95, Iqb95a, IWWY92, Izv99, JF99, Jaq90, Kal96, Kal94, KCC96, Kob94, Kor99, Kro92, Kud99, Kul97, KL92, KN95a, KN95b, LGZL98, Lem96, LBB91, Lin91, xLzSlLzH94, LG93a, MY92, Man91, Man93, MC97, MR99a, MD97, MSTH92, MIO96, MK96, MN97b, MN90, MT92b, MK91a, MK91b, MKPGS97, MH99, NL91, NK91, NKR99, Ole96, Ome98, PC90, PD96, PP99]. method [Pit97, Pit98, PW99, PG96, Pop98, PVR96, PSV99, RdC93, RDF^{+}97, RSW^{+}99, RHFR95, Rob98b, RH99, RW95, RSW96, SGF99a, SdCR97, SBGM91, SvdPS99, SSBA91, Sim98, SZG99, SKPP95, SAZ90, SGF99b, ST92, SZS95, Sul96, Szm95, Tan95, Tau95, TC93, TFY95, USW99, UKK98, VS95, VGK90, VS97, VES99, VVGI99, Wan99b, Wan99a, WP98, Wee92, Wee93, WZA^{+}98, WJ94, WJ95, WK90, WCJ92, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90, YY99, ZN93, ZZGS91, vMC91, Li95]. methodological [DPP^{+}99]. methodology [ZA99]. Methods [TB91, AM94, Art96, BS91a, BKMM93, BBMD99, BH99, BMT94, BBCM93, Bil91, BP93, BCP91, CRS90, CTDdM97, CTDdM98, CP95, CJ91, DEAM91, Deu99, DBVS99, DHL93, Duf96, FPV97, FZZ92, GHM93, Goe99, GPSW95, GPSM95, GJ99b, HZQ93, HB99b, Hut94, ICS95, IL91, IRV95, IVDV97, JD91a, JD91b, JF92, JN94, KM95, KML98, Kre99, LLF97, LB95, Les99, LLM91, LP99, MV97, MMW91, Nad90a, NBJK91, OPP96, PVRR97, RB99, Reb93, STJ90, Sim92, Sim93, Sim97, Sim99, Siv91, SG91, SHNS98, Swe91, Tay99, TK91, US91, VV99, VDVV99, dCRSF99]. metrics [CGK^{+}93]. Metropolis [TFY95]. Mg [MSP^{+}98]. MHD [AMS96, CNO98, CZ92, DVH90, DMM^{+}97, DGK97, GA90, HS95, HPS92, LB95, Lee90b, LB99, LBR92, LBS96, Ott90, STJ90, YM92, Zie98, dFZLS95]. MHD-like [CZ92]. Micro [Fan90, Mic94]. micro-crystallites [Mic94]. Micro-VAX [Fan90]. Microcanonical [CK99, DM91, Ran90, DG93, FR95, dO99]. Microcomputer [Pes93, RSW92]. micromagnetic [Fer97]. microphases [BHM99]. microprocessor [MES90b]. Microscopic [BD99, CUVS95, LD91, Var97]. microstrip [BS91b]. microstructure [Nar96]. microtasked [MSTK93]. microwave [KMT^{+}99]. Mie [DS92]. million [Rap91a, Rap91b, Rap93]. Milne [YYK^{+}92, YKO^{+}94, YYOM90]. MIMD [KdLNV93, WLD95]. mind [Duc96]. mini [MH96]. mini-jet [MH96]. minimal [Sim93]. minimisation [LZS95]. minimization [Smi92a]. MINUIT [LBC^{+}93]. MIRACLE [TG99]. MIS [Fau91]. MIS-like [Fau91]. MISCOMP [Whi98]. Mixage3D [MLM96]. mixed [Bru95, Ein90, HT93]. mixing [JY97a, JY97b, MSR96]. mixture [GGJG97]. mixtures [Car94, Car96, LH92, MB94]. mKdV [FG90]. MLOG [Car96]. mo [ERB90, FZZ92, RDB^{+}90, WDB^{+}91]. MOCCT [HS95]. mode [CKW98, PCC^{+}95, Siv91, TCE99b, You99]. Model [ABH99, ABE^{+}94, BGK^{+}94, Bai91, BS96, BHM99, BKKS99, BJK92, CCM95, ÇYG99, CP92, CT97, CBB99, CHM91, CK99, DP91, DG93, DATL99, DER90, DJUM99, DM91, DPRS90, FMSdC99, FMH97, FZZ92, FV99, Gal99, GBWES97, GDRR94, Ged98, Gib92, GLF97, HL92, HN99, Hor91, HPS92, IBR98, IMA94, IDW99, Jan99, KD95, KK99, KIMS98, KZS95, KKH^{+}92, KKKK99, KR95a, Kni90, KH96, LL99a, LBB91, Lim96, LC93a, Lön92, LHL98, LC93b, MSD93, MBZM99a, MBZM99b, MCCC99, MPT98, MR95, MB91, MP99, NSE95, PCSD99, PMB99, PR93, PAC^{+}98, Pur90, RASS99, RVB99, ST95, SC99a, SDB96, SMR99, SC99b, SA99b, SPA99, Sou99, SS91b, SA99c, SVV99, Sug99, SH97, Sun97, Swa91, Szo99, VS95, VB96, Var97, VB92b, WP96, DKS98, SH97]. model-based [MB91]. Modeling [AGIS99, CPM99, Rho97, WNS99, Aut99, FKKJ91, GY99, MAB^{+}94, MTH^{+}99, NKV94, PMUV90, RJS99, RBP^{+}90, RMMUJ99]. Modelling [MSR96, AS95, Art94, Art95b, BC91a, CMSK99, DJMT91, For91, Gie90, HE99, MLB93, MD96, MMGS93, MLM96, McD90, MS91b, Nar96, Rai91, Rob97, SF98, TSJ91, WBB99, You99]. models [ADK^{+}99, AAD92a, App99, BGCRRO99, BP90b, BB92, BD99, BC99c, CVZ99, CMV99, Cor99, FMH93, GLV99, HCO99, Hof99a, IJV92, Lam98, LH97, MDMR99, MMU99, Neu99, Ome97, PI99, PTU97, Sci91, Sel96, Smi92a, TMG99, VG99, ZK99, vVF96, Deb92]. Modern [Nad90a, Deu99]. modes [GW98, JWJK92, Mac91, TCE99a]. MODFETT [Fau91]. modification [Cop91]. Modified [Zha98c, BGK^{+}94, BSB99, CGP94, GC95, HAR91, MR99a, SdCR97, WC90]. modular [BvG97, CC99]. modulational [Riz99]. module [CDG^{+}90]. modules [DLLR96]. MOIL [ERS^{+}95]. Molecular [BST95, BSB99, CP91, DCR99, DW99, FKKJ91, ISS91, IK99, KdLNV93, LMM^{+}97, LBC91, LMW96, Luo96, MKL99, RJ90, Ryc90b, Smi91, TMS^{+}99, Ale93, Alf99, AF90, AM90, BKDMM91, BDKM91, BWNR99, BvdSvD95, BBG94, BvG97, BKNS97, BT92a, BC91a, BCOY93, BCOY94, BCOY95, BMM97, BF90b, BF90c, Bru95, CGM^{+}96, CD90, CF95, CDR91, CGHB98, CSS94b, DSD91, EK94, FHSO99, FJP90, Fuk94, GSS97, Gup92, HE94, Han90, HNR90, IOM95, JAT97, JBSZ97, JF92, JF99, JH91a, KB99, KLS91, KF99, Kut99, LDOO95, LHSD95, Lil94, MC95, Mar97, MTM99, MPRS97, MSFR96, MS95a, MHHL99, MT92b, MTG98, ME91, MP90b, MPB91, MP93, MPSvG94, NVK93, NKV94, NVK94, Nak97a, Nak97b, OMC93, OPP96, PBL94, PCC^{+}95, PLC91, Pes93, PS98b, PMM98, PK98, RT98, Rap91a, Rap91b]. molecular [Rap93, RJS99, RLP^{+}99, RF94, RLPR91, RLPR94, RLER97, RLF96, Ryc90a, RJ91a, RJ91b, Sal91, SSPD92, Sat97b, SGB^{+}93, SA92, SSH94, Ská90, SZG99, Smi92b, SAJ^{+}97a, SAJ^{+}97b, SM91b, TC97, TSJ91, TK91, TD99, TvG91, TOSM91, TRSM98, Tox99, ThLL^{+}93, VW95, WR99, WF90, YBS91, Zyb99, dM94, KKKS95]. Molecular-dynamics [KdLNV93]. molecule [BKMM93, EHKK93, Jac91, MK91b, Pop98, SKST91, Ten97]. molecule-corrugated [MK91b]. molecule-molecule [SKST91]. molecule-surface [Jac91, Lem96]. molecules [BMSW91, CKW98, CP91, DMLM91, Eng91, GW98, HD96, HE91, JWM^{+}96, KLS96, LGZL98, LMW94, MBW^{+}95, Mal90, Mit96, MSW92b, MSW92a, NP96, Nor94, PCC^{+}95, Pop96, SG98a, Sch91b, Sch92, SvNF92, TML93, THF95, WENG95, dlVM98]. MOLFDIR [VVA^{+}94]. Moment [KKK91]. momenta [Fri97, FVGF98, Ran90]. moments [AMS96, NK91]. momentum [FS92, Koi92, KPD97, OC98, SS90b, Tak92, VS97, Wei99, dlVM98]. Mon [Fin90, Fin91]. monitoring [CLM98, SW91a]. monoenergetic [MFG^{+}98]. monolayers [RLP^{+}99]. monomial [BEFJ99]. monotonical [Wei93]. Monte [LR92a, RP95, Tru93b, ALR99, Alb99, AL94, AAD92b, APN^{+}91, ADM^{+}92, AMM^{+}94, ABH^{+}94, ALR97, BN92, BvEW91, BW94, BB93b, BM99b, BS91c, BSW97, BC99a, BPK95, Ber92b, BMT94, BS90, BHM99, BBCM93, Blo92, BC99b, BKW97, BKS99, BRM99, BVVJA99, BLLM94, BG94, CMNP95, CC97a, CDG^{+}90, CSS94a, CMNP97, CHM91, CF95, CvN99, Cop91, DB93, DS90, ES92a, ES92b, Fan94, FR95, FF97a, FKKJ91, FR96, FBGV99, FCS99, GP95, GJ96, GE99, GGMS95, GW94, HP93, HMH99, HSP93, HCVM99, Heu90, HKL93, HN99, HK96a, HK96b, HK97b, IRS97, IER97, IMG99, JW90, JWW91, JKW91, JRWWW92, JP92, JWW94, JPRW^{+}97, JPS^{+}99, Jan95, JLFS91, Jun95, KM91, KM98, KTT98, KB94, KP94, KF99, KR99, LMF97, LPS99, LM95, MMH99, MSV92, MWA^{+}92, MBC99a]. Monte [MC97, MPB96, MEA^{+}90, Mit96, MH96, MNP95, MNP96, MSJ99, NP96, Odo90, Odo92, Ohl99, Pai93, PI99, Pan93, PP99, PK99, Pes93, Poz91, RB99, RI97, RKV91, RH99, SGB99, SMV90, Sci91, Sel96, Sik99, Sin90, SWT91, SJPW96, SL99, SF98, Sti94, Swe91, TLK94, TFY95, VS95, VVGI99, VVIBA99, Wag92, WM99, WW90, Wan99b, Wan99a, WCK99, WS95b, WBB99, Zhe99, dCRSF99]. Moose [Bos98]. MORATE [HLL^{+}95, ThLL^{+}93]. Moriarty [MSTK93]. morphological [HB99b]. morphology [CDR91, HNS99]. MORPHY [Pop96]. Morse [LM91b, LPSM92]. MORSMATEL [LPSM92]. mortality [RdMP99]. MOSFET [RGLV^{+}99]. MOSPLV [Chi90]. Mössbauer [BNSW92, Chi90]. motion [CMC^{+}99, FG90, GP95, HTI99, KH97, MSTH92, Ome98, VV99, WM99, vGHM^{+}95]. motivated [MDSF^{+}99, dSTC99]. Movement [CK91b]. moves [Sin90]. moving [DGK97, MR95, Ris96, ŠČ91]. MP [MSTK93, Len98, Por90]. MPI [FH97]. MPPs [HSSB98]. MS [Top96]. MS-DOS [Top96]. MSSM [KM98]. MSW [KCK99]. Muller [Bis91]. Mullges [Hop97]. Multi [JBB99, MPB91, Rap91a, Rap91b, Rap93, BK93, BF90c, BHL92, Car96, CJ91, Fro91b, GHE99, GRS90, Kaw95, KA95, Kup99, LLW98, MK95, MSW92a, MP90b, PI95, Pit97, Pit98, PM91, RL98, SGB^{+}93, SSW94, Szm95, XYNI96]. multi-channel [Szm95]. multi-color [CJ91]. Multi-colour [MPB91]. multi-configuration [Fro91b, PI95]. multi-dimensional [GRS90, Kaw95, Kup99, XYNI96]. multi-electron [GHE99]. multi-grid [BK93, SSW94]. multi-leg [Pit97, Pit98]. Multi-million [Rap91a, Rap91b, Rap93]. multi-node [BHL92]. multi-nuclide [Car96]. Multi-overlap [JBB99]. multi-phase [MK95]. multi-processor [MSW92a, MP90b]. multi-purpose [PM91]. multi-session [RL98]. multi-step [KA95]. multi-term [LLW98]. multi-transputer [BF90c]. Multichannel [Zha91, Du93a, Du93b, KP94, MF95b]. multicolour [Kob94]. multicomponent [vVF96]. Multiconfiguration [KNN91]. Multicontact [Rad99a]. Multicriteria [Tur99]. Multidimensional [CT97, JHK97, Ves98, All93, APaI99, BM95, GGSR94, Kle92, Neu99, PDL98a, RBS92]. multielectron [MKWH91]. multifaceted [CR98]. multifractality [dSVTM99]. multifractals [STI99]. multifragmentation [LR92b, Ran93]. Multigrid [LLM91, Siv91]. multilayer [DP91]. multilayers [ALR99].

Multilevel [Bra91, Nak97b]. multinode [MLM96]. Multiparticle [Sin90, Chi92, DCM99, HP93]. multiphase [Zlo95]. Multiphonon [Man94]. multiphoton [CJ91, CGHB98, KR91, LM91b, PMM98, Pot98, TYJ92]. Multiple [KPD97, BR95, Ber99, CDG^{+}90, CD92, Cop91, Far98, FG95b, JW90, KK96a, MMS93, NVK93, Por90, VP96, HUCS98]. multiple-processor [MMS93]. multiple-time-step [NVK93]. multiplication [KNK95, PDZ92]. multipliers [DH97b, Kre99, SL99]. multiply [YK95]. Multipoint [RBB98]. multipole [KN95a, KN95b, Lud91, MR96, SKPP95]. Multiprocessor [Len98, AL94, SWT91]. multiprocessors [MTM92, Pet92]. Multiquadric [All93, HHR98]. Multiresolution [Leh99, NKV94, CNPT91]. multiscale [Nak97a]. multiscaling [CZ99]. multispecies [MMTK94]. Multistate [Sal91]. multistep [ICS95]. multivariant [YJ92]. Multivariate [RCVJ99, HSM95]. muon [GJ96, OPT99, RM90]. muon-catalyzed [RM90]. muonic [NP96]. Mutation [RdMP99, MBZM99a].

N [Art95a, Gie90, Hoc92, DCR99, Tau92, BAAD^{+}97, Gra98, HP97a]. N-body [BAAD^{+}97]. Nagumo [IDW99]. NaI [NCC^{+}96]. nanoelectronics [NVK94]. Nanosecond [BME96, DBME96]. narrowband [KSA^{+}91]. Nath [CZ93]. Natural [Hor90, NN93, Tau92]. Navier [HS99, OGG91, QV95, WZA^{+}98]. Near [MR99a, FR96, Gri97, IMG99, KS92a, MC99, Pur99, Rie99, YB99]. Near-neighbor [MR99a]. need [BJL^{+}96]. needs [Bow99]. negative [CMC^{+}99]. neighbor [CD90, CD91, CD92, MR99a, MT92b]. neighbour [Ryc90a, Ryc90b]. neighbourhood [MN92]. neighbours [AM90]. neoclassical [Pai93]. NERO [TGG97]. nested [YK95, ZY97]. net [CO92, KK96a, Odo96, SAY93]. nets [SA95]. network [BGK^{+}94, BCD^{+}93, BR97b, BP90b, DL90, Frü93, GH91c, HTZ^{+}98, LLF97, PRL94, Pod99, SA91, TMG99, Ves98, dSTC99, dSB99]. Networking [GN98]. networks [BB93a, BR95, BR97a, Ber96, Bis91, BS99b, Den99, DD94, FNFBdlN99, GC95, Hum90a, Hum90c, Hum90b, JK99, Kin99, LVC^{+}94, LPPR91, LPR92, SAG91b, TVH^{+}99, dMP99, GJ98, Deb92]. Neumann [BMOF92, BK91]. Neural [Den99, DD94, Odo96, SCS^{+}96, BGK^{+}94, BB93a, BR97a, BCD^{+}93, Ber96, BR97b, CO92, Deb92, DL90, Frü93, GC95, GH91c, Hum90a, Hum90c, Hum90b, Kin99, LLF97, LVC^{+}94, LPR92, PRL94, Pod99, SAG91b, SA91, SAY93, SA95, TMG99, TVH^{+}99, dSTC99, dMP99, GJ98, Bis91, Cyb90]. neural-net [SAY93]. neural-nets [SA95]. neurons [TVH^{+}99]. neutral [ABB^{+}96, FT99a, FWH95]. neutral-to-ionic [FT99a]. neutrino [KCK99]. neutrinos [RI97]. neutron [Cop91, GJ96, Glü97, GRS90, KKKS95, PP99, WBB99, dBPB^{+}93]. neutron-proton [dBPB^{+}93]. neutrons [MFG^{+}98]. Newton [AL93, BSSH99, KM95]. Newtonian [CMM92, AP97, MY92, MMCH93]. Next [CGL^{+}98, AM90, KM91, RGLV^{+}99, Sig97]. next-to-leading [KM91, Sig97]. NFTC [PL99]. Ni [PI95]. nine [PS98a]. nine-point [PS98a]. NIRVANA [Zie98]. nitrogen [FKFE99]. NLLjet [KM91]. NLO [Pöt99]. NLOM [KKH^{+}92]. nMOLDYN [KKKS95]. NMR [TvG91]. no [BBS92, DMvG99]. no-exchange [BBS92]. noble [LC93a]. node [BHL92, HTZ^{+}98]. Noise [CGdlR99, BRWN96, CNPT91, Frü95, Frü97, FS99a, GV99, RKV91, SA99a, vVF96]. noise-removal [CNPT91]. noisy [Fog99, GKP93a, GKP93b, ZKHG99]. NoLLS [LH97]. nome [Lee92]. Non [CP95, FB99, MBC^{+}99b, AAD92a, BK91, BvG97, CMM92, CZ92, EOR94, Frü97, GST99a, HKG90, HF91, IDW99, Lam96a, Lam96b, LL96, MY92, Mic94, PD96, PB96, QAd^{+}95, Rod92, ŠČ91, SGF99a, SMH98, SGF99b, VB91, WP98, dMCT99]. non- [SMH98]. non-abrupt [Lam96a, Lam96b]. non-adiabatic [BvG97]. non-dedicated [QAd^{+}95]. non-dissipative [BK91]. non-equilibrium [IDW99, VB91]. non-extensivity [dMCT99]. non-Gaussian [Frü97]. non-ideal [CZ92, LL96]. non-integrable [SGF99a]. non-isothermal [ŠČ91]. Non-isotropic [CP95]. non-linear [PD96, Rod92]. non-magnetic [PB96]. non-Newtonian [CMM92, MY92]. non-orthogonal [HF91, SGF99b, WP98]. non-periodic [Mic94]. non-potential [GST99a]. non-relativistic [HF91]. Non-sequential [FB99]. non-spherical [EOR94]. non-trivial [AAD92a]. non-uniform [HKG90]. Nonadiabatic [WRF91, ZN93]. Nonclassical [MRC98]. nonconcentric [NVC96]. nonconservative [Cor99]. nonequilibrium [SC99a, TD99]. nonholonomic [Kut99]. noniterative [ST92]. nonlinear [AH91, Bes98, FS99a, Her91, HGCM98, Ito94, IDVV97, IVDV97, JMSW99, KKMS99, LH97, Mal97, MRC98, MMTK94, Nun90, PD90, PAC^{+}98, PL99, Ren92, Riz99, Rob98b, Swa91, TG96, TGG97, USW99, Wee93, WZKH98b]. nonlinear-condensation [JMSW99]. nonlocal [HB90, Izv99, KKH^{+}92, Pap92a, Pap92b, USW99]. nonnegative [SG98b]. nonorthogonal [Zat96]. Nonperturbative [CJ91]. nonpotential [GST99b]. Nonreactive [Lem96]. nonrectangular [GJ99a]. nonrelativistic [Sin94]. nonseparable [RU92]. nonsinglet [KKK95]. nonsymmorphic [DPD96c]. nonuniform [SV94]. NOON [SGF99b]. Nordsieck [GM92, Str91b]. normal [BGT95, GW98, HP97b, Kal93, PCC^{+}95, RU92, Ren92, TCE99b, YK90]. normal-mode [TCE99b]. normalization [BCM^{+}95, SS93a]. Note [TFY95, Ano94-32, Ano96-31, Art91, Du90, FJ91, Jam96c, LC90b, NM97]. Notes [Art96, Jon90, MT92a, Whi93]. notice [Ano90q, Ano91w, Ano91x, Ano91y, Ano92q, Ano92r, Ano92s, Ano92t, Ano93n, Ano94y]. Novel [Mic91, Szo99, FMSdC99, KK96b, OM93, POA92, RVCF99]. novo [Sik99]. NP [STI99]. NP-completeness [STI99]. NPTOOLS [CC98]. NT [AO98, FP98]. Nuclear [Sun97, AA98b, BMMMMR99, BZO92, BDF^{+}97, BGG91, BVR95, Ber99, Bon93, CUVS95, DS90, DPP^{+}99, Glü97, GW94, Hni94, HS93b, Joó97, LR92b, LC93b, PVR97, Ran93, TG99, US91]. nucleation [IST99, LS93]. nuclei [BM92b, DG93, GRRG95, GLF97, Mal90, Pi92, Pur90, RGL97, Sat97b, SH97, VST94]. nucleic [LZS95, LHSD95]. nucleon [Di 95, GGMS95, IER97, LHL98, PB93, UTI94]. nucleon-nucleon [GGMS95, UTI94]. nucleus [Cap93, NC96, PVRR97]. nuclide [Car96]. nuclides [Gal99]. null [ZS91]. null-field [ZS91]. number [Art91, CBTL97, DH97a, DH97b, GW95, GP99, Gut95, Ham93a, Ham93b, Ham94, Ham95a, Ham95b, Hen94, Hen95, HDF97, Jam90, Jam94, Jam96b, LB95, Lüs94, Mak93, MNZ90, Shc99, VKSAN95, VRS^{+}95, VRS^{+}99, WW90]. numbers [BS91c, BSW97, Fer99, IWWY92, JHK97, MM91, MTM92, TBCH90, TC93]. numeric [Fit90]. Numerical [BBB99, BM99c, CMSK99, CS99b, DEAM91, DdMdO^{+}99, Dub91a, Fij99b, Fij99c, FMH93, FMIM93, GDRR94, Has99, HAR91, HKM98b, HKM98a, HE91, KI99, KCK99, Kot96, Les99, Lin90, MR96, MK96, MS91b, NSE95, Ome98, PBDZ94, PGK90a, PdA99, RHFR95, RVB99, ŠČ91, Sec99, SP93c, SSK97, SSF92, SPT98, Sta96, Sug91, Tan90, TKS91, TS92, US91, WJ91, ZN93, dQRdS99, AG99, AB99, Bal95, BCSS99, BP93, BHP97, Cal96, CRS90, CBD^{+}95, CGP94, Fij99a, Fij00, FJP90, FJ95b, Fri97, Gib92, HD92, HS95, IBR98, ITMI95, Iqb95b, IDVV97, KMz91, KH97, Kle92, KV99, KKK95, KLS96, KNN91, Kul97, KL92, Lee90a, Li95, Lin91, Lio97, LC93a, LP99, Luc94b, Mac92b, MY92, Mer90, MB91, MN97b, Nak91, NK93, NM97, NL91, NN93, Nun90, OY99, POA92]. numerical [PJSFM92, PJSF93, PBN93, PSV99, RW97, RLWW97, Rei90, RW95, RSW96, SM91a, SFVW95, Sch91a, STJ90, SSHT91, SS93b, Sim92, Sim93, Sim98, Sim99, SAZ90, Str93a, TC93, UOMO94, UKK98, VGK90, Var93, VRS^{+}95, VRS^{+}99, Hop97, Hop97, Jon90]. numerically [CW92, DDK99]. Numerische [Weg91]. Numerov [JD91a, JF92]. NUNUGPV [MNP96].

O [HP97a]. ob [Cra98]. Object [GF98, KDH^{+}98, BJL^{+}96, CSC^{+}97, CI98, Deg99, HAC^{+}99, Hol96, MC97, RFH^{+}95, TEJJ93, VLG95]. Object-Oriented [KDH^{+}98, GF98, BJL^{+}96, CSC^{+}97, Deg99, HAC^{+}99, Hol96, RFH^{+}95, VLG95]. objects [BL99b, Cab90, MS91b, XYIT97, ZS91]. oblate [GS98a]. oblique [YB92]. Obrechkoff [Sta96]. observables [BA94, MPN^{+}93b, MNPP96, MNPP99]. occam [AHZ90]. occupancy [Pae91]. occuring [DM90]. occurring [OPP96, RW97, RLWW97]. Ocean [SW91a, PGJ91]. octagonal [LL99a]. octonionic [DM98]. octree [LS93]. ODBMS [Asa98]. odd [Pur90]. odd-A [Pur90]. ODE [Aro96, CTR98]. ODEs [AR95, CTDdM97, CTDdM98]. ODMG [MM98]. ODMG-compatible [MM98]. OEDIPUS [Sal97]. Off [Gut92, SA99c, ABH^{+}94, BHM99, Kre99, Sto98]. off-diagonal [Kre99]. Off-equilibrium [SA99c]. off-lattice [BHM99]. Off-line [Gut92]. off-shell [ABH^{+}94]. offset [RHGDM99]. older [PRdM99]. oligonucleotide [BSB99]. on-line [DHK^{+}98, KDH^{+}98]. on-shell [FT92]. One [BHS^{+}99, Jaq90, LPC94, RBP^{+}90, YA91, ABF^{+}99, AG91, ACA99, BN92, BBR^{+}90, BA94, BTC99b, BFK95, BFK97, CRS90, DM90, EHK^{+}92, FJKvO95, Fuk94, GS97, HPV99, HD92, HK96b, IY95, JN94, Kal94, Kor99, KK95, KK97, LC93a, MÖ93, NE93, NSE95, Pöt99, Pot98, RLPR91, RLPR94, Sim93, SGT90, TGG97, UKK98, WZKH98a, vO91]. one- [ACA99, HD92, Pöt99]. One-dimensional [BHS^{+}99, LPC94, RBP^{+}90, YA91, AG91, CRS90, JN94, NE93, NSE95, Sim93, UKK98]. one-dimensionally [BN92]. one-electron [BA94, Fuk94, KK95, KK97, Pot98, RLPR91, RLPR94]. one-flavour [ABF^{+}99]. one-loop [BBR^{+}90, BFK95, BFK97, EHK^{+}92, FJKvO95, HPV99, SGT90, vO91]. one-photon [DM90]. one-turn [TGG97]. oneloop [BFK97]. Onium [Sal97]. Online [CB99, Kru98, SW99]. onto [GG96]. OOP [Bla94a]. OOPS [RFH^{+}95]. Opacity [DS96]. ope.math [Fuj94b]. open [BCP91, Kre99, RDB^{+}90]. open-ended [BCP91]. open-shell [Kre99, RDB^{+}90]. Opening [Gen98]. operate [Sci93b]. Operations [Ixa97]. operator [Aok97, Bes98, DBME96, Fuj94b, HL92, JH91a, KS96, MOC91, RW95, RSW96, BME96]. operators [FGH91, Koo93, LPSM92, Pul97, RNR^{+}99, Sei94]. optical [BRWN96, BWNR99, CHC91, FMSdC99, IRV95, KKH^{+}92, Kot92, LD91, PHN99, PTMUJdFM99, RMMUJ99]. optics [Aut99, KSZ97]. Optimal [GGJG97, Lem97, SR91, SW99, Sug99, Frü93, Goe93, HS96, OPT99, Rho97, RBG99, SSS91]. optimisation [TBA^{+}98]. Optimization [Krö96, PK98, VFO94, BBMD99, CT97, Hof99b, KSZ97, KTT98, KP94, KWS^{+}96, LMS94, MDSF^{+}99, PDL98b, PDL98a, TH95, Tur99]. optimize [CGM99, KK96b]. optimized [Jaq90, Luc94a, Mon98]. optimizing [For91]. option [Pac91]. options [Len98]. orbital [CGM^{+}96, FF97b, Fuk94, SGF99b, ThLL^{+}93, VS97, WJ94]. orbitals [BSKO96, HF91, MSFR96, Zat96]. orbits [DB93, Far98, LEBB99]. Order [Glü97, IOM95, Joó97, SPA99, AAKP95b, AAKP95a, AAKP98, ABHR99, ADM^{+}92, ABH^{+}94, AB90, BMSW91, BC99a, BM95, CZ99, CTDdM97, CTDdM98, CTR98, DS92, FJ91, HRK99, HMG92, IMG99, IDVV97, JW90, Kan95, KM91, KRVZ98, LMW94, MIO96, MES90b, NE93, PSV99, RdC93, Ryc90a, RJ90, Ryc90b, RJ91a, RJ91b, Sig97, Sim93, Sim97, Sim99, SL97, SGF99b, YB99, dT93]. Order- [Glü97, Joó97, IOM95, SGF99b]. order-disorder [IMG99]. Ordered [AS95, IK99, SL97]. Ordering [BEFJ99, HLPT93]. orders [SG98b, SG99]. ordinary [ABHR99, LK96, Rod92, SP93b]. organic [CKS91]. organization [Ano91-35]. Organizing [Ano99o, LPPR91]. orientation [Rei93]. orientations [Lil93]. Oriented [KDH^{+}98, BJL^{+}96, CSC^{+}97, CI98, Deg99, GF98, HAC^{+}99, Hol96, Joó97, KKKS95, MPS99, RFH^{+}95, TEJJ93, VLG95]. Ornstein [HRK95]. orthogonal [Eas91a, HF91, Koo93, MD96, SGF99b, WP98]. Orthogonalising [MR99b]. orthogonality [KTM96]. orthonormal [EHK^{+}92]. ORTHOVEC [Eas91a]. oscillating [Sim98, Sim99]. oscillation [XYNI96, XYI96, XYE99]. oscillator [DD97a, DD97b, FF97b, LPC94, LM91b, MR96, TS92]. oscillators [Kal94, LPSM92, RBB98]. oscillatory [Alb99, Bra91, Ixa97].

OSIPE [CJP^{+}94]. Osservatorio [MRPC94]. other [AFP99, MSP^{+}98, PT92, VVIBA99]. output [Jan99, OA95, OA96, RASS99]. ovals [GCP95]. overdamped [GV99]. overhunting [MBZM99a]. overlap [AH92, HWS92, JBB99, PC94a]. overrelaxation [Kob94]. overstable [HKG90]. overview [Mar96]. Oxford [Art93, Eas93, Tru93a]. oxides [MOPR99, MSTH92, Mic94]. oxygen [HHV99, PvBMN99, PRRL99, VS95]. OZ [CP95].

P [Art96, Cra98, Fin90, Fin91, Gie90, Par99, Rai91, WW93, Whi93, PG96]. P3M [The94]. pA [ST95]. package [AI90a, BD94, BNSW92, BFK97, BN95, Can94a, Can94b, CBW97, Cha92, CY99, CC98, FFB99, FT92, Flo97, Fro91c, Fro91d, Gro91, Har97, HS93a, Hen94, Hol96, HKSV90, Ito94, JL93, Kad96, KT91, Kaw95, KKKS95, KRVZ98, KLS91, LP93, Luč95, MMTK94, MMSM93, MP93, PFG96, PCC^{+}95, PRL94, Por90, PM91, Ren92, Sci93a, Sci93b, Sei91, Sem98, Var97, VVA^{+}94, VRS^{+}95, VRS^{+}99, Wyb94, vO91]. packages [BBP^{+}90, CP99, DLS90, Rei90]. packet [BME96, DBME96, Lem96, MK91b, Zha91]. Packing [LC96b, BM91]. PACS [Uka98]. PAD1 [CS98b]. PAD2 [HUCS98]. PAD3 [CHUS98]. Padé [HZQ93]. pages [Art93, Art95a, Art96, Bis91, Bon93, Bro92, CT96a, Cyb90, Deb92, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Rai91, Reb93, Top96, Tru93a, Tru93b, Weg91, Whi93]. Painlevé [Ren92]. Pair [Str93b, TOSM91, ABH^{+}94, FR96, FJ95b, ISS91, JW90, JWW91, JWW94, LL99c, SJPW96, TRSM98]. Pair-Distribution [Str93b]. pairing [Sun97]. pairwise [CKS91]. PAKPDF [Cha92]. Palermo [MRPC94]. Pao [Tay99]. paper [Met96, Nar92]. paperback [Art93, Art95a, Top96]. Parabolic [SG98b, Tau92, BNS^{+}95]. Paragon [Gut95]. PARALLACS [MPSvG94]. Parallel [AHZ90, BMPS99, BHL92, DBVS99, FVR92, FHSO99, FBH90, JAT97, KB99, MC97, MD97, NVK93, Nak97b, RBS94, SGB^{+}93, SF94a, SF94b, SAJ^{+}97b, The94, AS95, AH99, BJL^{+}96, BMSW91, BBL^{+}99, BADP96, BAAD^{+}97, BTC99a, BvdSvD95, BT92a, BT92b, BCOY94, BCOY95, BMM97, Bru95, CGM^{+}96, CC99, CUVS95, CSP92, CA92, Dec95, DDK99, DG99, Din91, EABH95, ERB90, FFG^{+}96, GG96, GPSM95, GHGB97, HCVM99, IOM95, JTAW97b, Kah91, KKH^{+}98, LMM^{+}97, Lag92, LGZL98, MMGS93, MLM96, MPRS97, Mic94, MTR94, MPSvG94, MTNA98, NKV94, NVK94, Pai93, PS98b, PK98, QAd^{+}95, Rap93, RBS92, RDB^{+}90, SFM^{+}96, SBGM91, Shu94, SWT91, SSH94, Smi91, Smi92b, SG91, SAJ^{+}97a, Str93a, TR93, TGAW96, Uka98, Van96, WLD95, WS95a, WDB^{+}91, WJ94, WJ95, XZ91, ZX91, Hoc92]. parallelisation [MS96, MPB91]. Parallelising [MP90b, Yeu91]. Parallelism [LMS94, Mar97]. Parallelization [CDM99, FMH97, Kre97, RP95, SHNS98, WS95b, Bab92, BCOY93, Sek92]. Parallelized [PT95, Mak93]. paramagnetic [PB96]. parameter [AO97, CZ99, MB91, Sal91, vdB93]. Parameterised [MCMD91]. parameterization [ST98]. parameterizing [NCC^{+}96]. parameters [HP97b, KK95, LH97, LHSD95, Tur99, Wes93, Wie96, KK97]. parametric [DPP^{+}99, SWG99]. parametrization [Dik97, GCP95, vMC91]. parametrizations [Cha92]. parentage [Kag92, xLzHjW95]. Paris [Fly90]. Parisi [KKK95, KL92]. Parker [Rai91]. Parrinello [CC99, WJ94, WJ95]. parsimony [VPE94]. Part [Lam96a, Lam96b, MN97b, FPV94, FPV95]. Partial [AA98b, Aok97, BKDMM91, CT94a, CTvB95, Deg99, DHL93, MDSF^{+}99, RW95, RSW96, SM93, SP93c, Sty90]. partially [GF96, MP90a, SBK98]. Particle [AS95, Art94, FBdR^{+}98, HL92, LB91, Pur90, SZ99, ADK^{+}99, ADS92, Bas99a, BA95, BV99, BBG94, Bla94b, BT92b, Bra91, DL90, DB93, DBVS99, Dub91b, DPRS90, Eas91b, FWH95, FBVdlN99, GCP95, Gei97, GW94, HS99, JN94, KZS95, KL93, Lan92, LB95, xLzSlLzH94, Man91, Mar97, Meg94, MTH^{+}99, Mor92, NE93, NC96, Nun90, PAC^{+}98, PFBdlN99, Ran90, Rap91a, Rap91b, Rap93, RZE99, RFH^{+}95, Rho92, RCL98, RHFR95, RVB99, SLGB96, SEZ99b, SEZ99a, Ste97, SZS95, Sul96, Tan95, UOMO94, VS95, VP96, WLD95, WS95a]. Particle-fluid [LB91]. Particle-In-Cell [FBdR^{+}98, Lan92, MTH^{+}99, SZS95, Sul96, VS95]. Particle-mesh [Art94, Eas91b]. particle-tracking [ADS92]. Particle-vibration [Pur90]. Particles [Ris96, AIS^{+}97, BDKM91, BKS99, BG94, EOR94, GLF97, HTI99, HN99, Kag92, LB95, Lia95, LS93, MMH99, MEA^{+}90, PSR99, VB96, VVGI99, WW90, WCL90]. particular [MN97a]. parton [ABE^{+}94, Cha92, Gei97, GW94, MKPGS97, PB93, Sey95, Zha98a, Zha98b]. parton-cascades [Gei97]. parton-hadron [Gei97]. Pascal [HHK92, Kae95]. passage [SP93b]. passing [BvdSvD95, CA92, MAB^{+}94]. patch [Bha91, CS99a]. Pater [Tru93a]. path [Mak91, RDF^{+}97, SM91b, VVGI99, dCRSF99]. path-integration [SM91b]. paths [GPM95]. pathways [Gal99]. Pattern [CGRV99, LPR92, BE90, CNPT91, CO92, DL90, GH91c, MBMM99, Odo96, RCVJ99, RNR^{+}99, dMP99]. patterns [CTR98, GP95]. Pauli [HGF91]. PAW [Zan90]. PC [GRS90, Hea93, Hea96, MNZ90, RM90]. PC-Fortran [RM90]. PCs [AO98, Ham97]. PDE [BB92, ST92]. PDETWO [CZ92]. PDFLIB [PB93]. peak [BBR^{+}90, FR96, MPN^{+}93b, MNPP96]. pedagogical [Don99]. Penna [MBZM99a, MBZM99b]. PEPSI [MSV92]. Peptide [DMvG99]. percolation [BP90a, Bab92, HÓ99, SA99c, Sto98]. perfect [BEF^{+}99, DL98, Lia95]. Performance [DCR99, SSH94, BCD^{+}93, Bow99, CD91, CDD^{+}96, DG99, Hel99, Mar96, MLM96, Sek92, CLM98]. performing [FGH91]. periodic [BT91, BC91b, CJ91, Eva90, Far98, GJ99a, LEBB99, MGMR97, Mic94, PD90, PG96, RF96, SFFDR94, Sim98, Sim99, TZC96]. periodically [PGK90a]. periodicities [Zlo90]. Permeability [SLB99]. permutation [Ber95]. personal [BNSW92, CT96b, Fly90, Fly93, Ham93a, Ham93b, Ham95a, Ham95b, LC96a]. perspective [Den99, TB96]. perspectives [dlSG99]. perturbation [BMSW91, BDF^{+}97, FGR98, IRV95, LMW94, MSW92b, MSW92a].

Perturbative [Sal97, Gra98, MBC^{+}99b]. perturbed [BDF^{+}97, FGM98, LPR^{+}93, PCNO94, PCC94, PC94b, PC94c, PC94a, BPL97]. perturbed-cluster [PCNO94, PCC94, PC94b, PC94c, PC94a]. PEST [PMJ93]. Phase [ABF^{+}99, BSKO96, Bas99a, BL99b, BAS99b, Ten97, Alb99, BL99a, BC99a, Blo92, BF90b, ÇYG99, CS98a, FT99a, FNFBdlN99, HP93, HLMM99, HKLM99, IMG99, MCO99, MK95, NY99, Neu99, Pur99, RP91, Sim93, SS91b, SVV99, Szm95, Tox99, UTI94, UKK98, vVF96]. Phase-amplitude [BSKO96]. phase-lag [Sim93]. phase-shifts [UTI94]. phase-space [HP93, SS91b]. phases [FKFE99, HMR99]. PHENIX [DHK^{+}98, KDH^{+}98]. phenomena [Nak97a, SDB96, Zag99]. phenomenological [VPE94]. phenomenology [VP96]. phenomenon [OGG91]. Philip [Ano94-33]. PHIPHI [MPN93a]. Phonon [GP95]. phosphate [BSB99]. phosphate/phosphonate [BSB99]. phosphoglycerate [GPM95]. phosphonate [BSB99]. photo [HUCS98, Zei94]. photo-electron [HUCS98]. photo-induced [Zei94]. photodissociation [Ale93, BKDMM91, BKMM93, Hea91]. photoelectron [CS98b, CHUS98]. photoemission [CT96b]. photoionization [Bar98]. photomultipliers [OA96]. Photon [BM92b, Bra98, BLLM94, DM90, GJ97, Kas91a, MH96, MKPGS97, PB93, Sch98, SRR95, TP99]. photon-impact [Bra98]. photonic [BPMW95, FJKvO95, HHH99, SYM98, WP98]. photons [JW90, Pöt99]. photorefractive [OBRR92]. PHOTOS [BW94, BvEW91]. Phys [BCOY95, DBME96, Ham95a, Jam96b, J̈K97a, KK97, RLPR94, Stu95b, Sul96, VRS^{+}99, Zha98a, Nar92]. Physical [CGP94, DL98, Oln98, Ano96k, BBP^{+}90, Cab90, DS96, DH97a, Gie90, IVDV97, LH97, MB91]. physically [MDSF^{+}99]. physicists [Art95a]. Physics [Ano96k, Art95a, Art96, Don99, Fij00, Fle92, Gie90, JY97a, Jon90, Par99, Tru93b, Whi93, AB97, AFP99, All93, ACG^{+}99, Ano90g, Ano91h, Ano94j, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano97j, Ano97k, AH99, Art93, AO98, BH92, BH99, BK92, Bon93, CS91, Den99, DB93, DLS90, DD94, Duc96, Fin90, Fin91, FT99b, FP98, Hum90b, Kin99, Kli98, KH95, Lan99a, Lan99b, Lob99, LPR92, Lön94, MM98, MNS99, Ohl92, Ohl97, Ott90, PMUV90, Pas96, QTR99, Sch98, Sjö94, Top96, UTI94, VP96, Oln98, Tru93b, Art96]. physics-based [DB93]. Physlets [Chr99]. PIC [AG91, ADS92, Dec95, DBVS99, EABH95, GLSW95, ITW97, JTAW97b, LHL95, MV97, MAH98, VLG95]. PIC-circuit [AG91]. picture [Zha91]. Piecewise [IRV95, BNS^{+}95]. PIES [JMR^{+}93]. Pilot [Hel99]. PiN [LHL98]. pion [Di 95, Glü97, LHL98, PB93]. pion-nucleon [LHL98]. pitching [CC91a]. place [PJ98]. Planar [CL96, LLM91, PD90, TD99, WA97]. Planar-basis [CL96]. Planck [KH97, MGMR97, MOC91, MMTK94, PS98a, SSF92, SS93b, SS94a, WP96]. plane [BSST90, DPD96a, DPD96b, DPD96c, DPD96d, JKSC99, KWS^{+}96, LVL90, SJ91, YK90, YB92]. plane-stratified [LVL90]. plane-wave [JKSC99, KWS^{+}96]. plant [CPM99, Hel99]. plasma [AS95, BB99, CTE95, DBVS99, ECT93, FG95b, Gib92, Hew94, HE99, IKTY99, KB98, KI99, KRT95, KR95b, LB91, MD96, MC97, MTH^{+}99, MAH98, PG94, PdA99, QTR99, Rea95, RCL98, SV94, SAP^{+}95, SS93b, Sto99, VB92b, WLD95]. plasma-based [MTH^{+}99]. plasma-surface [HE99]. plasmas [DMM^{+}97, DB93, DJUM99, HL92, HKG90, JAVV95, LLW98, MSD93, MRPC94, MKWH91, MOC91, MP90a, Nun90, Pai93, PR93, PL99, RBP^{+}90, VGK90, WBB99]. plastic [KF99, RP91]. platform [CUVS95, Rob98b]. PLEC [For91]. plots [SW91b]. PMHT [SZ99]. PMTs [OA95]. Poincaré [CTdO96]. point [CHW91, CP99, CK91b, Cle93, Dik97, Fra92, FT95, FBdC98, HS99, MSFR96, MJ90, PS98a, RB99]. point-force [HS99]. pointers [MKS^{+}96]. points [Luc94a, MG93, Rie99, YJ92]. pointwise [HLT93, HT93, THF95]. pointwise/basis [HT93]. Poisson [HCO99, Hol96, HKC98, IBR98, NY99, TDV96, UKK98]. Poissonian [JV99]. polar [Ome97]. polarised [GGMS95]. polarizable [KN95a]. polarization [Hni94, VW95]. polarized [AIS^{+}97, BHY92, HKM98b, JKW91, MSV92, dBPB^{+}93, BKW97]. Polarons [Izv99]. POLDIS [BKW97]. pollution [LABV90]. poloidal [AHH94, LB99]. POLRAD [AIS^{+}97]. poly [BKNS97, FS99b]. poly-atomic [BKNS97, FS99b]. polyampholytes [SUYH99]. polyatomic [CP91, JWM^{+}96, KC98, SG98a, dlVM98]. polyatomics [hLTM^{+}92, Ome98, SHL^{+}95]. polycarbonates [HTK99]. polycrystalline [Chi90, Wie95]. polycrystals [Zlo95]. polyethylene [MPB96]. polyhedra [SAZ90]. polyhedral [RJS99]. polymer [AS99, BC91a, FNFBdlN99, FBH90, MB94, WS95b]. polymer-polymer [AS99]. polymers [BVVJA99, BCOY94, FR90, JAT97, Rai91, SFFDR94, VVIBA99, WK90, BCOY95]. polymorphism [DNS98]. polynomial [AL93, BCM^{+}95, FFB99, GFAALS98, PZH^{+}96, YA91, YIW^{+}91, ZL99]. polynomials [Bil92, BEFJ99, Mon98, POA92]. polynoms [BTC99b]. polypeptide [Der99, For91]. POLYRATE [hLTM^{+}92, SHL^{+}95]. POMULT [Far98]. POP [Mor92]. population [BMeS99, MBZM99a]. porous [ADK^{+}99]. portable [CDD^{+}96, FG95b, Lüs94, VRS^{+}95, VRS^{+}99, WS95a]. POS [RGL^{+}98]. POS-SPRITE [RGL^{+}98]. posed [Abd90, Wee93]. position [Goe93]. positions [Kar91, Wie95]. positive [Tau92]. positron [AA99, Bar93, RGL^{+}98, SG98a]. positrons [HM92, SM93]. possible [Eva90, RGLV^{+}99]. post [PMJ93]. post-processor [PMJ93]. Potential [LHSD95, AAS95, ATP98, Ale93, BSST90, Boy95, BVVJA99, Din91, FMIM93, FJ95b, GST99a, GCP95, IRV95, IDW99, Izv99, KWS^{+}96, OE99, PC94c, Rou95, SFFDR94, SvNF92, TOSM91, TRSM98]. Potentials [MR99b, BMOF94, CC97b, DDDL98, Hni94, HB90, ISS91, KR91, LPC94, Lio97, LD91, LL99c, MCMD91, Mon91, Pac91, Pap92a, Pap92b, Rit92, WR99, YKO^{+}94]. Potts [BJK92, ÇYG99, CBB99, CHM91, CK99, DM91, LL99a, PI99]. powder [BLGT94]. POWDERSPEC [BLGT92, BLGT94]. Power [Tar99, Jak98, KMT^{+}99, Lee92, Lio97, MP95, SL97]. power-law [MP95]. power-series [Lee92]. powers [Str91b]. POWEV [Sci93a]. pp [TP99]. practical [DLS90, Kar91]. prandtl [TBCH90]. Precision [Swa91, PL90, Por90]. Preconditioned [AR95, SG91, Nak97b]. preconditioner [FFG^{+}96]. Preconditioning [BEF^{+}99]. predicting [CTJ^{+}95, CDR91]. prediction [Der99, Lin90, MMU99]. predictions [BBL^{+}97]. predictor [AH91, Kro92, YY99]. predictor-corrector [AH91, Kro92, YY99]. Preface [Ano91-28, Ano94z, Ano99p, BMP91, BE95, CTM98, FGH98, GK90, GW91, Nad90b, Nad92, Nad96, PS95, SCS99, Sha91]. Preparation [CS98b]. PRESCOLD [GDGR97]. presheath [MC97]. Press [Art93, Cra98, Eas93, Tay99]. pressure [BC91a, CHMG92, FF97a, KSSB99, LC93a, RARF96, SZG99, Tar99, WT99]. pressures [GBWES97]. preventing [XYNI96, XYI96]. Price [Art95a, Art96, Bis91, Bon93, Bro92, CT96a, Deb92, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Rai91, Reb93, Tru93b, Whi93]. pricing [Deu99]. primitive [BTC99b, OY99]. principal [CW92]. principle [GKP93a, GKP93b, LK96]. principles [CY93, FT99a]. printed [DP91, Nak91]. priority [MC95]. probabilistic [CvN99, Art96]. probabilities [BS91a, KL93, MF94, VST94]. probability [GJ98, POA92]. PROBFIS [VST94]. problem [AP98, AHH94, BS95, CMM92, FS91, xLzSlLzH94, MMCH93, Moo93, Ple96, Pul97, Ryb98, ŠČ91, SFM^{+}96, SMB^{+}93, SJB^{+}94, Sim93, XZ91, ZX91]. Problems [Oln98, Par99, AB99, AK91, CTS99, Chr99, Du93a, DD94, FGM98, HHR98, Hum90c, IDVV97, IVDV97, IDV99, Jaq90, KNN91, KH95, Lin91, Lob99, Moo94, PdA99, RW97, RLWW97, Rho97, SS90a, SS93b, SSBA91, Sim98, Sim99, TH95, VS97, Wee93, WZKH98b, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90, dCRSF99]. procedure [Hea91, MSH91, PS98a, Scr95]. Procedures [HD92, MHN99, CT94a, Fri97, FVGF98]. Proceedings [Cyb90, Fin90, Fin91, Gie90, Reb93, Whi93]. process [BBL^{+}97, FBGV99, MH96]. processes [BPK95, CGHB98, DM90, FHI^{+}98, GW95, GDGR97, KR91, KSM92, MFG^{+}98, Pot98, Tan90, TMG99, WRF91, Zei94]. Processing [Ano94a, Ano94b, Ano94c, BMSW91, BCOY94, BCOY95, BMM97, Bro94, Coo99, ERB90, Goe93, Hor90, LC96b, Rap91a, Rap91b, Rap93, RTLM92, RDB^{+}90, WDB^{+}91, Won90, Bro92]. processor [CD91, Fan90, FT95, MMS93, MSW92a, MP90b, PMJ93, RBS92, SGB^{+}93]. processors [Kah91, LG93b, MM91]. produced [ML92]. product [AC92, BKDMM91, BEFJ99, Fuj94b, HK94, SPP^{+}98]. production [ABH^{+}94, ALR97, Bra90, CMNP97, Chi92, FR96, GW94, HKL93, JW90, JWW91, JWW94, KM98, MNP95, Pap97, Pap99, SJPW96, TP99]. products [Can94a, Can94b, HMG92, Wyb94]. profile [AI90b, Lam96a, SRB94]. profiles [HMH99, KKM^{+}99]. Program [Ano90s, Ano90t, Ano90u, Ano90v, Ano90w, Ano91-29, Ano91-30, Ano91-31, Ano91-32, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano93o, Ano93p, Ano93q, Ano93r, Ano94-27, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano95z, Ano95y, Ano95-27, Ano95-28, Ano95-29, Ano95-30, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano96-29, Ano96-30, Ano97-30, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, CP99, GKP93b, KW96, Kal93, Mar91, YN91, Abd90, AA94a, AAKP95a, AAS95, AA98a, AAKP98, AB97, AI90a, AI90b, ABH99, APN^{+}91, Ano94m, Ano96r, Ano96k, ABB^{+}96, AH97, BKMM93, BBL^{+}97, BDKM91, BM91, Bar93, BCM^{+}95, BGT95, BPMW95, BLGT92, BPK95]. program [BVR95, BS90, Ber90, Ber99, Bil92, BvG97, BPL97, Boc92, BFK97, BBS92, CRS90, CDG^{+}90, CY99, CO92, Chi90, CR98, CS91, CDR91, CT96b, CGR93, CM92b, Deg99, DS90, DKS98, DD97b, Eas91a, ERS^{+}95, ERB90, FS92, Far98, Flo97, For91, FG95a, FF97b, Fro91a, Fro91b, FL91, FGH91, FB93, FSMVM93, GF96, GLAT98, Gei97, GW95, GW98, GGMS95, GW94, HRK99, HK97a, HHCS95, Hea93, Hea96, HFD90, HN97a, Heu90, HF91, HGF91, HHK92, HLL^{+}95, IK96, IDVV97, IDV99, JW90, JWW91, JRWWW92, JP92, JWW94, JPRW^{+}97, JPS^{+}99, Jan90, JL97, JN94, JWF93, JPF96, JF97, Kae95, Kag92, KW90, KKH^{+}92, KR95a, KKKS95, KS92a, KKK95, KLS96, KA95, KK95, KK97, KL92, LH97, LDOO95, Lim91a, Lim91b, LR92a, LSZ92, LC96a, LH92]. program [Lön92, LMS^{+}93, LP90, hLTM^{+}92, LPR^{+}93, LC93b, Mal90, Mal97, MF95b, ML92, MS98, MSFR96, MS95a, MH96, MOSM93b, MOSM93a, MOSM96, MPN93a, MPN^{+}93b, MNPP96, MNPP99, MSTK93, MÖ93, MNS99, NE93, Nor94, NN93, Odo90, Odo92, Odo96, OMC93, Pac91, PWG93, PP94, PCNO94, Pop96, Pot98, QBP98, RT98, RGL97, Rob98b, Rod92, RDB^{+}90, RSW92, RJ90, Ryc90b, RJ91a, RJ91b, Sat97a, SMV90, STR91a, STR92, Sci91, SvNF92, Sig97, SSH94, Ská90, SJPW96, Sor98, Sta99, SF94c, SHL^{+}95, Ste97, ST98, Str91b, SF93, TMR94, TRSM98, ThLL^{+}93, VB96, Var97, VVA^{+}94, VST94, Wag92, WENG95, jWzHxL95, WP98, WS95a, Wie95, Wie96, WDB^{+}91, Yeu91, Zat96, Zlo95, DMLM91, MBW^{+}95]. program-package [AI90a]. Programming [Oln98, CSC^{+}97, CJP^{+}94, Ein90, LMS94, MMSM93, Pet92, Ryb98]. Programs [FG91, GKP93a, GKP93b, HWS92, HS93b, NCC^{+}96, NRG90, RG92, Tak92, BKWM91, BLGT94, BPP98, BSST90, BN94, FPV94, FPV95, GS98b, HLT93, HT93, Jac99, JKW91, PCC^{+}95, RM90, SAP^{+}95, SMH98, Tay99, TSJ91, TML93, Wes93]. project [AFP99, Ano98a, Ano98-27, Bos98, CBD^{+}95, FT99b, Lön94, Men99]. Projected [SH97]. projection [GGJG97, GW98]. projective [Dik94, NKR99]. PROJET [Gra95]. prolate [GS98a]. Prolog [Cab90]. prompt [VST94]. propagating [Hea91]. Propagation [IY95, TCE99b, BME96, CK91a, DBME96, FMSdC99, Goz91, HSSB98, HKA^{+}94, KMT^{+}99, NY99, SV94, SZG99, SHNS98, TYJ92, TMS^{+}99, Zha91, UKA97]. propagator [CG99, JGP93, MS98, SM97]. propagators [PT95]. Properties [DBGY99, Sik99, BKNS97, CDUT99, CBG99, CTJ^{+}95, Flo97, FB93, FJ94, GSS97, Gra94, HHH99, Lag99, MT91, POA92, PHN99, PCC^{+}95, RGL97, SGB99, SP95a, SH91, VBS96, VVIBA99, WT99]. property [SUYH99]. proportions [GGJG97]. prospective [HN97a]. protein [BJL^{+}96, GPSW95, Han99, LP90, RS99, vGHM^{+}95]. protein-folding [Han99]. protein-like [RS99]. protein-solvent [GPSW95]. proteins [LHSD95, Sik99, SP95a]. proton [ALR97, ABE^{+}94, BvG97, Cha92, CSS94a, Gra95, IK99, dBPB^{+}93]. proton-ordered [IK99]. proton-proton [ABE^{+}94]. protostellar [BM95]. prototype [Ano98a, BJM^{+}98]. prototypes [Avd96]. Pseudo [MR99b, Sei94, Bar98, CKS91, HK94, PZH^{+}96, dFZLS95]. pseudo-pairwise [CKS91]. Pseudo-Potentials [MR99b]. pseudo-random [HK94]. pseudo-spectral [dFZLS95]. pseudo-spectroscopy [PZH^{+}96]. pseudo-states [Bar98]. pseudopotential [CL96, GHGB97, HP97a, PvBMN99]. pseudopotentials [FS99b]. Pseudorandom [Ham93a, Ham93b, Ham94, DH97b, HDF97, Jam90, Jam94, Jam96b, VKSAN95]. pseudoresonant [Sch93]. Publ [Par99]. Publisher [Ano94-32, Ano96-31]. Publishing [Art95a, Tru93a]. pulsating [DF95]. pulses [Eng91, GDRR94]. Punctuated [BS99b]. pure [Gal99, HD96, ML92]. purpose [Hal90, IMES90, IYOY94, IEY^{+}96, MP95, PM91, Sta99]. PVM [FHSO99, JAT97]. pyramid [RJ90, Ryc90b, RJ91a]. PYTHIA [Mre97, Sjö94, Lön99].

QB [QBP98]. QCD [ABF^{+}99, ABB^{+}96, Aya92, BMT94, Bow99, BBJ91, CSS94a, CS99b, DLLR96, DL91, FFG^{+}96, Gei97, HK97a, HvSA98a, JL97, KS96, KM91, Kni90, KS92a, Luo96, Pöt99]. QCD-SVZ [Aya92]. QCDF90 [DLLR96]. QDC [Lön92]. QED [ABB^{+}96, BvEW91, BW94, CSS94a, JP92, JWJK92, PRH90, Wag92]. QPSI [FFB99]. Quadratically [Mon98]. quadrature [HS96, Sag91a, SSS91]. quadrupole [JWF93]. quality [DATL99, Jam94, Jam96b, JTAW97a, Lüs94, Shc99, Sor98, Sor98]. quanta [GLF97]. quantitatively [CTS99]. Quantum [BPL97, GSS97, Jac94, LPR^{+}93, MTM99, PLC91, QAd^{+}95, RLP^{+}99, WM99, WCK99, Zha94, BKDMM91, BKMM93, BM99b, BGG91, BBCM93, BBJ95, BvG97, BC99b, CG99, CPCF99, Coh92, Eng91, Gil99, GG96, GJ99b, HE94, JGP93, Kal94, KML98, KNN91, KSG91, LLF97, Lam96a, Lam96b, LM95, Lob99, LD91, Mak91, Mit96, NVK94, NL91, NND98, Pac91, RDF^{+}97, Ric91, Rie99, RH99, RBB98, SB97, SGF99a, Scr95, SM97, SS91b, SF98, Sug91, TB91, Ver99, VVA^{+}94, Wha94, Zha91, ZK99, Fly90, Fly93, Top96]. quantum-mechanical [JGP93]. quark [Din91, HSM95, JWW91, JWW94, KKK91, LHL98, WA94]. Quasi [Aly90, HK96a, HK96b, HK97b, Wan99b, BdT97, BA94, BWNR99, BS91c, BSW97, FFB99, JHK97, KK91, MD96, Shu94]. quasi-axisymmetric [KK91]. quasi-harmonic [BdT97]. Quasi-Monte [HK96a, HK96b, HK97b, Wan99b]. quasi-orthogonal [MD96]. quasi-polynomial [FFB99]. quasi-random [BS91c, BSW97, JHK97, Shu94]. Quasi-static [Aly90, BWNR99]. quasicrystals [TMS^{+}99]. Quasienergy [Pot98]. quasilattices [BP90a]. quasiperiodic [CBG99, LL99a, SGRZ99]. quenching [AO97]. questions [SW98]. queues [MC95]. QuickBasic [Tay99].

R [CT96a, Gie90, Rai91, Top96, LPSM92, Bar93, Bar98, BBS92, BN94, MTG98, QBP98, YM92]. r-dependent [LPSM92]. R-matrix [Bar93, Bar98, BBS92, BN94, MTG98, QBP98]. R-MHD [YM92]. RACHEL [DS96]. RAD [MOSM93b, MOSM93a]. RAD-TRAP [MOSM93b, MOSM93a]. Radial [TON95, AAS95, All93, BSKO96, LHL95, LL99c, SFVW95]. radiation [BCD^{+}93, CDG^{+}90, CM92a, DF95, GM92, JW90, ML92, MSH91, MOSM93b, MOSM93a, MOSM96, MSJ99, Ste97, YM92, YK95]. radiation-magnetohydrodynamics [YM92]. radiation-trapping [MOSM93b]. radiative [AIS^{+}97, ADM^{+}92, ABH^{+}94, BvEW91, BW94, BM95, CC97a, CSS94a, DM90, GJ96, Glü97, Joó97, KB94, KS92a, KSM92, MPN93a, SH91, VSS91, Wag92]. radiative-transition [VSS91]. radii [Iri96]. radioactivity [GDGR97, ML92]. radiofrequency [BDH99]. radionuclides [Ste97]. radius [Str91b]. Raman [CZ93]. random [BJW90, BS91c, BSW97, BTC99a, BJK92, ÇYG99, CBTL97, DH97a, Fer99, FV99, GP99, Goz91, Gut95, Ham95a, Ham95b, Hen94, Hen95, HK94, IJV92, IWWY92, JHK97, JV99, Lüs94, MM91, MTM92, Mak93, MNZ90, PI99, SGRZ99, SA99b, SPA99, SBGM91, Shc99, Shu94, SSW94, Sou99, TC93, TFY95, VS99, YJ92, ZL99]. randomness [BC99a]. RANEXP [Hen94]. range [Ess95, EK94, MN92, RJ90, Ryc90b, RJ91a, RJ91b, SC99b, SAJ^{+}97b, Str93b, TR93, The94, dMCT99]. ranged [WCL90]. rank [Wol98b]. RANLUX [HJ97, Ham97, Jam94, Jam96b]. RAPGAP [Jun95]. rapid [LPSM92]. rapidly [AK99, CC91a]. rate [AGIS99, HLL^{+}95, Jan95]. rates [BM92a, FWH95, GLAT98, hLTM^{+}92, PMM98, SHL^{+}95, ThLL^{+}93]. ratio [BH90, KTM96]. Rational [BRB97, RBB93, RBB96, WC90, XYNI96, XYI96]. ray [BCP91, AI90b, CT97, CT96b, CGR93, KSSB99, MS95a, Wie95]. RBS [HHK92]. RD [SF94b]. RD-SHAKE [SF94b]. reaction [BS91a, CBW97, GLAT98, HLL^{+}95, Jan95, KM95, KA95, Les99, hLTM^{+}92, MEA^{+}90, NK96, RFMM99, SHL^{+}95, ThLL^{+}93]. reaction-diffusion [Les99]. reaction-rate [Jan95]. reactions [Alb99, BZO92, BLLM94, Cap93, Di 95, HRK99, HHV99, Jan95, Lag99, MWA^{+}92, SR91, Tox99]. reactivation [RM90]. reactive [KC98, Lag92, LMS^{+}93, NBJK91, TK91]. reactor [DB93, IMA94]. real [BV99, CT94b, JGP93, Kah91, KRVZ98, SLGB96, SG99, SF94c]. real-time [BV99]. realistic [Ber92b, MNPP96, MNPP99, NKV94, Sol99]. reality [ASJ98, FT99b]. realizations [Fuj94b]. rearrangement [Gal99]. reasoning [Hum90c]. Recognition [SA91, dSTC99, BE90, CNPT91, CO92, DL90, GH91c, LPR92, Odo96, RNR^{+}99, TG99, dMP99]. reconfigurable [SWT91]. Reconnection [STJ90, Lee90b, Tan95]. reconnective [PD90]. Reconstruction [BG95, LL99c, Yat99, Bla94b, CGK^{+}93, CI98, FKMR96, IYOY94, LG93a, TG99]. recoupling [FVR92, FPV94, FPV95, FPV97, Lim91b, Wei99]. recovery [GKP93a, GKP93b]. rectangular [NN93]. Recurrence [Avd96]. recursion [AH99, GH91b, MIO96]. Recursive [HK95, KDH^{+}98, SS90b]. reduce [Can94a, Can94b, Wyb94, ZmSjFyJ97, BS98, BCM^{+}95, Bil92, Boc92, Eas91a, Har97, IK96, Ito94, Kad96, Ren92, Sei91, SMH98, Eas93, Art95a]. REDUCE-based [Boc92]. Reduced [SY95, BD94, LC93b]. reducing [Kas91a, SL99, ZYQ94]. Reduction [SA92, EHK^{+}92, Gut92, MS98, SGT90, Stu95a, Stu95b]. reductions [ABHR99, MRC98]. Reents [Oln98]. Reference [ES92b, Ale93, ES92a]. references [PJ98]. Refined [WBB99]. refinement [TvG91, ZY97, Zie98]. reflection [BT91, MS91a, Wie95]. reflections [Zlo95]. reflective [Fuj94a]. reflectivity [CGR93]. reflex [AG91]. region [CW92]. regions [Lin90, QV95]. register [BTC99a, HDF97, MM91, MTM92]. Regular [ABHR99, IDV99, Ple96]. regularization [Gra92, Iqb95a, RVCF99, Wee92, Wee93]. regularized [Iqb95b]. Reinhardt [Bis91]. rejection [Wan99b]. related [AK99]. relation [ITW97]. relations [Avd96]. relationships [RAD99b]. Relativistic [Gra94, PK91, PI95, PVR97, VVA^{+}94, APaI99, ABH99, BVR95, Ber99, Dub91a, HF91, JPF96, JF97, JTAW97b, KK95, KK97, LN93, Man91, Man93, MSFR96, PWG93, PFG96, PS98a, PVR96, PVRR97, Pös98, RGL97, Szm95, YN94]. relativity [AB99, CC98, FMIM93, JOP98, PS97, Sal94, d’I98]. relaxation [BM99a, CZ92, HD96, Hof99b, PC94a]. relaxed [FF97b, HS93a]. relaxed-orbital [FF97b]. reliability [CvN99, CLM98]. reliable [PJSFM92, PJSF93, Wee92, WZKH98b]. remainder [HW95]. Remote [CDdG^{+}98]. removal [CNPT91, HDF97]. renormalization [Gra98, HS93a, MBC^{+}99b, RF95]. reordering [WDB^{+}91]. REOS [FF97b]. repetitive [UP90]. replicated [Smi92b, SF94a, SF94b]. Replicator [OD99]. Reply [Can94a]. report [Ano98-27]. represent [Cab90]. representation [AUG92, BEFJ99, CY95, HLPT93, OY99, Pod99, Pop98, Sch91b]. representations [Abl98, Ber95, BdM93, Can94a, Can94b, CK91a, HE91, NC96, NRG90, RASS99, Wyb94]. representing [All93]. repulsive [BKMM93]. required [DMvG99]. requirements [MC99]. rescattering [ABH99, ST95]. Resistive [Sto99, DGK97, HPS92, JAVV95, Ott90]. Resolution [KCC96, BK93, MMPT94]. resolvability [vdB93]. resolved [CHUS98]. resolvent [GH91b, WCJ92]. resonance [GDRR94, KR95a, KS92a, ZS91]. Resonances [GFS99, BBB^{+}98, QBP98, ST98, TGG97]. resonant [BV95, Pap92a, PMB99, PGK90b]. respect [VFO94]. respectively [Fle92]. response [DATL99, IMG99, JBH99, KT91, Kot92, KSA^{+}91, NCC^{+}96, RVB99, Sto99, vGSB99]. restarted [LGZL98, Zha98c]. restoration [ZA99]. restricted [VVIBA99, Kre99]. result [Str91b]. Results [HTZ^{+}98, Bor99, CDUT99, FBGV99, FR90, Han99, JH91b, vHKH97, HK96b]. Retrieval [BR95, MPT98, TMG99]. returns [KKKT99]. Review [Art93, Art95a, Art96, Bis91, Bon93, Bro92, CT96a, Cra98, Cyb90, Deb92, Eas93, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Hop97, Jon90, Mon99, Oln98, Par99, Rai91, Reb93, Ryb98, Tay99, Top96, Tru93a, Tru93b, Weg91, Whi93, Duf96, Jam90, Ker90a, Kre99, MMW91]. revised [FG97]. Revision [BPL97, Ano91h, Ano94j]. revisited [Bla94a, ITMI95, Ohl99, dCRSF99]. reweighting [BC99b]. REX [CGR93]. Reynolds [Art91]. RFSFNS [VRS^{+}95, VRS^{+}99]. rheology [MMCH93]. RHF [Kre99]. rhodopseudomonas [HB99a]. Riccati [SL95]. Ricci [Kad96]. Richardson [AA94b, AA94a, AA98a]. Riemann [Abg91, SPP^{+}98, vHK99]. Riemannian [SMH98]. rigid [CP91, HSW99, Ome98, Sta96]. Rigorous [VB92a, SKPP95]. ring [CvN99]. rings [CGRV99]. ripple [DB93]. risk [Deu99, DPP^{+}99]. RLFI [DLS90]. RMATRX [Bar93]. RMATRX-ION [Bar93]. RMATRX1 [BEN95]. RNA [FCS99]. RNGs [BTC99a]. ro [HT93, Sch92, TML93, THF95]. ro-vibrational [HT93, Sch92, TML93, THF95]. road [Kro92]. Robust [OPT99, FKMR96, WZKH98b]. role [Aya92]. room [HCVM99]. roots [BRB92, VRS^{+}95, VRS^{+}99]. Rostoker [MCV^{+}90]. Rotating [Goe93, KW99, LM91b, Nor94, Sta96]. Rotation [MTNA98, FLV99]. Rotation-vibration [MTNA98]. rotational [JWM^{+}96, LPSM92, RT98]. ROTIONS [RT98]. rotor [BGK^{+}94]. roughening [SPA99]. Rounding [ÇYG99, CBTL97]. routine [Dub91b, Lev90, NVFNP93, Sat97b, YJ92]. routines [Mon91]. Routing [dCRSF99]. royal [Kro92]. RPA [Ber90]. rule [CS91, FVGF98]. rule-learning [CS91]. rules [Aya92, JBH99, Sem98, VP96]. Run [Lam98, DNS98]. run-time [DNS98]. runaway [HSP93]. Runge [Sim98, VDVV99]. runtime [SAJ^{+}97b]. Rydberg [Jan90].

S [Bon93, Rai91, MMS93, WW93]. SABSPV [CMNP95]. saddle [Cle93]. salesman [SFM^{+}96]. Salford [Ham94]. Salpeter [WA94]. sample [GGJG97, Wie95]. samples [BB93b, Chi90, CR98]. Sampling [ABE^{+}94, BC99b, CF95, Iqb95a, JHK97, Luc94a, Ran90, SGB99, Wan99b]. sandpile [CVZ99, CMV99]. SAT [STI99]. satellite [VV99]. satisfying [Lem97]. Saxon [GCP95]. SBGBBG [Wie95]. scalable [MM98]. ScaLAPACK [CDD^{+}96]. scalar [Boy95, CZ99, EHK^{+}92, Mar91, RJ90, SGT90, Stu95a, Stu95b, SJ91]. scale [BG98, Chr98, CSP92, FGR98, Fer97, Goz91, KH95, MY92, MHHL99, OF92, PGJ91, PFG96, SW91a]. Scaling [Mid92, TH95, BG98, CZ99, GHGB97, KKKT99, PC99, TP99]. scanning [Ohl92]. SCATTAMPDIAMOL [Nor94]. scattered [Gri97, KK96a]. Scattering [CC91b, EOR94, FJ95b, AIS^{+}97, ADM^{+}92, AMM^{+}94, ABB^{+}96, AM94, ALR97, BA94, Bar98, BS99a, BM92b, BN94, Bur98, BCP91, CMNP95, CC97a, CSS94a, CPCF99, CR98, Cop91, CB99, DS92, FR96, FB94, Gra95, HUCS98, HTI99, HKA^{+}94, IER97, Jac91, JRWWW92, JP92, JPRW^{+}97, Jun95, KC98, KB94, KKKS95, KML98, KS93, Lag92, Lem96, LM95, LBB91, Lin91, LHL98, MF95a, MF95b, Man94, MCV98, MS95a, MSH91, MPN93a, MK91b, NBJK91, NVC96, Nor94, Pap92b, PP99, PMB99, PLC91, RW97, RLWW97, Ric91, SL94a, SM93, SSPD92, SG98a, SR98, SS91a, SSBA91, ST92, SKST91, Szm95, TC97, UTI94, Wha94, XZ91, YN94, YK90, YB92, Zha91, dBPB^{+}93, BKW97]. Scenario [DPP^{+}99]. SCF [Kre99, FZZ92, RDB^{+}90]. Scharfetter [ZYQ94]. scheme [AP97, Aok97, BF91, FT95, Gra92, HS99, IMA94, KKMS99, KV99, MAH98, NY99, POA92, PJSFM92, PJSF93, Ple96, UKA97, Var93, VB91, XYNI96, XYI96, YY99, ZmSjFyJ97, ZYQ94]. schemes [Art91, BDL92, Mic91, Moo94, NL91]. Schladming [Reb93]. Schloss [Cyb90]. Schönung [Weg91]. Schottky [BBB99]. Schrödinger [AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, AH91, AE99, BKWM91, BV95, CRS90, CK91a, DEAM91, Du93b, Hut94, ICS95, IIM92, ITMI95, KKMS99, Kul97, LPC94, LM91b, MRC98, Mic91, PT95, PM91, PSV99, SM91a, SM93, SFVW95, Sim92, Sim93, Sim97, SPT98, WF93, ZZGS91]. Schrödinger-type [AE99]. Schulten [Deb92]. Schüttler [Fin90, Fin91]. Schwinger [HvSA98a, HvSA98b]. science [ACG^{+}99, RLP^{+}99]. Sciences [Hoc92, AFP99, Hoc92]. Scientific [Tru93a, AHZ90, Bra90, CSC^{+}97, CJP^{+}94, Gli90, Ryb98, Ano99o]. scintillation [OA96]. SCOT [Sti94]. Scotia [MMSM93]. SCT89 [SSPD92]. search [EK94, HSM95, PBDZ94, SAY93]. searches [HN97a]. Searching [SFM^{+}96]. Second [BM95, CTDdM98, FJ91, IDVV97, JBSZ97, JW90, Ker90a, Lee90a, McD90, MIO96, MES90b, SCS^{+}96, Sim93, WC90, WCK99]. second-generation [Ker90a, McD90]. second-level [SCS^{+}96]. second-order [FJ91, IDVV97, JW90, MIO96, MES90b, Sim93]. Second-order-accurate [BM95]. section [BKMM93, LN93, MMC95, MJ90]. sections [BKDMM91, BM92a, BM92b, Cap93, CTdO96, CM92b, DM90, Gra95, HD96, KA95, LC96a, Mal90, MPN93a, MPN^{+}93b, NOC98, Pöt99, RT98, SG98a]. segregation [Bas99a, KKM^{+}99]. Seitz [SAZ90]. selected [JN94, Nor94, SF94c]. Selection [Frü93, VP96]. selectivity [SR91]. Self [AW94, Kre99, LPPR91, SLGB96, TAMSW90, ZAW94, AI90a, BR95, BR97a, BS90, BWS94, CDM99, Kni90, KNN91, MS95b, PJ98, PAC^{+}98, PVR96, RJS99, RSW^{+}99, TK91, WF93, Wuc95]. self-adaptive [Wuc95]. self-assembly [RJS99]. self-avoiding [BS90]. Self-Consistent [Kre99, AW94, SLGB96, TAMSW90, ZAW94, BWS94, KNN91, PAC^{+}98, PVR96, TK91, WF93]. self-contained [AI90a]. self-energies [Kni90]. self-energy [RSW^{+}99]. self-gravitating [CDM99, MS95b]. self-interacting [BR97a]. self-interaction [BR95]. Self-organizing [LPPR91]. self-sorting [PJ98]. semi [BBL^{+}97, BZO92, BWS94, BKW97, CS98a, HS96, MB94, SSS91, WRF91, XYE99, Li95]. semi-analytic [BBL^{+}97]. semi-classical [WRF91]. semi-grand-canonical [MB94]. semi-heavy [BZO92]. semi-implicit [CS98a, Li95]. semi-inclusive [BKW97]. semi-infinite [BWS94, HS96, SSS91]. semi-Lagrangian [XYE99]. semiclassical [CPCF99, GH91a, Hor91, Jac94]. Semiconductor [Gus91, BMW91, FKKJ91, JLFS91, Kot92, Pes93, SSHT91, ZmSjFyJ97]. semiconductors [HBH91, RKV91, STR91a]. semicontinuous [HNS99]. semiempirical [FZZ92, ThLL^{+}93]. semileptonic [GJ96, GJ97]. semisimple [NRG90, RG92]. sensitivities [DATL99]. Sensitivity [Hof99a, ZA99, CvN99, Hel99, PGJ91, PCSD99, Pod99, SL99, Sta99, Tur99, You99]. separation [FBGV99, KKK91, NK99, Pur99, vVF96]. separations [Tox99]. separatrix [DMM^{+}97]. sequence [BGCRRO99, Der99, HW95, MSP^{+}98, Wen91]. sequences [BTC99a, Gro91, Mid92]. sequential [Deg99, FB99, LR92b]. SERENA [MS95a]. serial [AS95, GS98b]. Series [Fle92, Hoc92, Tru93b, AB90, BMOF92, BH92, DVH90, DGKS99, Gro91, Lee92, Lio97, Mac92a, Ole96, PBN93, Rod92, Sat97a, VGK90, Wen91]. servers [FP98]. Session [GUW98, Lan98, RL98]. set [Bil91, DLLR96, EHKK93, HvSA98a, HP97a, HT93, Mon91, Sor98]. sets [Pac91, TZC96, Wil90]. settling [HSW99]. setup [Ern98]. SEXIE [RSW92, TMR94]. SG [ZmSjFyJ97]. SHAKE [HR91, SF94b]. shallow [PCSD99]. shallow-water [PCSD99]. Sham [J̈K97a, AK99, BDF^{+}97, J̈K97b]. shape [GCP95, HHH99, SAZ90]. shapes [BRWN96, BWNR99]. sharing [BADP96]. sharp [FBB99]. shear [EH91, JAT97, Mac91]. sheared [CR99]. SHEEP [d’I98]. sheet [PD90, vHK99]. Shell [SH97, ABH^{+}94, CMM92, CCM95, FT92, Gal99, Kre99, LC96a, MMCH93, RJS99, RDB^{+}90, Sun97, TBA^{+}98]. shell-model [Sun97]. SHELL2 [FT92]. shells [Boc92, Kre99, RSW92, TMR94]. Shift [CNO98, BTC99a, HDF97, JF97, MM91, MTM92]. Shift-and-invert [CNO98]. shift-register [HDF97]. shifts [FSMVM93, PI95, SL97, UTI94]. Shlaer [KDH^{+}98]. Shock [Cor99, OY99, BB99, Man91, Zyb99]. shock-like [BB99]. Shock-waves [Cor99]. shocks [HFD90]. shooting [BCP91, EHKK93, Far98, IIM92, ITMI95, Kor99]. short [EK94, LHA95, RJ90, Ryc90b, RJ91a, RJ91b, SAJ^{+}97b, TR93, The94, Zhe99]. short-range [EK94, RJ90, Ryc90b, RJ91a, RJ91b, SAJ^{+}97b]. short-time [Zhe99]. shower [Kas91b, Kas91a, MKPGS97, Sey95]. showers [KT91]. shown [KR99]. Si [RGLV^{+}99]. Siegmund [Fly90]. Sigmund [Fly93]. sign [FBVdlN99]. Signal [BR97b, GKP93a, GKP93b, KD95, SGB^{+}93]. signals [ZKHG99]. Signature [AA99]. silicate [RARF96]. silicon [PHN99, PvBMN99, TAMSW90, Len98]. sillenite [MSR96]. silo [KKM^{+}99]. SIMD [Ols96]. similar [CR98]. SIMLYS [KLS91, KS92b]. Simple [Jor95, JH91b, KNK95, Alf99, BKKS99, BC91b, BLLM94, CC97b, KRVZ98, PJSFM92, PJSF93, Pit97, Pit98]. Simplicial [BBJ95]. simplification [IK96]. simplifications [Di 95]. simplified [EHKK93, FWH95, Sun97, Wan93]. simplify [Por98]. simulate [JKW91, PCC^{+}95]. Simulated [FH97, YBS91, AK91, ST99, Tox99]. Simulating [HS99, LHA95, MS95b, AG91, BJW90, DB93, GHM93, MWA^{+}92, Ohl97, PMB99, RHFR95, TB91, TF98, VES99, XYIT97]. Simulation [ADK^{+}99, AS95, BdT97, BWNR99, BDH99, BMW91, CDUT99, CHC91, CKS91, EH91, HTI99, HSW99, KMT^{+}99, MBZM99b, MFG^{+}98, MCO99, OGG91, PRRL99, ALR99, AF90, Asa98, AS99, BDKM91, BRWN96, BAAD^{+}97, BHY92, BLGT92, BLGT94, BM99c, BCSS99, BS90, BHM99, BP90b, Bin99, BKS99, BP93, BRM99, BVVJA99, BF90b, BF90c, BC99c, BG94, CDM99, CD92, Chi90, CBD^{+}95, CGR93, DCR99, DBVS99, DJUM99, DM91, DPRS90, Fan90, Fan94, FHSO99, Fau91, FMSdC99, FKKJ91, Fin90, Fin91, FMH93, FBH90, FR90, GP95, GDRR94, Gei97, GW95, GLV99, GFAALS98, Gri97, Gup92, Gus91, HTK99, HL92, HSP93, Heu90, Hew94, HN99, ITW97, IMG99, JAT97, JHK97, JLFS91, KT91, Kas91a, KB94, KR95b, KH96, Kud99, KR99, LMF97, Lam96a, Lam96b, LL99b, Lön92]. simulation [MSD93, MC93, MMW96, MC95, MC97, MPB96, MPRS97, MEA^{+}90, MMU99, MKL99, MHHL99, MK95, MSJ99, MSTK93, ME91, MAH98, MT91, MPB91, MP93, MB94, Nak97a, Nun90, Odo90, Odo92, Pai93, PD90, PP99, PG94, PZ91, PGK90a, PL99, PdA99, Poz91, PFBdlN99, RFH^{+}95, RCL98, Rit92, RNGR99, Ryc90a, RJ90, Ryc90b, RJ91a, RJ91b, SMV90, SMBV90, SA92, SY95, SWT91, Sta99, Sta96, Str93a, Sug91, Tan95, TP99, TvG91, TLK94, VB92b, WNS99, WM99, WW90, WCL90, WCK99, WS95b, WT99, YY99, Zha94, Zyb99, dFZLS95, vGHM^{+}95, Sal97, Top96]. Simulations [All99, BA95, Cat95, Alb99, AL94, AAD92b, AM90, Bab92, BB99, BSB99, BS96, BBG94, BT92b, BCOY93, BCOY94, BCOY95, BMM97, Bru95, BMPS99, CF95, DSB91, DLLR96, DMvG99, DMLM91, DSD91, DW99, Dub91a, ERS^{+}95, EK94, FBGV99, FBdR^{+}98, FMIM93, GE99, GJ99b, Han99, HÓ99, HMR99, JBB99, Jan95, JL97, KdLNV93, KB99, KKKS95, KL99b, Kot96, KLS91, KF99, Kut99, LPS99, Lee90b, LB91, LHSD95, LBC91, Lil93, Lil94, LMW96, Luo96, Lüs94, MMH99, MBW^{+}95, MS96, MY92, MD97, MR95, Mon98, MT92b, NS90, PI99, Pes93, PP92, QAd^{+}95, Ran90, RLP^{+}99, RF94, Rho92, RS99, Rou95, RARF96, Sci91, Sel96, SW91a, SP95a, SF94a, SF94b, SSW94, Sol99, Swe91, TC97, TSJ91, Tob99, TD99, TMS^{+}99, UOMO94, VW95, VB91, WLD95, WR99, WJ91, Zha96, Zhe99]. simulations [ZmSjFyJ97]. Simultaneous [SS94b, AP97, Car94, Car96, LR92b]. sine [BT91]. sine/cosine [BT91]. Singapore [Cra98]. Single [AZR96, GCP95, Pap99, AI90b, BTC99b, CKW98, FJ95a, KL93, MMC95, MCO99, NOC98, OC98, Ste97]. single-mode [CKW98]. Single-particle [GCP95, KL93]. single-phase [MCO99]. singlet [JWM^{+}96]. singlet-triplet [JWM^{+}96]. singular [CS99a, UP90]. singularities [MP95]. singularity [BMOF94, Ple96]. site [BR95, BR97a, Ome97, SA99c]. six [RLWW97]. sixth [Ano91h]. size [AI90b, HHH99, PC99]. Skeleton [Dec95]. Skyrme [DD97a, DD97b]. Slab [RJ91b]. slabs [SV94]. Slater [FG95a]. SLC [JW90, JWW91, JWW94]. SLCPM12 [IDV99]. slender [KK91]. slices [YB99]. Slow [Hof99b, NKR99, BC91b, MF94]. slowing [IDW99]. slowing-down [IDW99]. small [CMNP95, DJMT91, HN99, MEA^{+}90, PGJ91]. small-angle [CMNP95, MEA^{+}90]. small-scale [PGJ91]. smart [RCL98]. smart-particle [RCL98]. SMATASY [KR95a]. SMI [GF98]. Smooth [Kup99, BA95, BK92, PD90, Rho92]. smoothed [Man91, Meg94, RHFR95]. smoothers [SZ99]. SMS92 [JF97]. SnO [PRRL99]. Society [Tru93a]. soft [BAS99b, JW90]. soft-committee [BAS99b]. Softcover [Bro92, Fle92]. SoftSciences [AFP99]. Software [BJM^{+}98, AFP99, Ano98-27, BJL^{+}96, BH99, BBP^{+}90, Don99, EABH95, FJ95b, KDH^{+}98, KRVZ98, KLS91, Kru98, Luč95, MXR98, PM91, RTLM92, Top96, Tru93a, Van96]. Solar [VGM90, GBWES97, Goe90, KCK99, Lin90, NS90, PGK90b]. solid [Ask94, FKFE99, MNS99, SZS95, Sul96, XYIT97]. solidification [CP92]. solids [BA95, BBG94, DKT96, Mit96, Sin90, Tar99, WT99, WF90, Rai91]. solitary [Her91, PD96]. Soliton [Zag99, WA94]. Solitons [Fog99, BB99]. Solution [BDD^{+}92, CZ93, DD97a, DD97b, FS91, KSG91, LM91b, MOC91, Moo93, Moo94, VS97, YN94, Abd90, AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, AP98, AW94, AH91, AHH94, BS91b, BK91, Bes98, BvG97, BS99a, BSSH99, Boy95, CTE95, CP95, DMLM91, Fij99a, Fij99b, Fij99c, Fij00, FJP90, HK95, HKM98b, HKM98a, HRK95, ITMI95, IMA94, IDVV97, KKK95, KNN91, Kul97, KL92, LHSD95, LLW98, Lin91, Mac92a, MBW^{+}95, MK96, MOSM93a, PVR97, QV95, Rho97, Rod92, RW95, RSW96, SM91a, SFVW95, ŠČ91, SSHT91, SSF92, SS93b, SL97, TDV96, Tur99, Wee92, YK95, ZAW94]. Solutions [HKC98, DEAM91, DL98, DHL93, FGM98, GUW98, Her91, IBR98, Krö99, PD96, SS91a, Sim98, Sim99, Swa91]. solvated [LDOO95]. solvation [IBR98]. solve [Bin99, DD94, SB97, USW99, Var93, WF93]. solvent [DMvG99, GPSW95]. solver [Aro96, BPG97, BV99, Cam95, LBR92, MP90a, PK91, PHT93, PS98a, SS94a, TGAW96, XYE99, YA91, YIW^{+}91]. solvers [Abg91, VB92a, YIW^{+}91]. Solving [HvSA98a, HvSA98b, Hum90c, Aok97, AK99, BKWM91, CTvB95, CTDdM97, CTDdM98, Du93b, Duf96, FJ91, Hut94, IDV99, JN94, Jor95, KH97, KKY99, NY99, OPP96, PS98a, Pös98, PSV99, Ryb98, ST92, Szm95, TG96, UKK98, WZA^{+}98, XZ91, YYOM90]. Some [Bor99, GM92, KV90, BS91c, CDUT99, UTI94, VKSAN95]. sonoluminescing [MKL99]. Sons [Top96]. sorting [MP99, PJ98]. SOS [LR92b]. soul [Sol99]. source [BM99c, FWH95, KI99, NK99]. sources [KB98, Mak93, NK99]. Sp [NC96]. space [AW94, Blo92, CP99, Coo99, DPD96a, DPD96c, DD95, Gei97, GJ99b, HP93, HLMM99, HKLM99, KS98, LM91a, Man93, MT92a, MES90b, NY99, PVR97, RSW^{+}99, SPP^{+}98, SMH98, SS91b, UKK98, WENG95, WZKH98a, ZAW94]. spaces [PWG93, Rut98]. spacing [KTM96]. spallation [ML92]. sparse [AH99, KNK95, PDZ92, Sci93a, Sci93b, SF94c, TGAW96]. SPARSEM [Sci93b]. sparticle [KM98]. Spatial [DL90, TVH^{+}99, NE93]. spatially [Hea91, Ris96]. SPECFIT [BNSW92]. Special [Fra92, IYOY94, IEY^{+}96, Dub91a, IMES90, Sim93]. Special-purpose [IYOY94, IEY^{+}96, IMES90]. species [RdMP99]. species-dependent [RdMP99]. specific [GUW98, GFAALS98, PI95]. specification [KH96, MKS^{+}96, SMB^{+}93, SKM94, SJB^{+}94]. specified [Lil93, SRB94]. specified-profile [SRB94]. spectra [BNSW92, BMMMMR99, BLGT92, BLGT94, Chi90, FSMVM93, FBdC98, HLT93, HT93, HHK92, JWM^{+}96, LBC^{+}93, LC93b, NCC^{+}96, NvdHPG97, Ohl97, OA96, Pot98, Ste97, Sto99, TC97, TML93, THF95]. Spectral [CMM92, CBG99, CS98a, GHM93, LABV90, MV97, PD90, QV95, SGRZ99, dFZLS95, Mon99]. spectrometer [GRS90, TLK94]. spectroscopic [CT96b, VdN92, ZFZ92]. spectroscopy [AA99, DL91, FJP90, OMC93, PZH^{+}96, PBL94, SAP^{+}95]. Spectrum [dSVTM99, HIK90, LVL90]. specular [LVL90]. speed [IWWY92, YYK^{+}92, YKO^{+}94, YKM^{+}96, YYOM90]. SPEFY [LMS94]. SPH [HI93, MK95, TR93]. Sphere [Str93b, LL90, NVC96, Pan99]. spheres [BM91, Ber90, HSW99, Iri96, LH92, Nie94]. Spherical [CTE95, AC92, CMM92, Cop91, DEAM91, EOR94, HP97a, Lem97, LM91b, Man93, Mar91, MMCH93, NVC96, PVRR97, PVR97, QV95]. spherical-wave [HP97a]. spherically [DL98]. spheroidal [GS98a]. spheroids [CC91b]. Spices [RCL98]. spiking [TVH^{+}99]. spiking-bursting [TVH^{+}99]. spin [BHY92, BDH99, DER90, Flo97, HHCS95, Hof99b, JBB99, JV99, Kno90, Kre99, KBL98, MCMD91, PC99, PTU97, PvBMN99]. spin-12 [Flo97]. spin-density [PvBMN99]. spin-exchange [DER90]. spin-Restricted [Kre99]. SPINDIS [Sun97]. spinless [MJ90]. spinodal [BF90b]. Spinor [Abl98, PRH90]. SPINSGA [Flo97]. spline [Iqb95a, PT92, Pös98, SKG91, UTI94, YYOM90]. splines [BDL92]. split [DBME96, GHE99, BME96]. split-domain [GHE99]. split-operator [DBME96, BME96]. splitting [Hea91, MOC91, RW95, RSW96]. Sponsors [Ano90x]. spread [CPM99, MXR98]. spreading [VS99]. SPRING [Kaw95]. Springer [Art96, Bis91, Bon93, Bro92, CT96a, Cyb90, Deb92, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Hop97, Jon90, Mon99, Oln98, Reb93, Ryb98, Tru93b, Weg91, Whi93]. Springer-Verlag [Art96, Bis91, Bon93, CT96a, Deb92, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Jon90, Reb93, Tru93b, Weg91, Whi93]. SPRITE [RGL^{+}98]. sputtered [MMH99]. sputtering [IKTY99]. SPYTHIA [Mre97]. square [Ber92b]. squares [BDL92, Bru91, CGR93, Dem97, HHK92, LH97, Mar91, MK91a, Nie94, Rob98b, Swa91]. Sr [Car94, Car94]. SRLOG [Car94]. SSGF [BWS94]. SSM [EHKK93]. SSNTD [MFG^{+}98]. SSOR [FFG^{+}96]. Stability [Goe90, ICS95, SS90a, UKA97, AZM92, CNO98, DSB91, DVH90, DMM^{+}97, FKFE99, KCH92, LB99, VGK90, Win91]. stabilized [MT92a]. stack [SRFN92]. stage [SMB^{+}93, SJB^{+}94]. stagnation [CK91b]. stagnation-point [CK91b]. stall [CC91a, JH91b, OGG91]. standard [Bor99, CZ99, HHK92, Kni90, Krö90, SS91b, DKS98]. standardization [Car94, Car96]. standards [Krö90]. standing [Tru94]. Stark [MKWH91]. stars [DF95]. starting [BTC99b]. State [Coo99, BKDMM91, BKWM91, BKMM93, BS96, Bin99, BSSH99, ÇYG99, CBB99, CHM91, CY99, CK99, DM91, FL91, HHR98, Hut94, JW90, LL99a, Lem97, LLW98, LC93a, MMC95, MES90b, MNS99, Nor94, PGK90a, RDF^{+}97, RGL97, RBB98, SS94b, SF93, Szo99, ZN93, dCdJDvL99]. state-space [MES90b]. states [APN^{+}91, BR95, BR97a, BMSW91, Bar98, BV95, CBG99, Du93b, HHT95, JPS^{+}99, Kag92, LHL98, Pap92a, Pap92b, SMR99, SvdPS99, TAMSW90, WA94]. static [Aly90, BWNR99, BKNS97, DL98, Ver99, WENG95]. Stationary [JH91a, PGK90a, Szo99]. Statistical [Kin99, ACG^{+}99, CO92, GLF97, Lob99, LC96b, MR95, Odo96, Ran93, TSJ91]. statistics [KKKT99, ST99, SC99a]. status [Ano98-27]. steady [LLW98, LC93a]. steady-state [LLW98]. stellar [NS90]. stellarator [JMR^{+}93]. step [AH91, BKKS99, CD92, IVDV97, JD91b, KA95, KPD97, NVK93, WF93]. step-forward [WF93]. sticking [Jac94]. stiff [AR95, Aro96]. STO [RLER97]. Stochastic [BHP97, KML98, MMGS93, MLM96, AGIS99, BP93, CT97, CVZ99, Hof99b, KSZ97, SP93b, SP93c, VS97]. stock [KKKT99, Sol99]. stokes [CMM92, MMCH93, HS99, OGG91, QV95, WZA^{+}98]. storage [BF98, OBRR92]. STOs [RLPR91, RLPR94]. straggling [HM92]. strain [Wie95]. strain/stress [Wie95]. strained [Lam96a, Lam96b, RGLV^{+}99]. strained-Si [RGLV^{+}99]. strains [Wie96]. Strategic [Wol98a, Ern98, PDL98b]. Strategies [KH96, BV99, Lön99, PJ99, YBS91]. strategy [BCOY93, Krö96, LGZL98, Rob98a, Smi92b, SF94a, SF94b]. stratified [LVL90, XYIT97]. streamline [CK91b]. strength [FF97b]. stress [Lil94, Wie95]. stressed [STR91a]. STRFLO [Pot98]. string [IJV92]. strings [Str93a]. strip [OPT99]. strips [dQRdS99]. Strömungen [Weg91]. Strömungsmechanik [Weg91]. strong [AB90, PR93]. Strongly [dGT99]. Structural [Wol98b, BM99a, CTJ^{+}95, NHK^{+}95, PCC^{+}95, SWG99]. Structure [Cra98, RARF96, ABF^{+}99, AK99, BL99a, BKNS97, BG99, CUVS95, Der99, DKT96, Fro91c, Fro91d, FS99b, HFD90, HKM98b, JKSC99, JPF96, KWS^{+}96, KK95, KK97, LM91a, Lim91a, LH92, LS93, MCV^{+}90, MCMD91, Mak93, Man99, MOPR99, MSTH92, Mic94, Mit96, MR99b, Pag94, PGJ91, PWG93, PFG96, PBL94, Sti94, SS94b, TBA^{+}98, Yeu91, Zat96, Zei97, Bon93]. Structured [Won90, JAT97]. structures [BN92, BB99, BS91b, BPMW95, CY93, Eic90, FBVdlN99, For91, Fri97, MSJ99, RTLM92, SMV90, SS91a, SW91a, TCE99a, TCE99b, TAMSW90, TvG91, WP98, vHK99]. students [Tob99, WY99]. studies [BBS91, Fin90, Fin91, FV99, HN97a, IKTY99, IK99, LS93, PBL94, PC94a, YAD99, Zei94, dQRdS99]. Study [Kal96, TP99, WW93, Zei97, Ale93, BP90a, CC91a, CHM91, CJ91, CK99, DPP^{+}99, DCM99, DH97b, FT99a, FS99a, Gei97, Goe93, Has99, HZQ93, HHV99, Hum90a, JTAW97b, KS96, KMz91, Li95, NL91, RHGDM99, RGLV^{+}99, RLF96, Sci94, Sik99, Sti94, Tan95, UOMO94, VGK90, VKSAN95]. Sturm [YKM^{+}96, IDV99, PM91]. Styer [Top96]. sub [OPT99]. sub-optimal [OPT99]. subbarrier [VST94]. Subgrid [DJMT91]. subgroup [CP99]. Subject [Ano90y, Ano91-33, Ano91-34]. sublayer [HSW91]. submatrices [GH91b]. subpicosecond [GDRR94]. subprogram [TDV96]. subprograms [EHKK93]. Subroutine [Sty90, Dem97, HM92, HS93a, KT91, LBC^{+}93, LG93b, Sci93a, Sci93b]. Subroutines [DM90, Han90]. subsamples [GGJG97]. subsets [Frü93]. subshell [Ste97]. subshells [GF96]. substance [ML92]. substitutional [HMH99]. substrates [Bha91]. successive [Kob94]. suite [SF98]. suited [FT95]. Sum [FVGF98, Aya92, HK94, JBH99, MG93, Smi92b]. Summary [GUW98, Lan98]. Summation [GJ99a, FPV94, FPV95, Ome97, SLGB96, TB96, Wen91, WA97]. sums [Mon91, Nie95]. superalgebra [CT90]. SUPERCALC [Sei91]. supercomputer [Ano98-27, Din91, Gut95, PS98b, Ric91]. supercomputers [BMW91, RWL^{+}98]. superconductors [Str93a]. superfluid [MBC99a]. superlattices [CT97, PMB99]. Superparamagnetic [DBGW99]. superradiation [BDH99]. supersonic [Mac91]. superstable [Kar91]. supersymmetric [AAH98, DKS98, FHI^{+}98, KKS90, Mre97]. supersymmetries [AH97, AAH98]. supersymmetry [Luč95]. supply [CGM99]. support [DNS98, DH97a, Fro91d, MXR98]. supported [AI90b, TSL^{+}99]. Supporting [Ano91-35, FP98]. suppressing [XYE99]. suppression [DH97a, TLK94]. supralinear [RVB99]. Supramolecular [RJS99]. surface [AA99, Alb99, AM94, Ask94, BHS^{+}99, BWS94, BJK92, CMC^{+}99, DATL99, HE99, HKA^{+}94, IJV92, Jac91, Jac94, Jan95, KT91, LVL90, Lem96, MMW96, Man94, MBC99a, MDMR99, MF94, MK91b, MBC^{+}99b, OE99, Pop98, PRRL99, SW91b, Zei94, Zha91]. surface-dynamic [MDMR99]. Surfaces [Cra98, ATP98, BJW90, CL96, DKT96, FB94, HE94, HS93a, Kis91, PHN99, Pur99, SL94a, SSPD92, SvNF92, SSW94, VG99, WBS95, Wha94, WA97]. survey [TB96]. Susceptibilities [CY93]. suspended [XYIT97]. suspensions [Hin95, HSW99, HS99]. Susycal [KKS90]. SUSYGEN [KM98]. Suura [JRWWW92, SJPW96]. SVZ [Aya92]. Swendsen [BJK92, Min91]. switching [HTZ^{+}98, HN99]. Symanzik [JL97]. Symbolic [AUG92, Aut99, CT90, KKS90, NND98, QTR99, RF95, Sha92, Str93b, Abl98, CC98, Fit90, Gro90, Luč95, MR96, RU92]. symbolic-numeric [Fit90]. symbols [LC90a, LC92, TBS98, Wei11]. SYMCD [Ito94]. symmetric [DL98, Flo97, HE91, ICS95, NRG90, SKWH99, SF94c, TDV96]. Symmetries [CTR98, CHW91, FFB99, Ito94, JOP98]. Symmetrized [DPD96a, DPD96b, DPD96c, DPD96d]. symmetry [BT91, CTDdM97, CTDdM98, CP99, EH91, FG95a, FG97, KTT98, LM91b, MCV98, MSFR96, NRG90, PC94b, PVR97, RG92, Ská90, TEJJ93, YB99]. symmetry-adapted [Ská90]. symmorphic [DPD96a]. Symplectic [Ple96, BGT95]. Symposium [Cyb90]. synaptic [TMG99]. synchronization [GM99, TVH^{+}99]. synchronous [AL94, Mar97]. synergetic [Cyb90, SS93b, SS94a]. system [BO98, BAAD^{+}97, BH92, BVR95, BCSS99, BC91b, CPCF99, CSS94b, CB99, DHK^{+}98, DGK97, DD95, ERB90, GST99a, GST99b, GRS90, Hal90, HHCS95, Heu90, HM99, Ker90a, KCC96, Kli98, KM95, KKY99, Kot92, LSZ92, xLzHjW95, MRPC94, MRC98, MMS93, Mer90, MAB^{+}94, Mor92, PBDZ94, PVR97, RDB^{+}90, Sat97a, SGB^{+}93, SCS^{+}96, SWT91, SAJ^{+}97b, jWzHxL95, Whi98, WDB^{+}91, WGH^{+}98, dFZLS95, Rho97]. Systematic [VPE94, FGM98]. Systems [Str93b, Abd90, AAS95, ATP98, AL94, BBMD99, BHY92, BDH99, BSST90, BKNS97, BT92b, BKS99, BMM97, BF90c, CDM99, CZ99, CHW91, CTdO96, CG99, CD91, CD92, CC91b, CGM99, CSS94b, CZ92, DS96, DdMdO^{+}99, Duf96, Far98, Fer97, Flo97, FBH90, FS99b, FG95b, GUW98, GF98, GLV99, GM99, HHT95, Har97, HB99a, Hum90c, Ito94, IL91, IDVV97, JK99, Jan90, JV99, KdLNV93, KKH^{+}98, KMz91, KSA^{+}91, KDH^{+}98, Kre99, KBL98, KN95a, KN95b, LMF97, Lan98, LDOO95, Lia95, LHA95, xLzSlLzH94, LMS^{+}93, LEBB99, LD91, McD90, MMTK94, MBMM99, MH99, PG96, Pot98, Rad99a, Ran90, RDF^{+}97, Rie99, RSW^{+}99, Rod92, RAD99b, Ryc90b, Sat97b, SGF99a, Sei94, SS93a, SvdPS99, SM97, SP95b, Sou99, SM91b, SH91, TB91, TZC96, Van96, VS99, VVGI99, Zag99]. systems [Zha94, ZK99, dMCT99, dQRdS99, dVdRL93, Gie90]. systolic [Bru95].

T3D [BAAD^{+}97]. T9000 [MLM96]. table [MT92b, Wei11]. tables [DDDL98, Kae95, KW96, MN92, PP94, TBS98]. Tabulated [WR99]. Tackling [CTS99]. tailed [Frü95]. tails [SEZ99a]. Talus [SF98]. tanh [PD96]. tanh-function [PD96]. taps [Shc99]. TARA [Deg99]. TARCER [MS98]. target [MJ90, RMMUJ99]. targets [BA94]. task [MAB^{+}94]. TAUOLA [JW91, JKW91, JWDK93, JWJK92]. taxonomically [MPT98]. Teaching [PMUV90, Tob99, Gil99, MNS99]. technique [AH99, BN92, BM99a, FWH95, HS95, HSM95, Kas91a, KR95b, LLW98, Lio97, Lud91, MOC91, MN90, Nun90, Ome97, OPT99, PRH90, SZG99, TSK^{+}97, ZY97, dCdJDvL99, BME96, DBME96]. Techniques [WF90, Abg91, DR90, Fle92, Hol96, Isk91, PK98, SKG91, SP95a, SA91, TB96, WA97]. technologies [BJM^{+}98, ZA99]. Tecolote [HAC^{+}99]. Teilchenphysik [Reb93]. Temperature [LL99a, AG99, AF90, FKFE99, FCS99, HCVM99, Lil94, MCMD91, Szo99, VW95]. Temperature-driven [LL99a]. temperatures [Jac91]. templates [Bla94a, OPY92]. Temporal [PGK90b]. ten [Den99]. tension [CMC^{+}99, IJV92]. tensor [Can94a, Can94b, Chr98, IK96, Por98, SPP^{+}98, Wie95, Wyb94]. tensors [EHK^{+}92, Wol98b]. teraflop [Ano98-27, MMS93]. teraflop/s [MMS93]. term [CMC^{+}99, LLW98]. terminal [Luc94a]. terms [BMSW91, KM91]. ternary [PFVC99]. Test [GGSR94, LP99, NSE95, Pac91]. testbed [MM98]. tested [Man91]. tests [DL91, FJ95b, HGCM98]. tetrad [CC98]. tetrahedron [CY95]. tetrathiafulvalene [FT99a]. tetrathiafulvalene- [FT99a]. textual [Ker90b]. their [BdM93, DKT96, HDF97, IIM92, MKS^{+}96, Pös98, RBB98, SS90a, SKM94, Sim93]. them [BP90a]. themes [Tru94]. Theorems [Rut98, Lio97]. theoretic [Hol96]. Theoretical [Zei94, BA94, Pag94, ZFZ92, dGT99]. theories [All99, Coh92, GM92, JH91b, KKS90, LM91a, RF95, SKG91, TKS91].

Theory [Bar98, AUG92, AM94, AK99, AB90, BKDMM91, BVR95, BKNS97, BG99, FGR98, FS99b, Fuj94b, J̈K97b, Kot96, LC96a, Lüs94, MF95b, MSTK93, MNS99, PHN99, PVRR97, PVR97, QBP98, RdMP99, Reb93, SGRZ99, SS94b, TZC96, ThLL^{+}93, VV99, WRF91, Wha94, J̈K97a]. Thermal [Rea95, ALR99, CDUT99, Cop91, GFAALS98, MY92, MDSF^{+}99, WF90]. thermodynamic [BK93, KL99b, LMF97]. thermodynamics [LH92, PBL94, dSVTM99]. thermometric [HP97b]. theta [Lee90a]. thick [Bha91]. thin [HB99b, Wie96]. third [AI90a]. Thomas [Mac92a, Moo93, PR93, ZZGS91]. those [ZFZ92]. Three [Art95b, BS98, DSB91, PL99, SJ91, TBCH90, AN97, AAS91, AC92, Avd96, BR97a, BS95, BF90a, DER90, DM91, DPRS90, FMH93, Fuj94a, Goe93, Gri97, Heu90, IYOY94, IMA94, KH97, LP93, Lee90b, LP90, Mac91, MOC91, NVK93, Nar96, QV95, Rit92, RDMS95, SBK98, SSHT91, YIW^{+}91, Zie99]. three-body [AAS91, BS95, NVK93, Rit92]. three-denominator [RDMS95]. Three-dimensional [Art95b, DSB91, PL99, SJ91, TBCH90, AN97, DER90, DPRS90, Fuj94a, Goe93, Gri97, Heu90, IMA94, KH97, LP93, Lee90b, LP90, MOC91, QV95, SBK98, SSHT91, YIW^{+}91, Zie99]. three-dimensionality [Mac91]. Three-loop [BS98, Avd96]. three-state [DM91]. tied [DVH90, VGK90]. tight [CGM^{+}96, IOM95, DCR99]. tight-binding [CGM^{+}96, IOM95, DCR99]. Tikhonov [Wee92]. tiling [LL99a]. Tilted [Ber95]. Time [BS91a, BGG91, BK91, BBS91, GGH91, Jac91, KC98, Lee90b, SDB96, SKST91, Wha94, Wuc95, AR95, BKDMM91, BKMM93, BV99, BVR95, Bil91, CGdlR99, Cam95, CK91a, CD92, CKW98, DEAM91, DeV91, DNS98, Eng91, Eva92, GH91a, Gei97, GHE99, GRS90, HKA^{+}94, ICS95, IL91, Jan95, JH91a, Kas91a, KPD97, KNN91, KS98, KR91, Lag99, LK96, LM91b, Lin91, LD91, Man93, Mic91, MT92a, MSH91, MP90a, NVK93, NBJK91, PMB99, PLC91, Pot98, PSV99, RSW^{+}99, Sat97a, Sch91a, SP93b, SEZ99a, SPT98, SM91b, ST98, TB91, TK91, Zhe99, ZZGS91, vGSB99]. time-advance [LK96]. Time-asymptotic [GGH91]. time-delay [ST98]. Time-dependent [BS91a, BGG91, BK91, BBS91, Jac91, KC98, Lee90b, SKST91, Wuc95, BKDMM91, Bil91, CK91a, CKW98, DEAM91, Eng91, Eva92, GH91a, GHE99, HKA^{+}94, ICS95, Jan95, JH91a, KNN91, KR91, LM91b, Lin91, LD91, Mic91, MP90a, NBJK91, PLC91, PSV99, Sch91a, SPT98, SM91b, TK91, ZZGS91, vGSB99]. time-domain [MSH91]. time-evolution [BVR95]. Time-independent [Wha94, Pot98]. time-step [CD92]. TIMEDEL [ST98]. times [JGP93, SPP^{+}98, SMH98]. timescales [LHA95]. timestep [SSK97]. tin [MOPR99]. tiny [SG99]. TJ [JMR^{+}93]. TJ-II [JMR^{+}93]. TL [RVB99]. TLM [SSBA91]. TOF [GRS90]. tokamak [AMS96, CBD^{+}95, DB93, JAVV95, KCH92, MOC91, Sto99]. tokamaks [AZM92, AHH94, Eva92, FWH95, KH97, MLB93, SRB94]. Tokyo [Fly90]. Tomographic [HLMM99, HKLM99]. tool [BBB^{+}98, CJP^{+}94, KH95, MSK98, Mor92, Whi98]. tool-kit [Whi98]. toolkit [LN92]. Tools [Kas91b, BJL^{+}96, DD94, SP95a]. top [HSM95, HE91].

TOPAZ0 [MPN^{+}93b, MNPP99]. TOPAZO [MNPP96]. Topics [Tru93b]. topological [Hoe90]. topology [AAD92a]. toroidal [CKS91, DB93, HSP93, KRT95, LBS96, PL99, TSK^{+}97, vMF94]. toroidal-field [DB93]. total [BKMM93, CSP92, HP97a, MMC95, SSK97]. total-energy [HP97a]. trace [Wes93]. TRACER [JL93]. traces [Gra92]. Track [CGK^{+}93, Frü95, Frü97, FS99a, MS96, OPY92, SRF94, TG99, Bla94b, Dik94, FKMR96, FFS98, LVC^{+}94, OPT99, SAG91b]. Tracking [BGK^{+}94, ADS92, FS99a, GH91c, SZ99]. tracks [BO98, BE90, Frü93, KK96a, SA91]. traders [Sol99]. Traffic [BC99c, Sug99, WNS99]. trained [SA95]. training [GC95]. trajectories [BC91b, MS95a, SB97]. Trajectory [YN91, ZZGS91, KLS91]. transcription [MDSF^{+}99]. Transfer [SA99a, BRWN96, CK99, FT99a, HB99a, Jac94, MF94, NN93, SRR95]. transfers [BvG97]. transform [AL93, BT91, DEAM91, FS91, GH91a, Goe93, Iqb95a, Iqb95b, PJ98, RBB93, Sec99, SL95]. Transformation [Ker90b, AH92, Bal95, Kul97, WDB^{+}91]. transformations [HW95, JMSW99, Mid92, NND98, Wen91]. transformer [AZM92]. transforming [Gro91]. transforms [BRB97, Bra91, Dik97, Lem97, RBB96, SL95]. transient [GFAALS98]. Transition [AM94, Wan99a, BKMM93, BHS^{+}99, BCD^{+}93, Bin99, BF90b, CDUT99, CDG^{+}90, ÇYG99, FT99a, FNFBdlN99, FGH91, GPM95, IMG99, LL99a, LLM91, MHN99, MSTH92, Mic94, SVV99, VSS91, ZFZ92]. transitional [PZ91]. transitions [Alb99, BL99a, BC99a, BAS99b, CMC94, FG91, PD90, STR91a]. Transmission [SSBA91, BPMW95, SYM98]. Transmission-line [SSBA91]. transonic [WJ91]. transparent [AE99]. Transport [HD96, Mon99, BBL^{+}99, BV99, CM92a, CBD^{+}95, Fuj94a, GSS97, HBH91, HHC90, JN94, Mac92b, MMH99, MLB93, Mor92, NE93, Pai93, SA99a]. Transport-Dominated [Mon99]. transputer [BF90b, BF90c, CKS91, CP91, MLM96, MPRS97, RBP^{+}90, Yeu91]. transputers [CA92, FVR92]. transversity [HKM98a]. trap [DJUM99, MOSM93b, MOSM93a]. TrapCAD [VB96]. trapping [MOSM93b, MOSM93a, MOSM96]. traps [VB96]. Traveling [Pan99, BD99, SFM^{+}96]. travelling [TG96]. Trease [Whi93]. treating [Fan94]. treatment [Bla94b, FFS98, Fuj94a, HP93, PLC91, PCNO94, PCC94, PC94b, PC94c, PC94a, PB96, Sal91, ZKHG99]. tree [BADP96, BAAD^{+}97, CDM99, NK96, Ols96, PBDZ94, PG94, SL94b]. tree-code [CDM99]. trees [FLV99, KK96b, Krö96, MSK98, SS90b]. Trends [Lan98, Ern98, Rob98a, TSL^{+}99]. triangular [Has99]. triangulated [BJW90, BJK92, Cat95, SSW94]. Triatom [TML93]. Triatomic [ATP98, OMC93, Sch92, SvNF92, TML93, THF95]. triaxial [BM92b, MR96]. tribute [Ano94-33]. tridiagonal [DHL93]. trigger [SCS^{+}96]. triode [AG91]. triple [LC90b]. triplet [JWM^{+}96]. triply [BMSW91]. trivial [AAD92a]. tropospheric [LP99]. TRS [STR91a, STR92]. TRSS [STR92]. truncated [DVH90, HvSA98a]. truncation [SAZ90]. Tsallis [ST99]. TSYM [MSFR96]. tube [BO98, Man91]. tubes [BPG97, Str93a]. Tunable [BF91]. tunnel [HP97b]. tunneling [Gil99, PMB99, ZN93]. turbulence [AG99, Art94, Art95b, GA90, GHM93, KR95b, VGM90, Art93]. turbulent [CBD^{+}95, DJMT91, MY92, dFZLS95]. Turek [Cra98]. turn [TGG97]. tutorial [Cab90]. twist [Bor99]. Two [BKWM91, BPG97, BK92, BCP91, CY99, DLS90, Fle92, Goz91, HN97c, Lee90b, MDSF^{+}99, RB99, Sch98, SAP^{+}95, SG99, YIW^{+}91, AAS95, AH91, AAD92a, ACA99, ABE^{+}94, AAH98, BR97a, BH90, BS91b, BPP98, BLLM94, CS99a, Deg99, DGK97, DCM99, Dub91b, FT92, FJKvO95, FBH90, FG95b, GBWES97, GGSR94, Han90, HD92, HSSB98, HN97b, JW90, JD91b, JTAW97b, KR95b, LM91a, LPSM92, MC93, MSR96, MS98, MMTK94, MH96, MKPGS97, NVFNP93, OPP96, PC90, PI99, PFVC99, Pöt99, RLPR91, RLPR94, RLER97, SGB99, Sat97b, Sci94, SVBD92, SKPP95, Szo99, Ver99, WZKH98a, Wei93, Wil90, YB99, YK90, ZN93, ZFZ92]. Two- [Lee90b, YIW^{+}91]. two-body [ACA99]. Two-center [HN97c, HN97b, RLPR91, RLPR94]. two-centre [Dub91b, Han90, HD92]. two-component [ABE^{+}94]. Two-dimensional [BPG97, BS91b, CS99a, DGK97, DCM99, FBH90, FG95b, HSSB98, JTAW97b, KR95b, LM91a, MC93, MMTK94, PC90, PI99, PFVC99, Ver99, Wei93, YB99]. two-electron [AAS95, RLER97, Wil90, ZFZ92]. two-ion [GBWES97]. two-jet [Pöt99]. two-loop [FT92, FJKvO95, MS98]. two-nuclei [Sat97b]. Two-photon [Sch98, BLLM94, MH96, MKPGS97]. Two-program [CY99]. Two-scale [Goz91]. two-state [ZN93]. two-step [AH91, JD91b]. two-temperature [Szo99]. TWOGEN [BLLM94]. TWOGEN-a [BLLM94]. type [BMOF92, BP90b, CTE95, GJ96, HW95, HZQ93, KCC96, LP93, Moo94, Rut98, VD94, YKM^{+}96, Zag99, AE99]. types [FR90, MKS^{+}96, SKM94].

UCL [Eis98]. ugly [Lan99b]. Uhlig [Hop97]. UK [Tru93a, MTG98]. UK£ [Top96]. ULF [KV90]. ultrarelativistic [Zha98a, Zha98b]. ultrashort [Eng91]. unbinding [HHGCM99]. Uncertainty [Hel99, DPP^{+}99, Hof99a, MMU99, RAD99b, SWG99, ZA99]. uncertainty-like [RAD99b]. Understanding [Bin99]. unfolding [NCC^{+}96, SGRZ99]. uniaxially [STR91a]. UNIBAB [AMM^{+}94]. Unified [Wei99, Bla94b, YY99]. uniform [HKG90, OF92]. uniform-scale [OF92]. uniformity [KTM96]. unimolecular [SR91]. Unitarisation [Sal97]. unitary [DG99, NND98]. units [BWNR99]. Universal [Wil90, BvEW91, BW94, DD95, Ham95a, Ham95b, YA91, YIW^{+}91]. Universality [App99, Sou99]. Universitätswochen [Reb93]. University [Art93, Eas93, MG90, DMLM91, MBW^{+}95]. UNIX [MMTK94, RBS94]. unoriented [Glü97]. unpredictable [GP99]. Unrestricted [PB96, TD99]. unsteady [CK91b, MT92a, OGG91]. unstructured [ADS92, KZS95]. unsymmetrical [YKO^{+}94]. update [JWJK92]. updated [TMR94].

Upgrade [JPRW^{+}97, JW91, JW95]. upon [Ren92]. upper [Top96]. upper-level [Top96]. upwind [Art91, ZmSjFyJ97]. urban [MSJ99]. USA [Whi93]. usage [BLS96]. Use [UP90, YK95, Art91, BVVJA99, BC91a, CD90, DHL93, EABH95, FZZ92, HE94, Hum90b, LK96, MKS^{+}96, MSW92a, PC94b, Siv91, SL99, YYOM90]. useful [ES92a]. User [GLSW95, Gli90, JN94, MG90, Zan90]. User-configurable [GLSW95]. Using [MB91, Oln98, AA94b, AA94a, AAKP95b, AAKP95a, AA98a, AAKP98, ADK^{+}99, BMOF92, BB93a, BKWM91, BKMM93, BM99a, BB93b, BdT97, BA95, BPP98, BRB92, BRB97, BS99a, BSSH99, BMW91, BC99c, CPM99, CRS90, CT90, CTDdM97, CTDdM98, CGK^{+}93, CKS91, CGHB98, CKW98, CA92, DL90, Dik97, DGK97, DM98, EHKK93, FVR92, FPV97, FMSdC99, Fer97, FBdR^{+}98, Flo97, FKMR96, FS99b, FT95, GRRG95, GJ98, GH91a, GKP93a, GKP93b, GH91b, Her91, Hol96, HRK95, Jaq90, KDH^{+}98, KN95a, Lag99, LBB91, LC96a, LS93, MY92, Mar97, MMC95, MR99a, MS91a, MOC91, MAB^{+}94, MCO99, MSH91, MS91b, Moo93, Moo94, MK91b, MTNA98, Nun90, OGG91, PZH^{+}96, PP99, PI95, PK98, QBP98, Rob98b, Rou95, RBB96, Ryb98, ST99, SB97, SAJ^{+}97b, ST98, SS94b, TP99]. using [TC93, TZC96, VB91, WP98, WCJ92, Zei97]. usy23 [FHI^{+}98]. utilities [Fro91d]. utilizing [KSA^{+}91].

V [Cra98, BBL^{+}97, Fau91, Hol96, Lam96a, Lam96b, Pes93, PC94a]. v1.0 [CA92]. v1.1 [FIK^{+}97]. v2.0 [FHI^{+}98]. Va [RP95]. vacancy [HMH99]. Vacuum [Hni94, BS98, vMF94]. Vacuum-polarization [Hni94]. Validation [SA95]. validity [CMC^{+}99]. value [CW92, FS91, IDVV97, KRT95, Sim93, Sim98, Sim99, UP90]. valued [AH97]. values [dlVM98]. variability [AGIS99]. variable [GFAALS98, HLPT93, Jor95, Nak97b, SGF99a, Szm95]. variable-charge [Nak97b]. variables [BPP98, NVFNP93, SVBD92]. variance [Jan99, SL99]. variant [FT95]. variates [HK94]. variation [Tau95]. Variational [SGB99, CMM92, FS91, VS97]. variations [AP98, IIM92]. Various [HE99, HRK95, LPC94, NL91]. VAX [BHL92, Fan90]. Vector [RJ91a, AS95, ALR97, BMSW91, CD91, Eas91a, ERB90, FT95, Goe93, GW98, HRK95, HHGCM99, Kni90, LG93b, MM91, Mid92, QTR99, Rap91a, RDB^{+}90, SSHT91, WDB^{+}91]. vector-concurrent [SSHT91]. vector-parallel [AS95]. vector-serial [AS95]. vectorisable [AAS91]. vectorisation [MPB91]. vectorised [GS98b]. Vectorizable [Kob94, TYJ92]. Vectorization [Pae91, Sek92]. Vectorized [BS90, DER90, RJ91a, RJ91b, WK90, Heu90, Min91, TOSM91, WW90, WCL90]. vectors [BTC99a, Lil93, SS90a, Shu94]. Vegas [Ohl99]. velocity [App99, BG95, PW99, Sug99]. Verification [Meg94]. Verlag [Art96, Bis91, Bon93, Bro92, CT96a, Cyb90, Deb92, Fin90, Fin91, Fle92, Fly90, Fly93, Gie90, Hoc92, Jon90, Reb93, Tru93b, Weg91, Whi93]. Verlet [CD90, CD91, CD92]. versatile [KD95, PRL94]. Version [Ohl97, AA98a, AAKP98, AI90a, ABH99, AMM^{+}94, BW94, Eas91a, FG97, HLL^{+}95, JW91, JWW91, JWDK93, JWW94, JW95, JPRW^{+}97, JL93, JN94, Kaw95, KL99a, KS92b, xLzHjW95, Lön92, LN92, Lön99, hLTM^{+}92, LPR^{+}93, MSFR96, MNPP99, MSTK93, Pi92, ST95, STR92, Ská90, SHL^{+}95, TOSM91, TRSM98, WK90, Zau94, Gra95]. versus [BHP97, Lam96b, MKL99]. Vertex [FKMR96, LG93a, CGK^{+}93]. vertical [NN93]. very [KB94, LHA95]. VI [PB96]. via [Bas99a, Bra98, JMSW99, MTM99, Rob97, RARF96, Smi92a]. VIBR3AT [OMC93]. vibration [LPSM92, MTNA98, Pur90]. vibration-rotational [LPSM92]. Vibrational [LGZL98, VBS96, CBG99, GW98, HLT93, HT93, JF92, JF99, Sch91b, Sch92, SVV99, TCE99a, TCE99b, TML93, THF95]. vibrationally [Lim96]. vibrationally-adiabatic [Lim96]. vibrations [IK99, PvBMN99]. Vichniac [Gie90]. videoconferencing [GN98]. VIDSIM [SMV90]. VII [BMSW91]. viii [Fin91, MSW92b]. violating [GW95]. violation [dAASZ97]. ViPA [GW98]. Virtual [FT99b, Eas91b, IYOY94, Pöt99]. viscosity [FR90, OY99]. viscous [Ris96]. vision [DBGW99]. VisRes [BBB^{+}98]. visual [CNPT91, NHK^{+}95]. visualization [KB99]. Visualizing [Ric91]. Vlasov [Fij00, BBL^{+}99, Bas99a, CTE95, ECT93, Fij99a, Fij99b, Fij99c, NY99, Nun90, UKK98]. VLF [Nun90]. VMEC [JMR^{+}93]. VNI [Gei97]. Voigt [LG93b]. VOIGTL [LG93b]. Vol [Hoc92, Whi93, Fin90, Fin91, Gie90]. Volume [FNFBdlN99, Tru93b, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90i, Ano90j, Ano90k, Ano90l, Ano90m, Ano90s, Ano90t, Ano90u, Ano90v, Ano90w, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91m, Ano91n, Ano91o, Ano91-29, Ano91-30, Ano91-31, Ano91-32, Ano91-34, Ano92b, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93o, Ano93p, Ano93q, Ano93r, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94m, Ano94l, Ano94n, Ano94o, Ano94p, Ano94q, Ano94-27, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95k, Ano95l, Ano95m]. volume [Ano95n, Ano95o, Ano95p, Ano95z, Ano95y, Ano95-27, Ano95-28, Ano95-29, Ano95-30, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96l, Ano96m, Ano96n, Ano96o, Ano96p, Ano96q, Ano96y, Ano96z, Ano96-27, Ano96-28, Ano96-29, Ano96-30, Ano97i, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97q, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97-30, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, CF95, Iri96, KCC96, Kud99]. volume [LL90, MSH91]. Volumes [Ano99g, Fle92]. Vortex [BF90a, CS99a, HSW91, KK91, SBGM91]. Vortex-in-cell [BF90a]. Vries [HAR91, Var93]. vs [BA94, IMG99].

W [Art93, Art96, Hoc92, Whi93]. W. [Oln98]. walks [BS90]. wall [PD90]. walls [JAT97, SA99b]. Wang [BJK92, Min91]. Wannier [CL96]. warped [SPP^{+}98]. Warr [Tru93a]. waste [DPP^{+}99, Hel99]. water [FT99b, HCVM99, KN95a, MT91, Pan99, PCSD99]. Wave [DL91, NOC98, dVdRL93, AS92, Art94, Art95b, BSST90, BME96, BHP97, BD99, CY99, Deg99, DBME96, Eis98, FG91, Fuk94, Gib92, Goz91, GD93, GHE99, HP97a, Her91, JKSC99, JH91a, KB98, KWS^{+}96, Lem96, Lin91, LMS^{+}93, MSR96, MMC95, MK91b, Nun90, PD96, PAC^{+}98, Riz99, SM91a, SM93, SFVW95, Sat97b, Sch91a, Sch91b, SZG99, TYJ92, Zha91, Zyb99]. Wave-function [DL91, BHP97]. wave-operator [JH91a]. wave-packet [BME96, DBME96, MK91b]. wave-particle [Nun90, PAC^{+}98]. wavefunction [DeV91, DG99, Hea91, PK91, PHT93]. wavefunctions [BDL92, CT94b, Hea91, JWF93]. waveguide [CHC91]. wavelengths [DJMT91]. Wavelet [WZA^{+}98]. Wavelet-distributed [WZA^{+}98]. Wavelets [WKH98, SKWH99]. wavepacket [BS91a, GH91a, HKA^{+}94, KC98, MS91a, NBJK91, SKST91]. waves [Cor99, DPD96a, DPD96b, DPD96c, DPD96d, JAVV95, KV90, NVC96, Pan99, SV94, TG96, YK90, YB92]. way [BTC99a, Tur99, WF93, YB99]. Wbase [Can94a, Wyb94, Can94b]. weak [LB91]. weakly [LLW98, PdA99]. weather [Ker90b]. Web [CDdG^{+}98, Lan99b, Lan99a]. Web-enhanced [Lan99a]. Wedig [Art96]. Weight [KP94, OPT99]. weighted [JBH99, PBDZ94]. Weinberger [Cra98]. well [FT95]. wells [Lam96a, Lam96b]. Wendy [Tru93a]. Weniger [BRB97, RBB96, RBB98]. WEXTER [BPP98]. where [ES92a]. whether [KS98]. Whittaker [BMOF92]. Whole [Str93b]. wide [FR96, CDdG^{+}98]. wide-angle [FR96]. wideband [KSA^{+}91]. WIEN [KWS^{+}96]. Wigner [SAZ90, CPCF99, Sha92]. Wigner-Seitz [SAZ90]. Wiley [Top96]. Willett [Tru93a]. Willoughby [Kal96]. Wilson [JL97, Kal96]. wind [GBWES97, VGM90]. Windows [AO98, FP98]. WIRED [CDdG^{+}98]. within [AAS95, Bar93, KH96, Lem96, VV99, ZA99]. without [JKSC99, Pae91, Pop98]. WIX [Ran93]. Woods [GCP95]. WOPPER [ABH^{+}94]. work [BADP96, FBB99]. work- [BADP96]. Workflow [BBE^{+}98]. working [MXR98]. workstation [MP90b, RBS94]. workstations [FP98]. World [BS96, CDdG^{+}98]. World-Wide [CDdG^{+}98]. wormlike [Krö99]. Woźniakowski [Kle92]. WPHACT [AB97]. Wright [Eas93]. write [Tob99]. WTO [Pas96]. WVM [Wie96]. WW [AB97]. WWF [vOFB94]. WWGENPV [CMNP97, MNP95]. WWW [MXR98]. WY [Whi93].

x [Art96, AI90b, Art93, CT97, CT96b, CGR93, KSSB99, LMW94, MS95a, Oln98, Tay99, Wie95]. X-ray [AI90b, CT97, CT96b, CGR93, KSSB99, MS95a, Wie95]. XANES [AI90a]. Xfey [Lai98]. xi [Whi93, IK99]. xii [Art95a, Bon93, Reb93]. xiii [Bro92, Hoc92, Weg91]. xloops [BFK98]. XRLINE [AI90b]. xvi [Deb92, Tru93b]. xxiv [Art93].

Y-MP [MSTK93]. Y. [Tay99]. Yamamoto [BP90b]. YASP [MP93]. year [Den99]. Yennie [JRWWW92, SJPW96]. YFS2 [JW90]. YMn [GE99]. York [Fly90, Oln98].

Z0POLE [KS92a]. Zachary [Ryb98]. ZEBEC [KRVZ98]. Zernike [HRK95]. Zero [MCMD91, KML98, LM95, VS97]. zeros [KRVZ98, SG99]. ZhuKè [Lim96]. zinc [SMV90]. zinc-blende [SMV90]. ZnSe [RHGDM99]. ZnSe/BeTe [RHGDM99]. ZPC [Zha98b, Zha98a].

References

Abrashkevich:1994:FPF

[AA94a] A. G. Abrashkevich and D. G. Abrashkevich. FDEXTR, a program for the finite-difference solution of the coupled-channel Schrödinger equation using Richardson extrapolation. Computer Physics Communications, 82(2–3):209–220, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901694.

Abrashkevich:1994:FDS

[AA94b] A. G. Abrashkevich and D. G. Abrashkevich. Finitedifference solution of the coupled-channel Schrödinger equation using Richardson extrapolation. Computer Physics Communications, 82(2–3):193–208, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901686.

Abrashkevich:1998:FNVa

[AA98a] A. G. Abrashkevich and D. G. Abrashkevich. FDEXTR 2.1: a new version of a program for the finite-difference solution of the coupled-channel Schrödinger equation using the Richardson extrapolation. Computer Physics Communications, 113(1):105–107, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000332.

Avrigeanu:1998:PLD

[AA98b] M. Avrigeanu and V. Avrigeanu. Partial level densities for nuclear data calculations. Computer Physics Communications, 112(2–3):191–226, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000502.

Adam:1999:SFS

[AA99] Gh. Adam and S. Adam. Signature of Fermi surface jumps in positron spectroscopy data. Computer Physics Communications, 120(2–3):215–221, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002209.

Amir-Azizi:1992:ALF

[AAD92a] S. Amir-Azizi and G. J. Daniell. Application of the linear filtering algorithm to two models with non-trivial topology. Computer Physics Communications, 70(3):451–458, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465592901069.

Amir-Azizi:1992:LFA

[AAD92b] S. Amir-Azizi and G. J. Daniell. Linear filtering algorithms for Monte Carlo simulations. Computer Physics Communications, 67(3):443–452, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290052Z.

Ayari:1998:CLS

[AAH98] M. A. Ayari, M. I. Ayari, and V. Hussin. Computation of Lie supersymmetries for the supersymmetric two bosons equations. Computer Physics Communications, 115(2–3):416–427, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001192.

Abrashkevich:1995:FPF

[AAKP95a] A. G. Abrashkevich, D. G. Abrashkevich, M. S. Kaschiev, and I. V. Puzynin. FESSDE, a program for the finiteelement solution of the coupled-channel Schrödinger equation using high-order accuracy approximations. Computer Physics Communications, 85(1):65–81, January 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400107D.

Abrashkevich:1995:FES

[AAKP95b] A. G. Abrashkevich, D. G. Abrashkevich, M. S. Kaschiev, and I. V. Puzynin. Finite-element solution of the coupledchannel Schrödinger equation using high-order accuracy approximations. Computer Physics Communications, 85(1):40– 64, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400106C.

Abrashkevich:1998:FNVb

[AAKP98] A. G. Abrashkevich, D. G. Abrashkevich, M. S. Kaschiev, and I. V. Puzynin. FESSDE 2.2: a new version of a program for the finite-element solution of the coupled-channel Schrödinger equation using high-order accuracy approximations. Computer Physics Communications, 115(1):90–92, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559800099X.

Alvarez:1991:VAC

[AAS91] Luis Javier Alvarez, Ali Alavi, and Jörn Ilja Siepmann. A vectorisable algorithm for calculating three-body interactions. Computer Physics Communications, 62(2–3):179–186, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190093Z.

Abrashkevich:1995:HPC

[AAS95] A. G. Abrashkevich, D. G. Abrashkevich, and M. Shapiro. HSTERM — a program to calculate potential curves and radial matrix elements for two-electron systems within the hyperspherical adiabatic approach. Computer Physics Communications, 90(2–3):311–339, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500084S.

Ayala:1990:GAH

[AB90] Carles Ayala and Marià Baig. Generation and analysis of high order strong coupling series for SU(2) lattice gauge theory. Computer Physics Communications, 58(1–2):199–209, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090145Q.

Accomando:1997:WPW

[AB97] Elena Accomando and Alessandro Ballestrero. WPHACT 1.0 — a program for WW, Higgs and four-fermion physics at e^{+}e^{−} colliders. Computer Physics Communications, 99 (2–3):270–296, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001245.

Arbona:1999:DCB

[AB99] A. Arbona and C. Bona. Dealing with the center and boundary problems in 1D numerical relativity. Computer Physics Communications, 118(2–3):229–235, May 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001915.

Arbuzov:1996:HPC

[ABB^{+}96] A. Arbuzov, D. Bardin, J. Blümlein, L. Kalinovskaya, and T. Riemann. HECTOR 1.00: a program for the calculation of QED, QCD and electroweak corrections to ep and l^{±}N deep inelastic neutral and charged current scattering. Computer Physics Communications, 94(2–3):128–184, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000057.

Abdelmalek:1990:PSI

[Abd90] Nabih N. Abdelmalek. A program for the solution of illposed linear systems arising from the discretization of the Fredholm integral equation of the first kind. Computer Physics Communications, 58(3):285–292, April 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900648.

Aurenche:1994:DSI

[ABE^{+}94] P. Aurenche, F. W. Bopp, R. Engel, D. Pertermann, J. Ranft, and S. Roesler. DTUJET-93: Sampling inelastic proton-proton and antiproton-proton collisions according to the two-component dual parton model. Computer Physics Communications, 83(1):107–123, October 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559490037X.

Alexandrou:1999:PSO

[ABF^{+}99] C. Alexandrou, A. Borici, A. Feo, Ph. de Forcrand, A. Galli, F. Jegerlehner, and T. Takaishi. Phase structure of one-flavour QCD. Computer Physics Communications, 121–122:595, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700030.

Abgrall:1991:HCR

[Abg91] Rémi Abgrall. Hypersonic calculations by Riemann solvers techniques. Computer Physics Communications, 65(1–3):1–7, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190148E.

Anlauf:1994:WMC

[ABH^{+}94] H. Anlauf, J. Biebel, A. Himmler, P. Manakos, T. Mannel, W. Schönau, and H. D. Dahmen. WOPPER: a Monte Carlo event generator for W off-shell pair production including higher order electromagnetic radiative corrections. Computer Physics Communications, 79(3):487–502, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901899.

An:1999:LNV

[ABH99] Tai An and Sa Ben-Hao. LUCIAE 3.0: a new version of a computer program for the firecracker model and rescattering in relativistic heavy-ion collisions. Computer Physics Communications, 116(2–3):353–365, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001386.

Aguirregabiria:1999:ROR

[ABHR99] J. M. Aguirregabiria, Ll. Bel, A. Hernández, and M. Rivas. Regular order reductions of ordinary and delay-differential equations. Computer Physics Communications, 116(1): 95–106, January 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001593.

Ablamowicz:1998:SRC

[Abl98] Rafal Ablamowicz. Spinor representations of Clifford algebras: a symbolic approach. Computer Physics Communications, 115(2–3):510–535, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001180.

Amaya:1992:IPT

[AC92] A. Amaya and E. Chacón. Integral of a product of three 6-dimensional spherical harmonics. Computer Physics Communications, 71(1–2):159–172, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900819.

Angulo:1999:MEE

[ACA99] J. C. Angulo, J. C. Cuchí, and J. C. Antolín. Maximumentropy estimation of atomic one- and two-body densities. Computer Physics Communications, 121–122:597, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700054.

Amaral:1999:ECS

[ACG^{+}99] L. A. N. Amaral, P. Cizeau, P. Gopikrishnan, Y. Liu, M. Meyer, C.-K. Peng, and H. E. Stanley. Econophysics: can statistical physics contribute to the science of economics? Computer Physics Communications, 121–122:145–152, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559900301X.

Agarwal:1991:RAC

[AD91] Ramesh K. Agarwal and Jerry E. Desse. Recent advances in computational aerodynamics. Computer Physics Communications, 65(1–3):8–16, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190149F.

Alda:1999:SFT

[ADK^{+}99] W. Alda, W. Dzwinel, J. Kitowski, J. Mościński, J. Rybicki, and K. Boryczko. Simulation of flow through porous media using particle models. Computer Physics Communications, 121–122:594, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700029.

Anlauf:1992:KMC

[ADM^{+}92] Harald Anlauf, Hans D. Dahmen, Panagiotis Manakos, Thomas Mannel, and Thorsten Ohl. KRONOS — a Monte Carlo event generator for higher order electromagnetic radiative corrections to deep inelastic scattering at HERA. Computer Physics Communications, 70(1):97–119, May 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290095G.

Assous:1992:PTM

[ADS92] Franck Assous, Pierre Degond, and Jacques Segre. A particle-tracking method for 3D electromagnetic PIC codes on unstructured meshes. Computer Physics Communications, 72(2–3):105–114, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290142L.

Arnold:1999:DTB

[AE99] A. Arnold and M. Ehrhardt. Discrete transparent boundary conditions for general Schrödinger-type equations. Computer Physics Communications, 121–122:598, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700066.

Amini:1990:ETM

[AF90] M. Amini and D. Fincham. Evaluation of temperature in molecular dynamics simulation. Computer Physics Communications, 56(3):313–324, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090017U.

Avdeev:1997:TAA

[AFKT97] Leo V. Avdeev, J. Fleischer, M. Yu. Kalmykov, and M. N. Tentyukov. Towards automatic analytic evaluation of diagrams with masses. Computer Physics Communications, 107(1–3):155–166, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001124.

Alberto:1999:SPS

[AFP99] Pedro Vieira Alberto, Carlos Fiolhais, and João Paiva. The SoftSciences project — software for learning physics and other sciences. Computer Physics Communications, 121–122: 593, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700017.

Ambrosiano:1991:ODP

[AG91] John J. Ambrosiano and James L. Geary. A one-dimensional PIC-circuit code for simulating a reflex triode. Computer Physics Communications, 67(2):210–222, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190017F.

Alvarez:1999:ITG

[AG99] J. D. Alvarez and L. Garcia. Ion temperature gradient turbulence numerical calculations. Computer Physics Communications, 121–122:596, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700042.

Amaral:1999:MHR

[AGIS99] Luís A. Nunes Amaral, Ary L. Goldberger, Plamen Ch. Ivanov, and H. Eugene Stanley. Modeling heart rate variability by stochastic feedback. Computer Physics Communications, 121–122:126–128, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002957.

Allison:1991:EFT

[AH91] A. C. Allison and X. Huang. An exponentially fitted two-step method with predictor-corrector form for the solution of nonlinear or coupled Schrödinger equations. Computer Physics Communications, 66(2–3):211–218, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900702.

Aiani:1992:EOM

[AH92] Karen E. Aiani and John S. Hutchinson. Evaluation of overlap matrix elements by the transformation method (HEG). Computer Physics Communications, 69(1):46–52, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290126J.

Ayari:1997:GMP

[AH97] M. A. Ayari and V. Hussin. GLie: a Maple program for Lie supersymmetries of Grassmann-valued differential equations. Computer Physics Communications, 100(1–2):157–176, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001294.

Arnold:1999:PSM

[AH99] Wolfram T. Arnold and Roger Haydock. A parallel sparse matrix technique, “dynamic recursion method”, in condensed matter physics. Computer Physics Communications, 121–122:599, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700078.

Attenberger:1994:MCC

[AHH94] S. E. Attenberger, S. P. Hirshman, and W. A. Houlberg. MACH — a computer code for solution of the poloidal asymmetry eigenvalue problem in tokamaks. Computer Physics Communications, 79(2):341–350, April 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900779.

Allan:1990:PFS

[AHZ90] R. J. Allan, L. Heck, and S. Zurek. Parallel FORTRAN in scientific computing: a new occam harness called fortnet. Computer Physics Communications, 59(2):325–344, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090182Z.

Aldea:1990:FAE

[AI90a] Nicolae Aldea and Emil Indrea. Fourier analysis of EX-AFS and XANES data — a self-contained FORTRAN program-package — the third version. Computer Physics Communications, 60(1):145–154, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090083D.

Aldea:1990:XPE

[AI90b] Nicolae Aldea and Emil Indrea. XRLINE, a program to evaluate the crystallite size of supported metal catalysts by single X-ray profile Fourier analysis. Computer Physics Communications, 60(1):155–163, August 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090084E.

Akushevich:1997:PFC

[AIS^{+}97] I. Akushevich, A. Ilyichev, N. Shumeiko, A. Soroko, and A. Tolkachev. POLRAD 2.0. FORTRAN code for the radiative corrections calculation to deep inelastic scattering of polarized particles. Computer Physics Communications, 104(1–3):201–244, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000623.

Arjunwadkar:1991:SAB

[AK91] Mihir Arjunwadkar and D. G. Kanhere. A simulated annealing based algorithm for eigenvalue problems. Computer Physics Communications, 62(1):8–15, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901163.

Auer:1999:RCA

[AK99] J. Auer and E. Krotscheck. A rapidly converging algorithm for solving the Kohn–Sham and related equations in electronic structure theory. Computer Physics Communications, 118(2–3):139–144, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598002021.

Auerbach:1993:NCA

[AL93] Scott M. Auerbach and Claude Leforestier. A new computational algorithm for Green’s functions: Fourier transform of the Newton polynomial expansion. Computer Physics Communications, 78(1–2):55–66, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390142Y.

Altevogt:1994:ADL

[AL94] Peter Altevogt and Andreas Linke. An algorithm for dynamic load balancing of synchronous Monte Carlo simulations on multiprocessor systems. Computer Physics Communications, 79(3):373–380, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559490183X.

Albano:1999:MCS

[Alb99] Ezequiel V. Albano. Monte Carlo simulations of surface chemical reactions: Irreversible phase transitions and oscillatory behaviour. Computer Physics Communications, 121–122:388–391, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599003616.

Alexander:1993:LRP

[Ale93] Millard H. Alexander. A linear reference potential algorithm for the study of molecular photodissociation. Computer Physics Communications, 75(1–2):87–97, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390166A.

Alfe:1999:IMD

[Alf99] Dario Alfè. Ab initio molecular dynamics, a simple algorithm for charge extrapolation. Computer Physics Communications, 118(1):31–33, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001957.

Allison:1993:MRB

[All93] John Allison. Multiquadric radial basis functions for representing multidimensional high energy physics data. Computer Physics Communications, 77(3):377–395, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390184E.

Allen:1999:STL

[All99] Michael P. Allen. Simulations and theories of liquid crystals. Computer Physics Communications, 121–122:219–224, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003173.

Arneodo:1997:DMC

[ALR97] M. Arneodo, L. Lamberti, and M. Ryskin. DIPSI: a Monte Carlo generator for elastic vector meson production in charged lepton-proton scattering. Computer Physics Communications, 100(1–2):195–214, February 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001026.

Adam:1999:KMC

[ALR99] E. Adam, F. Lançon, and B. Rodmacq. Kinetic Monte Carlo simulation of the thermal evolution of multilayers. Computer Physics Communications, 121–122:545–546, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004038.

Aly:1990:QSE

[Aly90] J. J. Aly. Quasi-static evolution of a force-free magnetic field. Computer Physics Communications, 59(1):13–20, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090152Q.

Arnold:1990:EMB

[AM90] Anton Arnold and Norbert Mauser. An efficient method of bookkeeping next neighbours in molecular dynamics simulations. Computer Physics Communications, 59(2):267–275, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090175Z.

Armand:1994:TMM

[AM94] G. Armand and J. R. Manson. Transition matrix methods in atom-surface scattering theory. Computer Physics Communications, 80(1–3):53–89, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900981.

Anlauf:1994:UVM

[AMM^{+}94] H. Anlauf, P. Manakos, T. Mannel, T. Ohl, H. Meinhard, and H.-D. Dahmen. UNIBAB, version 2.0: Monte Carlo event generation for large angle Bhabha scattering at LEP. Computer Physics Communications, 79(3):466–486, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901880.

Atanasiu:1996:IMM

[AMS96] C. V. Atanasiu, A. Moraru, and A. A. Subbotin. Improvement of the moments method for MHD equilibrium calculation in a divertor tokamak configuration. Computer Physics Communications, 98(1–2):137–152, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559600094X.

Allison:1997:CTD

[AN97] D. C. S. Allison and M. T. Noga. Computing the threedimensional convex hull. Computer Physics Communications, 103(1):74–82, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597000271.

Anonymous:1990:AI

[Ano90a] Anonymous. Author index. Computer Physics Communications, 59(1):197–198, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465590901692.

Anonymous:1990:AIVa

[Ano90b] Anonymous. Author index to volume 56. Computer Physics Communications, 56(3):413–415, January 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090025V.

Anonymous:1990:AIVb

[Ano90c] Anonymous. Author index to volume 58. Computer Physics Communications, 58(3):349–350, April 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900718.

Anonymous:1990:AIVc

[Ano90d] Anonymous. Author index to volume 59. Computer Physics Communications, 59(3):548–551, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090096J.

Anonymous:1990:AIVd

[Ano90e] Anonymous. Author index to volume 60. Computer Physics Communications, 60(3):422–424, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900408.

Anonymous:1990:AIVe

[Ano90f] Anonymous. Author index to volume 61. Computer Physics Communications, 61(3):444–446, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090059A.

Anonymous:1990:CPC

[Ano90g] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 59(1):v–vii, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090148T.

Anonymous:1990:C

[Ano90h] Anonymous. Contents. Computer Physics Communications, 59(1):xiii–xiv, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090150Y.

Anonymous:1990:CVa

[Ano90i] Anonymous. Contents to volume 56. Computer Physics Communications, 56(3):411–412, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090024U.

Anonymous:1990:CVb

[Ano90j] Anonymous. Contents to volume 58. Computer Physics Communications, 58(3):347–348, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090070H.

Anonymous:1990:CVc

[Ano90k] Anonymous. Contents to volume 59. Computer Physics Communications, 59(3):545–547, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090095I.

Anonymous:1990:CVd

[Ano90l] Anonymous. Contents to volume 60. Computer Physics Communications, 60(3):419–421, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900394.

Anonymous:1990:CVe

[Ano90m] Anonymous. Contents to volume 61. Computer Physics Communications, 61(3):441–443, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900589.

Anonymous:1990:EBa

[Ano90n] Anonymous. Editorial Board. Computer Physics Communications, 58(1–2):v–vii, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090130S.

Anonymous:1990:EBb

[Ano90o] Anonymous. Editorial Board. Computer Physics Communications, 60(1):v–vii, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090073A.

Anonymous:1990:EBc

[Ano90p] Anonymous. Editorial Board. Computer Physics Communications, 61(1–2):v–vii, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090098L.

Anonymous:1990:EN

[Ano90q] Anonymous. Erratum notice. Computer Physics Communications, 58(3):345, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090069D.

Anonymous:1990:LC

[Ano90r] Anonymous. List of contributors. Computer Physics Communications, 61(1–2):275, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090128N.

Anonymous:1990:PIVa

[Ano90s] Anonymous. Program index to volume 56. Computer Physics Communications, 56(3):416–417, January 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090026W.

Anonymous:1990:PIVb

[Ano90t] Anonymous. Program index to volume 58. Computer Physics Communications, 58(3):351–352, April 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900729.

Anonymous:1990:PIVc

[Ano90u] Anonymous. Program index to volume 59. Computer Physics Communications, 59(3):552–553, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090097K.

Anonymous:1990:PIVd

[Ano90v] Anonymous. Program index to volume 60. Computer Physics Communications, 60(3):425–426, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090041X.

Anonymous:1990:PIVe

[Ano90w] Anonymous. Program index to volume 61. Computer Physics Communications, 61(3):447, December 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090060E.

Anonymous:1990:S

[Ano90x] Anonymous. Sponsors. Computer Physics Communications, 61(1–2):xii, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090100F.

Anonymous:1990:SI

[Ano90y] Anonymous. Subject index. Computer Physics Communications, 61(1–2):277–278, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090129O.

Anonymous:1991:AIa

[Ano91a] Anonymous. Author index. Computer Physics Communications, 65(1–3):333–335, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190187P.

Anonymous:1991:AIb

[Ano91b] Anonymous. Author index. Computer Physics Communications, 67(1):157, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190228D.

Anonymous:1991:AIVa

[Ano91c] Anonymous. Author index to volume 62. Computer Physics Communications, 62(2–3):386–388, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190111W.

Anonymous:1991:AIVb

[Ano91d] Anonymous. Author index to volume 63. Computer Physics Communications, 63(1–3):579–582, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190277R.

Anonymous:1991:AIVc

[Ano91e] Anonymous. Author index to volume 64. Computer Physics Communications, 64(3):524–526, June 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901439.

Anonymous:1991:AIVd

[Ano91f] Anonymous. Author index to volume 66. Computer Physics Communications, 66(2–3):410–412, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900883.

Anonymous:1991:AIVe

[Ano91g] Anonymous. Author index to volume 68. Computer Physics Communications, 68(1–3):497–498, November 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591902157.

Anonymous:1991:CPC

[Ano91h] Anonymous. Computer physics communications instructions to authors (sixth revision). Computer Physics Communications, 62(1):ix–xviii, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190114Z.

Anonymous:1991:Ca

[Ano91i] Anonymous. Contents. Computer Physics Communications, 62(2–3):vii–viii, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190091X.

Anonymous:1991:Cb

[Ano91j] Anonymous. Contents. Computer Physics Communications, 63(1–3):xiii–xv, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190232A.

Anonymous:1991:Cc

[Ano91k] Anonymous. Contents. Computer Physics Communications, 65(1–3):xiii–xv, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190147D.

Anonymous:1991:Cd

[Ano91l] Anonymous. Contents. Computer Physics Communications, 67(1):xiii, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190218A.

Anonymous:1991:CVa

[Ano91m] Anonymous. Contents to volume 62. Computer Physics Communications, 62(2–3):383–385, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591901107.

Anonymous:1991:CVb

[Ano91n] Anonymous. Contents to volume 64. Computer Physics Communications, 64(3):521–523, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591901428.

Anonymous:1991:CVc

[Ano91o] Anonymous. Contents to volume 66. Computer Physics Communications, 66(2–3):407–409, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900872.

Anonymous:1991:EBa

[Ano91p] Anonymous. Editorial Board. Computer Physics Communications, 62(1):v–vii, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190113Y.

Anonymous:1991:EBb

[Ano91q] Anonymous. Editorial Board. Computer Physics Communications, 63(1–3):v–vii, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190229E.

Anonymous:1991:EBc

[Ano91r] Anonymous. Editorial Board. Computer Physics Communications, 64(1):v–vii, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190043K.

Anonymous:1991:EBd

[Ano91s] Anonymous. Editorial Board. Computer Physics Communications, 65(1–3):v–vii, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190145B.

Anonymous:1991:EBe

[Ano91t] Anonymous. Editorial Board. Computer Physics Communications, 66(1):v–vii, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591900012.

Anonymous:1991:EBf

[Ano91u] Anonymous. Editorial Board. Computer Physics Communications, 67(1):v–vii, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591902168.

Anonymous:1991:EBg

[Ano91v] Anonymous. Editorial Board. Computer Physics Communications, 68(1–3):v–vii, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190189R.

Anonymous:1991:ENa

[Ano91w] Anonymous. Erratum notice. Computer Physics Communications, 62(1):162–165, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591901288.

Anonymous:1991:ENb

[Ano91x] Anonymous. Erratum notice. Computer Physics Communications, 64(2):343, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190042J.

Anonymous:1991:ENc

[Ano91y] Anonymous. Erratum notice. Computer Physics Communications, 67(2):356, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190028J.

Anonymous:1991:Fb

[Ano91z] Anonymous. Foreword. Computer Physics Communications, 67(1):xi–xii, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591902179.

Anonymous:1991:Fa

[Ano91-27] Anonymous. Forward. Computer Physics Communications, 63(1–3):xi, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190230I.

Anonymous:1991:P

[Ano91-28] Anonymous. Preface. Computer Physics Communications, 63 (1–3):xii, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591902319.

Anonymous:1991:PIVa

[Ano91-29] Anonymous. Program index to volume 62. Computer Physics Communications, 62(2–3):389–390, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190112X.

Anonymous:1991:PIVb

[Ano91-30] Anonymous. Program index to volume 63. Computer Physics Communications, 63(1–3):584, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190279T.

Anonymous:1991:PIVc

[Ano91-31] Anonymous. Program index to volume 64. Computer Physics Communications, 64(3):527–529, June 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190144A.

Anonymous:1991:PIVd

[Ano91-32] Anonymous. Program index to volume 66. Computer Physics Communications, 66(2–3):413–414, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900894.

Anonymous:1991:SI

[Ano91-33] Anonymous. Subject index. Computer Physics Communications, 65(1–3):336–341, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190188Q.

Anonymous:1991:SIV

[Ano91-34] Anonymous. Subject index to volume 63. Computer Physics Communications, 63(1–3):583, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190278S.

Anonymous:1991:SO

[Ano91-35] Anonymous. Supporting organization. Computer Physics Communications, 67(1):xiv, August 1991. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046559190219B.

Anonymous:1992:AI

[Ano92a] Anonymous. Author index. Computer Physics Communications, 73(1–3):221–222, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290044Y.

Anonymous:1992:AIVa

[Ano92b] Anonymous. Author index to volume 67. Computer Physics Communications, 67(3):552–554, January 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900613.

Anonymous:1992:AIVb

[Ano92c] Anonymous. Author index to volume 69. Computer Physics Communications, 69(2–3):492–494, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901885.

Anonymous:1992:AIVc

[Ano92d] Anonymous. Author index to volume 70. Computer Physics Communications, 70(3):613–616, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290120N.

Anonymous:1992:AIVd

[Ano92e] Anonymous. Author index to volume 71. Computer Physics Communications, 71(3):348–350, September 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290021P.

Anonymous:1992:AIVe

[Ano92f] Anonymous. Author index to volume 72. Computer Physics Communications, 72(2–3):309–310, November 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290159V.

Anonymous:1992:CVa

[Ano92g] Anonymous. Contents to volume 67. Computer Physics Communications, 67(3):549–551, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290060C.

Anonymous:1992:CVb

[Ano92h] Anonymous. Contents to volume 69. Computer Physics Communications, 69(2–3):489–491, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901874.

Anonymous:1992:CVc

[Ano92i] Anonymous. Contents to volume 70. Computer Physics Communications, 70(3):609–612, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290119J.

Anonymous:1992:CVd

[Ano92j] Anonymous. Contents to volume 71. Computer Physics Communications, 71(3):346–347, September 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290020Y.

Anonymous:1992:CVe

[Ano92k] Anonymous. Contents to volume 72. Computer Physics Communications, 72(2–3):307–308, November 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290158U.

Anonymous:1992:EBa

[Ano92l] Anonymous. Editorial Board. Computer Physics Communications, 69(1):v–vii, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290122F.

Anonymous:1992:EBb

[Ano92m] Anonymous. Editorial Board. Computer Physics Communications, 70(1):v–vii, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290085D.

Anonymous:1992:EBc

[Ano92n] Anonymous. Editorial Board. Computer Physics Communications, 71(1–2):v–vii, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900668.

Anonymous:1992:EBd

[Ano92o] Anonymous. Editorial Board. Computer Physics Communications, 72(1):v–vii, October 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290001F.

Anonymous:1992:EBe

[Ano92p] Anonymous. Editorial Board. Computer Physics Communications, 73(1–3):??, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290023R.

Anonymous:1992:ENa

[Ano92q] Anonymous. Erratum notice. Computer Physics Communications, 69(2–3):486, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592901852.

Anonymous:1992:ENb

[Ano92r] Anonymous. Erratum notice. Computer Physics Communications, 71(1–2):206, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290084C.

Anonymous:1992:ENc

[Ano92s] Anonymous. Erratum notice. Computer Physics Communications, 71(3):343, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290018T.

Anonymous:1992:ENd

[Ano92t] Anonymous. Erratum notice. Computer Physics Communications, 72(2–3):304–306, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290157T.

Anonymous:1992:PIVa

[Ano92u] Anonymous. Program index to volume 67. Computer Physics Communications, 67(3):555–556, January 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900624.

Anonymous:1992:PIVb

[Ano92v] Anonymous. Program index to volume 69. Computer Physics Communications, 69(2–3):495–497, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901896.

Anonymous:1992:PIVc

[Ano92w] Anonymous. Program index to volume 70. Computer Physics Communications, 70(3):617–619, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290121E.

Anonymous:1992:PIVd

[Ano92x] Anonymous. Program index to volume 71. Computer Physics Communications, 71(3):351–352, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290022Q.

Anonymous:1992:PIVe

[Ano92y] Anonymous. Program index to volume 72. Computer Physics Communications, 72(2–3):312–313, November 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290160Z.

Anonymous:1993:AIVa

[Ano93a] Anonymous. Author index to volume 74. Computer Physics Communications, 74(3):455–458, March 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390028B.

Anonymous:1993:AIVb

[Ano93b] Anonymous. Author index to volume 75. Computer Physics Communications, 75(3):419–421, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390053F.

Anonymous:1993:AIVc

[Ano93c] Anonymous. Author index to volume 76. Computer Physics Communications, 76(3):413–415, August 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390065K.

Anonymous:1993:AIVd

[Ano93d] Anonymous. Author index to volume 77. Computer Physics Communications, 77(3):453–455, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390190N.

Anonymous:1993:CVa

[Ano93e] Anonymous. Contents to volume 74. Computer Physics Communications, 74(3):452–454, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390027A.

Anonymous:1993:CVb

[Ano93f] Anonymous. Contents to volume 75. Computer Physics Communications, 75(3):417–418, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390052E.

Anonymous:1993:CVc

[Ano93g] Anonymous. Contents to volume 76. Computer Physics Communications, 76(3):411–412, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390064J.

Anonymous:1993:CVd

[Ano93h] Anonymous. Contents to volume 77. Computer Physics Communications, 77(3):450–452, November 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390189J.

Anonymous:1993:EBa

[Ano93i] Anonymous. Editorial Board. Computer Physics Communications, 74(1):v–vii, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390100Q.

Anonymous:1993:EBb

[Ano93j] Anonymous. Editorial Board. Computer Physics Communications, 75(1–2):v–vii, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390159A.

Anonymous:1993:EBc

[Ano93k] Anonymous. Editorial Board. Computer Physics Communications, 76(1):v–vii, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390115S.

Anonymous:1993:EBd

[Ano93l] Anonymous. Editorial Board. Computer Physics Communications, 77(1):v–vii, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390030G.

Anonymous:1993:EBe

[Ano93m] Anonymous. Editorial board. Computer Physics Communications, 78(1–2):v–vii, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901372.

Anonymous:1993:EN

[Ano93n] Anonymous. Erratum notice. Computer Physics Communications, 77(2):299, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593900122.

Anonymous:1993:PIVa

[Ano93o] Anonymous. Program index to volume 74. Computer Physics Communications, 74(3):459–460, March 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390029C.

Anonymous:1993:PIVb

[Ano93p] Anonymous. Program index to volume 75. Computer Physics Communications, 75(3):422–423, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390054G.

Anonymous:1993:PIVc

[Ano93q] Anonymous. Program index to volume 76. Computer Physics Communications, 76(3):416–417, August 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390066L.

Anonymous:1993:PIVd

[Ano93r] Anonymous. Program index to volume 77. Computer Physics Communications, 77(3):456–457, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390191E.

Anonymous:1994:APLa

[Ano94a] Anonymous. Announcement — processing of L^{A}TEX manuscripts. Computer Physics Communications, 79(1):ix, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902240.

Anonymous:1994:APLb

[Ano94b] Anonymous. Announcement — processing of L^{A}TEX manuscripts. Computer Physics Communications, 79(3):??, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901805.

Anonymous:1994:APL

[Ano94c] Anonymous. Announcement-processing of L^{A}TEX manuscripts. Computer Physics Communications, 79(2):??, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900639.

Anonymous:1994:AIVa

[Ano94d] Anonymous. Author index to volume 78. Computer Physics Communications, 78(3):303–304, January 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900108.

Anonymous:1994:AIVb

[Ano94e] Anonymous. Author index to volume 80. Computer Physics Communications, 80(1–3):275, March 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901066.

Anonymous:1994:AIVc

[Ano94f] Anonymous. Author index to volume 81. Computer Physics Communications, 81(3):427–429, July 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900906.

Anonymous:1994:AIVd

[Ano94g] Anonymous. Author index to volume 82. Computer Physics Communications, 82(2–3):311–313, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901783.

Anonymous:1994:AIVe

[Ano94h] Anonymous. Author index to volume 83. Computer Physics Communications, 83(2–3):337–338, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900612.

Anonymous:1994:AIVf

[Ano94i] Anonymous. Author index to volume 84. Computer Physics Communications, 84(1–3):348–349, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902216.

Anonymous:1994:CPC

[Ano94j] Anonymous. Computer physics communications instructions to authors (1994 revision). Computer Physics Communications, 83(1):ix–xix, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900299.

Anonymous:1994:C

[Ano94k] Anonymous. Contents. Computer Physics Communications, 80(1–3):xiii, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900949.

Anonymous:1994:CVaa

[Ano94l] Anonymous. Contents to volume 78. Computer Physics Communications, 78(3):301–302, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900094.

Anonymous:1994:CVA

[Ano94m] Anonymous. Contents to volume 79, author index to volume 79, program index to volume 79. Computer Physics Communications, 79(3):561–570, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901953.

Anonymous:1994:CVb

[Ano94n] Anonymous. Contents to volume 81. Computer Physics Communications, 81(3):425–426, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900892.

Anonymous:1994:CVc

[Ano94o] Anonymous. Contents to volume 82. Computer Physics Communications, 82(2–3):309–310, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901775.

Anonymous:1994:CVd

[Ano94p] Anonymous. Contents to volume 83. Computer Physics Communications, 83(2–3):335–336, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900604.

Anonymous:1994:CVe

[Ano94q] Anonymous. Contents to volume 84. Computer Physics Communications, 84(1–3):346–347, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902208.

Anonymous:1994:CB

[Ano94r] Anonymous. CPC — the beginning. Computer Physics Communications, 84(1–3):x, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901988.

Anonymous:1994:EBa

[Ano94s] Anonymous. Editorial Board. Computer Physics Communications, 79(1):??, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594902232.

Anonymous:1994:EBb

[Ano94t] Anonymous. Editorial Board. Computer Physics Communications, 80(1–3):v–vii, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900922.

Anonymous:1994:EBc

[Ano94u] Anonymous. Editorial Board. Computer Physics Communications, 81(1–2):iii–vii, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901074.

Anonymous:1994:EBd

[Ano94v] Anonymous. Editorial Board. Computer Physics Communications, 82(1):v–vii, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901252.

Anonymous:1994:EBe

[Ano94w] Anonymous. Editorial Board. Computer Physics Communications, 83(1):v–vii, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900280.

Anonymous:1994:EBf

[Ano94x] Anonymous. Editorial Board. Computer Physics Communications, 84(1–3):??, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901961.

Anonymous:1994:EN

[Ano94y] Anonymous. Erratum notice. Computer Physics Communications, 79(2):351, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900787.

Anonymous:1994:P

[Ano94z] Anonymous. Preface. Computer Physics Communications, 80 (1–3):xi–xii, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900930.

Anonymous:1994:PIVa

[Ano94-27] Anonymous. Program index to volume 78. Computer Physics Communications, 78(3):305, January 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900116.

Anonymous:1994:PIVb

[Ano94-28] Anonymous. Program index to volume 81. Computer Physics Communications, 81(3):430, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900914.

Anonymous:1994:PIVc

[Ano94-29] Anonymous. Program index to volume 82. Computer Physics Communications, 82(2–3):314–315, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901791.

Anonymous:1994:PIVd

[Ano94-30] Anonymous. Program index to volume 83. Computer Physics Communications, 83(2–3):339, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900620.

Anonymous:1994:PIVe

[Ano94-31] Anonymous. Program index to volume 84. Computer Physics Communications, 84(1–3):350, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902224.

Anonymous:1994:PN

[Ano94-32] Anonymous. Publisher’s note. Computer Physics Communications, 84(1–3):ix, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559490197X.

Anonymous:1994:TPG

[Ano94-33] Anonymous. A tribute to Philip George Burke. Computer Physics Communications, 84(1–3):xi–xiii, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901996.

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index to volume 85. Computer Physics Communications, 85(3):504–506, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900772.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index to volume 86. Computer Physics Communications, 86(3):315–317, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900837.

Anonymous:1995:AIVc

[Ano95c] Anonymous. Author index to volume 87. Computer Physics Communications, 87(3):440–442, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900179.

Anonymous:1995:AIVd

[Ano95d] Anonymous. Author index to volume 88. Computer Physics Communications, 88(2–3):350–352, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900675.

Anonymous:1995:AIVe

[Ano95e] Anonymous. Author index to volume 89. Computer Physics Communications, 89(1–3):169, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900632.

Anonymous:1995:AIVf

[Ano95f] Anonymous. Author index to volume 90. Computer Physics Communications, 90(2–3):390–392, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559590039X.

Anonymous:1995:AIVg

[Ano95g] Anonymous. Author index to volume 91. Computer Physics Communications, 91(1–3):345–347, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900136.

Anonymous:1995:AIVh

[Ano95h] Anonymous. Author index to volume 92. Computer Physics Communications, 92(2–3):423–424, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900535.

Anonymous:1995:CPCa

[Ano95i] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 91(1–3):v– vii, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900071.

Anonymous:1995:CPCb

[Ano95j] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 92(1):v–vii, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559590073X.

Anonymous:1995:CVa

[Ano95k] Anonymous. Contents to volume 85. Computer Physics Communications, 85(3):501–503, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900756.

Anonymous:1995:CVb

[Ano95l] Anonymous. Contents to volume 86. Computer Physics Communications, 86(3):313–314, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900829.

Anonymous:1995:CVc

[Ano95m] Anonymous. Contents to volume 87. Computer Physics Communications, 87(3):438–439, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900152.

Anonymous:1995:CVd

[Ano95n] Anonymous. Contents to volume 88. Computer Physics Communications, 88(2–3):348–349, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900659.

Anonymous:1995:CVe

[Ano95o] Anonymous. Contents to volume 90. Computer Physics Communications, 90(2–3):388–389, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900373.

Anonymous:1995:CVf

[Ano95p] Anonymous. Contents to volume 92. Computer Physics Communications, 92(2–3):421–422, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900519.

Anonymous:1995:EBa

[Ano95q] Anonymous. Editorial Board. Computer Physics Communications, 85(1):??, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900357.

Anonymous:1995:EBb

[Ano95r] Anonymous. Editorial Board. Computer Physics Communications, 86(1–2):??, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900810.

Anonymous:1995:EBc

[Ano95s] Anonymous. Editorial Board. Computer Physics Communications, 87(1–2):??, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900012.

Anonymous:1995:EBd

[Ano95t] Anonymous. Editorial Board. Computer Physics Communications, 88(1):??, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900217.

Anonymous:1995:EBe

[Ano95u] Anonymous. Editorial Board. Computer Physics Communications, 89(1–3):??, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900594.

Anonymous:1995:EBf

[Ano95v] Anonymous. Editorial Board. Computer Physics Communications, 90(1):??, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900438.

Anonymous:1995:EBg

[Ano95w] Anonymous. Editorial Board. Computer Physics Communications, 91(1–3):??, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900055.

Anonymous:1995:EBh

[Ano95x] Anonymous. Editorial Board. Computer Physics Communications, 92(1):??, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595900713.

Anonymous:1995:PIVb

[Ano95y] Anonymous. Program index to volume 8. Computer Physics Communications, 86(3):318–319, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900845.

Anonymous:1995:PIVa

[Ano95z] Anonymous. Program index to volume 85. Computer Physics Communications, 85(3):507–508, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900799.

Anonymous:1995:PIVc

[Ano95-27] Anonymous. Program index to volume 87. Computer Physics Communications, 87(3):443, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900195.

Anonymous:1995:PIVd

[Ano95-28] Anonymous. Program index to volume 88. Computer Physics Communications, 88(2–3):353–354, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900691.

Anonymous:1995:PIVe

[Ano95-29] Anonymous. Program index to volume 90. Computer Physics Communications, 90(2–3):393–394, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900411.

Anonymous:1995:PIVf

[Ano95-30] Anonymous. Program index to volume 92. Computer Physics Communications, 92(2–3):425, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900551.

Anonymous:1996:AI

[Ano96a] Anonymous. Author index. Computer Physics Communications, 97(1–2):205–206, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596900958.

Anonymous:1996:AIVa

[Ano96b] Anonymous. Author index to volume 93. Computer Physics Communications, 93(2–3):324–326, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900193.

Anonymous:1996:AIVb

[Ano96c] Anonymous. Author index to volume 94. Computer Physics Communications, 94(2–3):276–277, April 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900272.

Anonymous:1996:AIVc

[Ano96d] Anonymous. Author index to volume 95. Computer Physics Communications, 95(2–3):223–224, June 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900132.

Anonymous:1996:AIVd

[Ano96e] Anonymous. Author index to volume 96. Computer Physics Communications, 96(2–3):333–335, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900399.

Anonymous:1996:AIVe

[Ano96f] Anonymous. Author index to volume 97. Computer Physics Communications, 97(3):361–363, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900454.

Anonymous:1996:AIVf

[Ano96g] Anonymous. Author index to volume 98. Computer Physics Communications, 98(3):417–419, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900338.

Anonymous:1996:CPCb

[Ano96h] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 95(1):v–vii, May 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596900077.

Anonymous:1996:CPCc

[Ano96i] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 96(1):v–vii, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596901071.

Anonymous:1996:CPCd

[Ano96j] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 98(1–2):v–vii, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559690003X.

Anonymous:1996:CPCa

[Ano96k] Anonymous. Computer Physics Communications: An international journal and program library for computational physics and physical chemistry. Computer Physics Communications, 95(1):??, May 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596900053.

Anonymous:1996:CVa

[Ano96l] Anonymous. Contents to volume 93. Computer Physics Communications, 93(2–3):322–323, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559690017X.

Anonymous:1996:CVb

[Ano96m] Anonymous. Contents to volume 94. Computer Physics Communications, 94(2–3):274–275, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596900259.

Anonymous:1996:CVc

[Ano96n] Anonymous. Contents to volume 95. Computer Physics Communications, 95(2–3):221–222, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596900119.

Anonymous:1996:CVd

[Ano96o] Anonymous. Contents to volume 96. Computer Physics Communications, 96(2–3):331–332, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900375.

Anonymous:1996:CVe

[Ano96p] Anonymous. Contents to volume 97. Computer Physics Communications, 97(3):359–360, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900430.

Anonymous:1996:CVf

[Ano96q] Anonymous. Contents to volume 98. Computer Physics Communications, 98(3):415–416, November 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900314.

Anonymous:1996:CIP

[Ano96r] Anonymous. CPC international program library. Computer Physics Communications, 93(1):ix–xvi, January 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596901034.

Anonymous:1996:EBa

[Ano96s] Anonymous. Editorial Board. Computer Physics Communications, 93(1):??, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596901010.

Anonymous:1996:EBb

[Ano96t] Anonymous. Editorial Board. Computer Physics Communications, 94(1):v–vii, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596901101.

Anonymous:1996:EBc

[Ano96u] Anonymous. Editorial Board. Computer Physics Communications, 96(1):??, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596901058.

Anonymous:1996:EBd

[Ano96v] Anonymous. Editorial Board. Computer Physics Communications, 97(1–2):v–vii, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596900892.

Anonymous:1996:EBe

[Ano96w] Anonymous. Editorial Board. Computer Physics Communications, 98(1–2):??, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596900016.

Anonymous:1996:EBf

[Ano96x] Anonymous. Editorial Board. Computer Physics Communications, 99(1):??, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596900090.

Anonymous:1996:PIVa

[Ano96y] Anonymous. Program index to volume 93. Computer Physics Communications, 93(2–3):327–328, February 1996.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900211.

Anonymous:1996:PIVb

[Ano96z] Anonymous. Program index, to volume 94. Computer Physics Communications, 94(2–3):278, April 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900296.

Anonymous:1996:PIVc

[Ano96-27] Anonymous. Program index to volume 95. Computer Physics Communications, 95(2–3):225, June 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900156.

Anonymous:1996:PIVd

[Ano96-28] Anonymous. Program index to volume 96. Computer Physics Communications, 96(2–3):336–337, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900417.

Anonymous:1996:PIVe

[Ano96-29] Anonymous. Program index to volume 97. Computer Physics Communications, 97(3):364, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900478.

Anonymous:1996:PIVf

[Ano96-30] Anonymous. Program index to volume 98. Computer Physics Communications, 98(3):421, November 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900351.

Anonymous:1996:PNC

[Ano96-31] Anonymous. Publisher’s note: classification and keywording of articles. Computer Physics Communications, 94(1):ix, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596901113.

Anonymous:1997:AI

[Ano97a] Anonymous. Author index. Computer Physics Communications, 104(1–3):275–276, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597811009.

Anonymous:1997:AIVb

[Ano97b] Anonymous. Author index to volume 100. Computer Physics Communications, 100(3):324–325, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597814087.

Anonymous:1997:AIVc

[Ano97c] Anonymous. Author index to volume 101. Computer Physics Communications, 101(3):291–292, May 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597834613.

Anonymous:1997:AIVd

[Ano97d] Anonymous. Author index to volume 102. Computer Physics Communications, 102(1–3):269–270, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597825337.

Anonymous:1997:AIVe

[Ano97e] Anonymous. Author index to volume 103. Computer Physics Communications, 103(2–3):305–306, July 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597841021.

Anonymous:1997:AIVf

[Ano97f] Anonymous. Author index to volume 105. Computer Physics Communications, 105(2–3):287–288, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597806030.

Anonymous:1997:AIVg

[Ano97g] Anonymous. Author index to volume 106. Computer Physics Communications, 106(3):269–270, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559785391X.

Anonymous:1997:AIVh

[Ano97h] Anonymous. Author index to volume 107. Computer Physics Communications, 107(1–3):293–294, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597853829.

Anonymous:1997:AIVa

[Ano97i] Anonymous. Author index to volume 99. Computer Physics Communications, 99(2–3):373–374, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597810351.

Anonymous:1997:CPCa

[Ano97j] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 103(1):v–vii, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597900096.

Anonymous:1997:CPCb

[Ano97k] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 107(1–3):v–vii, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597900229.

Anonymous:1997:CVb

[Ano97l] Anonymous. Content of volume 100. Computer Physics Communications, 100(3):322–323, March 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597814075.

Anonymous:1997:CVc

[Ano97m] Anonymous. Contents to volume 101. Computer Physics Communications, 101(3):289–290, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597834601.

Anonymous:1997:CVd

[Ano97n] Anonymous. Contents to volume 103. Computer Physics Communications, 103(2–3):303–304, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559784101X.

Anonymous:1997:CVe

[Ano97o] Anonymous. Contents to volume 105. Computer Physics Communications, 105(2–3):285–286, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597806029.

Anonymous:1997:CVf

[Ano97p] Anonymous. Contents to volume 106. Computer Physics Communications, 106(3):267–268, November 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597853908.

Anonymous:1997:CVa

[Ano97q] Anonymous. Contents to volume 99. Computer Physics Communications, 99(2–3):371–372, January 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559781034X.

Anonymous:1997:EBa

[Ano97r] Anonymous. Editorial Board. Computer Physics Communications, 100(1–2):v–vii, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597900059.

Anonymous:1997:EBb

[Ano97s] Anonymous. Editorial Board. Computer Physics Communications, 101(1–2):v–vii, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597900060.

Anonymous:1997:EBc

[Ano97t] Anonymous. Editorial Board. Computer Physics Communications, 104(1–3):??, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559790028X.

Anonymous:1997:EBd

[Ano97u] Anonymous. Editorial Board. Computer Physics Communications, 105(1):??, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597900011.

Anonymous:1997:EBe

[Ano97v] Anonymous. Editorial Board. Computer Physics Communications, 106(1–2):v–vii, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597900047.

Anonymous:1997:EBf

[Ano97w] Anonymous. Editorial Board. Computer Physics Communications, 107(1–3):??, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597900217.

Anonymous:1997:PIVb

[Ano97x] Anonymous. Program index to volume 100. Computer Physics Communications, 100(3):326–327, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597814099.

Anonymous:1997:PIVc

[Ano97y] Anonymous. Program index to volume 102. Computer Physics Communications, 102(1–3):271, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597825349.

Anonymous:1997:PIVd

[Ano97z] Anonymous. Program index to volume 104. Computer Physics Communications, 104(1–3):277, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597811010.

Anonymous:1997:PIVe

[Ano97-27] Anonymous. Program index to volume 105. Computer Physics Communications, 105(2–3):289–290, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597806042.

Anonymous:1997:PIVf

[Ano97-28] Anonymous. Program index to volume 106. Computer Physics Communications, 106(3):271, November 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597853921.

Anonymous:1997:PIVg

[Ano97-29] Anonymous. Program index to volume 107. Computer Physics Communications, 107(1–3):295, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597853830.

Anonymous:1997:PIVa

[Ano97-30] Anonymous. Program index to volume 99. Computer Physics Communications, 99(2–3):375–376, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597900035.

Anonymous:1998:ADE

[Ano98a] Anonymous. The ATLAS DAQ and event filter prototype “- 1” project. Computer Physics Communications, 110(1–3):95– 102, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001604.

Anonymous:1998:AI

[Ano98b] Anonymous. Auteur index. Computer Physics Communications, 110(1–3):259–263, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000472.

Anonymous:1998:AIVa

[Ano98c] Anonymous. Author index to volume 108. Computer Physics Communications, 108(2–3):306–307, February 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597838398.

Anonymous:1998:AIVb

[Ano98d] Anonymous. Author index to volume 109. Computer Physics Communications, 109(2–3):278–279, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597875595.

Anonymous:1998:AIVc

[Ano98e] Anonymous. Author index to volume 111. Computer Physics Communications, 111(1–3):277–279, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000538.

Anonymous:1998:AIVd

[Ano98f] Anonymous. Author index to volume 112. Computer Physics Communications, 112(2–3):275–276, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800085X.

Anonymous:1998:AIVe

[Ano98g] Anonymous. Author index to volume 113. Computer Physics Communications, 113(2–3):263–264, October 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598800027.

Anonymous:1998:AIVf

[Ano98h] Anonymous. Author index to volume 114. Computer Physics Communications, 114(1–3):401–402, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001829.

Anonymous:1998:AIVg

[Ano98i] Anonymous. Author index to volume 115. Computer Physics Communications, 115(2–3):566–568, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001908.

Anonymous:1998:CVa

[Ano98j] Anonymous. Contents to volume 108. Computer Physics Communications, 108(2–3):304–305, February 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597838386.

Anonymous:1998:CVb

[Ano98k] Anonymous. Contents to volume 109. Computer Physics Communications, 109(2–3):276–277, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597875583.

Anonymous:1998:CVc

[Ano98l] Anonymous. Contents to volume 112. Computer Physics Communications, 112(2–3):273–274, August 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598800052.

Anonymous:1998:CVd

[Ano98m] Anonymous. Contents to volume 113. Computer Physics Communications, 113(2–3):261–262, October 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598800015.

Anonymous:1998:CVe

[Ano98n] Anonymous. Contents to volume 115. Computer Physics Communications, 115(2–3):563–565, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599800010.

Anonymous:1998:EBa

[Ano98o] Anonymous. Editorial Board. Computer Physics Communications, 108(1):v–vii, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598900546.

Anonymous:1998:EBb

[Ano98p] Anonymous. Editorial Board. Computer Physics Communications, 109(1):??, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598900595.

Anonymous:1998:EBc

[Ano98q] Anonymous. Editorial Board. Computer Physics Communications, 112(1):??, July 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598900017.

Anonymous:1998:EBd

[Ano98r] Anonymous. Editorial Board. Computer Physics Communications, 113(1):??, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598900558.

Anonymous:1998:EBe

[Ano98s] Anonymous. Editorial Board. Computer Physics Communications, 115(1):??, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598900625.

Anonymous:1998:PIVa

[Ano98t] Anonymous. Program index to volume 108. Computer Physics Communications, 108(2–3):308, February 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597838404.

Anonymous:1998:PIVb

[Ano98u] Anonymous. Program index to volume 109. Computer Physics Communications, 109(2–3):280, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597875601.

Anonymous:1998:PIVc

[Ano98v] Anonymous. Program index to volume 111. Computer Physics Communications, 111(1–3):280, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800054X.

Anonymous:1998:PIVd

[Ano98w] Anonymous. Program index to volume 112. Computer Physics Communications, 112(2–3):277, August 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000861.

Anonymous:1998:PIVe

[Ano98x] Anonymous. Program index to volume 113. Computer Physics Communications, 113(2–3):265, October 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598800064.

Anonymous:1998:PIVf

[Ano98y] Anonymous. Program index to volume 114. Computer Physics Communications, 114(1–3):403–404, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001830.

Anonymous:1998:PIVg

[Ano98z] Anonymous. Program index to volume 115. Computer Physics Communications, 115(2–3):569–570, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800191X.

Anonymous:1998:TSA

[Ano98-27] Anonymous. The teraflop supercomputer APEmille: architecture, software and project status report. Computer Physics Communications, 110(1–3):216–219, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700180X.

Anonymous:1999:AI

[Ano99a] Anonymous. Author index. Computer Physics Communications, 117(1–2):189–190, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599800022.

Anonymous:1999:AIVa

[Ano99b] Anonymous. Author index to volume 116. Computer Physics Communications, 116(2–3):368–369, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001964.

Anonymous:1999:AIVb

[Ano99c] Anonymous. Author index to volume 117. Computer Physics Communications, 117(3):293–295, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900212X.

Anonymous:1999:AIVc

[Ano99d] Anonymous. Author index to volume 118. Computer Physics Communications, 118(2–3):301–302, May 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002404.

Anonymous:1999:AIVd

[Ano99e] Anonymous. Author index to volume 119. Computer Physics Communications, 119(2–3):314–315, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003343.

Anonymous:1999:AIVe

[Ano99f] Anonymous. Author index to volume 120. Computer Physics Communications, 120(2–3):275–276, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599800174.

Anonymous:1999:AIVf

[Ano99g] Anonymous. Author index to volumes 121–122. Computer Physics Communications, 121–122:757–778, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004828.

Anonymous:1999:C

[Ano99h] Anonymous. Contents. Computer Physics Communications, 121–122:xxi–xxxvi, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004816.

Anonymous:1999:CVa

[Ano99i] Anonymous. Contents to volume 116. Computer Physics Communications, 116(2–3):366–367, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002143.

Anonymous:1999:CVb

[Ano99j] Anonymous. Contents to volume 117. Computer Physics Communications, 117(3):291–292, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002118.

Anonymous:1999:CVc

[Ano99k] Anonymous. Contents to volume 118. Computer Physics Communications, 118(2–3):299–300, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002398.

Anonymous:1999:CVd

[Ano99l] Anonymous. Contents to volume 119. Computer Physics Communications, 119(2–3):312–313, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003331.

Anonymous:1999:CVe

[Ano99m] Anonymous. Contents to volume 120. Computer Physics Communications, 120(2–3):273–274, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599800162.

Anonymous:1999:I

[Ano99n] Anonymous. Index. Computer Physics Communications, 123(1–3):166, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004981.

Anonymous:1999:MSA

[Ano99o] Anonymous. Members of Scientific Advisory and Organizing Committees. Computer Physics Communications, 121–122:xx, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004142.

Anonymous:1999:P

[Ano99p] Anonymous. Preface. Computer Physics Communications, 121–122:xix, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002659.

Anonymous:1999:PIVa

[Ano99q] Anonymous. Program index to volume 116. Computer Physics Communications, 116(2–3):370, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001976.

Anonymous:1999:PIVb

[Ano99r] Anonymous. Program index to volume 117. Computer Physics Communications, 117(3):296, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002131.

Anonymous:1999:PIVc

[Ano99s] Anonymous. Program index to volume 118. Computer Physics Communications, 118(2–3):303, May 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599800034.

Anonymous:1999:PIVd

[Ano99t] Anonymous. Program index to volume 119. Computer Physics Communications, 119(2–3):316, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599800046.

Anonymous:1999:PIVe

[Ano99u] Anonymous. Program index to volume 120. Computer Physics Communications, 120(2–3):277, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004440.

Arcos:1997:KCD

[AO97] JoséM. Los Arcos and Félix Ortiz. kB: a code to determine the ionization quenching function Q (E ) as a function of the kB parameter. Computer Physics Communications, 103(1):83– 94, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559700012X.

Athanas:1998:EPH

[AO98] Michael Athanas and Michael Ogg. An evaluation of PCs for high energy physics under windows NT and Linux. Computer Physics Communications, 110(1–3):225–229, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001823.

Aoki:1997:IDO

[Aok97] Takayuki Aoki. Interpolated differential operator (IDO) scheme for solving partial differential equations. Computer Physics Communications, 102(1–3):132–146, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000209.

Airapetyan:1997:NIS

[AP97] R. G. Airapetyan and I. V. Puzynin. Newtonian iterative scheme with simultaneous iterations of inverse derivative. Computer Physics Communications, 102(1–3):97–108, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559700009X.

Aldridge:1998:CAS

[AP98] J. E. Aldridge and G. E. Prince. Computer algebra solution of the inverse problem in the calculus of variations. Computer Physics Communications, 115(2–3):489–509, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001283.

Aloy:1999:EIF

[APaI99] M. A. Aloy, J. A. Pons, and J. M. a. Ibáñez. An efficient implementation of flux formulae in multidimensional relativistic hydrodynamical codes. Computer Physics Communications, 120(2–3):115–121, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002362.

Angelini:1991:MCP

[APN^{+}91] L. Angelini, M. Pellicoro, L. Nitti, G. Preparata, and G. Valenti. A Monte Carlo program for generating hadronic final states. Computer Physics Communications, 67(2): 293–308, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190024F.

Appert:1999:UGV

[App99] C. Appert. Universality of the growth velocity distribution in 1 + 1 dimensional growth models. Computer Physics Communications, 121–122:363–365, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003549.

Aro:1995:PTD

[AR95] Colin J. Aro and Garry H. Rodrigue. Preconditioned time differencing for stiff ODEs in diurnal atmospheric kinetics. Computer Physics Communications, 92(1):27–53, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500089X.

Aro:1996:CSO

[Aro96] Colin J. Aro. CHEMSODE: a stiff ODE solver for the equations of chemical kinetics. Computer Physics Communications, 97(3):304–314, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000719.

Arter:1991:UUS

[Art91] Wayne Arter. The use of upwind schemes at high Reynolds number — a cautionary note. Computer Physics Communications, 66(2–3):207–210, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190069W.

Arter:1993:BRB

[Art93] W. Arter. Book review: The physics of fluid turbulence: W. D. McComb, Oxford University Press, 1991. xxiv + 572 pages, £26 (paperback). ISBN 0-19-856256-X. Computer Physics Communications, 78(1–2):218–219, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465593901578.

Arter:1994:PMM

[Art94] Wayne Arter. Particle-mesh modelling of drift-wave turbulence. Computer Physics Communications, 79(3):381–408, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901848.

Arter:1995:BRB

[Art95a] W. Arter. Book review: REDUCE for physicists: By N. MacDonald. Institute of Physics Publishing, 1994. xii + 167 pages. Price £25 (paperback). ISBN 0-7503-0277-1. Computer Physics Communications, 85(2):323, February 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900578.

Arter:1995:TDM

[Art95b] Wayne Arter. Three-dimensional modelling of drift-wave turbulence. Computer Physics Communications, 88(1):59–75, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500030J.

Arter:1996:BRB

[Art96] W. Arter. Book review: Probabilistic methods in applied physics: P. Kree and W. Wedig, eds. Lecture Notes in Physics 451, Springer-Verlag, Berlin, 1995. x + 393 pages. Hardcover price DM 116,00. ISBN 3-540-60214-3. Computer Physics Communications, 99(1):149–150, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001063.

Abbiendi:1991:NHF

[AS91] G. Abbiendi and L. Stanco. A new heavy flavour generator in e–p collisions. Computer Physics Communications, 66(1): 16–24, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591900034.

Abad:1992:CCW

[AS92] J. Abad and J. Sesma. Computation of Coulomb wave functions at low energies. Computer Physics Communications, 71(1–2):110–124, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290076B.

Anderson:1995:HOP

[AS95] David V. Anderson and Dan E. Shumaker. Hybrid Ordered Particle Simulation (HOPS) code for plasma modelling on vector-serial, vector-parallel, and massively parallel computers. Computer Physics Communications, 87(1–2):16–34, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594001752.

Auhl:1999:CSC

[AS99] R. Auhl and E. Straube. Computer-simulation of copolymers on polymer-polymer interfaces. Computer Physics Communications, 121–122:600, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670008X.

Asai:1998:OLD

[Asa98] Makoto Asai. ODBMS for LHC detector simulation. Computer Physics Communications, 110(1–3):125–130, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001653.

Agarwal:1998:RC

[ASJ98] D. A. Agarwal, S. R. Sachs, and W. E. Johnston. The reality of collaboratories. Computer Physics Communications, 110(1–3):134–141, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597001677.

Askar:1994:CSC

[Ask94] Attila Askar. The continuum solid and compliance functions in gas-surface low energy collisions. Computer Physics Communications, 80(1–3):168–199, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901023.

Aguado:1998:GFI

[ATP98] Alfredo Aguado, César Tablero, and Miguel Paniagua. Global fit of ab initio potential energy surfaces I. triatomic systems. Computer Physics Communications, 108(2–3):259–266, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001355.

Armand-Ugon:1992:SML

[AUG92] Daniel Armand-Ugon and Rodolfo Gambini. Symbolic manipulation of loops and lattice gauge theory in the loop representation. Computer Physics Communications, 72(1):29–38, October 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290004I.

Autin:1999:SMH

[Aut99] B. Autin. Symbolic modeling of high energy beam optics. Computer Physics Communications, 121–122:67–73, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002830.

Avdeev:1996:RRT

[Avd96] Leo. V. Avdeev. Recurrence relations for three-loop prototypes of bubble diagrams with a mass. Computer Physics Communications, 98(1–2):15–19, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000902.

Alkofer:1994:SCS

[AW94] R. Alkofer and H. Weigel. Self-consistent solution to a fermion determinant with space dependent fields. Computer Physics Communications, 82(1):30–41, August 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901295.

Ayala:1992:FRA

[Aya92] Carles Ayala. The μ(x, β, α) function and its role in the analysis of the QCD-SVZ sum rules. Computer Physics Communications, 70(2):401–408, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290202A.

Atanasiu:1992:IEM

[AZM92] C. V. Atanasiu, L. E. Zakharov, and A. Moraru. Integral equations method in equilibrium and stability computation for iron-core transformer tokamaks. Computer Physics Communications, 70(3):483–494, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290110K.

Ankudinov:1996:SCD

[AZR96] A. L. Ankudinov, S. I. Zabinsky, and J. J. Rehr. Single configuration Dirac–Fock atom code. Computer Physics Communications, 98(3):359–364, November 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000975.

Bartschat:1994:ESQ

[BA94] Klaus Bartschat and Nils Andersen. Electron scattering from quasi one-electron targets: experimental observables vs. theoretical scattering amplitudes. Computer Physics Communications, 84(1–3):335–345, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902194.

Benz:1995:SBS

[BA95] W. Benz and E. Asphaug. Simulations of brittle solids using smooth particle hydrodynamics. Computer Physics Communications, 87(1–2):253–265, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001763.

Becciani:1997:PTC

[BAAD^{+}97] U. Becciani, R. Ansaloni, V. Antonuccio-Delogu, G. Erbacci, M. Gambera, and A. Pagliaro. A parallel tree code for large N-body simulation: dynamic load balance and data distribution on a CRAY T3D system. Computer Physics Communications, 106(1–2):105–113, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001021.

Babalievski:1992:PPC

[Bab92] Filip V. Babalievski. On the parallelization of percolation cluster simulations. Computer Physics Communications, 67 (3):453–455, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900532.

Becciani:1996:WDS

[BADP96] U. Becciani, V. Antonuccio-Delogu, and A. Pagliaro. A work- and data-sharing parallel tree N-body code. Computer Physics Communications, 99(1):9–20, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559600121X.

Baillie:1991:NMC

[Bai91] Clive F. Baillie. A new MCRG calculation of the critical behavior of the 3d Ising model. Computer Physics Communications, 65(1–3):17–23, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190150J.

Balluch:1995:ANA

[Bal95] Martin Balluch. Adaptive numerical advection. the coordinate transformation equation method. Computer Physics Communications, 89(1–3):91–117, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001888.

Bartschat:1993:RIP

[Bar93] Klaus Bartschat. RMATRX-ION — a program to calculate electron and positron impact ionization within the R-matrix method. Computer Physics Communications, 75(1–2):219– 258, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390176D.

Bartschat:1998:RMP

[Bar98] Klaus Bartschat. The R-matrix with pseudo-states method: Theory and applications to electron scattering and photoionization. Computer Physics Communications, 114(1–3): 168–182, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000575.

Bastea:1999:PSV

[Bas99a] S. Bastea. Phase segregation via Vlasov–Boltzmann particle dynamics. Computer Physics Communications, 121–122: 270–273, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559900329X.

Biehl:1999:PTS

[BAS99b] Michael Biehl, Martin Ahr, and Enno Schlösser. Phase transitions in soft-committee machines. Computer Physics Communications, 121–122:614, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700212.

Brieger:1992:CAL

[BB92] Leesa Brieger and Ernesto Bonomi. Cellular automata — lattice gas models for PDE’s. Computer Physics Communications, 73(1–3):47–60, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290028W.

Bahan:1993:IJU

[BB93a] Graham Bahan and Roger Barlow. Identification of b jets using neural networks. Computer Physics Communications, 74(2):199–216, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390090Y.

Barlow:1993:FUF

[BB93b] Roger Barlow and Christine Beeston. Fitting using finite Monte Carlo samples. Computer Physics Communications, 77(2):219–228, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390005W.

Baboolal:1999:FSS

[BB99] S. Baboolal and R. Bharuthram. Fluid simulations of solitons and shock-like structures in an electrostatic plasma. Computer Physics Communications, 121–122:601, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700091.

Busby:1998:VAT

[BBB^{+}98] D. W. Busby, P. G. Burke, V. M. Burke, C. J. Noble, and N. S. Scott. VisRes: a GRACE tool for displaying and analysing resonances. Computer Physics Communications, 114(1–3):243–270, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000678.

Bardi:1999:NIS

[BBB99] J. Bardi, N. Binggeli, and A. Baldereschi. Numerical investigation of Schottky barriers at Al/Ga Al As interfaces. Computer Physics Communications, 121–122:602, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700108.

Bianchi:1993:AQM

[BBCM93] R. Bianchi, D. Bressanini, P. Cremaschi, and G. Morosi. Antisymmetry in quantum Monte Carlo methods. Computer Physics Communications, 74(2):153–163, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390086R.

Baker:1998:WMA

[BBE^{+}98] N. Baker, A. Bazan, F. Estrella, Z. Kovacs, T. Le Flour, J.- M. Le Goff, E. Leonardi, S. Lieunard, R. McClatchey, and J.-P. Vialle. Workflow management in the assembly of CMS ECAL. Computer Physics Communications, 110(1–3):170– 176, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001732.

Bernardo:1994:KPB

[BBG94] Dan N. Bernardo, Reena Bhatia, and Barbara J. Garrison. keV particle bombardment of solids: molecular dynamics simulations and beyond. Computer Physics Communications, 80(1–3):259–273, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901058.

Brickner:1991:QCM

[BBJ91] Ralph G. Brickner, Clive F. Baillie, and S. Lennart Johnsson. QCD on the connection machine: beyond LISP. Computer Physics Communications, 65(1–3):39–51, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190152B.

Bilke:1995:SQG

[BBJ95] S. Bilke, Z. Burda, and J. Jurkiewicz. Simplicial quantum gravity on a computer. Computer Physics Communications, 85(2):278–292, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400083E.

Bardin:1997:GVP

[BBL^{+}97] D. Bardin, J. Biebel, D. Lehner, A. Leike, A. Olchevski, and T. Riemann. GENTLE/4fan v. 2.0: a program for the semianalytic calculation of predictions for the process e^{+}e^{−} → 4f. Computer Physics Communications, 104(1–3):161–187, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000519.

Bassi:1999:PCK

[BBL^{+}99] M. Bassi, U. Becciani, U. Lombardo, G. Russo, N. Sandulescu, and W. Zuo. A parallel code for the kinetic Landau–Vlasov transport equation. Computer Physics Communications, 118(2–3):110–118, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599001952.

Bastea:1999:COM

[BBMD99] S. Bastea, A. Burkov, C. Moukarzel, and P. M. Duxbury. Combinatorial optimization methods in disordered systems. Computer Physics Communications, 121–122:199– 205, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003136.

Bok:1990:ISP

[BBP^{+}90] J. Bok, I. Barvík, P. Praus, P. Herman, and D. Cermáková. Integrated software packages in the physical laboratory. Computer Physics Communications, 61(1–2):219–224, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090119L.

Bardin:1990:DEO

[BBR^{+}90] D. Yu. Bardin, M. S. Bilenky, T. Riemann, M. Sachwitz, H. Vogt, and P. Ch. Christova. DIZET — electroweak one-loop corrections for e^{+}e^{−} → f^{+} + f^{−} around the Z^{0} peak. Computer Physics Communications, 59(2):303–312, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590901795.

Bottcher:1991:TDH

[BBS91] C. Bottcher, G. J. Bottrell, and M. R. Strayer. Timedependent Hartree–Fock studies of atomic collisions. Computer Physics Communications, 63(1–3):63–70, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190238G.

Burke:1992:NNE

[BBS92] V. M. Burke, P. G. Burke, and N. S. Scott. A new noexchange R-matrix program. Computer Physics Communications, 69(1):76–98, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290131H.

Brown:1991:LCC

[BC91a] D. Brown and J. H. R. Clarke. A loose-coupling, constantpressure, molecular dynamics algorithm for use in the modelling of polymer materials. Computer Physics Communications, 62(2–3):360–369, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190107V.

Bruhwiler:1991:DTS

[BC91b] David L. Bruhwiler and John R. Cary. Diffusion of trajectories in a simple Hamiltonian system with slow periodic forcing. Computer Physics Communications, 65(1–3):52–56, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190153C.

Berche:1999:MCI

[BC99a] Bertrand Berche and Christophe Chatelain. Monte Carlo investigation of the influence of randomness at firstorder phase transitions. Computer Physics Communications, 121–122:191–193, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599003100.

Boronat:1999:SDQ

[BC99b] J. Boronat and J. Casulleras. Sampling differences in quantum Monte Carlo: a generalized reweighting method. Computer Physics Communications, 121–122:466–467, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003835.

Bubak:1999:TSU

[BC99c] Marian Bubak and Przemysław Czerwiński. Traffic simulation using cellular automata and continuous models. Computer Physics Communications, 121–122:395–398, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559900363X.

Bellotti:1993:CBN

[BCD^{+}93] R. Bellotti, M. Castellano, C. De Marzo, N. Giglietto, G. Pasquariello, and P. Spinelli. A comparison between a neural network and the likelihood method to evaluate the performance of a transition radiation detector. Computer Physics Communications, 78(1–2):17–22, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901394.

Basios:1995:GRP

[BCM^{+}95] V. Basios, N. A. Chekanov, B. L. Markovski, V. A. Rostovtsev, and S. I. Vinitsky. GITA: a REDUCE program for the normalization of polynomial Hamiltonians. Computer Physics Communications, 90(2–3):355–368, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500080Y.

Brown:1993:DDP

[BCOY93] David Brown, Julian H. R. Clarke, Motoi Okuda, and Takao Yamazaki. A domain decomposition parallelization strategy for molecular dynamics simulations on distributed memory machines. Computer Physics Communications, 74(1):67–80, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390107N.

Brown:1994:DDP

[BCOY94] David Brown, Julian H. R. Clarke, Motoi Okuda, and Takao Yamazaki. A domain decomposition parallel processing algorithm for molecular dynamics simulations of polymers. Computer Physics Communications, 83(1):1–13, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900302. See erratum [BCOY95].

Brown:1995:EBD

[BCOY95] David Brown, Julian H. R. Clarke, Motoi Okuda, and Takao Yamazaki. Erratum: A domain decomposition parallel processing algorithm for molecular dynamics simulations of polymers [Comput. Phys. Commun. 83 (1994) 1–13]. Computer Physics Communications, 86(3):312, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000073. See [BCOY94].

Burkholder:1991:TRS

[BCP91] R. J. Burkholder, R.-C. Chou, and P. H. Pathak. Two ray shooting methods for computing the EM scattering by large open-ended cavities. Computer Physics Communications, 68(1–3):353–365, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591902094.

Bernaschi:1999:LCI

[BCSS99] M. Bernaschi, F. Castiglione, P. Seiden, and S. Succi. Learning cascade in the immune system dynamics: a numerical simulation. Computer Physics Communications, 121–122:122–125, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002945.

Bahri:1994:RME

[BD94] C. Bahri and J. P. Draayer. SU(3) reduced matrix element package. Computer Physics Communications, 83(1):59–94, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900353.

Brunet:1999:MMT

[BD99] Eric Brunet and Bernard Derrida. Microscopic models of traveling wave equations. Computer Physics Communications, 121–122:376–381, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003586.

Bartsch:1992:SME

[BDD^{+}92] Michael Bartsch, Micha Dehler, Martin Dohlus, Frank Ebeling, Peter Hahne, Reinhard Klatt, Frank Krawczyk, Michaela Marx, Zhang Min, Thomas Pröpper, Dietmar Schmitt, Petra Schütt, Bernhard Steffen, Bernhard Wagner, Thomas Weiland, Susan G. Wipf, and Heike Wolter. Solution of Maxwell’s equations. Computer Physics Communications, 73(1–3):22–39, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290026U.

Berces:1997:ICP

[BDF^{+}97] Attila Bérces, Ross M. Dickson, Liangyou Fan, Heiko Jacobsen, David Swerhone, and Tom Ziegler. An implementation of the coupled perturbed Kohn–Sham equations: perturbation due to nuclear displacements. Computer Physics Communications, 100(3):247–262, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001208.

Belozerova:1999:SCR

[BDH99] T. S. Belozerova, C. L. Davis, and V. K. Henner. Simulation of coherent radiofrequency superradiation from disordered spin systems. Computer Physics Communications, 121–122:214–218, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003161.

Bargiel:1991:CLM

[BDKM91] Monika Bargiel, Witold Dzwinel, Jacek Kitowski, and Jacek Mościński. C-language molecular dynamics program for the simulation of Lennard-Jones particles. Computer Physics Communications, 64(1):193–205, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190061O.

Brosolo:1992:CWC

[BDL92] M. Brosolo, P. Decleva, and A. Lisini. Continuum wavefunctions calculations with least-squares schemes in a B-splines basis. Computer Physics Communications, 71(3):207–214, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290009N.

Berube:1993:CCG

[BdM93] D. Bérubé and M. de Montigny. The computer calculation of graded contractions of Lie algebras and their representations. Computer Physics Communications, 76(3):389–410, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390063I.

Barrera:1997:SMA

[BdT97] G. D. Barrera and R. H. de Tendler. Simulation of metals and alloys using quasi-harmonic lattice dynamics. Computer Physics Communications, 105(2–3):159–168, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000775.

Budinich:1990:EAC

[BE90] M. Budinich and S. Esquivel. An efficient algorithm for charged tracks pattern recognition. Computer Physics Communications, 58(1–2):83–87, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090137P.

Brackbill:1995:P

[BE95] J. U. Brackbill and J. W. Eastwood. Preface. Computer Physics Communications, 87(1–2):xi, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595900039.

Bietenholz:1999:PIP

[BEF^{+}99] W. Bietenholz, N. Eicker, A. Frommer, Th. Lippert, B. Medeke, K. Schilling, and G. Weuffen. Preconditioning of improved and “perfect” fermion actions. Computer Physics Communications, 119(1):1–18, June 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001878.

Bunk:1999:OMF

[BEFJ99] B. Bunk, S. Elser, R. Frezzotti, and K. Jansen. Ordering monomial factors of polynomials in the product representation. Computer Physics Communications, 118(2–3):95–109, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599001988.

Berrington:1995:RBA

[BEN95] Keith A. Berrington, Werner B. Eissner, and Patrick H. Norrington. RMATRX1: Belfast atomic R-matrix codes. Computer Physics Communications, 92(2–3):290–420, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001238.

Bertsch:1990:RPJ

[Ber90] G. Bertsch. An RPA program for jellium spheres. Computer Physics Communications, 60(2):247–255, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090011O.

Berg:1992:DJB

[Ber92a] Bernd A. Berg. Double jackknife bias-corrected estimators. Computer Physics Communications, 69(1):7–14, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290124H.

Berg:1992:MCC

[Ber92b] Bernd A. Berg. Monte Carlo calculation of confidence limits for realistic least square fitting. Computer Physics Communications, 69(1):65–72, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290129M.

Bergdolt:1995:TIR

[Ber95] G. Bergdolt. Tilted irreducible representations of the permutation group. Computer Physics Communications, 86 (1–2):97–104, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595000095.

Berg:1996:NAI

[Ber96] Bernd A. Berg. New algorithm to investigate neural networks. Computer Physics Communications, 98(1–2):35–44, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000860.

Bertulani:1999:CPR

[Ber99] C. A. Bertulani. A computer program for relativistic multiple Coulomb and nuclear excitation. Computer Physics Communications, 116(2–3):345–352, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001416.

Beskrovnyi:1998:AMA

[Bes98] V. N. Beskrovnyi. Applying Mathematica to the analytical solution of the nonlinear Heisenberg operator equations. Computer Physics Communications, 111(1–3):76–86, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000290.

Brecht:1990:VCC

[BF90a] Stephen H. Brecht and John R. Ferrante. Vortex-in-cell calculations in three dimensions. Computer Physics Communications, 58(1–2):25–54, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090133L.

Bruge:1990:CMD

[BF90b] F. Brugè and S. L. Fornili. Concurrent molecular dynamics simulation of spinodal phase transition on transputer arrays. Computer Physics Communications, 60(1):31–38, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090076D.

Bruge:1990:DDL

[BF90c] F. Brugè and S. L. Fornili. A distributed dynamic load balancer and its implementation on multi-transputer systems for molecular dynamics simulation. Computer Physics Communications, 60(1):39–45, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090077E.

Bondeson:1991:TIS

[BF91] A. Bondeson and G. Y. Fu. Tunable integration scheme for the finite element method. Computer Physics Communications, 66(2–3):167–176, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190065S.

Brand:1998:DDL

[BF98] Susanne Brand and Patrick Fuhrmann. A distributed disk layer for mass storage at DESY. Computer Physics Communications, 110(1–3):131–133, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001665.

Brucher:1995:NMC

[BFK95] L. Brücher, J. Franzkowski, and D. Kreimer. A new method for computing one-loop integrals. Computer Physics Communications, 85(1):153–165, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400112F.

Brucher:1997:OPP

[BFK97] L. Brücher, J. Franzkowski, and D. Kreimer. oneloop 2.0 — a program package calculating one-loop integrals. Computer Physics Communications, 107(1–3):281–283, 286–292, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000982.

Brucher:1998:XAF

[BFK98] L. Brücher, J. Franzkowski, and D. Kreimer. xloops — automated Feynman diagram calculation. Computer Physics Communications, 115(2–3):140–160, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001623.

Bukin:1994:MCS

[BG94] A. D. Bukin and N. A. Grozina. Monte Carlo simulation of fluctuations of the ionization losses of heavy charged particles. Computer Physics Communications, 78(3):287–290, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900078.

Berkov:1995:RVD

[BG95] D. V. Berkov and N. L. Gorn. Reconstruction of the velocity distribution in conducting melts from induced magnetic field measurements. Computer Physics Communications, 86(3):255–263, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595000114.

Bowler:1998:LSI

[BG98] D. R. Bowler and M. J. Gillan. Length-scale ill conditioning in linear-scaling DFT. Computer Physics Communications, 112(2–3):103–111, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000617.

Bowler:1999:DME

[BG99] D. R. Bowler and M. J. Gillan. Density matrices in O(N) electronic structure calculations: theory and applications.

Computer Physics Communications, 120(2–3):95–108, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002210.

Bernaola-Galvan:1999:CCD

[BGCRRO99] P. Bernaola-Galván, P. Carpena, R. Román-Roldán, and J. L. Oliver. Compositional complexity of DNA sequence models. Computer Physics Communications, 121–122:136– 138, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002982.

Berger:1991:TDQ

[BGG91] J. F. Berger, M. Girod, and D. Gogny. Time-dependent quantum collective dynamics applied to nuclear fission. Computer Physics Communications, 63(1–3):365–374, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190263K.

Baginyan:1994:TMR

[BGK^{+}94] S. Baginyan, A. Glazov, I. Kisel, E. Konotopskaya, V. Neskoromnyi, and G. Ososkov. Tracking by a modified rotor model of neural network. Computer Physics Communications, 79(2):165–178, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900655.

Bazzani:1995:PCB

[BGT95] A. Bazzani, M. Giovannozzi, and E. Todesco. A program to compute Birkhoff normal forms of symplectic maps in R^{4}. Computer Physics Communications, 86(1–2):199–207, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400140W.

Barlow:1990:FRT

[BH90] Roger Barlow and Peter Hinde. Fitting the ratio of two distributions. Computer Physics Communications, 56(3):325–336, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090018V.

Belozerova:1992:ISA

[BH92] T. S. Belozerova and V. K. Henner. An interactive system for the application of generalized Fourier series in physics. Computer Physics Communications, 73(1–3):145–150, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290034V.

Belozerova:1999:GFM

[BH99] T. Belozerova and V. Henner. Generalized Fourier methods in mathematical physics — a software for education. Computer Physics Communications, 121–122:608, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700169.

Bhattacharyya:1991:ACP

[Bha91] Arun K. Bhattacharyya. Analysis of circular patch antennas on electrically thick substrates. Computer Physics Communications, 68(1–3):485–495, November 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591902146.

Buijs:1992:PAD

[BHL92] A. Buijs, B. Holl, and A. M. Lee. Parallel analysis of data on a multi-node VAX cluster. Computer Physics Communications, 67(3):407–411, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900495.

Besold:1999:MCS

[BHM99] G. Besold, O. Hassager, and O. G. Mouritsen. Monte Carlo simulation of diblock copolymer microphases by means of a fast off-lattice model. Computer Physics Communications, 121–122:542–544, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004026.

Breuer:1997:SWF

[BHP97] Heinz-Peter Breuer, Wolfgang Huber, and Francesco Petruccione. Stochastic wave-function method versus density matrix: a numerical comparison. Computer Physics Communications, 104(1–3):46–58, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000507.

Bazhanov:1999:ODM

[BHS^{+}99] D. I. Bazhanov, W. Hergert, V. S. Stepanyuk, N. A. Levanov, A. A. Katsnelson, P. Rennert, and C. Demangeat. One-dimensional magnetism of transition metal chains on Ag(001) surface. Computer Physics Communications, 121–122:605, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700133.

Belozerova:1992:CSC

[BHY92] T. S. Belozerova, V. K. Henner, and V. I. Yukalov. Computer simulation of coherent effects in polarized spin systems. Computer Physics Communications, 73(1–3):151–160, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290035W.

Burkitt:1990:IFM

[BI90] A. N. Burkitt and A. C. Irving. Inversion of the fermion matrix and the equivalence of the conjugate gradient and Lanczos algorithms. Computer Physics Communications, 59(3):447– 454, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090086G.

Billing:1991:BSI

[Bil91] Gert Due Billing. Basis set independent methods in timedependent dynamics. Computer Physics Communications, 63(1–3):38–50, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190236E.

Bilge:1992:RPI

[Bil92] Ayşe Hümeyra Bilge. A REDUCE program for the integration of differential polynomials. Computer Physics Communications, 71(3):263–268, September 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290013O.

Binder:1999:UGT

[Bin99] Kurt Binder. Understanding the glass transition and the amorphous state of matter: can computer simulation solve the challenge? Computer Physics Communications, 121–122:168–175, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003057.

Bishop:1991:BRB

[Bis91] C. M. Bishop. Book review: Neural networks — an introduction: B. Muller and J. Reinhardt, Springer-Verlag, Berlin, 1990. 266 pages, 5.25” diskette. Hardcover price £28.00. ISBN 3-540-52380-4. Computer Physics Communications, 67(2):357–359, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190029K.

Burda:1992:SWD

[BJK92] Z. Burda, J. Jurkiewicz, and L. Kärkkäinen. Swendsen– Wang dynamics for the Potts model on a dynamically triangulated random surface. Computer Physics Communications, 70(3):510–520, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290113D.

Baekdal:1996:OOD

[BJL^{+}96] Lars Bækdal, Wouter Joosen, Thomas Larsen, Jiri Kolafa, Jens H. Ovesen, John W. Perram, Henrik G. Petersen, Robert Bywater, and Mark Ratner. The object-oriented development of a parallel application in protein dynamics: why we need software tools for HPCN applications. Computer Physics Communications, 97(1–2):124–135, August 2, 1996.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000264.

Burckhart:1998:STP

[BJM^{+}98] D. Burckhart, R. Jones, L. Mapelli, M. Michelotto, A. Patel, M. Skiadelli, I. Soloviev, P-Y. Duval, A. Le Van Suu, R. Nacasch, Z. Qian, F. Touchard, M. Caprini, S. Kolos, K. Nurdan, and S. Wheeler. Software technologies for a prototype ATLAS DAQ. Computer Physics Communications, 110(1–3):113–119, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559700163X.

Baillie:1990:CAS

[BJW90] Clive F. Baillie, Desmond A. Johnston, and Roy D. Williams. Computational aspects of simulating dynamically triangulated random surfaces. Computer Physics Communications, 58(1–2):105–117, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090139R.

Berman:1991:TDS

[BK91] Michael Berman and Ronnie Kosloff. Time-dependent solution of the Liouville–von Neumann equation: non-dissipative evolution. Computer Physics Communications, 63(1–3):1–20, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190233B.

Bogdanova:1992:TAS

[BK92] N. Bogdanova and T. Kupenova. Two algorithms for smooth many-dimensional approximation in physics. Computer Physics Communications, 73(1–3):170–178, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290037Y.

Borzi:1993:MGM

[BK93] Alfio Borzì and Anni Koubek. A multi-grid method for the resolution of thermodynamic Bethe ansatz equations. Computer Physics Communications, 75(1–2):118–125, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390169D.

Balint-Kurti:1991:CPQ

[BKDMM91] Gabriel G. Balint-Kurti, Richard N. Dixon, C. Clay Marston, and A. J. Mulholland. The calculation of product quantum state distributions and partial cross-sections in timedependent molecular collision and photodissociation theory. Computer Physics Communications, 63(1–3):126–134, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190244F.

Biehl:1999:SME

[BKKS99] M. Biehl, M. Kinne, W. Kinzel, and S. Schinzer. A simple model of epitaxial growth: the influence of step edge diffusion. Computer Physics Communications, 121–122:347– 352, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003513.

Balint-Kurti:1993:CPC

[BKMM93] Gabriel G. Balint-Kurti, Steven P. Mort, and C. Clay Marston. A computer program to calculate the total energy absorption cross-section for the photodissociation of a diatomic molecule arising from a bound state → repulsive state transition using time dependent quantum dynamical methods. Computer Physics Communications, 74(2):289–296, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390098W.

Bockstedte:1997:DFT

[BKNS97] Michel Bockstedte, Alexander Kley, Jörg Neugebauer, and Matthias Scheffler. Density-functional theory calculations for poly-atomic systems: electronic structure, static and elastic properties and ab initio molecular dynamics. Computer Physics Communications, 107(1–3):187–222, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001173.

Bren:1999:MCS

[BKS99] D. Bren, P. Kulhánek, and D. Skandera. Monte Carlo simulation of the systems with many particles. Computer Physics Communications, 121–122:616, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700236.

Bravar:1997:PMC

[BKW97] Alessandro Bravar, Krzysztof Kurek, and Roland Windmolders. POLDIS: a Monte Carlo for POLarized (semiinclusive) Deep Inelastic Scattering. Computer Physics Communications, 105(1):42–61, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000635.

Balint-Kurti:1991:TCP

[BKWM91] Gabriel G. Balint-Kurti, Christopher L. Ward, and C. Clay Marston. Two computer programs for solving the Schrödinger equation for bound-state eigenvalues and eigenfunctions using the Fourier grid Hamiltonian method. Computer Physics Communications, 67(2):285–292, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190023E.

Belousov:1999:MMD

[BL99a] A. I. Belousov and Yu. E. Lozovik. Mesoscopic and macroscopic dipole clusters: structure and phase transitions. Computer Physics Communications, 121–122:607, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506700157.

Belousov:1999:PDF

[BL99b] A. I. Belousov and Yu. E. Lozovik. Phase diagram of a 2D frustrated array of mesoscopic objects. Computer Physics Communications, 121–122:510–511, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900394X.

Blankenbecler:1994:DTR

[Bla94a] Richard Blankenbecler. Deformable templates — revisited and extended — with an OOP implementation. Computer Physics Communications, 81(3):318–334, July 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900809.

Blankenbecler:1994:UTT

[Bla94b] Richard Blankenbeclér. A unified treatment of track reconstruction and particle identification. Computer Physics Communications, 81(3):335–342, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900817.

Beltran-Lopez:1992:PPE

[BLGT92] Virgilio Beltrán-López and L. González-Tovany. POWDER-SPEC, a program for efficient simulation of isotropic EPR spectra. Computer Physics Communications, 69(2–3):397– 405, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290177Z.

Beltran-Lopez:1994:PLN

[BLGT94] Virgilio Beltrán-López and L. González-Tovany. POW-DERSPEC 2, a library of new programs for efficient simulation of powder EPR spectra. Computer Physics Communications, 79(3):533–546, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901929.

Buijs:1994:TSM

[BLLM94] A. Buijs, W. G. J. Langeveld, M. H. Lehto, and D. J. Miller. TWOGEN-a simple Monte Carlo generator for two-photon reactions. Computer Physics Communications, 79(3):523–532, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901910.

Block:1992:MCP

[Blo92] Martin M. Block. Monte Carlo phase space evaluation. Computer Physics Communications, 69(2–3):459–476, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290183Y.

Belkov:1996:CHK

[BLS96] A. A. Bel’kov, A. V. Lanyov, and A. Schaale. Calculation of heat-kernel coefficients and usage of computer algebra. Computer Physics Communications, 95(2–3):123–130, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000318.

Bargiel:1991:CLP

[BM91] Monika Bargiel and Jacek Mościński. C-language program for the irregular close packing of hard spheres. Computer Physics Communications, 64(1):183–192, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190060X.

Bonnin:1992:ARF

[BM92a] X. Bonnin and R. Marchand. Analytic rates for fitted ionization cross-sections. Computer Physics Communications, 71(1–2):147–149, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290079E.

Bortolani:1992:PAS

[BM92b] E. Bortolani and G. Maino. Photon absorption and scattering cross-sections by triaxial nuclei. Computer Physics Communications, 70(1):207–218, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592901014.

Boss:1995:SOA

[BM95] Alan P. Boss and Elizabeth A. Myhill. Second-order-accurate radiative hydrodynamics and multidimensional protostellar collapse. Computer Physics Communications, 89(1–3):59–67, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594001866.

Barkema:1999:ESM

[BM99a]  G. T. Barkema and Normand Mousseau.  Exploring

structural mechanisms in disordered materials using the activation-relaxation technique. Computer Physics Communications, 121–122:206–209, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003148.

Batrouni:1999:DQM

[BM99b] G. G. Batrouni and H. Mabilat. Dual quantum Monte Carlo algorithm for hardcore bosons. Computer Physics Communications, 121–122:468–476, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003847.

Benmakhlouf:1999:NSD

[BM99c] A. Benmakhlouf and A. Menai. Numerical simulation of diffusion by dissociative mechanism in the case of finite foreign atom source. Computer Physics Communications, 121–122:609–610, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700170.

Braun:1996:NWP

[BME96] Michael Braun, Christoph Meier, and Volker Engel. Nanosecond wave-packet propagation with the Split-Operator Technique. Computer Physics Communications, 93(2–3):152–158, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001328.

Bernardes:1999:CBP

[BMeS99] A. T. Bernardes, J.-G. Moreira, and A. Castro e Silva. Chaotic behaviour on population dynamics. Computer Physics Communications, 121–122:613, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700200.

Brown:1997:DDP

[BMM97] David Brown, Herve’ Minoux, and Bernard Maigret. A domain decomposition parallel processing algorithm for molecular dynamics simulations of systems of arbitrary connectivity. Computer Physics Communications, 103(2–3):170–186, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000404.

Belmont-Moreno:1999:IEN

[BMMMMR99] E. Belmont-Moreno, K. Michaelian, A. Martinez, and A. Menchaca-Rocha. Information extraction from nuclear spectra with an evolutive algorithm. Computer Physics Communications, 121–122:606, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700145.

Badralexe:1992:CKF

[BMOF92] E. Badralexe, P. Marksteiner, Y. Oh, and A. J. Freeman. Computation of the Kummer functions and Whittaker functions by using Neumann type series expansions. Computer Physics Communications, 71(1–2):47–55, August 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900716.

Badralexe:1994:CCE

[BMOF94] E. Badralexe, P. Marksteiner, Yoonsik Oh, and A. J. Freeman. Coupled channel equation for potentials with a Coulomb singularity. Computer Physics Communications, 82(2–3): 120–128, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901619.

Bishop:1991:P

[BMP91] Marvin Bishop and Florian Müller-Plathe. Preface. Computer Physics Communications, 62(2–3):v, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900908.

Bubak:1999:PIL

[BMPS99] M. Bubak, J. Mościński, M. Pogoda, and R. Slota. Parallel implementation of 2D LGA and MD simulations. Computer Physics Communications, 121–122:618, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550670025X.

Baker:1991:DMB

[BMSW91] D. J. Baker, D. Moncrieff, V. R. Saunders, and S. Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules: VII experiments in vector and parallel processing for fourth-order energy terms involving triply excited intermediate states. Computer Physics Communications, 62(1):25–41, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901185.

Bernardson:1994:MCM

[BMT94] Shannon Bernardson, Paul McCarty, and Chris Thron. Monte Carlo methods for estimating linear combinations of inverse matrix entries in lattice QCD. Computer Physics Communications, 78(3):256–264, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900043.

Brennan:1991:SAS

[BMW91] Kevin F. Brennan, Nabil Mansour, and Yang Wang. Simulation of advanced semiconductor devices using supercomputers. Computer Physics Communications, 67(1):73–92, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591902227.

Babkevich:1992:AMC

[BN92] A. Yu. Babkevich and B. I. Nikolin. Application of the Monte Carlo technique to the investigation of onedimensionally disordered structures. Computer Physics Communications, 73(1–3):139–144, December 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290033U.

Burke:1994:RDR

[BN94] V. M. Burke and C. J. Noble. Recent developments in R-matrix programs for electron-atom scattering. Computer Physics Communications, 84(1–3):19–36, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902011.

Burke:1995:FFA

[BN95] V. M. Burke and C. J. Noble. Farm — a flexible asymptotic R-matrix package. Computer Physics Communications, 85(3):471–500, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594001785.

Bryan:1995:PPM

[BNS^{+}95] Greg L. Bryan, Michael L. Norman, James M. Stone, Renyue Cen, and Jeremiah P. Ostriker. A piecewise parabolic method for cosmological hydrodynamics. Computer Physics Communications, 89(1–3):149–168, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001914.

Barb:1992:SIP

[BNSW92] Florin Dan Barb, Oprea Netoiu, Monica Sorescu, and Martin Weiss. SPECFIT — an interactive package for Mössbauer spectra fitting with personal computers. Computer Physics Communications, 69(1):182–186, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290138O.

Baginyan:1998:FTD

[BO98] S. Baginyan and G. Ososkov. Finding tracks detected by a drift tube system. Computer Physics Communications, 108 (1):20–28, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597001288.

Bocko:1992:ERB

[Boc92] J. Bocko. EQSHELL — a REDUCE-based program for generation of equations of equilibrium for shells. Computer Physics Communications, 69(1):215–222, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290141K.

Bondorf:1993:BRB

[Bon93] Jakob B. Bondorf. Book review: Computational nuclear physics 1. nuclear structure: edited by K. Langanke, J. A. Maruhn and S. E. Koonin. Springer-Verlag, Berlin, 1991. 209 + xii pages. Hardcover price DM 88.00 ISBN 3-540-53571-

3. Computer Physics Communications, 74(3):450–451, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465593900269.

Borcherds:1999:SRS

[Bor99] P. H. Borcherds. Some results for the standard map with twist. Computer Physics Communications, 121–122:432– 436, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003768.

Bos:1998:MP

[Bos98]  Kors Bos.  The moose project.  Computer Physics

Communications, 110(1–3):160–163, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001719.

Bowler:1999:WQN

[Bow99] K. C. Bowler. Why QCD needs high performance computing. Computer Physics Communications, 121–122:512– 516, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003951.

Boyd:1995:ESP

[Boy95] John K. Boyd. Electrostatic scalar potential and electric field solution of an axisymmetric accelerator gap. Computer Physics Communications, 88(1):23–42, July 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500033C.

Babalievski:1990:ACQ

[BP90a] F. Babalievski and O. Peshev. An algorithm to construct quasilattices and study percolation on them. Computer Physics Communications, 60(1):27–30, August 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090075C.

Biller:1990:ASN

[BP90b] Peter Biller and Francesco Petruccione. Algorithms for the simulation of network models of the yamamoto type. Computer Physics Communications, 61(3):279–284, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090042Y.

Breuer:1993:NIB

[BP93] Heinz-Peter Breuer and Francesco Petruccione. On the numerical integration of Burgers’ equation by stochastic simulation methods. Computer Physics Communications, 77(2):207–218, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390004V.

Belien:1997:TDE

[BPG97] A. J. C. Beliën, S. Poedts, and J. P. Goedbloed. Twodimensional equilibrium in coronal magnetostatic flux tubes: an accurate equilibrium solver. Computer Physics Communications, 106(1–2):21–38, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000842.

Berends:1995:EMC

[BPK95] F. A. Berends, R. Pittau, and R. Kleiss. EXCALIBUR — a Monte Carlo program to evaluate all four-fermion processes at LEP 200 and beyond. Computer Physics Communications, 85(3):437–452, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400138R.

Blanco:1997:QMC

[BPL97] Miguel A. Blanco, A. Martiń Pendaś, and V. Luaña. Quantum mechanical cluster calculations of ionic materials: Revision 10 of the ab initio Perturbed Ion program. Computer Physics Communications, 103(2–3):287–302, July 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000301.

Bell:1995:PCP

[BPMW95] P. M. Bell, J. B. Pendry, L. Martín Moreno, and A. J. Ward. A program for calculating photonic band structures and transmission coefficients of complex structures. Computer Physics Communications, 85(2):306–322, February 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400131K.

Berends:1998:WET

[BPP98] F. A. Berends, C. G. Papadopoulos, and R. Pittau. WEX-TER and ERAFITTER: two programs to fit M_{W} at LEP2 using the best measurable kinematical variables. Computer Physics Communications, 115(1):32–44, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001040.

Bajpai:1995:RCM

[BR95] R. P. Bajpai and S. N. Rai. Retrieval characteristics of multiple site states networks with self-interaction. Computer Physics Communications, 88(1):51–58, July 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400130T.

Bajpai:1997:BTT

[BR97a] R. P. Bajpai and S. N. Rai. Behavior of two, three and four site states self-interacting neural networks. Computer Physics Communications, 100(1–2):17–25, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001427.

Berg:1997:SCL

[BR97b] Bernd A. Berg and Jürgen Riedler. Signal confidence limits from a neural network data analysis. Computer Physics Communications, 107(1–3):39–48, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001112.

Brandt:1990:CAP

[Bra90] Josef Brandt. Computer-aided production of scientific documents. Computer Physics Communications, 61(1–2):163–176, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090114G.

Brandt:1991:MCI

[Bra91] Achi Brandt. Multilevel computations of integral transforms and particle interactions with oscillatory kernels. Computer Physics Communications, 65(1–3):24–38, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190151A.

Bray:1998:CEP

[Bra98] Igor Bray. Calculation of electron- and photon-impact ionization via a close-coupling approach. Computer Physics Communications, 114(1–3):356–367, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000587.

Bhattacharya:1992:FRU

[BRB92] Ranjan Bhattacharya, Dhiranjan Roy, and Siddhartha Bhowmick. Finding roots using divergent functional iteration. Computer Physics Communications, 69(2–3):339– 346, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290172U.

Bhattacharya:1997:RIU

[BRB97] Ranjan Bhattacharya, Dhiranjan Roy, and Siddhartha Bhowmick. Rational interpolation using Levin–Weniger transforms. Computer Physics Communications, 101(3):213–222, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001695.

Brey:1999:DMC

[BRM99] J. Javier Brey and M. J. Ruiz-Montero. Direct Monte Carlo simulation of dilute granular flow. Computer Physics Communications, 121–122:278–283, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003318.

Brownrigg:1992:BRB

[Bro92] D. R. K. Brownrigg. Book review: Digital image processing: Bernd Jahne, Springer Verlag, Berlin, 1991. 383 + xiii pages. Softcover price DM98.00. ISBN 3-540-53782-1. Computer Physics Communications, 71(3):344–345, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290019U.

Brownrigg:1994:DIP

[Bro94] D. R. K. Brownrigg. Digital image processing. Computer Physics Communications, 82(2–3):308, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901767.

Bruneau:1991:LSC

[Bru91] Charles-Henri Bruneau. Least-squares computation of hypersonic flows. Computer Physics Communications, 65(1–3):57– 61, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190154D.

Bruge:1995:MGS

[Bru95] F. Brugè. A mixed geometric-systolic approach to parallel molecular dynamics simulations. Computer Physics Communications, 90(1):59–65, September 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500070V.

Barvik:1996:CSE

[BRWN96] I. Barvík, P. Reineker, Ch. Warns, and Th. Neidlinger. Computer simulation of excitonic optical line shapes: transfer integral fluctuations with dichotomic coloured noise. Computer Physics Communications, 97(1–2):154–162, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000288.

Berretti:1990:VPM

[BS90] Alberto Berretti and Alan D. Sokal. Vectorized program for Monte Carlo simulation of self-avoiding walks. Computer Physics Communications, 58(1–2):1–16, February/ March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090131J.

Balakrishnan:1991:TDW

[BS91a] N. Balakrishnan and N. Sathyamurthy. Time-dependent wavepacket methods for the calculation of collinear atomdiatom exchange reaction probabilities. Computer Physics Communications, 63(1–3):209–215, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190250O.

Bayard:1991:FDS

[BS91b] Jean-Pierre R. Bayard and Daniel H. Schaubert. Finite difference solution of infinite arrays of two-dimensional microstrip structures. Computer Physics Communications, 68 (1–3):366–392, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190210C.

Berblinger:1991:MCI

[BS91c] Michael Berblinger and Christoph Schlier. Monte Carlo integration with quasi-random numbers: some experience. Computer Physics Communications, 66(2–3):157–166, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190064R.

Bray:1995:CCC

[BS95] Igor Bray and Andris T. Stelbovics. The convergent closecoupling method for a Coulomb three-body problem. Computer Physics Communications, 85(1):1–17, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400134N.

Batrouni:1996:WLS

[BS96] G. G. Batrouni and R. T. Scalettar. World line simulations of the bosonic Hubbard model in the ground state. Computer Physics Communications, 97(1–2):63–81, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000227.

Baikov:1998:TLV

[BS98] P. A. Baikov and M. Steinhauser. Three-loop vacuum integrals in FORM and REDUCE. Computer Physics Communications, 115(2–3):161–169, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001258.

Blackett:1999:AFB

[BS99a] Anthony J. Blackett and Andris T. Stelbovics. Analytic first Born atomic scattering: a computer algebra solution using Maple V. Computer Physics Communications, 116(1):78–94, January 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001854.

Bornholdt:1999:PEE

[BS99b] Stefan Bornholdt and Kim Sneppen. Punctuated equilibrium in evolving genetic networks. Computer Physics Communications, 121–122:615, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700224.

Barvik:1999:MDS

[BSB99] I. Barvík, Jr., J. Stepánek, and J. Bok. Molecular dynamics simulations of the oligonucleotide with the modified phosphate/phosphonate internucleotide linkage. Computer Physics Communications, 121–122:603, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550670011X.

Bar-Shalom:1996:PAA

[BSKO96] A. Bar-Shalom, M. Klapisch, and J. Oreg. Phase-amplitude algorithms for atomic continuum orbitals and radial integrals. Computer Physics Communications, 93(1):21–32, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500100X.

Booth:1999:ESL

[BSSH99] Michael J. Booth, A. G. Schlijper, L. E. Scales, and A. D. J. Haymet. Efficient solution of liquid state integral equations using the Newton–GMRES algorithm. Computer Physics Communications, 119(2–3):122–134, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001861.

Blaha:1990:FPL

[BSST90] P. Blaha, K. Schwarz, P. Sorantin, and S. B. Trickey. Fullpotential, linearized augmented plane wave programs for crystalline systems. Computer Physics Communications, 59 (2):399–415, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465590901876.

Barone:1995:MDA

[BST95] Luciano Maria Barone, Riccardo Simonazzi, and Alexander Tenenbaum. Molecular dynamics on APE100. Computer Physics Communications, 90(1):44–58, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500076R.

Berblinger:1997:MCI

[BSW97] M. Berblinger, Ch. Schlier, and T. Weiss. Monte Carlo integration with quasi-random numbers: experience with discontinuous integrands. Computer Physics Communications, 99(2–3):151–162, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001312.

Besprozvannaya:1991:FSC

[BT91] Anna Besprozvannaya and David J. Tannor. Fast sine/cosine transform for periodic functions with reflection symmetry. Computer Physics Communications, 63(1–3):569–577, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190276Q.

Bonomi:1992:DPM

[BT92a] Ernesto Bonomi and Marco Tomassini. Data-parallel molecular dynamics: 1-D hard-core fluid. Computer Physics Communications, 70(1):32–40, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290088G.

Bonomi:1992:DPS

[BT92b] Ernesto Bonomi and Marco Tomassini. Data-parallel simulations of hard-particle systems. Computer Physics Communications, 73(1–3):40–46, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290027V.

Berdnikov:1999:CLV

[BTC99a] A. S. Berdnikov, S. B. Turtia, and A. Compagner. The combination of LP _{τ} vectors and shift register RNGs as a way to generate independent random sequences for parallel computations. Computer Physics Communications, 121–122:612, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700194.

Berdnikov:1999:FME

[BTC99b] A. S. Berdnikov, S. B. Turtia, and A. Compagner. The fast and memory-economical algorithm to generate the primitive polynoms for GF(2^{n}) algebras starting from a single one. Computer Physics Communications, 121–122:611, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700182.

Bunn:1998:CCE

[Bun98] Julian J. Bunn. Collaborative computing environments for HEP. Computer Physics Communications, 110(1–3):51–58, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001537.

Burke:1998:NMC

[Bur98] V. M. Burke. A new method for calculating angular integrals in electron-atom scattering. Computer Physics Communications, 114(1–3):210–224, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001027.

Borbely:1995:ISE

[BV95] I. Borbély and T. Vertse. Inhomogeneous Schrödinger equation for anti-bound and resonant states. Computer Physics Communications, 86(1–2):61–66, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400158X.

Bergeron:1999:ISR

[BV99] Stephen Bergeron and Alain Vincent. Implementation strategies for real-time particle transport solver. Computer Physics Communications, 120(2–3):177–184, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002349.

Berendsen:1995:GMP

[BvdSvD95] H. J. C. Berendsen, D. van der Spoel, and R. van Drunen. GROMACS: a message-passing parallel molecular dynamics implementation. Computer Physics Communications, 91(1–3):43–56, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500042E.

Barberio:1991:PUM

[BvEW91] Elisabetta Barberio, Bob van Eijk, and Zbigniew Was. Photos — a universal Monte Carlo for QED radiative corrections in decays. Computer Physics Communications, 66(1):115– 128, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190012A.

Billeter:1997:MMD

[BvG97] S. R. Billeter and W. F. van Gunsteren. A modular molecular dynamics/quantum dynamics program for nonadiabatic proton transfers in solution. Computer Physics Communications, 107(1–3):61–91, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000891.

Berghammer:1995:CPT

[BVR95] H. Berghammer, D. Vretenar, and P. Ring. Computer program for the time-evolution of a nuclear system in relativistic mean-field theory. Computer Physics Communications, 88(2–3):293–308, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500018B.

Broukhno:1999:MCS

[BVVJA99] A. V. Broukhno, P. N. Vorontsov-Velyaminov, B. Jonsson, and T. Akesson. Monte Carlo simulation of confined polymers at constant chemical potential with the use of expanded ensemble approach. Computer Physics Communications, 121–122:617, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700248.

Barberio:1994:PUM

[BW94] Elisabetta Barberio and Zbigniew Was. PHOTOS — a universal Monte Carlo for QED radiative corrections: version 2.0. Computer Physics Communications, 79(2):291–308, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900744.

Barvik:1999:SEO

[BWNR99] I. Barvík, Ch. Warns, Th. Neidlinger, and P. Reineker. Simulation of excitonic optical line shapes of cyclic molecular aggregates with 9 and 18 units: influence of quasistatic and dynamic disorder. Computer Physics Communications, 121–122:604, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700121.

Bormet:1994:SCS

[BWS94] Jörg Bormet, Bernd Wenzien, and Matthias Scheffler. A selfconsistent surface green-function (SSGF) method for the calculation of isolated adsorbate atoms on a semi-infinite crystal. Computer Physics Communications, 79(1):124–142, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902356.

Belyaeva:1992:CCC

[BZO92] T. L. Belyaeva, N. S. Zelenskaya, and N. V. Odintsov. Computation of correlation characteristics of nuclear reactions induced by semi-heavy ions. Computer Physics Communications, 73(1–3):161–169, December 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290036X.

Cooper:1992:FVI

[CA92] R. K. Cooper and R. J. Allan. Fortnet (3L) v1.0: Implementation and extensions of a message-passing harness for transputers using 3L parallel Fortran. Computer Physics Communications, 70(3):521–543, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290114E.

Cabrol:1990:APR

[Cab90] D. Cabrol. Applications of prolog to represent physical and chemical objects — a tutorial introduction. Computer Physics Communications, 61(1–2):34–57, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090105A.

Califano:1996:NAG

[Cal96] Francesco Califano. A numerical algorithm for geophysical and astrophysical inhomogeneous fluid flows. Computer Physics Communications, 99(1):29–42, December 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001178.

Camenzind:1995:IFT

[Cam95] Max Camenzind. An implicit FE-time solver for axisymmetric magnetic dynamo equations. Computer Physics Communications, 89(1–3):17–27, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500183G.

Candiello:1994:RCW

[Can94a] Antonio Candiello. Reply on the comment on “Wbase: a C package to reduce tensor products of Lie algebra representations”. Computer Physics Communications, 83(2–3):334, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900590. See [Can94b,

Wyb94].

Candiello:1994:WCP

[Can94b] Antonio Candiello. WBase: a C package to reduce tensor products of Lie algebra representations. Computer Physics Communications, 81(1–2):248–260, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901236. See comment [Wyb94] and

reply [Can94a].

Capurro:1993:ACS

[Cap93] O. A. Capurro. Absolute cross sections of compound nucleus reactions. Computer Physics Communications, 77(3):396–402, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390185F.

Carles:1994:SSS

[Car94] Agustín Grau Carles. SRLOG, the simultaneous standardization of ^{90}Sr + ^{90}Y + ^{89}Sr mixtures. Computer Physics Communications, 82(1):17–22, August 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901279.

Carles:1996:MSS

[Car96] Agustín Grau Carles. MLOG, the simultaneous standardization of multi-nuclide mixtures. Computer Physics Communications, 93(1):48–52, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001107.

Catterall:1995:SDT

[Cat95] S. Catterall. Simulations of dynamically triangulated gravity — an algorithm for arbitrary dimension. Computer Physics Communications, 87(3):409–415, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400117K.

Cvjetkovic:1999:OES

[CB99] V. Cvjetković and V. Bočvarski. Online expert system for analysis of inelastic electron scattering by metal atoms. Computer Physics Communications, 120(2–3):185–196, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002465.

Chatelain:1999:IDF

[CBB99] Christophe Chatelain, Pierre Emmanuel Berche, and Bertrand Berche. Influence of deterministic fluctuations on the 8-state Potts model. Computer Physics Communications, 121–122:197–198, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599003124.

Cohen:1995:NTP

[CBD^{+}95] B. I. Cohen, D. C. Barnes, J. M. Dawson, G. W. Hammett, W. W. Lee, G. D. Kerbel, J.-N. Leboeuf, P. C. Liewer, T. Tajima, and R. E. Waltz. The numerical tokamak project: simulation of turbulent transport. Computer Physics Communications, 87(1–2):1–15, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400166Y.

Carpena:1999:SPL

[CBG99] Pedro Carpena and Pedro Bernaola-Galván. Spectral properties and localization of vibrational states in quasiperiodic chains. Computer Physics Communications, 121–122: 620, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700273.

Compagner:1997:RER

[CBTL97] A. Compagner, A. S. Berdnikov, S. B. Turtia, and A. Larionov. Rounding errors in random number generators. Computer Physics Communications, 106(3):207–218, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000702.

Carver:1997:AAC

[CBW97] Glenn D. Carver, Paul D. Brown, and Oliver Wild. The ASAD atmospheric chemistry integration package and chemical reaction database. Computer Physics Communications, 105(2–3):197–215, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000568.

Carr:1991:SCE

[CC91a] Lawrence W. Carr and M. S. Chandrasekhara. A study of compressibility effects on dynamic stall of rapidly pitching airfoils. Computer Physics Communications, 65(1–3):62–68, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190155E.

Cooray:1991:SSS

[CC91b] M. F. R. Cooray and I. R. Ciric. Scattering by systems of spheroids in arbitrary configurations. Computer Physics Communications, 68(1–3):279–305, November 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190204X.

Caffo:1997:BMC

[CC97a] M. Caffo and H. Czyz. BHAGEN-1PH: a Monte Carlo event generator for radiative Bhabha scattering. Computer Physics Communications, 100(1–2):99–118, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001348.

Coelho:1997:FSM

[CC97b] A. A. Coelho and R. W. Cheary. A fast and simple method for calculating electrostatic potentials. Computer Physics Communications, 104(1–3):15–22, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000398.

Cyganowski:1998:MPN

[CC98] Sasha Cyganowski and John Carminati. The Maple package NPTOOLS; a symbolic algebra package for tetrad formalisms in general relativity. Computer Physics Communications, 115(2–3):200–214, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001635.

Cavazzoni:1999:PMD

[CC99] Carlo Cavazzoni and Guido L. Chiarotti. A parallel and modular deformable cell car–parrinello code. Computer Physics Communications, 123(1–3):56–76, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900418X.

Casas:1995:CCE

[CCM95] Gonzalo Galiano Casas, Agustín Grau Carles, and Agustín Grau Malonda. CAPMULT, the counting efficiency for electron capture by a KLMN four-shell model. Computer Physics Communications, 87(3):432–437, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500017A.

Chialvo:1990:UVN

[CD90] Ariel A. Chialvo and Pablo G. Debenedetti. On the use of the Verlet neighbor list in molecular dynamics. Computer Physics Communications, 60(2):215–224, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090007N.

Chialvo:1991:PAV

[CD91] Ariel A. Chialvo and Pablo G. Debenedetti. On the performance of an automated Verlet neighbor list algorithm for large systems on a vector processor. Computer Physics Communications, 64(1):15–18, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190046N.

Chialvo:1992:AVN

[CD92] Ariel A. Chialvo and Pablo G. Debenedetti. An automated Verlet neighbor list algorithm with a multiple timestep approach for the simulation of large systems. Computer Physics Communications, 70(3):467–477, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290108B.

Choi:1996:SPL

[CDD^{+}96] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: a portable linear algebra library for distributed memory computers — design issues and performance. Computer Physics Communications, 97(1–2):1–15, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000173.

Coperchio:1998:WWW

[CDdG^{+}98] M. C. Coperchio, M. Dönszelmann, N. de Groot, P. Gunnarsson, M. Litmaath, D. McNally, and N. Smirnov. WIRED — World-Wide Web Interactive Remote Event Display. Computer Physics Communications, 110(1–3):155–159, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001707.

Castellano:1990:MCP

[CDG^{+}90] M. Castellano, G. De Cataldo, N. Giglietto, E. Nappi, and P. Spinelli. A Monte Carlo program to design a multiple module transition radiation detector. Computer Physics Communications, 61(3):395–409, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900534.

Cohen:1996:LAC

[CdG96] A. M. Cohen and W. A. de Graaf. Lie algebraic computation. Computer Physics Communications, 97(1–2):53–62, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000215.

Capuzzo-Dolcetta:1999:PTC

[CDM99] R. Capuzzo-Dolcetta and P. Miocchi. Parallelization of a tree-code for the simulation of self-gravitating astrophysical systems. Computer Physics Communications, 121–122:423– 424, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003732.

Clydesdale:1991:HPP

[CDR91] G. Clydesdale, R. Docherty, and K. J. Roberts. HABIT — a program for predicting the morphology of molecular crystals. Computer Physics Communications, 64(2):311–328, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190040R.

Cagin:1999:SRT

[CDUT99] T. Cagin, G. Dereli, M. Uludogan, and M. Tomak. Simulation results of thermal and mechanical properties of some fcc transition metals. Computer Physics Communications, 121–122:619, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700261.

Chow:1995:IIM

[CF95] Kim-Hung Chow and David M. Ferguson. Isothermal-isobaric molecular dynamics simulations with Monte Carlo volume sampling. Computer Physics Communications, 91(1–3):283– 289, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500059O.

Chen:1999:CPQ

[CG99] Rongqing Chen and Hua Guo. The Chebyshev propagator for quantum systems. Computer Physics Communications, 119(1):19–31, June 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001799.

Cabrera:1999:NCT

[CGdlR99] Juan L. Cabrera, J. Gorroñogoitia, and F. J. de la Rubia. Noise correlation time controlled localization. Computer Physics Communications, 121–122:371–373, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003562.

Colgan:1998:CMM

[CGHB98] J. Colgan, D. H. Glass, K. Higgins, and P. G. Burke. The calculation of molecular multiphoton processes using the R-matrix-Floquet method. Computer Physics Communications, 114(1–3):27–41, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001064.

Chernov:1993:TVR

[CGK^{+}93] N. Chernov, A. Glazov, I. Kisel, E. Konotopskaya, S. Korenchenko, and G. Ososkov. Track and vertex reconstruction in discrete detectors using Chebyshev metrics. Computer Physics Communications, 74(2):217–227, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390091P.

Cudzewicz:1998:NGF

[CGL^{+}98] Ron Cudzewicz, Lisa Giacchetti, Mark Leininger, Tanya Levshina, Ramon Pasetes, Marilyn Schweitzer, and Stephen Wolbers. Next generation farms at Fermilab. Computer Physics Communications, 110(1–3):211–215, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001793.

Canning:1996:TBM

[CGM^{+}96] A. Canning, G. Galli, F. Mauri, A. De Vita, and R. Car. O(N) tight-binding molecular dynamics on massively parallel computers: an orbital decomposition approach. Computer Physics Communications, 94(2–3):89–102, April 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000094.

Cordoba:1999:GAO

[CGM99] A. Córdoba and L. I. González-Monroy. Genetic algorithms to optimize energy supply systems. Computer Physics Communications, 121–122:43–45, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002763.

Cordelli:1994:PNA

[CGP94] Alessandro Cordelli, Giuseppe Grosso, and Giuseppe Pastori Parravicini. Physical and numerical aspects in Lanczos and modified Lanczos calculations. Computer Physics Communications, 83(2–3):255–265, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900531.

Crabb:1993:RLS

[CGR93] T. A. Crabb, P. N. Gibson, and K. J. Roberts. REX — a least-squares fitting program for the simulation and analysis of X-ray reflectivity data. Computer Physics Communications, 77(3):441–449, November 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390188I.

Cartwright:1999:PFC

[CGRV99] Julyan H. E. Cartwright, Juan Manuel García-Ruiz, and Ana I. Villacampa. Pattern formation in crystal growth: Liesegang rings. Computer Physics Communications, 121– 122:411–413, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465599003707.

Charchula:1992:PPP

[Cha92] K. Charchula. The package PAKPDF 1.1 of parametrizations of parton distribution functions in the proton. Computer Physics Communications, 69(2–3):360–368, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290174W.

Chu:1991:SAW

[CHC91] S. T. Chu, W. P. Huang, and S. K. Chaudhuri. Simulation and analysis of waveguide based optical integrated circuits. Computer Physics Communications, 68(1–3):451–484, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591902135.

Chipaux:1990:MPS

[Chi90] R. Chipaux. MOSPLV, a program for simulation of complex Mössbauer spectra in polycrystalline samples. Computer Physics Communications, 60(3):405–415, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590900372.

Chilingarian:1992:DAM

[Chi92] A. A. Chilingarian. Dimensionality analysis of multiparticle production at high energies. Computer Physics Communications, 69(2–3):347–359, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290173V.

Chen:1991:FCC

[CHM91] Li Jen Chen, Chin Kun Hu, and Kit Sing Mak. FOR-TRAN code for the cluster Monte Carlo study of the q-state Potts model on D-dimensional hypercubic lattices. Computer Physics Communications, 66(2–3):377–382, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190083W.

Cooper:1992:MEA

[CHMG92] W. A. Cooper, S. P. Hirshman, S. Merazzi, and R. Gruber. 3D magnetohydrodynamic equilibria with anisotropic pressure. Computer Physics Communications, 72(1):1–13, October 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290002G.

Christensen:1998:LST

[Chr98] Steven M. Christensen. Large scale tensor analysis by computer. Computer Physics Communications, 115(2–3):245– 263, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001428.

Christian:1999:PDM

[Chr99] Wolfgang Christian. Physlets: delivering media-focused problems anytime anywhere. Computer Physics Communications, 121–122:569–572, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004087.

Chen:1998:CPAb

[CHUS98] X. Chen, G. R. Harp, Y. Ueda, and D. K. Saldin. Computation of photoelectron and Auger-electron diffraction III. Evaluation of angle-resolved intensities PAD3. Computer Physics Communications, 112(1):91–101, July 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000320.

Champagne:1991:CCL

[CHW91] B. Champagne, W. Hereman, and P. Winternitz. The computer calculation of Lie point symmetries of large systems of differential equations. Computer Physics Communications, 66(2–3):319–340, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900805.

Collaboration:1998:CRA

[CI98] CMS Collaboration and Vincenzo Innocente. CMS reconstruction and analysis: an object oriented approach. Computer Physics Communications, 110(1–3):192–197, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001768.

Chu:1991:NMS

[CJ91] Shih-I Chu and Tsin-Fu Jiang. Nonperturbative methods for the study of multiphoton dynamics in intense periodic and multi-color laser fields. Computer Physics Communications, 63(1–3):482–493, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190271L.

Colonna:1994:OTS

[CJP^{+}94] François Colonna, Luc-Henri Jolly, Raymond A. Poirier, János G. Ángyán, and Georg Jansen. OSIPE — a tool for scientific programming in FORTRAN. Computer Physics Communications, 81(3):293–317, July 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900795.

Cerjan:1991:ETP

[CK91a] C. Cerjan and K. C. Kulander. Efficient time propagation for finite-difference representations of the timedependent Schrödinger equation. Computer Physics Communications, 63(1–3):529–537, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190274O.

Chow:1991:MDS

[CK91b] C. Y. Chow and K. H. Kao. Movement of the dividing streamline in unsteady stagnation-point flows. Computer Physics Communications, 65(1–3):69–75, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190156F.

Creswick:1999:MTM

[CK99] Richard J. Creswick and Seung-Yeon Kim. Microcanonical transfer matrix study of the Q-state Potts model. Computer Physics Communications, 121–122:26–29, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002714.

Chynoweth:1991:SOL

[CKS91] S. Chynoweth, U. C. Klomp, and L. E. Scales. Simulation of organic liquids using pseudo-pairwise interatomic forces on a toroidal transputer array. Computer Physics Communications, 62(2–3):297–306, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190102Q.

Collins:1998:EIM

[CKW98] L. A. Collins, J. D. Kress, and R. B. Walker. Excitation and ionization of molecules by a single-mode laser field using a time-dependent approach. Computer Physics Communications, 114(1–3):15–26, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000551.

Chang:1996:PBP

[CL96] Yia-Chung Chang and Guangwei Li. Planar-basis pseudopotential method and planar Wannier functions for surfaces and heterostructures. Computer Physics Communications, 95(2–3):158–170, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000562.

Clejan:1993:SPF

[Cle93] I. Clejan. A saddle point finding algorithm for functionals. Computer Physics Communications, 77(1):57–63, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390036C.

Cottrell:1998:WID

[CLM98] R. L. A. Cottrell, Connie A. Logg, and David E. Martin. What is the Internet doing? Performance and reliability monitoring for the HEP community. Computer Physics Communications, 110(1–3):142–148, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001689.

Chaturvedi:1992:ECE

[CM92a] S. Chaturvedi and R. G. Mills. ECRCYL: a code for electron cyclotron radiation transport. Computer Physics Communications, 70(1):183–206, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290100D.

Crothers:1992:IPE

[CM92b] D. S. F. Crothers and M. McCartney. ION — a program to evaluate cross-sections for ionisation in ion-atom collisions. Computer Physics Communications, 72(2–3):288–294, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290155R.

Carles:1998:CCC

[CM98]  A. Grau Carles and A. Grau Malonda.  CHEREN,

the Cherenkov counting efficiency. Computer Physics Communications, 111(1–3):258–264, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000393.

Carles:1994:ECE

[CMC94] Agustín Grau Carles, Agustín Grau Malonda, and Pilar Grau Carles. EMI, the counting efficiency for electron capture, electron capture-gamma and isomeric transitions. Computer Physics Communications, 79(1):115–123, February 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902348.

Creus:1999:VLI

[CMC^{+}99] A. Hernández Creus, H. Martín, P. Carro, R. C. Salvarezza, and A. J. Arvia. The validity of the linear interface motion equation containing a negative surface tension term. The dealloying of β-brass. Computer Physics Communications, 121–122:654, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700595.

Cadek:1992:SVA

[CMM92] Ondrej Cadek, Zdenek Martinec, and Ctirad Matyska. Spectral variational approach to the non-newtonian stokes problem in a spherical shell. Computer Physics Communications, 71(1–2):56–70, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900727.

Cacciari:1995:SMC

[CMNP95] Matteo Cacciari, Guido Montagna, Oreste Nicrosini, and Fulvio Piccinini. SABSPV — a Monte Carlo integrator for small-angle Bhabha scattering. Computer Physics Communications, 90(2–3):301–310, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500102L.

Charlton:1997:WMC

[CMNP97] David G. Charlton, Guido Montagna, Oreste Nicrosini, and Fulvio Piccinini. WWGENPV 2.0- A Monte Carlo event generator for four-fermion production at e^{+}e^{−} colliders. Computer Physics Communications, 99(2–3):355–370, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001221.

Charafi:1999:NMC

[CMSK99] M. M. Charafi, A. Menai, A. Sadok, and A. Kamal. Numerical modelling of change bottom elevation in a cohesive bed channel. Computer Physics Communications, 121–122:621, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700285.

Chessa:1999:ECS

[CMV99] Alessandro Chessa, Enzo Marinari, and Alessandro Vespignani. Energy constrained sandpile models. Computer Physics Communications, 121–122:622, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700297.

Chen:1998:SIL

[CNO98] J. Chen, N. Nakajima, and M. Okamoto. Shift-and-invert Lanczos algorithm for ideal MHD stability analysis. Computer Physics Communications, 113(1):1–9, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000782.

Castellano:1991:MNR

[CNPT91] M. Castellano, E. Nappi, F. Posa, and G. Tomasicchio. A multiresolution noise-removal algorithm for visual pattern recognition in imaging detectors. Computer Physics Communications, 66(2–3):293–307, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190078Y.

Cherubini:1992:LPN

[CO92] A. Cherubini and R. Odorico. LVQNET 1.10- a program for neural net and statistical pattern recognition. Computer Physics Communications, 72(2–3):249–264, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290152O.

Cohen:1992:CAQ

[Coh92] Eyal Cohen. Computer animations of quantum field theories. Computer Physics Communications, 70(3):441–446, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465592901047.

Cooper:1999:SSP

[Coo99] G. R. J. Cooper. State space processing of geophysical data. Computer Physics Communications, 121–122:623–624, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700303.

Copley:1991:MCC

[Cop91] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: modification to spherical geometry. Computer Physics Communications, 66(2–3):403–406, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190086Z.

Cornille:1999:SWN

[Cor99] H. Cornille. Shock-waves for nonconservative except mass models. Computer Physics Communications, 121–122:625– 626, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700315.

Craven:1991:MDR

[CP91] C. J. Craven and G. S. Pawley. Molecular dynamics of rigid polyatomic molecules on transputer arrays. Computer Physics Communications, 62(2–3):169–178, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190092Y.

Cerny:1992:CML

[CP92] R. Cerný and P. Prikryl. A computational model of laserinduced melting and solidification with density change. Computer Physics Communications, 73(1–3):179–191, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290038Z.

Chen:1995:NIS

[CP95] Zhuo-Min Chen and B. Montgomery Pettitt. Non-isotropic solution of an OZ equation: matrix methods for integral equations. Computer Physics Communications, 85(2):239–250, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400146S.

Chen:1999:PPP

[CP99] Jin-Quan Chen and Jia-Lung Ping. Program packages for point groups and space groups with subgroup chain symmetry adaptation. Computer Physics Communications, 120(1):71– 93, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002015.

Ciccotti:1999:WAS

[CPCF99] Giovanni Ciccotti, Carlo Pierleoni, Fabrizio Capuani, and Vladimir S. Filinov. Wigner approach to the semiclassical dynamics of a quantum many-body system: the dynamic scattering function of ^{4}He. Computer Physics Communications, 121–122:452–459, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003811.

Cannas:1999:MPS

[CPM99] Sergio A. Cannas, Sergio A. Páez, and Diana E. Marco. Modeling plant spread in forest ecology using cellular automata. Computer Physics Communications, 121–122:131– 135, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002970.

Clark:1998:APC

[CR98] Roger C. Clark and John S. Reid. ABSFAC: a program for the calculation of the absorption during scattering in multifaceted crystals and similar samples. Computer Physics Communications, 111(1–3):243–257, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000150.

Cordero:1999:DSG

[CR99] Patricio Cordero and Dino Risso. Dynamics of sheared gases. Computer Physics Communications, 121–122:225– 230, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003185.

Crampin:1998:BRT

[Cra98] S. Crampin. Book review: I. Turek, V. Drchal, J. Kudrnovský, M. ob, P. Weinberger, Electronic Structure of Disordered Alloys, Surfaces and Interfaces (1997) Kluwer Academic Press,Singapore 0-7923-9798-3. Computer Physics Communications, 111(1–3):274–275, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000368.

Cash:1990:FPN

[CRS90] J. R. Cash, A. D. Raptis, and T. E. Simos. A Fortran program for the numerical integration of the one-dimensional Schrödinger equation using exponential and Bessel fitting methods. Computer Physics Communications, 56(3):391–407, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090022S.

Clearwater:1991:RLP

[CS91] S. H. Clearwater and E. G. Stern. A rule-learning program in high energy physics event classification. Computer Physics Communications, 67(2):159–182, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190014C.

Chen:1998:ASI

[CS98a] L. Q. Chen and Jie Shen. Applications of semi-implicit fourier-spectral method to phase field equations. Computer Physics Communications, 108(2–3):147–158, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700115X.

Chen:1998:CPAa

[CS98b] X. Chen and D. K. Saldin. Computation of photoelectron and Auger-electron diffraction I. Preparation of input data for the cluster calculation PAD1. Computer Physics Communications, 112(1):67–79, July 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000307.

Carrillo:1999:ESV

[CS99a] José A. Carrillo and Juan Soler. On the evolution of a singular vortex patch in a two-dimensional incompressible fluid flow. Computer Physics Communications, 121–122:244– 250, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003227.

Coriano:1999:QEE

[CS99b] Claudio Corianò and Çetin Savkli. QCD evolution equations: Numerical algorithms from the Laguerre expansion. Computer Physics Communications, 118(2–3):236–258, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001581.

Cary:1997:CCF

[CSC^{+}97] John R. Cary, Svetlana G. Shasharina, Julian C. Cummings, John V. W. Reynders, and Paul J. Hinker. Comparison of C++ and Fortran 90 for object-oriented scientific programming. Computer Physics Communications, 105(1):20–36, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559700043X.

Clarke:1992:LSI

[CSP92] L. J. Clarke, I. Stich, and M. C. Payne. Large-scale ab initio total energy calculations on parallel computers. Computer Physics Communications, 72(1):14–28, October 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290003H.

Charchula:1994:CQQ

[CSS94a] K. Charchula, G. A. Schuler, and H. Spiesberger. Combined QED and QCD radiative effects in deep inelastic lepton-proton scattering: the Monte Carlo generator DJANGO6. Computer Physics Communications, 81(3):381– 402, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900868.

Cox:1994:ACA

[CSS94b] Hazel Cox, Stephen J. Smith, and Brian T. Sutcliffe. An application of the computer algebra system Maple for calculations on atomic and molecular systems. Computer Physics Communications, 84(1–3):186–200, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902119.

Cecchini:1990:SSM

[CT90] R. Cecchini and M. Tarlini. Symbolic superalgebra manipulations using COMMON LISP. Computer Physics Communications, 60(2):265–270, September 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090013Q.

Cheb-Terrab:1994:MPP

[CT94a] E. S. Cheb-Terrab. Maple procedures for partial and functional derivatives. Computer Physics Communications, 79 (3):409–424, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901856.

Christley:1994:CCR

[CT94b] J. A. Christley and I. J. Thompson. CRCWFN: coupled real Coulomb wavefunctions. Computer Physics Communications, 79(1):143–155, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902364.

Cheb-Terrab:1996:BRB

[CT96a] E. S. Cheb-Terrab. Book review: Classical mechanics with Maple: By R. L. Greene. Springer-Verlag, Berlin, 1995. Approx. 230 pages, 55 figures. Hardcover price DM 48,00. ISBN 0-387-94512-1. Computer Physics Communications, 97 (3):358, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000483.

Contini:1996:CPX

[CT96b] G. Contini and S. Turchini. CONTUR: a program for X-ray photoemission spectroscopic personal computer-based data analysis. Computer Physics Communications, 94(1):49–52, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001417.

Chassapis:1997:MOS

[CT97] Constantine S. Chassapis and Thomas Tsakalakos. Multidimensional optimization of a stochastic model for Xray diffraction from superlattices. Computer Physics Communications, 99(2–3):163–179, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001270.

Cheb-Terrab:1997:CAS

[CTDdM97] E. S. Cheb-Terrab, L. G. S. Duarte, and L. A. C. P. da Mota. Computer algebra solving of first order ODEs using symmetry methods. Computer Physics Communications, 101(3):254– 268, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000180.

Cheb-Terrab:1998:CAS

[CTDdM98] E. S. Cheb-Terrab, L. G. S. Duarte, and L. A. C. P. da Mota. Computer algebra solving of second order ODEs using symmetry methods. Computer Physics Communications, 108 (1):90–114, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559700132X.

Cheb-Terrab:1996:PSH

[CTdO96] E. S. Cheb-Terrab and H. P. de Oliveira. Poincaré sections of Hamiltonian systems. Computer Physics Communications, 95(2–3):171–189, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559600032X.

Cheb-Terrab:1995:SVE

[CTE95] E. S. Cheb-Terrab and A. G. Elfimov. The solution of Vlasov’s equation for complicated plasma geometry. I. spherical type. Computer Physics Communications, 85(2):251–266, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400144Q.

Chelikowsky:1995:APS

[CTJ^{+}95] James R. Chelikowsky, N. Troullier, X. Jing, D. Dean, N. Binggeli, K. Wu, and Y. Saad. Algorithms for predicting the structural properties of clusters. Computer Physics Communications, 85(3):325–335, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400147T.

Cheb-Terrab:1998:P

[CTM98] E. S. Cheb-Terrab and R. G. McLenaghan. Preface. Computer Physics Communications, 115(2–3):v, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598900893.

Cheb-Terrab:1998:SFO

[CTR98] E. S. Cheb-Terrab and A. D. Roche. Symmetries and first order ODE patterns. Computer Physics Communications, 113(2–3):239–260, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559800071X.

Campolongo:1999:TQL

[CTS99] F. Campolongo, S. Tarantola, and A. Saltelli. Tackling quantitatively large dimensionality problems. Computer Physics Communications, 117(1–2):75–85, March 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001659.

Cheb-Terrab:1995:CAA

[CTvB95] E. S. Cheb-Terrab and K. von Bülow. A computational approach for the analytical solving of partial differential equations. Computer Physics Communications, 90(1):102– 116, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500083R.

Chinn:1995:MNS

[CUVS95] C. R. Chinn, A. S. Umar, M. Vallières, and M. R. Strayer. Microscopic nuclear structure on a parallel platform. Computer Physics Communications, 86(1–2):40–60, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000062.

Cooke:1999:LPS

[CvN99] Roger M. Cooke and Jan M. van Noortwijk. Local probabilistic sensitivity measures for comparing FORM and Monte Carlo calculations illustrated with dike ring reliability calculations. Computer Physics Communications, 117(1–2):86–98, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001660.

Chessa:1999:CES

[CVZ99] Alessandro Chessa, Alessandro Vespignani, and Stefano Zapperi. Critical exponents in stochastic sandpile models. Computer Physics Communications, 121–122:299–302, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003380.

Chang:1992:GMC

[CW92] Chao-Hsi Chang and Jian-Xiong Wang. A general method for calculating principal value integrals numerically in an N-dimensional region. Computer Physics Communications, 69(2–3):330–338, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290171T.

Crockford:1993:SFP

[CY93] D. J. Crockford and W. Yeung. Susceptibilities for first principles band structures. Computer Physics Communications, 75(1–2):55–64, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593901648.

Charlesworth:1995:ARL

[CY95] J. P. A. Charlesworth and W. Yeung. An algebraic representation of the linear-analytic tetrahedron method. Computer Physics Communications, 88(2–3):186–194, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595000228.

Chernysheva:1999:TPP

[CY99] L. V. Chernysheva and V. L. Yakhontov. Two-program package to calculate the ground and excited state wave functions in the Hartree–Fock–Dirac approximation. Computer Physics Communications, 119(2–3):232–255, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001970.

Cybenko:1990:BRB

[Cyb90] G. Cybenko. Book review: Neural and synergetic computers: Hermann Haken, editor, Proceedings of the International Symposium at Schloss Elmau, Bavaria, 13-17 June 1988, Springer Verlag, Berlin, 263 pages. DM 88 (hard cover). ISBN 3-540-50339-0. Computer Physics Communications, 60(3):417, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465590900383.

Celik:1999:RPT

[ÇYG99] Tarik Çelik, Fatih Yasar, and Yigit Gündüç. Rounding of phase transition in the 2D 8-state random bond Potts model. Computer Physics Communications, 121–122:194– 196, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003112.

Cuperman:1992:IPC

[CZ92] S. Cuperman and D. Zoler. Implementation of PDETWO code to non-ideal MHD-like systems of equations: relaxation, accuracy and boundary conditions. Computer Physics Communications, 67(3):435–442, January 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290051Y.

Coisson:1993:SRN

[CZ93] R. Coïsson and D. Zavattaro. Solution of Raman–Nath equations. Computer Physics Communications, 76(2):231–234, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390134X.

Castellano:1999:CMS

[CZ99] Claudio Castellano and Marco Zannetti. Crossover from multiscaling to standard scaling in systems with conserved scalar order parameter. Computer Physics Communications, 121–122:317–320, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003434.

delAguila:1997:IAC

[dAASZ97] F. del Aguila, J. A. Aguilar-Saavedra, and M. Zralek. Invariant analysis of CP violation. Computer Physics Communications, 100(3):231–246, March 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001592.

Dennis:1999:NRS

[DATL99] Robin L. Dennis, J. R. Arnold, Gail S. Tonnesen, and Yonghong Li. A new response surface approach for interpreting Eulerian air quality model sensitivities. Computer Physics Communications, 117(1–2):99–112, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001672.

Dhaeseleer:1993:EPB

[DB93] William D. D’haeseleer and Craig D. Beidler. An efficient physics-based Monte Carlo algorithm for simulating alpha-particle orbits in tokamak fusion-reactor plasmas with toroidal-field ripple. Computer Physics Communications, 76(1):1–38, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390116T.

Domany:1999:SCD

[DBGW99] Eytan Domany, Marcelo Blatt, Yoram Gdalyahu, and Daphna Weinshall. Superparamagnetic clustering of data: application to computer vision. Computer Physics Communications, 121–122:5–12, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002672.

Dmitruk:1999:PDE

[DBGY99] I. Dmitruk, M. Bilyi, I. Gorban, and O. Yeshchenko. Properties of dielectric excitonic liquid in crystals. Computer Physics Communications, 121–122:628, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700339.

Dietz:1996:ABW

[DBME96] H. Dietz, M. Braun, C. Meier, and V. Engel. Addendum to: Nanosecond wave-packet propagation with the split-operator technique [Comput. Phys. Commun. 93 (1996) 152]. Computer Physics Communications, 98(1–2):265, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000665.

denBok:1993:GEP

[dBPB^{+}93] H. W. den Bok, Z. Papandreou, Th. S. Bauer, E. Voutier, and M. E. Brandan. A GEANT extension for polarized neutron-proton scattering. Computer Physics Communications, 74(3):375–380, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390020D.

DiMartino:1999:PPP

[DBVS99] B. Di Martino, S. Briguglio, G. Vlad, and P. Sguazzero. Parallel PIC plasma simulation through particle decomposition methods. Computer Physics Communications, 121–122: 696, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550670100X.

deJongh:1999:DMT

[dCdJDvL99] M. S. L. du Croo de Jongh, J. M. Doumen, and J. M. J. van Leeuwen. Density matrix technique for ground state calculations. Computer Physics Communications, 121–122: 483–488, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003872.

Droz:1999:KTD

[DCM99] M. Droz, Bastien Chopard, and Alexandre Masselot. Kinetics of two-dimensional ballistic annihilation: a multiparticle lattice gas study. Computer Physics Communications, 121–122:366–370, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003550.

DiMartino:1999:HPF

[DCR99] B. Di Martino, M. Celino, and V. Rosato. A[n] High Performance Fortran implementation of a Tight-Binding Molecular Dynamics simulation. Computer Physics Communications, 120(2–3):255–268, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002519.

deCordoba:1999:RPR

[dCRSF99] P. Fernández de Córdoba, L. M. García Raffi, J. M. Sanchis, and A. Ferrando. Routing problems revisited: path integral formalism and Monte Carlo methods. Computer Physics Communications, 121–122:634, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550670039X.

Duch:1994:NNT

[DD94] Wlodzisław Duch and Geerd H. F. Diercksen. Neural networks as tools to solve problems in physics and chemistry. Computer Physics Communications, 82(2–3):91–103, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901589.

Duch:1995:FSM

[DD95] Wlodzisław Duch and Geerd H. F. Diercksen. Feature space mapping as a universal adaptive system. Computer Physics Communications, 87(3):341–371, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000239.

Dobaczewski:1997:SSHa

[DD97a] J. Dobaczewski and J. Dudek. Solution of the Skyrme– Hartree–Fock equations in the Cartesian deformed harmonic oscillator basis I. The method. Computer Physics Communications, 102(1–3):166–182, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000040.

Dobaczewski:1997:SSHb

[DD97b] J. Dobaczewski and J. Dudek. Solution of the Skyrme– Hartree–Fock equations in the Cartesian deformed harmonic oscillator basis II. The program HFODD. Computer Physics Communications, 102(1–3):183–209, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000052.

Danese:1998:EPE

[DDDL98] G. Danese, I. De Lotto, D. Dotti, and F. Leporati. Ewald potentials evaluated through look-up tables. Computer Physics Communications, 108(2–3):211–217, February 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001409.

Decyk:1999:APM

[DDK99] Viktor K. Decyk, Dean E. Dauger, and Pieter R. Kokelaar. Appleseed: a parallel Macintosh cluster for numerically intensive computing. Computer Physics Communications, 121–122:627, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700327.

Duarte:1999:NAD

[DdMdO^{+}99] L. G. S. Duarte, L. A. C. P. da Mota, H. P. de Oliveira, R. O. Ramos, and J. E. F. Skea. Numerical analysis of dynamical systems and the fractal dimension of boundaries. Computer Physics Communications, 119(2–3):256–271, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002040.

Dateo:1991:NST

[DEAM91] Christopher E. Dateo, Volker Engel, Raphael Almeida, and Horia Metiu. Numerical solutions of the timedependent Schrödinger equation in spherical coordinates by Fourier transform methods. Computer Physics Communications, 63(1–3):435–445, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190268P.

Debenham:1992:BRB

[Deb92] R. Debenham. Book review: Models of neural networks: E. Domany, J. L. van Hemmen and K. Schulten, eds., Springer-Verlag, Berlin, 1991. 345 + xvi pages, 78 figures. Hardcover price DM78.00 ISBN 3-540-51109-1. Computer Physics Communications, 70(1):219–220, May 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901025.

Decyk:1995:SPC

[Dec95] Viktor K. Decyk. Skeleton PIC codes for parallel computers. Computer Physics Communications, 87(1–2):87–94, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594001693.

Degener:1999:TOO

[Deg99] Thomas F. Degener. TARA — an object-oriented program for a partial wave analysis of sequential two body decays. Computer Physics Communications, 118(1):34–48, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001945.

Demetriou:1997:CFS

[Dem97] I. C. Demetriou. CXFTV2: a Fortran subroutine for the discrete least squares convex approximation. Computer Physics Communications, 100(3):297–310, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001580.

Denby:1999:NNH

[Den99] Bruce Denby. Neural networks in high energy physics: a ten year perspective. Computer Physics Communications, 119(2–3):219–231, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001994.

Desalvo:1990:VCT

[DER90] A. Desalvo, G. Erbacci, and R. Rosa. Vectorized code for the three-dimensional spin-exchange kinetic Ising model on cubic and diamond lattices. Computer Physics Communications, 60(3):305–310, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090029Z.

Derreumaux:1999:IPP

[Der99] Philippe Derreumaux. Ab initio prediction of polypeptide structure from its sequence. Computer Physics Communications, 121–122:139–140, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002994.

Deutsch:1999:CMP

[Deu99] Hans-Peter Deutsch. Computational methods in the pricing and risk management of modern financial derivatives. Computer Physics Communications, 121–122:157– 160, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003033.

DeVries:1991:TEH

[DeV91] Paul L. DeVries. The time evolution of the hydrogen wavefunction in intense laser fields. Computer Physics Communications, 63(1–3):95–99, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190241C.

Dorfi:1995:ARH

[DF95] Ernst A. Dorfi and Michael U. Feuchtinger. Adaptive radiation hydrodynamics of pulsating stars. Computer Physics Communications, 89(1–3):69–90, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001877.

deFainchtein:1995:FES

[dFZLS95] Rosalinda de Fainchtein, Steven T. Zalesak, Rainald Löhner, and Daniel S. Spicer. Finite element simulation of a turbulent MHD system: comparison to a pseudo-spectral simulation. Computer Physics Communications, 86(1–2):25–39, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400143P.

DeAngelis:1993:MMH

[DG93] A. R. DeAngelis and D. H. E. Gross. A microcanonical model of hot nuclei. Computer Physics Communications, 76(1):113– 126, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390125V.

Dehnen:1999:HWI

[DG99] Richard Dehnen and Frank Grosßmann. Hydrogen wavefunction in intense laser fields: a unitary integrator for a high performance parallel computer. Computer Physics Communications, 120(1):33–40, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002052.

Diyankov:1997:MTD

[DGK97] Oleg V. Diyankov, Igor V. Glazyrin, and Serge V. Koshelev. MAG — two-dimensional resistive MHD code using an arbitrary moving coordinate system. Computer Physics Communications, 106(1–2):76–94, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000830.

Diggs:1999:BAS

[DGKS99] Brian Diggs, Christopher Genovese, Joseph B. Kadane, and Robert H. Swendsen. Bayesian analysis of series expansions. Computer Physics Communications, 121–122:1–4, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002660.

daGama:1999:SDF

[dGT99] M. M. Telo da Gama and J. M. Tavares. Strongly dipolar fluids: a theoretical and computational challenge. Computer Physics Communications, 121–122:256–258, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003240.

Dyadkin:1997:FEL

[DH97a] Iosif G. Dyadkin and Kenneth G. Hamilton. A family of enhanced Lehmer random number generators, with hyperplane suppression, and direct support for certain physical applications. Computer Physics Communications, 107(1–3):258–280, December 22, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.cpc.cs.

qub.ac.uk/cpc/;   http://www.cpc.cs.qub.ac.uk/cpc/
cgi-bin/list_summary.pl?CatNumber=ADGW;   http://www.
sciencedirect.com/science/article/pii/S001046559700101X.

Dyadkin:1997:SBM

[DH97b] Iosif G. Dyadkin and Kenneth G. Hamilton. A study of 64bit multipliers for Lehmer pseudorandom number generators. Computer Physics Communications, 103(2–3):103–130, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000520.

Desmond:1998:PLD

[DHK^{+}98] Edmond Desmond, John Haggerty, Hyon Joo Kehayias, Thomas Kozlowski, Martin L. Purschke, and Chris Witzig. PHENIX on-line distributed computing system architecture. Computer Physics Communications, 110(1–3):107–112, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001628.

DiPeso:1993:UBT

[DHL93] G. DiPeso, D. W. Hewett, and D. J. Larson. The use of block tridiagonal methods to improve convergence of DADI solutions to coupled elliptic partial differential equations. Computer Physics Communications, 77(1):33–45, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390034A.

DiMarzio:1995:CSP

[Di 95] Frank Di Marzio. Computational simplifications in pion and kaon induced nucleon knock-on reactions. Computer Physics Communications, 85(2):200–212, February 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400118L.

dInverno:1998:ASG

[d’I98] Ray A. d’Inverno. Applications of SHEEP in General Relativity. Computer Physics Communications, 115(2–3):330– 349, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001271.

Dikoussar:1994:APF

[Dik94] N. D. Dikoussar. Adaptive projective filters for track finding. Computer Physics Communications, 79(1):39–51, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902283.

Dikoussar:1997:FPU

[Dik97] N. D. Dikoussar. Function parametrization by using 4point transforms. Computer Physics Communications, 99 (2–3):235–254, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001105.

Ding:1991:CQP

[Din91] H. Q. Ding. Computing quark potential on a parallel supercomputer. Computer Physics Communications, 65(1–3):92– 99, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190159I.

Dubois:1991:SMI

[DJMT91] T. Dubois, F. Jauberteau, M. Marion, and R. Temam. Subgrid modelling and the interaction of small and large wavelengths in turbulent flows. Computer Physics Communications, 65(1–3):100–106, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190160M.

Dougar-Jabon:1999:SME

[DJUM99] V. D. Dougar-Jabon, A. M. Umnov, and F. A. Vivas Mejia. A 3D simulation model for electron cyclotron plasmas in a double cusp magnetic trap. Computer Physics Communications, 121–122:629, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700340.

Djouadi:1998:HPH

[DKS98] A. Djouadi, J. Kalinowski, and M. Spira. HDECAY: a program for Higgs boson decays in the Standard Model and its supersymmetric extension. Computer Physics Communications, 108(1):56–74, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001239.

Drchal:1996:ICE

[DKT96] V. Drchal, J. Kudrnovský, and I. Turek. Ab-initio calculations of the electronic and atomic structure of solids and their surfaces. Computer Physics Communications, 97(1–2):111–123, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000252.

Denby:1990:SPR

[DL90] Bruce Denby and Stephan L. Linn. Spatial pattern recognition in a high energy particle detector using a neural network algorithm. Computer Physics Communications, 56(3):293–297, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090015S.

DeGrand:1991:WFT

[DL91] Thomas A. DeGrand and Richard D. Loft. Wave-function tests for lattice QCD spectroscopy. Computer Physics Communications, 65(1–3):84–91, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190158H.

Delgaty:1998:PAI

[DL98] M. S. R. Delgaty and Kayll Lake. Physical acceptability of isolated, static, spherically symmetric, perfect fluid solutions of Einstein’s equations. Computer Physics Communications, 115(2–3):395–415, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001301.

Dasgupta:1996:QSF

[DLLR96] Indranil Dasgupta, Andrea Ruben Levi, Vittorio Lubicz, and Claudio Rebbi. QCDF90: a set of Fortran 90 modules for a high-level, efficient implementation of QCD simulations. Computer Physics Communications, 98(3):365–397, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596001038.

Drska:1990:TPP

[DLS90] Ladislav Drska, Richard Liska, and Milan Sinor. Two practical packages for computational physics — GCPM, RLFI. Computer Physics Communications, 61(1–2):225–230, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090120P.

delosSantos:1999:DLG

[dlSG99] F. de los Santos and P. L. Garrido. Driven lattice gases: new perspectives. Computer Physics Communications, 121–122:321–323, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003446.

delaVega:1998:DME

[dlVM98] JoséM. García de la Vega and Beatriz Miguel. Determination of momentum expectation values for polyatomic molecules. Computer Physics Communications, 109(1):34–46, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001343.

Dubois:1990:SEC

[DM90] Alain Dubois and Alfred Maquet. Subroutines for the evaluation of cross sections of one-photon radiative processes occuring in fast-electron H-atom collisions. Computer Physics Communications, 60(2):271–280, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090014R.

Drouffe:1991:MST

[DM91] J.-M. Drouffe and K. J. M. Moriarty. Microcanonical simulation for the three-state Potts model. Computer Physics Communications, 64(2):207–213, May 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190031F.

deMeyere:1994:CMD

[dM94] Arnout de Meyere. Computing the molecular distribution in ferroelectric liquid crystal displays. Computer Physics Communications, 79(3):353–363, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901813.

Dray:1998:FOE

[DM98] Tevian Dray and Corinne A. Manogue. Finding octonionic eigenvectors using Mathematica. Computer Physics Communications, 115(2–3):536–547, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800126X.

deMagalhaes:1999:LRI

[dMCT99] Aglaé C. N. de Magalhães, Sergio A. Cannas, and Francisco A. Tamarit. Long range interactions and nonextensivity in magnetic systems. Computer Physics Communications, 121–122:692, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700960.

Davis:1991:EDM

[DMLM91] Malcolm E. Davis, Jeffry D. Madura, Brock A. Luty, and J. Andrew McCammon. Electrostatics and diffusion of molecules in solution: simulations with the University of Houston Brownian Dynamics Program. Computer Physics Communications, 62(2–3):187–197, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900942.

Degtyarev:1997:KIM

[DMM^{+}97] L. Degtyarev, A. Martynov, S. Medvedev, F. Troyon, L. Villard, and R. Gruber. The KINX ideal MHD stability code for axisymmetric plasmas with separatrix. Computer Physics Communications, 103(1):10–27, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000374.

deMenezes:1999:CAI

[dMP99] Marcio Argollo de Menezes and T. J. P. Penna. Chaotic automata improving pattern recognition ability by attractor neural networks. Computer Physics Communications, 121–122:698, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506701023.

Daura:1999:PFS

[DMvG99] Xavier Daura, Alan E. Mark, and Wilfred F. van Gunsteren. Peptide folding simulations: no solvent required? Computer Physics Communications, 123(1–3):97–102, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002611.

Decyk:1998:HSI

[DNS98] Viktor K. Decyk, Charles D. Norton, and Bolesław K. Szymanski. How to support inheritance and run-time polymorphism in Fortran 90. Computer Physics Communications, 115(1):9–17, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001015.

Dumitru:1999:CAI

[DNSP99] Mihaela Dumitru, V. Ninulescu, P. E. Sterian, and M. Piscureanu. Computer aided instruction in laser dynamics. Computer Physics Communications, 121–122:583–585, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004117.

deOliveira:1999:EDM

[dO99] Paulo Murilo Castro de Oliveira. Energy degeneracies from microcanonical averages. Computer Physics Communications, 121–122:16–21, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002696.

Donnelly:1999:PPS

[Don99]  Denis Donnelly.  Physics pedagogical software: The

CIP software contest. Computer Physics Communications, 121–122:557–561, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004063.

Das:1991:GCM

[DP91] Nirod K. Das and David M. Pozar. A generalized CAD model for printed antennas and arrays with arbitrary multilayer geometries. Computer Physics Communications, 68(1–3):393–440, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591902113.

Dirl:1996:SPWa

[DPD96a] R. Dirl, K. Payer, and B. L. Davies. Symmetrized plane waves: 1. symmorphic space groups. Computer Physics Communications, 98(1–2):52–72, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000720.

Dirl:1996:SPWb

[DPD96b] R. Dirl, K. Payer, and B. L. Davies. Symmetrized plane waves: 2. symmetrized P m¯3m-plane waves. Computer Physics Communications, 98(1–2):73–82, October 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000732.

Dirl:1996:SPWc

[DPD96c] R. Dirl, K. Payer, and B. L. Davies. Symmetrized plane waves: 3. nonsymmorphic space groups. Computer Physics Communications, 98(1–2):83–97, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000744.

Dirl:1996:SPWd

[DPD96d] R. Dirl, K. Payer, and B. L. Davies. Symmetrized plane waves: 4. symmetrized P m¯3n-plane waves. Computer Physics Communications, 98(1–2):98–110, October 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000756.

Draper:1999:SPU

[DPP^{+}99] David Draper, Antonio Pereira, Pedro Prado, Andrea Saltelli, Ryan Cheal, Sonsoles Eguilior, Bruno Mendes, and Stefano Tarantola. Scenario and parametric uncertainty in GESAMAC: a methodological study in nuclear waste disposal risk assessment. Computer Physics Communications, 117(1–2):142–155, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001702.

Dyachenko:1990:PMT

[DPRS90] A. J. Dyachenko, A. N. Pushkarev, A. M. Rubenchik, and V. F. Shvets. A particle model for three-dimensional Langmuir collapse simulation. Computer Physics Communications, 60(2):239–245, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090010X.

deQueiroz:1999:NSD

[dQRdS99] Sergio L. A. de Queiroz, F. D. A. Aarão Reis, and R. R. dos Santos. Numerical studies of dilute Ising systems on strips. Computer Physics Communications, 121–122:210– 213, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559900315X.

Degrand:1990:CTD

[DR90] Thomas A. Degrand and Pietro Rossi. Conditioning techniques for dynamical fermions. Computer Physics Communications, 60(2):211–214, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090006M.

Ding:1990:MCP

[DS90] Linkai Ding and Evert Stenlund. A Monte Carlo program for nuclear collision geometry. Computer Physics Communications, 59(2):313–318, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590901809.

DiMarzio:1992:MSF

[DS92] Frank Di Marzio and Jakub Szajman. Mie scattering in the first-order corrected eikonal approximation. Computer Physics Communications, 70(2):297–304, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901955.

Drska:1996:OCE

[DS96] Ladislav Drska and Milan Sinor. Opacity calculations for extreme physical systems: code RACHEL. Computer Physics Communications, 97(1–2):163–174, August 2, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559600029X.

Dorso:1999:FHD

[DS99] C. O. Dorso and A. Strachan. Fragmentation of hot drops. Computer Physics Communications, 121–122:240– 243, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003215.

Danabasoglu:1991:TDS

[DSB91] G. Danabasoglu, A. Saati, and S. Biringen. Three-dimensional simulations of incompressible and compressible flow stability. Computer Physics Communications, 65(1–3):76–83, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190157G.

dosSantos:1999:HDI

[dSB99] Rita M. Zorzenon dos Santos and A. T. Bernardes. How does the immune network learn? Computer Physics Communications, 121–122:754, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701576.

Deitrick:1991:NME

[DSD91] G. L. Deitrick, L. E. Scriven, and H. T. Davis. A new method of error analysis for molecular simulations. Computer Physics Communications, 62(2–3):327–335, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190104S.

daSilva:1999:RAB

[dSTC99] Crisógono R. da Silva, Francisco A. Tamarit, and Evaldo M. F. Curado. Recognition ability in a biologically motivated neural network. Computer Physics Communications, 121–122:103–107, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002908.

daSilva:1999:SEM

[dSVTM99] Luciano R. da Silva, Raúl O. Vallejos, Constantino Tsallis, and Renio S. Mendes. Spectrum energy multifractality and equilibrium thermodynamics. Computer Physics Communications, 121–122:737, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701400.

duToit:1993:BFI

[dT93] C. F. du Toit. Bessel functions J_{n}(z) and Y_{n}(z) of integer order and complex argument. Computer Physics Communications, 78(1–2):181–189, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901534.

Duschl:1995:F

[DT95] Wolfgang J. Duschl and Werner M. Tscharnuter. Foreword. Computer Physics Communications, 89(1–3):xi, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595900616.

Du:1990:NCE

[Du90] M. L. Du. A note on Cooley’s energy correction formula. Computer Physics Communications, 61(3):294–296, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900442.

Du:1993:CME

[Du93a] M. L. Du. Computing matrix elements through inhomogeneous equations in multichannel problems. Computer Physics Communications, 77(2):229–240, October 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390006X.

Du:1993:ECS

[Du93b] M. L. Du. Eigenvalue corrector for solving bound states of multichannel Schrödinger equations. Computer Physics Communications, 76(1):39–47, June 1993. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046559390117U.

Dubal:1991:NSS

[Dub91a] Mark R. Dubal. Numerical simulations of special relativistic, magnetic gas flows. Computer Physics Communications, 64(2):221–234, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190033H.

Dubois:1991:FRE

[Dub91b] Alain Dubois. Fast routine for the evaluation of twocentre integrals in heavy-particle collisions. Computer Physics Communications, 64(2):300–310, May 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190039N.

Duch:1996:CPM

[Duc96] Wlodzisław Duch. Computational physics of the mind. Computer Physics Communications, 97(1–2):136–153, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000276.

Duff:1996:RFM

[Duf96] Iain S. Duff. A review of frontal methods for solving linear systems. Computer Physics Communications, 97(1–2):45–52, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000203.

Dunn:1996:CDM

[Dun96] Dennis Dunn. A Chebyshev differentiation method. Computer Physics Communications, 96(1):10–16, July 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000033.

deVries:1993:WLD

[dVdRL93] Pedro de Vries, Hans de Raedt, and Ad Lagendijk. Wave localization in disordered and fractal systems. Computer Physics Communications, 75(3):298–310, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390046F.

DeBruyne:1990:IMS

[DVH90] P. De Bruyne, M. Velli, and A. W. Hood. The ideal MHD stability of line-tied coronal loops: a truncated Fourier series approach. Computer Physics Communications, 59(1):55– 59, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090155T.

Dippel:1999:MDS

[DW99] S. Dippel and D. E. Wolf. Molecular dynamics simulations of granular chute flow. Computer Physics Communications, 121–122:284–289, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900332X.

Delisle:1991:CFE

[DWL91] Gilles Y. Delisle, Ke Li Wu, and John Litva. Coupled finite element and boundary element method in electromagnetics. Computer Physics Communications, 68(1–3):255–278, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190203W.

Eastwood:1995:BFE

[EABH95] J. W. Eastwood, W. Arter, N. J. Brealey, and R. W. Hockney. Body-fitted electromagnetic PIC software for use on parallel computers. Computer Physics Communications, 87(1–2):155– 178, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400165X.

Eastwood:1991:OVR

[Eas91a] James W. Eastwood. ORTHOVEC: version 2 of the REDUCE program for 3-D vector analysis in orthogonal curvilinear coordinates. Computer Physics Communications, 64(1):121–122, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190054O.

Eastwood:1991:VPE

[Eas91b] James W. Eastwood. The virtual particle electromagnetic particle-mesh method. Computer Physics Communications, 64(2):252–266, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190036K.

Eastwood:1993:BRB

[Eas93] J. W. Eastwood. Book review: Algebraic computing with Reduce: M. MacCallum and F. Wright. Oxford University Press, 1991. £15. ISBN 0-19-853443-4. Computer Physics Communications, 76(1):140, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390127X.

Elfimov:1993:VME

[ECT93] A. G. Elfimov and E. S. Cheb-Terrab. Vlasov–Maxwell equations for complicated plasma geometry. Computer Physics Communications, 77(3):357–373, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390182C.

Evans:1991:SSS

[EH91] M. W. Evans and D. M. Heyes. Simulation and symmetry of shear and elongational flow. Computer Physics Communications, 62(2–3):249–266, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591900986.

Ezawa:1992:BFR

[EHK^{+}92] Y. Ezawa, T. Hayashi, M. Kikugawa, J. Kodaira, T. Muta, R. Najima, J. Saito, S. Wakaizumi, T. Watanabe, T. Yano, and M. Yonezawa. Brown–Feynman reduction of one-loop Feynman diagrams to scalar integrals with orthonormal basis tensors. Computer Physics Communications, 69(1):15–45, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290125I.

El-Hajj:1993:SSS

[EHKK93] Ali El-Hajj, Hafez Kobeissi, and Mahmoud Korek. SSM: a set of subprograms for calculating eigenvalues for a diatomic molecule using a simplified shooting method. Computer Physics Communications, 74(2):297–302, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390099X.

Eickel:1990:LLS

[Eic90] Jürgen Eickel. Logical and layout structures of documents. Computer Physics Communications, 61(1–2):201–208, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090117J.

Einarsson:1990:AML

[Ein90] Bo Einarsson. Application of mixed language programming. Computer Physics Communications, 61(1–2):150–162, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090113F.

Eissner:1998:UDW

[Eis98] W. Eissner. The UCL distorted wave code. Computer Physics Communications, 114(1–3):295–341, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000824.

Everaers:1994:FGS

[EK94] Ralf Everaers and Kurt Kremer. A fast grid search algorithm for molecular dynamics simulations with short-range interactions. Computer Physics Communications, 81(1–2):19– 55, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594901090.

Engel:1991:EMU

[Eng91] Volker Engel. Excitation of molecules with ultrashort laser pulses: exact time-dependent quantum calculations. Computer Physics Communications, 63(1–3):228–242, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190252G.

Eremin:1994:SNS

[EOR94] Ju. A. Eremin, N. V. Orlov, and V. I. Rozenberg. Scattering by non-spherical particles. Computer Physics Communications, 79(2):201–214, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559490068X.

Ernenwein:1990:PSI

[ERB90] René Ernenwein, Marie-Madeleine Rohmer, and Marc Benard. A program system for ab initio mo calculations on vector and parallel processing machines I. evaluation of integrals. Computer Physics Communications, 58(3):305–328, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090066A.

Ernst:1998:CDC

[Ern98] Michael Ernst. Computing at DESY — current setup, trends and strategic directions. Computer Physics Communications, 110(1–3):12–17, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559700146X.

Elber:1995:MPS

[ERS^{+}95] Ron Elber, Adrian Roitberg, Carlos Simmerling, Robert Goldstein, Haiying Li, Gennady Verkhivker, Chen Keasar, Jing Zhang, and Alex Ulitsky. MOIL: a program for simulations of macromolecules. Computer Physics Communications, 91(1–3):159–189, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500047J.

Eberhard:1992:EMC

[ES92a] Philippe H. Eberhard and Olivier P. Schneider. Example of Monte Carlo integrals where reference functions are useful. Computer Physics Communications, 67(3):378–388, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465592900462.

Eberhard:1992:RFD

[ES92b] Philippe H. Eberhard and Olivier P. Schneider. Reference functions to decrease errors in Monte Carlo integrals. Computer Physics Communications, 67(3):363–377, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290045Z.

Esselink:1995:CAL

[Ess95] Klaas Esselink. A comparison of algorithms for long-range interactions. Computer Physics Communications, 87(3):375– 395, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500003X.

Evans:1990:IPA

[Eva90] M. W. Evans. On the isolation of possible artifacts due to cubic periodic boundary conditions. Computer Physics Communications, 59(3):495–497, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090090N.

Evans:1992:CTD

[Eva92] G. A. Evans. Computing time-dependent eddy currents in tokamaks. Computer Physics Communications, 69(2–3):243– 252, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290164T.

Fan:1990:EDS

[Fan90] G. Y. Fan. Electron diffraction simulation on micro-VAX II computers with the aid of an array processor. Computer Physics Communications, 59(2):429–438, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590901898.

Fang:1994:EMT

[Fan94] Hai-Ping Fang. An efficient method for treating conditional Monte Carlo simulation. Computer Physics Communications, 83(2–3):147–155, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900442.

Farantos:1998:PPC

[Far98] Stavros C. Farantos. POMULT: a program for computing periodic orbits in Hamiltonian systems based on multiple shooting algorithms. Computer Physics Communications, 108(2–3):240–258, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001318.

Fauquembergue:1991:CSI

[Fau91] R. Fauquembergue. Computer simulation of III-V MES-FET’s, MODFETT’s and MIS-like FET’s. Computer Physics Communications, 67(1):63–72, August 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591902216.

FroeseFischer:1993:PCA

[FB93] Charlotte Froese Fischer and Tomas Brage. A program for computing autoionization properties. Computer Physics Communications, 74(3):381–398, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593900214.

Franchini:1994:IAS

[FB94] A. Franchini and V. Bortolani. Inelastic he-atom scattering from metal surfaces. Computer Physics Communications, 80(1–3):90–118, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559490099X.

Faraudo:1999:NSA

[FB99] J. Faraudo and J. Bafaluy. Non-sequential adsorption of diffusing disks: an effective medium approach. Computer Physics Communications, 121–122:633, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700388.

Fall:1999:LWF

[FBB99] Caspar J. Fall, N. Binggeli, and A. Baldereschi. Local work function around sharp aluminum facet edges. Computer Physics Communications, 121–122:631, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700364.

Fulling:1998:CGG

[FBdC98] Stephen A. Fulling, Itshak Borosh, and Andrea da Conturbia. Cataloguing general graphs by point and line spectra. Computer Physics Communications, 115(2–3):93–112, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001155.

Fivaz:1998:FEA

[FBdR^{+}98] M. Fivaz, S. Brunner, G. de Ridder, O. Sauter, T. M. Tran, J. Vaclavik, L. Villard, and K. Appert. Finite element approach to global gyrokinetic particle-in-cell simulations using magnetic coordinates. Computer Physics Communications, 111(1–3):27–47, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559800023X.

Figueroa:1999:MCS

[FBGV99] Susana G. Figueroa, Felipe J. Blas, Keith E. Gubbins, and Lourdes F. Vega. Monte Carlo simulations of adsorption applied to process separation: a comparison with experimental results. Computer Physics Communications, 121–122:638, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700431.

Forrest:1990:PSD

[FBH90] B. M. Forrest, A. Baumgärtner, and D. W. Heerman. Parallel simulation of dense two-dimensional polymer systems. Computer Physics Communications, 59(3):455–462, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090087H.

Fernandez-Barbero:1999:EPS

[FBVdlN99] A. Fernández-Barbero, B. Vincent, and F. J. de las Nieves. Effect of the particle sign of charge on mesoscopic cluster fractal structures. Computer Physics Communications, 121–122:635, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700406.

Fischer:1999:HMC

[FCS99] A. Fischer, F. Cordes, and C. Schütte. Hybrid Monte Carlo with adaptive temperature choice: efficient conformational analysis of RNA. Computer Physics Communications, 121–122:37–39, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900274X.

Ferre:1997:LSM

[Fer97] Ricardo Ferré. Large scale micromagnetic calculations for finite and infinite 3D ferromagnetic systems using FFT. Computer Physics Communications, 105(2–3):169–186, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000726.

Fernandez:1999:AER

[Fer99] Julio F. Fernández. Algorithm for exponential random numbers. Computer Physics Communications, 121–122:78–82, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002854.

Fernandes:1997:HMC

[FF97a] Fernando M. S. Silva Fernandes and F. F. Martins Freitas. Hypervolume Monte Carlo method at constant pressure. Computer Physics Communications, 102(1–3):161–165, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000210.

Fritzsche:1997:RPR

[FF97b] S. Fritzsche and C. Froese Fischer. REOS — a program for relaxed-orbital oscillator strength calculations. Computer Physics Communications, 99(2–3):323–334, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001166.

Filho:1999:QMP

[FFB99] Tarcísio M. Rocha Filho, Annibal Figueiredo, and Léon Brenig. [QPSI] A Maple package for the determination of quasi-polynomial symmetries and invariants. Computer Physics Communications, 117(3):263–272, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001866.

Fischer:1996:PSP

[FFG^{+}96] S. Fischer, A. Frommer, U. Glässner, Th. Lippert, G. Ritzenhöfer, and K. Schilling. A parallel SSOR preconditioner for lattice QCD. Computer Physics Communications, 98(1–2):20–34, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000896.

Fruhwirth:1998:TEL

[FFS98] R. Frühwirth and S. Frühwirth-Schnatter. On the treatment of energy loss in track fitting. Computer Physics Communications, 110(1–3):80–86, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001574.

Fontenelle:1990:CMK

[FG90] Márcia T. Fontenelle and Jason A. C. Gallas. Constants of motion for the KdV and mKdV equations. Computer Physics Communications, 60(2):225–230, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090008O.

FroeseFischer:1991:PCL

[FG91] Charlotte Froese Fischer and Michel R. Godefroid. Programs for computing LS and LSJ transitions from MCHF wave functions. Computer Physics Communications, 64(3):501–519, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901417.

Fritzsche:1995:PCE

[FG95a] S. Fritzsche and I. P. Grant. A program for the complete expansion of jj-coupled symmetry functions into Slater determinants. Computer Physics Communications, 92(1):111–126, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500088W.

Fuentes:1995:EPT

[FG95b] Néstor O. Fuentes and Hebe O. Gavarini. ECMC, a portable two-dimensional code for plasma equilibrium computation on coaxial-multiple-coil systems. Computer Physics Communications, 90(1):169–188, September 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000084.

Fritzsche:1997:CRV

[FG97] S. Fritzsche and I. P. Grant. CESD97 — a revised version to expand jj-coupled symmetry functions into determinants. Computer Physics Communications, 103(2–3):277–286, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000350.

FroeseFischer:1991:PPA

[FGH91] Charlotte Froese Fischer, Michel R. Godefroid, and Alan Hibbert. A program for performing angular integrations for transition operators. Computer Physics Communications, 64(3):486–500, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190140G.

Friebel:1998:P

[FGH98] Wolfgang Friebel, Ulrich Gensch, and Lars Hagge. Preface. Computer Physics Communications, 110(1–3):xi–xii, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598800039.

Farto:1998:ASC

[FGM98] José-Miguel Farto, Ana-Belén Gon’Zalez, and Pablo Martin. An algorithm for the systematic construction of solutions to perturbed problems. Computer Physics Communications, 111(1–3):110–132, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559800037X.

Fernandez:1998:CAL

[FGR98] F. M. Fernández, R. Guardiola, and J. Ros. Computer algebra and large scale perturbation theory. Computer Physics Communications, 115(2–3):170–182, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001350.

Fachat:1997:IEB

[FH97] André Fachat and Karl Heinz Hoffmann. Implementation of Ensemble-Based Simulated Annealing with dynamic load balancing under MPI. Computer Physics Communications, 107(1–3):49–53, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000969.

Fujimoto:1998:UVE

[FHI^{+}98] J. Fujimoto, K. Hikasa, T. Ishikawa, M. Jimbo, T. Kaneko, K. Kato, S. Kawabata, T. Kon, M. Kuroda, Y. Kurihara, T. Munehisa, D. Perret-Gallix, Y. Shimizu, and H. Tanaka. S usy23 v2.0: An event generator for supersymmetric processes at e^{+}e^{−} colliders. Computer Physics Communications, 111(1–3):185–216, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000174.

Fang:1999:PMD

[FHSO99] Zhiwu Fang, A. D. J. Haymet, Wataru Shinoda, and Susumu Okazaki. Parallel molecular dynamics simulation: Implementation of PVM for a lipid membrane. Computer Physics Communications, 116(2–3):295–310, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000897.

Fijalkow:1999:NSVa

[Fij99a] Eric Fijalkow. A numerical solution to the Vlasov equation. Computer Physics Communications, 116(2–3):319–328, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001465. See erra-

tum [Fij00].

Fijalkow:1999:NSVb

[Fij99b] Eric Fijalkow. Numerical solution to the Vlasov equation: The 1D code. Computer Physics Communications, 116(2–3): 329–335, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001477.

Fijalkow:1999:NSVc

[Fij99c] Eric Fijalkow. Numerical solution to the Vlasov equation: The 2D code. Computer Physics Communications, 116(2–3): 336–344, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001489.

Fijalkow:2000:ENS

[Fij00] Eric Fijalkow. Erratum to “A numerical solution to the Vlasov equation” [Computer Physics Communications 116 (1999) 319–328]. Computer Physics Communications, 124(2–3):359, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004336. See [Fij99a].

Fujimoto:1997:GVF

[FIK^{+}97] J. Fujimoto, T. Ishikawa, T. Kaneko, K. Kato, S. Kawabata, Y. Kurihara, T. Munehisa, D. Perret-Gallix, Y. Shimizu, and H. Tanaka. grc4f v1.1: a four-fermion event generator for e^{+}e^{−} collisions. Computer Physics Communications, 100(1–2):128–156, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001269.

Fincham:1990:BRB

[Fin90] David Fincham. Book review: Computer simulation studies in condensed matter physics. Recent developments: D. P. Landau, K. K. Mon and H.-B. Schüttler, eds., Springer Proceedings in Physics, vol. 33, Springer-Verlag, Berlin, 1988, 233 + ix pages. DM98.00 (hardcover). ISBN 3-540-50449-4. Computer Physics Communications, 56(3):409–410, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090023T.

Fincham:1991:BRB

[Fin91] David Fincham. Book review: Computer simulation studies in condensed matter physics II: New directions: D. P. Landau, K. K. Mon and H.-B. Schüttler, eds., Springer Proceedings in Physics, vol. 45, Springer-Verlag, Berlin, 1990. 189 + viii pages. Hardcover price DM 89.00 ISBN: 3-540-51906-8 and 0-387-51906-8. Computer Physics Communications, 62(1):166–167, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591901299.

Fitch:1990:SNI

[Fit90] John Fitch. The symbolic-numeric interface. Computer Physics Communications, 61(1–2):22–33, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590901049.

Friedman:1991:NSC

[FJ91] R. S. Friedman and M. J. Jamieson. A note on solving coupled second-order linear differential equations. Computer Physics Communications, 62(1):53–57, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190120A.

FroeseFischer:1994:MCA

[FJ94] Charlotte Froese Fischer and Per Jönsson. MCHF calculations for atomic properties. Computer Physics Communications, 84(1–3):37–58, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559490202X.

Friedman:1995:JLD

[FJ95a] R. S. Friedman and M. J. Jamieson. On Johnson’s log-derivative algorithm for a single equation. Computer Physics Communications, 85(2):231–238, February 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400133M.

Friedman:1995:SPP

[FJ95b] Ronald S. Friedman and Michael J. Jamieson. Scattering by a pair of potential energy barriers: tests of numerical software. Computer Physics Communications, 85(3):382–388, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400148U.

Fleischer:1995:EGI

[FJKvO95] Jochem Fleischer, Fred Jegerlehner, Karol Kolodziej, and Geert Jan van Oldenborgh. EEWW: a generator for e^{+}e^{−} → W^{+}W^{−} including one-loop and leading photonic two-loop corrections. Computer Physics Communications, 85(1):29–39, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400113G.

Friedman:1990:NSC

[FJP90] R. S. Friedman, M. J. Jamieson, and S. C. Preston. On the numerical solution of coupled eigenvalue differential equations arising in molecular spectroscopy. Computer Physics Communications, 58(1–2):17–23, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090132K.

Fabianski:1999:CCS

[FKFE99] R. Fabiański, B. Kuchta, L. Firlej, and R. D. Etters. Computer calculations of a stability of low temperature phases of solid nitrogen. Computer Physics Communications, 121–122:630, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700352.

Ferry:1991:MDE

[FKKJ91] David K. Ferry, Alfred M. Kriman, Meng Jeng Kann, and Ravindra P. Joshi. Molecular dynamics extensions of Monte Carlo simulation in semiconductor device modeling. Computer Physics Communications, 67(1):119–134, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190225A.

Fruhwirth:1996:VRT

[FKMR96] R. Frühwirth, P. Kubinec, W. Mitaroff, and M. Regler. Vertex reconstruction and track bundling at the LEP collider using robust algorithms. Computer Physics Communications, 96(2–3):189–208, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000409.

FroeseFischer:1991:PGC

[FL91] Charlotte Froese Fischer and Bin Liu. A program to generate configuration-state lists. Computer Physics Communications, 64(3):406–416, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591901358.

Fletcher:1992:BRB

[Fle92] D. F. Fletcher. Book review: Computational techniques for fluid dynamics: C. A. J. Fletcher, Springer Series in Computational Physics, Springer-Verlag, Heidelberg, 2nd ed. 1991. Two volumes; 425 and 515 pages, respectively. Softcover price DM 65.00 and DM 75.00, respectively. ISBN 3-540-53058-4 and 3-540-53601-9. Computer Physics Communications, 70(1):221, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592901036.

Flocke:1997:SPP

[Flo97] Norbert Flocke. SPINSGA: a program package for the evaluation of properties of spin-12 Heisenberg systems using the symmetric group approach. Computer Physics Communications, 106(1–2):114–124, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000738.

Fack:1999:RDB

[FLV99] V. Fack, S. Lievens, and J. Van der Jeugt. On rotation distance between binary coupling trees and applications for 3njcoefficients. Computer Physics Communications, 119(2–3):99– 114, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002167.

Flybjerg:1990:BRB

[Fly90] Henrik Flybjerg. Book review: Quantum mechanics on the personal computer: Siegmund Brandt and Hans D. Dahmen, Springer-Verlag, Berlin, Heidelberg, New York, London, Paris, Tokyo, Hong Kong, 1989. 270 pages and 69 figures. Hardcover price DM98.00. ISBN 3-540-51541-0. Computer Physics Communications, 61(3):437–439, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900578.

Flyvbjerg:1993:BRB

[Fly93]  Henrik Flyvbjerg.  Book review: Quantum mechan-

ics on the personal computer: Sigmund Brandt and Hans D. Dhmen, Springer-Verlag, Berlin, Heidelberg, etc., 2nd edition, 1992. 269 pages and 69 figures. Hardcover price DM98.00. ISBN 3-540-55722-9. Computer Physics Communications, 77(2):300–301, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593900133.

Fletcher:1993:NSL

[FMH93] D. F. Fletcher, M. McCaughey, and R. W. Hall. Numerical simulation of a laminar jet flow: a comparison of three CFD models. Computer Physics Communications, 78(1–2):113–120, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901475.

Fettes:1997:PED

[FMH97] W. Fettes, I. Morgenstern, and T. Husslein. Parallelization of the exact diagonalization of the t- t ’-hubbard model. Computer Physics Communications, 106(1–2):1–9, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000970.

Font:1993:NSP

[FMIM93] JoséA. Font, JoséM. Martí, JoséM. Ibáñes, and Juan A. Miralles. Numerical simulations of 2D potential flows in general relativity. Computer Physics Communications, 75(1–2):31– 46, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593901626.

Ferrando:1999:CSE

[FMSdC99] A. Ferrando, J. J. Miret, E. Silvestre, and P. Fernández de Córdoba. Computer simulation of the electromagnetic propagation in optical fibers using a novel biorthogonalbasis model method. Computer Physics Communications, 121–122:637, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550670042X.

Fernandez-Nieves:1999:VPT

[FNFBdlN99] A. Fernández-Nieves, A. Fernández-Barbero, and F. J. de las Nieves. Volume phase transition in polymer networks. Computer Physics Communications, 121–122:636, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506700418.

Fogedby:1999:SNB

[Fog99] Hans Fogedby. Solitons in the noisy Burgers equation. Computer Physics Communications, 121–122:382–385, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003598.

Forcada:1991:PPB

[For91] Mikel L. Forcada. PLEC: a program for building, modelling and optimizing polypeptide structures. Computer Physics Communications, 64(1):131–139, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190056Q.

Foster:1998:SWN

[FP98] David Foster and Alberto Pace. Supporting windows NT as physics workstations and servers. Computer Physics Communications, 110(1–3):230–238, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001835.

Fack:1994:NEP

[FPV94] V. Fack, S. N. Pitre, and J. Van der Jeugt. New efficient programs to calculate general recoupling coefficients: Part I: Generation of a summation formula. Computer Physics Communications, 83(2–3):275–292, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900558.

Fack:1995:NEP

[FPV95] V. Fack, S. N. Pitre, and J. Van der Jeugt. New efficient programs to calculate general recoupling coefficients.

Part II: Evaluation of a summation formula. Computer Physics Communications, 86(1–2):105–122, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001796.

Fack:1997:CGR

[FPV97] V. Fack, S. N. Pitre, and J. Van der Jeugt. Calculation of general recoupling coefficients using graphical methods. Computer Physics Communications, 101(1–2):155–170, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001701.

Freire:1990:LBI

[FR90] Juan J. Freire and Antonio Rey. Lower bounds for the intrinsic viscosity of flexible polymers. comparison with Brownian dynamics simulation results for different types of chains. Computer Physics Communications, 61(3):297–303, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900453.

Fernandes:1995:HMM

[FR95] Fernando M. S. Silva Fernandes and J. P. Prates Ramalho. Hypervolumes in microcanonical Monte Carlo. Computer Physics Communications, 90(1):73–80, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500075Q.

Field:1996:BMC

[FR96] J. H. Field and T. Riemann. BHAGENE3, a Monte Carlo event generator for lepton pair production and wide-angle Bhabha scattering in e^{+}e^{−} collisions near the Z peak. Computer Physics Communications, 94(1):53–87, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500131X.

Fradkin:1992:SAC

[Fra92] M. A. Fradkin. Special algorithm for computer calculation of the point group invariants. Computer Physics Communications, 73(1–3):197–204, December 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900406.

Fritzsche:1997:MPC

[Fri97] S. Fritzsche. Maple procedures for the coupling of angular momenta I. data structures and numerical computations. Computer Physics Communications, 103(1):51–73, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000325.

FroeseFischer:1991:CIP

[Fro91a] Charlotte Froese Fischer. A configuration interaction program. Computer Physics Communications, 64(3):473–485, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190139C.

FroeseFischer:1991:GMC

[Fro91b] Charlotte Froese Fischer. A general multi-configuration Hartree–Fock program. Computer Physics Communications, 64(3):431–454, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190137A.

FroeseFischer:1991:MASa

[Fro91c] Charlotte Froese Fischer. The MCHF atomic-structure package. Computer Physics Communications, 64(3):369–398, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591901336.

FroeseFischer:1991:MASb

[Fro91d] Charlotte Froese Fischer. MCHF atomic-structure package: support libraries and utilities. Computer Physics Communications, 64(3):399–405, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591901347.

Fruhwirth:1993:SOS

[Frü93] Rudolf Frühwirth. Selection of optimal subsets of tracks with a feed-back neural network. Computer Physics Communications, 78(1–2):23–28, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901408.

Fruhwirth:1995:TFL

[Frü95] Rudolf Frühwirth. Track fitting with long-tailed noise: a Bayesian approach. Computer Physics Communications, 85 (2):189–199, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559400121H.

Fruhwirth:1997:TFN

[Frü97] R. Frühwirth. Track fitting with non-Gaussian noise. Computer Physics Communications, 100(1–2):1–16, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001555.

Faisal:1991:SIV

[FS91] F. H. M. Faisal and P. Scanzano. Solution of the initial-value problem in laser-atom interaction: a variational Fourier transform method. Computer Physics Communications, 63(1–3):265–278, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190255J.

Fang:1992:CPC

[FS92] D. F. Fang and J. F. Shriner, Jr. A computer program for the calculation of angular-momentum coupling coefficients. Computer Physics Communications, 70(1):147–153, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290097I.

Fruhwirth:1999:TFA

[FS99a] R. Frühwirth and A. Strandlie. Track fitting with ambiguities and noise: a study of elastic tracking and nonlinear filters. Computer Physics Communications, 120(2–3):197–214, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002313.

Fuchs:1999:IPE

[FS99b] Martin Fuchs and Matthias Scheffler. Ab initio pseudopotentials for electronic structure calculations of polyatomic systems using density-functional theory. Computer Physics Communications, 119(1):67–98, June 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800201X.

FroeseFischer:1993:PCI

[FSMVM93] Charlotte Froese Fischer, Lidia Smentek-Mielczarek, Nathalie Vaeck, and Gregory Miecznik. A program to compute isotope shifts in atomic spectra. Computer Physics Communications, 74(3):415–431, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593900236.

Fleischer:1992:SPC

[FT92] J. Fleischer and O. V. Tarasov. SHELL2: a package for the calculation of two-loop on-shell Feynman diagrams in FORM. Computer Physics Communications, 71(1–2):193–205, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290083B.

Fujino:1995:IFW

[FT95] Seiji Fujino and Toshiki Takeuchi. ILU factorization well suited to the vector processor using a variant of the 5-point difference scheme. Computer Physics Communications, 85 (3):371–381, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559400152R.

Fang:1999:FPS

[FT99a] Zhong Fang and Kiyoyuki Terakura. First-principles study on the neutral-to-ionic phase transition in charge-transfer complex tetrathiafulvalene- p-chloranil. Computer Physics Communications, 121–122:632, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700376.

Fiolhais:1999:VWV

[FT99b] Carlos Fiolhais and Jorge Alberto Trindade. Virtual water, a virtual reality project for learning physics and chemistry. Computer Physics Communications, 121–122:639, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700443.

Fujimura:1994:DFE

[Fuj94a]  Toichiro Fujimura.  A double finite element method

with accurate reflective boundary condition treatment for three-dimensional transport. Computer Physics Communications, 82(2–3):111–119, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901600.

Fujitsu:1994:OMO

[Fuj94b] Akira Fujitsu. ope.math: operator product expansions in free field realizations of conformal field theory. Computer Physics Communications, 79(1):78–99, February 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902313.

Fukuda:1994:AEW

[Fuk94] Hiroshi Fukuda. Asymptotic energy and wave function of one-electron molecular orbital. Computer Physics Communications, 78(3):291–300, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900086.

Frontera:1999:CSR

[FV99] Carlos Frontera and Eduard Vives. Computer studies of the 2D random field Ising model at T = 0. Computer Physics Communications, 121–122:188–190, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003094.

Fritzsche:1998:MPC

[FVGF98] S. Fritzsche, S. Varga, D. Geschke, and B. Fricke. Maple procedures for the coupling of angular momenta II. sum rule evaluation. Computer Physics Communications, 111(1–3):167– 184, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000216.

Fack:1992:PCR

[FVR92] V. Fack, J. Van der Jeugt, and K. Srinivasa Rao. Parallel computation of recoupling coefficients using transputers. Computer Physics Communications, 71(3):285–304, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290015Q.

Feng:1995:NST

[FWH95] Y. Feng, B. Wolle, and K. Hübner. New, simplified technique for calculating particle source rates due to neutral beam injection into tokamaks. Computer Physics Communications, 88(2–3):161–172, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000136.

Filatov:1992:UMF

[FZZ92] M. J. Filatov, I. L. Zilberberg, and G. M. Zhidomirov. On the use of a model function of Coulomb hole approximation for evaluation of the Coulomb integrals in semiempirical SCF MO methods. Computer Physics Communications, 73(1–3):192–196, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900392.

Geertsema:1990:MTA

[GA90] G. T. Geertsema and A. Achterberg. MHD turbulence and accretion disks. Computer Physics Communications, 59(1):145– 152, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090164V.

Galiano:1999:CAR

[Gal99] Gonzalo Galiano. Computing the atomic rearrangement pathways for pure electron nuclides capture by a five-shell model. Computer Physics Communications, 117(3):273–277, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599001897.

Gautschi:1993:CGF

[Gau93] Walter Gautschi. On the computation of generalized Fermi–Dirac and Bose–Einstein integrals. Computer Physics Communications, 74(2):233–238, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390093R.

Gan-Baruch:1997:TIF

[GBWES97] Zameret Gan-Baruch, Rudolf Wegmann, Aharon Eviatar, and Hermann U. Schmidt. A two-ion fluid model with gyrotropic pressures for the comet interaction with the solar wind. Computer Physics Communications, 103(2–3):131–144, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000337.

Gernoth:1995:MBA

[GC95] Klaus A. Gernoth and John W. Clark. A modified backpropagation algorithm for training neural networks on data with error bars. Computer Physics Communications, 88(1): 1–22, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595000125.

Garrote:1995:SPC

[GCP95] E. Garrote, R. Capote, and R. Pedrosa. Single-particle calculations in an axially deformed Woods–Saxon potential with cassinian ovals parametrization of the shape deformation. Computer Physics Communications, 92(2–3):267–276, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001172.

Grieves:1993:CCW

[GD93] Brian Grieves and Dennis Dunn. Calculation of continuum wave functions. Computer Physics Communications, 77(3):313–324, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390179G.

Goncalves:1997:PCH

[GDGR97] M. Gonçalves, S. B. Duarte, F. Garcia, and O. Rodriguez. PRESCOLD: Calculation of the half-life for alpha decay, cluster radioactivity and cold fission processes. Computer Physics Communications, 107(1–3):246–252, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001185.

Gavrilov:1994:NSM

[GDRR94] E. M. Gavrilov, D. F. DuBois, H. A. Rose, and A. M. Rubenchik. Numerical simulation of a model of resonance absorption for subpicosecond laser pulses. Computer Physics Communications, 81(1–2):65–73, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901112.

Gruner:1999:MCS

[GE99] M. E. Gruner and P. Entel. Monte Carlo simulations of magnetovolume effects in YMn_{2}. Computer Physics Communications, 121–122:646, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700510.

Geddes:1998:BCM

[Ged98] Neil Geddes. The babar computing model. Computer Physics Communications, 110(1–3):38–42, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001513.

Geiger:1997:VMS

[Gei97] Klaus Geiger. VNI 3.1 MC — simulation program to study high-energy particle collisions in QCD by space–time evolution of parton-cascades and parton-hadron conversion. Computer Physics Communications, 104(1–3):70–160, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000489.

Gensch:1998:OA

[Gen98] U. Gensch. Opening address. Computer Physics Communications, 110(1–3):xiii–xiv, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598800040.

Gaigalas:1996:EHP

[GF96] Gediminas Gaigalas and Charlotte Froese Fischer. Extension of the HF program to partially filled f-subshells. Computer Physics Communications, 98(1–2):255–264, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000926.

Gaspar:1998:SOO

[GF98] C. Gaspar and B. Franek. SMI++ — object-oriented framework for designing control systems for HEP experiments. Computer Physics Communications, 110(1–3):87–90, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001586.

Gonzalez-Fernandez:1998:DST

[GFAALS98] C. F. González-Fernández, F. Alhama, M. Alarcón, and J. F. López-Sánchez. Digital simulation of transient heat conduction with polynomial variable thermal conductivity and specific heat. Computer Physics Communications, 111(1–3):53– 58, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000265.

Gonzalez-Ferez:1999:RLA

[GFS99] R. González-Férez and W. Schweizer. Resonances of light atoms in electric and magnetic fields. Computer Physics Communications, 121–122:644, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700492.

Goldfield:1996:MCC

[GG96] E. M. Goldfield and S. K. Gray. Mapping coriolis-coupled quantum dynamics onto parallel computer architectures. Computer Physics Communications, 98(1–2):1–14, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000914.

Goodrich:1991:TAF

[GGH91] John W. Goodrich, Karl Gustafson, and Kadosa Halasi. Time-asymptotic flow calculation. Computer Physics Communications, 65(1–3):107–116, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190161D.

Garrido:1997:OPE

[GGJG97] Lluís Garrido, Sergio Gómez, Aurelio Juste, and Vicens Gaitan. Optimal projection to estimate the proportions of the different subsamples in a given mixture sample. Computer Physics Communications, 104(1–3):37–45, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000556.

Gullenstern:1995:MCP

[GGMS95] Stefan Güllenstern, Pawel Górnicki, Lech Mankiewicz, and Andreas Schäfer. Monte Carlo program for polarised nucleonnucleon collisions. Computer Physics Communications, 87 (3):416–431, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595000147.

Garrido:1994:TAB

[GGSR94] Lluís Garrido, Vicens Gaitan, and Miguel Serra-Ricart. Test of agreement between two multidimensional empirical distributions. Computer Physics Communications, 84(1–3):297–306, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559490216X.

Gauss:1991:TDS

[GH91a] Jürgen Gauss and Eric J. Heller. A time-dependent semiclassical wavepacket method using a fast Fourier transform (FFT) algorithm. Computer Physics Communications, 63 (1–3):375–388, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190264L.

Godin:1991:BRL

[GH91b] T. J. Godin and Roger Haydock. The block recursion library: accurate calculation of resolvent submatrices using the block recursion method. Computer Physics Communications, 64(1):123–130, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190055P.

Gyulassy:1991:ETN

[GH91c] Miklos Gyulassy and Magnus Harlander. Elastic tracking and neural network algorithms for complex pattern recognition. Computer Physics Communications, 66(1):31–46, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591900056.

Grobe:1999:SDA

[GHE99] R. Grobe, S. L. Haan, and J. H. Eberly. A split-domain algorithm for time-dependent multi-electron wave functions. Computer Physics Communications, 117(3):200–210, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559800112X.

Goringe:1997:LSD

[GHGB97] C. M. Goringe, E. Hernández, M. J. Gillan, and I. J. Bush. Linear-scaling DFT-pseudopotential calculations on parallel computers. Computer Physics Communications, 102(1–3):1– 16, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000295.

Ghosh:1993:ASM

[GHM93] Sanjoy Ghosh, Murshed Hossain, and William H. Matthaeus. The application of spectral methods in simulating compressible fluid and magnetofluid turbulence. Computer Physics Communications, 74(1):18–40, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390103J.

Gibbon:1992:NMP

[Gib92] Paul Gibbon. A numerical model of the plasma beat-wave accelerator. Computer Physics Communications, 69(2–3): 299–305, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290168X.

Gielen:1990:BRB

[Gie90] Stan Gielen. Book review: Cellular automata and modelling of complex physical systems: P. Manneville, N. Boccara, G. Y. Vichniac and R. Bidaux, eds., Springer Proceedings in Physics, vol. 46, Springer-Verlag, Berlin, 1990. 319 + ix pages. Hardcover price DM 96.00. ISBN 3-540-51933-5. Computer Physics Communications, 61(3):433–435, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590900556.

Gilfoyle:1999:NTA

[Gil99] G. P. Gilfoyle. A new teaching approach to quantum mechanical tunneling. Computer Physics Communications, 121–122:573–577, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004099.

Gluck:1996:MCT

[GJ96] F. Glück and I. Joó. Monte Carlo type radiative correction calculations for neutron, muon and hyperon semileptonic decays. Computer Physics Communications, 95(2–3):111–122, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559600015X.

Gluck:1997:HPB

[GJ97] F. Glück and I. Joó. Hard photon Bremsstrahlung effects in hyperon semileptonic decays. Computer Physics Communications, 107(1–3):92–112, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001100.

Garrido:1998:DPD

[GJ98] Lluís Garrido and Aurelio Juste. On the determination of probability density functions by using Neural Networks. Computer Physics Communications, 115(1):25–31, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001076.

Gronbech-Jensen:1999:SLI

[GJ99a] Niels Grønbech-Jensen. Summation of logarithmic interactions in nonrectangular periodic media. Computer Physics Communications, 119(2–3):115–121, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002003.

Guerin:1999:GMH

[GJ99b] S. Guérin and H. R. Jauslin. Grid methods and Hilbert space basis for simulations of quantum dynamics. Computer Physics Communications, 121–122:496–498, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003902.

Goossens:1990:P

[GK90] Marcel Goossens and Wolfgang Kerner. Preface. Computer Physics Communications, 59(1):xi, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090149U.

Gelfgat:1993:PSRa

[GKP93a] V. I. Gelfgat, E. L. Kosarev, and E. R. Podolyak. Programs for signal recovery from noisy data using the maximum likelihood principle: I. general description. Computer Physics Communications, 74(3):335–348, March 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900177.

Gelfgat:1993:PSRb

[GKP93b] V. I. Gelfgat, E. L. Kosarev, and E. R. Podolyak. Programs for signal recovery from noisy data using the maximum likelihood principle: II. program implementation. Computer Physics Communications, 74(3):349–357, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900188.

Gustafson:1991:CDA

[GL91] Karl Gustafson and Robert Leben. Computation of dragonfly aerodynamics. Computer Physics Communications, 65 (1–3):121–132, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190163F.

Garrett:1998:APC

[GLAT98] Bruce C. Garrett, Gillian C. Lynch, Thomas C. Allison, and Donald G. Truhlar. ABCRATE: a program for the calculation of atom-diatom reaction rates. Computer Physics Communications, 109(1):47–54, March 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001379.

Gontchar:1997:CCC

[GLF97] I. Gontchar, L. A. Litnevsky, and P. Fröbrich. A Ccode for combining a Langevin fission dynamics of hot nuclei with a statistical model including evaporation of light particles and giant dipole γ-quanta. Computer Physics Communications, 107(1–3):223–245, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001082.

Gliss:1990:DUI

[Gli90] B. Gliss. Database user interfaces in a scientific environment. Computer Physics Communications, 61(1–2):93–102, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090108D.

Goplen:1995:UCM

[GLSW95] Bruce Goplen, Larry Ludeking, David Smith, and Gary Warren. User-configurable MAGIC for electromagnetic PIC calculations. Computer Physics Communications, 87(1–2):54–86, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500010D.

Gluck:1997:ORC

[Glü97] F. Glück. Order-α radiative correction calculations for unoriented allowed nuclear, neutron and pion β decays. Computer Physics Communications, 101(3):223–231, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001683.

Ginovart:1999:DSB

[GLV99] M. Ginovart, D. López, and J. Valls. Discrete simulation of biological systems based on individual and environment models. Computer Physics Communications, 121–122:642, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700479.

Gravielle:1992:SNI

[GM92] M. S. Gravielle and J. E. Miraglia. Some Nordsieck integrals of interest in radiation and atomic collision theories. Computer Physics Communications, 69(1):53–58, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290127K.

Gonzalez-Miranda:1999:BGS

[GM99] J. M. González-Miranda. Bistable generalized synchronization of chaotic systems. Computer Physics Communications, 121–122:429–431, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003756.

Galvez:1998:NVC

[GN98] P. Galvez and H. Newman. Networking, videoconferencing and collaborative environments. Computer Physics Communications, 110(1–3):43–50, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001525.

Goedbloed:1990:SSC

[Goe90] J. P. Goedbloed. Stability of solar coronal loops. Computer Physics Communications, 59(1):39–53, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090154S.

Goedecker:1993:RTD

[Goe93] S. Goedecker. Rotating a three-dimensional array in an optimal position for vector processing: case study for a three-dimensional fast Fourier transform. Computer Physics Communications, 76(3):294–300, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390057J.

Goedecker:1999:CMM

[Goe99] S. Goedecker. A comparison of major O(N) methods. Computer Physics Communications, 121–122:643, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506700480.

Gozani:1991:TSE

[Goz91] Joseph Gozani. Two-scale expansion of wave propagation in a random medium. Computer Physics Communications, 65(1–3):117–120, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190162E.

Gancza:1995:PFP

[GP95] W. M. Gańcza and T. Paszkiewicz. Phonon focussing patterns: Monte Carlo simulation of the motion of ballistic phonon beams in cubic crystals. Computer Physics Communications, 85(3):423–436, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400149V.

Gonzalez:1999:RNG

[GP99] Jorge A. González and Ramiro Pino. A random number generator based on unpredictable chaotic functions. Computer Physics Communications, 120(2–3):109–114, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002337.

Guilbert:1995:MET

[GPM95] Christophe Guilbert, David Perahia, and Liliane Mouawad. A method to explore transition paths in macromolecules. applications to hemoglobin and phosphoglycerate kinase. Computer Physics Communications, 91(1–3):263–273, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500052H.

Gorczyca:1995:AAC

[GPSM95] T. W. Gorczyca, M. S. Pindzola, F. S. Shieh, and C. L. McCreary. Adaptation of asymptotic close-coupling methods to massively parallel computers. Computer Physics Communications, 88(2–3):211–216, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500068Q.

Goodfellow:1995:NMA

[GPSW95] Julia M. Goodfellow, William R. Pitt, Oliver S. Smart, and Mark A. Williams. New methods for the analysis of the protein-solvent interface. Computer Physics Communications, 91(1–3):321–329, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500056L.

Graudenz:1992:CLT

[Gra92] Dirk Graudenz. Calculation of long traces of γ-matrices in the dimensional regularization scheme. Computer Physics Communications, 69(1):173–181, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290137N.

Grant:1994:RCA

[Gra94] Ian P. Grant. Relativistic calculation of atomic properties. Computer Physics Communications, 84(1–3):59–77, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902038.

Graudenz:1995:PJC

[Gra95] Dirk Graudenz. PROJET: Jet cross sections in deeply inelastic electron proton scattering. Version 4.1. Computer Physics Communications, 92(1):65–89, November 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500097Y.

Gracey:1998:CPR

[Gra98] J. A. Gracey. Computation of perturbative renormalization group functions — the large N_{f} algorithm. Computer Physics Communications, 115(2–3):113–123, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001167.

Grinin:1997:IAF

[Gri97] S. V. Grinin. Integer-arithmetic FDTD codes for computer simulation of internal, near and far electromagnetic fields scattered by three-dimensional conductive complicated form bodies. Computer Physics Communications, 102(1–3):109–131, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000143.

Garcia:1999:BCC

[GRM^{+}99] F. Garcia, O. Rodriguez, J. Mesa, J. D. T. Arruda-Neto, V. P. Likhachev, E. Garrote, R. Capote, and F. Guzmán.

BARRIER code: Calculation of fission barriers. Computer Physics Communications, 120(1):57–70, July 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900199X.

Grotendorst:1990:AFM

[Gro90] J. Grotendorst. Approximating functions by means of symbolic computation and a general extrapolation method. Computer Physics Communications, 59(2):289–301, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590901784.

Grotendorst:1991:MPT

[Gro91] J. Grotendorst. A Maple package for transforming series, sequences and functions. Computer Physics Communications, 67(2):325–342, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190026H.

Garcia:1995:DCL

[GRRG95] F. Garcia, O. Rodriguez, V. A. Rubchenya, and E. Garrote. DENCOM: Code for level density calculations of deformed nuclei using a combined method. Computer Physics Communications, 86(1–2):129–146, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400137Q.

Grosz:1990:PCM

[GRS90] T. Grosz, Gy. Rubin, and A. Sandor. PC controlled multi-dimensional data-collection system for a double time of flight (TOF) neutron spectrometer. Computer Physics Communications, 61(1–2):234–238, November 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090122H.

Gil:1997:ELF

[GS97] A. Gil and J. Segura. Evaluation of Legendre functions of argument greater than one. Computer Physics Communications, 105(2–3):273–283, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000763.

Gil:1998:CEP

[GS98a] A. Gil and J. Segura. A code to evaluate prolate and oblate spheroidal harmonics. Computer Physics Communications, 108(2–3):267–278, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001264.

Gomez:1998:MPG

[GS98b] Claude Gomez and Tony Scott. Maple programs for generating efficient FORTRAN code for serial and vectorised machines. Computer Physics Communications, 115(2–3): 548–562, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001143.

Gianturco:1997:QCT

[GSS97] F. A. Gianturco, S. Serna, and A. V. Storozhev. Quantum calculations of transport properties in molecular gases. Computer Physics Communications, 103(2–3):251–276, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001506.

Gallego:1999:DGCa

[GST99a] R. Gallego, M. San Miguel, and R. Toral. Domain growth and coarsening inhibition in a non-potential system. Computer Physics Communications, 121–122:640, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700455.

Gallego:1999:DGCb

[GST99b] R. Gallego, M. San Miguel, and R. Toral. Domain growth and coarsening inhibition in a nonpotential system. Computer Physics Communications, 121–122:324–326, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003458.

Gupta:1992:CAM

[Gup92] Sumnesh Gupta. Computing aspects of molecular dynamics simulation. Computer Physics Communications, 70(2):243– 270, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290191Z.

Gustafson:1991:SDS

[Gus91] Karl Gustafson. Semiconductor device simulation. Computer Physics Communications, 65(1–3):133–136, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190164G.

Gutowski:1992:LDR

[Gut92] Marek W. Gutowski. Off-line data reduction. Computer Physics Communications, 73(1–3):209–216, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290042W.

Gutbrod:1995:FRN

[Gut95] F. Gutbrod. A fast random number generator for the Intel Paragon supercomputer. Computer Physics Communications, 87(3):291–306, June 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500005Z.

Gagliardi:1998:LSS

[GUW98] Fabrizio Gagliardi, Akira Ukawa, and Peter Wegner. Large systems and specific solutions: Summary of CHEP97 Session F. Computer Physics Communications, 110(1–3):76–79, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001562.

Genchev:1999:ECN

[GV99] Z. D. Genchev and V. I. Vaskivskii. Effect of coloured noise on an overdamped Josephson junction. Computer Physics Communications, 121–122:641, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700467.

Gustafson:1991:P

[GW91] Karl Gustafson and Walter Wyss. Preface. Computer Physics Communications, 65(1–3):xi, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190146C.

Gyulassy:1994:HMC

[GW94] Miklos Gyulassy and Xin-Nian Wang. HIJING 1.0: a Monte Carlo program for parton and particle production in high energy hadronic and nuclear collisions. Computer Physics Communications, 83(2–3):307–331, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900574.

Gibbs:1995:HPS

[GW95] M. J. Gibbs and B. R. Webber. HERBVI — a program for simulation of baryon and lepton number violating processes. Computer Physics Communications, 90(2–3):369–380, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500073O.

Grafton:1998:VCP

[GW98] Anthony K. Grafton and Ralph A. Wheeler. ViPA: a computer program for vector projection analysis of normal vibrational modes of molecules. Computer Physics Communications, 113(1):78–84, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000381.

Grubert:1999:MMC

[GY99] D. Grubert and J. M. Yeomans. Mesoscale modeling of contact line dynamics. Computer Physics Communications, 121–122:236–239, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003203.

Holian:1999:TOO

[HAC^{+}99] K. S. Holian, L. A. Ankeny, S. P. Clancy, J. H. Hall, S. R. Lee, J. C. Marshall, G. R. McNamara, J. W. Painter, and M. E. Zander. Tecolote: an object-oriented framework for hydrodynamics applications. Computer Physics Communications, 121–122:658, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700637.

Hall:1990:GPI

[Hal90] B. D. Hall. A general-purpose interface system for the laboratory. Computer Physics Communications, 61(1–2):239–245, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090123I.

Hamilton:1993:PNGa

[Ham93a] Kenneth G. Hamilton. Pseudorandom number generators for personal computers. Computer Physics Communications, 75(1–2):105–117, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390168C.

Hamilton:1993:PNGb

[Ham93b] Kenneth G. Hamilton. Pseudorandom number generators for personal computers II. Computer Physics Communications, 78(1–2):172–180, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901523.

Hamilton:1994:PNG

[Ham94] Kenneth G. Hamilton. Pseudorandom number generators for Salford FTN77. Computer Physics Communications, 81 (1–2):237–247, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901228.

Hamilton:1995:EBU

[Ham95a] Kenneth G. Hamilton. Erratum: A universal GFSR random number generator for personal computers [Comput. Phys. Commun. 85 (1995) 127–152]. Computer Physics Communications, 86(1–2):208, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500024A. See [Ham95b].

Hamilton:1995:UGR

[Ham95b] Kenneth G. Hamilton. A universal GFSR random number generator for personal computers. Computer Physics Communications, 85(1):127–152, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400114H. See erratum [Ham95a].

Hamilton:1997:ARP

[Ham97] Kenneth G. Hamilton. Assembler RANLUX for PCs. Computer Physics Communications, 101(3):249–253, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000179.

Hansen:1990:GSC

[Han90] J. P. Hansen. General subroutines for the calculation of atomic and molecular two-centre integrals. Computer Physics Communications, 58(1–2):217–221, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090147S.

Hansmann:1999:RRG

[Han99] Ulrich H. E. Hansmann. Recent results from generalizedensemble protein-folding simulations. Computer Physics Communications, 121–122:129–130, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002969.

Herbst:1991:NHI

[HAR91] B. M. Herbst, Mark J. Ablowitz, and E. Ryan. Numerical homoclinic instabilities and the complex modified Korteweg–de Vries equation. Computer Physics Communications, 65(1–3):137–142, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190165H.

Hartley:1997:ERP

[Har97] David Hartley. EDS: a REDUCE package for exterior differential systems. Computer Physics Communications, 100(1–2):177–194, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001385.

Hasegawa:1999:NSB

[Has99] M. Hasegawa. Numerical study on broken ergodicity in triangular billiards. Computer Physics Communications, 121–122:649, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700546.

Horacek:1990:MCG

[HB90] Jirí Horácek and Jirí Bok. K-matrix calculation for general nonlocal potentials. Computer Physics Communications, 59(2):319–323, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090181Y.

Herman:1999:ETA

[HB99a] P. Herman and I. Barvík. The exciton transfer in the antenna systems of rhodopseudomonas acidophila. Computer Physics Communications, 121–122:652, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700571.

Hrach:1999:NMM

[HB99b] R. Hrach and J. Boldys. New methods in the morphological analysis of thin metal films. Computer Physics Communications, 121–122:660, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700650.

Higman:1991:ETS

[HBH91] J. M. Higman, J. Bude, and K. Hess. Electronic transport in semiconductors at high energy. Computer Physics Communications, 67(1):93–104, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591902238.

Hirabayashi:1999:LBM

[HCO99] Miki Hirabayashi, Yu Chen, and Hirotada Ohashi. Lattice BGK models for the Poisson and KPZ equations. Computer Physics Communications, 121–122:656, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700613.

Hernandez-Cobos:1999:DFE

[HCVM99] J. Hernández-Cobos, L. F. Vega, and Allan D. Mackie. Determination of the free energy of water at room temperature by parallel grand canonical Monte Carlo. Computer Physics Communications, 121–122:653, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700583.

Hansen:1992:PAN

[HD92] Jan P. Hansen and Alain Dubois. Procedures for analytical and numerical calculation of Coulombic one- and two-centre integrals. Computer Physics Communications, 67(3):456–464, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465592900543.

Heck:1996:TRC

[HD96] E. L. Heck and A. S. Dickinson. Transport and relaxation cross-sections for pure gases of linear molecules. Computer Physics Communications, 95(2–3):190–220, June 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000331.

Heuer:1997:CCS

[HDF97] Andreas Heuer, Burkhard Dünweg, and Alan M. Ferrenberg. Considerations on correlations in shift-register pseudorandom number generators and their removal. Computer Physics Communications, 103(1):1–9, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700026X.

Hsieh:1991:NAR

[HE91] H. C. Hsieh and W. C. Ermler. Numerical and analytical representations of Eulerian angle matrices for symmetric top molecules. Computer Physics Communications, 66(2–3):266–270, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190075V.

Hagler:1994:UQE

[HE94] A. T. Hagler and C. S. Ewig. On the use of quantum energy surfaces in the derivation of molecular force fields. Computer Physics Communications, 84(1–3):131–155, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902089.

Hrach:1999:VAC

[HE99] R. Hrach and M. Entlicher. Various approaches in computer modelling of plasma-surface interaction. Computer Physics Communications, 121–122:661, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700662.

Heather:1991:AWS

[Hea91] Robert W. Heather. An asymptotic wavefunction splitting procedure for propagating spatially extended wavefunctions: application to intense field photodissociation of H_{2}^{+}. Computer Physics Communications, 63(1–3):446–459, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190269Q.

Head:1993:LPP

[Hea93] A. K. Head. LIE, a PC program for Lie analysis of differential equations. Computer Physics Communications, 77(2): 241–248, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390007Y.

Head:1996:LPP

[Hea96] A. K. Head. LIE, a PC program for Lie analysis of differential equations. Computer Physics Communications, 96(2–3):311–313, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000690.

Helton:1999:USA

[Hel99] J. C. Helton. Uncertainty and sensitivity analysis in performance assessment for the Waste Isolation Pilot Plant. Computer Physics Communications, 117(1–2):156–180, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001714.

Hendrikse:1995:CIE

[HEM95] Z. W. Hendrikse, M. O. Elout, and W. J. A. Maaskant. Computation of the independent elements of the dynamical matrix. Computer Physics Communications, 86(3):297–311, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400164W.

Hennecke:1994:RER

[Hen94] Michael Hennecke. RANEXP: experimental random number generator package. Computer Physics Communications, 79 (2):261–267, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900728.

Hennecke:1995:FIR

[Hen95] Michael Hennecke. A Fortran 90 interface to random number generation. Computer Physics Communications, 90(1): 117–120, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500065N.

Hereman:1991:ESW

[Her91] Willy Hereman. Exact solitary wave solutions of coupled nonlinear evolution equations using MACSYMA. Computer Physics Communications, 65(1–3):143–150, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190166I.

Heuer:1990:FVF

[Heu90] Hans-Otto Heuer. A fast vectorized Fortran 77 program for the Monte Carlo simulation of the three-dimensional Ising system. Computer Physics Communications, 59(2):387–398, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590901865.

Hewett:1994:LFE

[Hew94] D. W. Hewett. Low-frequency electromagnetic (darwin) applications in plasma simulation. Computer Physics Communications, 84(1–3):243–277, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902143.

Hibbert:1991:GPCa

[HF91] Alan Hibbert and Charlotte Froese Fischer. A general program for computing angular integrals of the nonrelativistic Hamiltonian with non-orthogonal orbitals. Computer Physics Communications, 64(3):417–430, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901369.

Heck:1990:CPC

[HFD90] L. Heck, D. R. Flower, and G. Pineau Des Forêts. A computer program for calculating the structure of magnetohydrodynamical shocks in interstellar clouds. Computer Physics Communications, 58(1–2):169–179, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090143O.

Hereman:1998:AIT

[HGCM98] Willy Hereman, Ünal Göktas, Michael D. Colagrosso, and Antonio J. Miller. Algorithmic integrability tests for nonlinear differential and lattice equations. Computer Physics Communications, 115(2–3):428–446, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001210.

Hibbert:1991:GPCb

[HGF91] Alan Hibbert, Robert Glass, and Charlotte Froese Fischer. A general program for computing angular integrals of the Breit–Pauli Hamiltonian. Computer Physics Communications, 64(3):455–472, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190138B.

Hogeweij:1990:HTA

[HHC90] G. M. D. Hogeweij, G. Hordósy, and N. J. Lopes Cardozo. Heat transport analysis with error calculation. Computer Physics Communications, 59(2):245–258, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090173X.

Han:1995:FPC

[HHCS95] Q. Z. Han, S. T. Hsieh, H. C. Chiang, and H. Z. Sun. A FORTRAN program for the CFPs of a Boson system with F spin. Computer Physics Communications, 85(3):463–470, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400136P.

Hoyuelos:1999:DFD

[HHGCM99] Miguel Hoyuelos, Emilio Hernández-García, Pere Colet, and Maxi San Miguel. Defect-freezing and defect-unbinding in the vector complex Ginzburg–Landau equation. Computer Physics Communications, 121–122:414–419, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003719.

Hergert:1999:PCS

[HHH99] W. Hergert, R. Hillebrand, and W. Harms. 2D photonic crystals — shape and size properties. Computer Physics Communications, 121–122:651, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670056X.

Hnatowicz:1992:PPL

[HHK92] V. Hnatowicz, V. Havránek, and J. Kvítek. A Pascal program for the least-squares evaluation of standard RBS spectra. Computer Physics Communications, 72(2–3):295–303, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290156S.

Hu:1998:CMB

[HHR98] Xu-Guang Hu, Tak-San Ho, and Herschel Rabitz. The collocation method based on a generalized inverse multiquadric basis for bound-state problems. Computer Physics Communications, 113(2–3):168–179, October 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000964.

Haerle:1995:EAD

[HHT95] Rainer Haerle, Roger Haydock, and Ronald L. Te. Estimating average densities of states for disordered systems. Computer Physics Communications, 90(1):81–86, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500079U.

Hrach:1999:SRD

[HHV99] R. Hrach, V. Hrachová, and M. Vicher. A study of reactions in DC glow discharge in oxygen. Computer Physics Communications, 121–122:662, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700674.

Henneken:1993:SFE

[HI93] Edwin A. C. Henneken and Vincent Icke. SPH faces Emery’s jump. Computer Physics Communications, 74(2):239–246, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390094S.

Hnatowicz:1990:GFC

[HIK90] Vladimír Hnatowicz, Valery Ilyushchenko, and Peter Kozma. GSAP: Fortran code for gamma-spectrum analysis. Computer Physics Communications, 60(1):111–125, August 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090081B.

Hinsen:1995:HLE

[Hin95] K. Hinsen. HYDROLIB: a library for the evaluation of hydrodynamic interactions in colloidal suspensions. Computer Physics Communications, 88(2–3):327–340, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500029F.

Hamilton:1997:AR

[HJ97] Kenneth G. Hamilton and F. James. Acceleration of RAN-LUX. Computer Physics Communications, 101(3):241–248, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000027.

Hoogland:1994:GPR

[HK94] Jiri Hoogland and Ronald Kleiss. Generation of pseudorandom variates with fixed sum and product. Computer Physics Communications, 79(2):179–189, April 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900663.

Haydock:1995:RSL

[HK95] Roger Haydock and David B. Kim. Recursive solution of Liouville’s equation. Computer Physics Communications, 87 (3):396–408, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595000217.

Hoogland:1996:DBEa

[HK96a] Jiri K. Hoogland and Ronald Kleiss. Discrepancy-based error estimates for quasi-Monte Carlo I. General formalism. Computer Physics Communications, 98(1–2):111–127, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000847.

Hoogland:1996:DBEb

[HK96b] Jiri K. Hoogland and Ronald Kleiss. Discrepancy-based error estimates for quasi-Monte Carlo II. Results in one dimension. Computer Physics Communications, 98(1–2):128–136, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000835.

Hakkinen:1997:CCP

[HK97a] Jari Häkkinen and Hamid Kharraziha. Colour: a computer program for QCD colour factor calculations. Computer Physics Communications, 100(3):311–321, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001403.

Hoogland:1997:DBE

[HK97b] Jiri Hoogland and Ronald Kleiss. Discrepancy-based error estimates for quasi-Monte Carlo III. Error distributions and central limits. Computer Physics Communications, 101(1–2):21– 30, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001543.

Huang:1994:DAF

[HKA^{+}94] Youhong Huang, Donald J. Kouri, Mark Arnold, Thomas L. Marchioro II, and David K. Hoffmann. Distributed approximating function approach to time-dependent wavepacket propagation in 3-dimensions: atom-surface scattering. Computer Physics Communications, 80(1–3):1–16, March 1994.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900957.

Hoyles:1998:SPE

[HKC98] Matthew Hoyles, Serdar Kuyucak, and Shin-Ho Chung. Solutions of Poisson’s equation in channel-like geometries. Computer Physics Communications, 115(1):45–68, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000903.

Hermans:1990:LOM

[HKG90] D. Hermans, W. Kerner, and M. Goossens. Linearly overstable magnetic convection in 1D compressible and non-uniform plasmas. Computer Physics Communications, 59(1):127–138, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090162T.

Hilgart:1993:EMC

[HKL93] John Hilgart, Ronald Kleiss, and François Le Diberder. An electroweak Monte Carlo for four-fermion production. Computer Physics Communications, 75(1–2):191–218, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390175C.

Hancock:1999:TMLb

[HKLM99] S. Hancock, P. Knaus, M. Lindroos, and Mike Metcalf. Tomographic measurements of longitudinal phase space density. Computer Physics Communications, 121–122:648, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700534.

Hirai:1998:NSEb

[HKM98a] M. Hirai, S. Kumano, and M. Miyama. Numerical solution of Q^{2} evolution equation for the transversity distribution Δ_{T} q. Computer Physics Communications, 111(1–3):150–166, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000289.

Hirai:1998:NSEa

[HKM98b] M. Hirai, S. Kumano, and M. Miyama. Numerical solution of Q^{2} evolution equations for polarized structure functions. Computer Physics Communications, 108(1):38–55, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559700129X.

Hrebicek:1990:IID

[HKSV90] Jirí Hrebícek, Jan Kucera, Petr Svenda, and Vladimír A. Vasilenko. IDA — interactive data approximation package. Computer Physics Communications, 61(1–2):231–233, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090121G.

Han:1992:PSM

[HL92] J. H. Han and J. N. Leboeuf. Particle simulation model of the Lorentz collision operator in guiding-center plasmas. Computer Physics Communications, 69(2–3):277–286, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290166V.

Hu:1995:MNV

[HLL^{+}95] Wei-Ping Hu, Gillian C. Lynch, Yi-Ping Liu, Ivan Rossi, James J. P. Stewart, Rozeanne Steckler, Bruce C. Garrett, Alan D. Isaacson, Da hong Lu, Vasilios S. Melissas, and Donald G. Truhlar. MORATE 6.5: a new version of a computer program for direct dynamics calculations of chemical reaction rate constants. Computer Physics Communications, 88(2–3):344–346, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500038H.

Hancock:1999:TMLa

[HLMM99] S. Hancock, M. Lindroos, E. McIntosh, and M. Metcalf. Tomographic measurements of longitudinal phase space density. Computer Physics Communications, 118(1):61–70, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599001940.

Henderson:1993:COD

[HLPT93] James R. Henderson, C. Ruth Le Sueur, Steven G. Pavett, and Jonathan Tennyson. Coordinate ordering in the discrete variable representation. Computer Physics Communications, 74(2):193–198, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390089U.

Henderson:1993:DPF

[HLT93] James R. Henderson, C. Ruth Le Sueur, and Jonathan Tennyson. DVR3D: programs for fully pointwise calculation of vibrational spectra. Computer Physics Communications, 75 (3):379–395, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390050M.

Lu:1992:PNV

[hLTM^{+}92] Da hong Lu, Thanh N. Truong, Vasilios S. Melissas, Gillian C. Lynch, Yi-Ping Liu, Bruce C. Garrett, Rozeanne Steckler, Alan D. Isaacson, Sachchida N. Rai, Gene C. Hancock, J. G. Lauderdale, Tomi Joseph, and Donald G. Truhlar. POLYRATE 4: a new version of a computer program for the calculation of chemical reaction rates for polyatomics. Computer Physics Communications, 71(3):235–262, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290012N.

Harrington:1991:ECT

[HM91] Roger F. Harrington and Joseph R. Mautz. Electromagnetic coupling through apertures by the generalized admittance approach. Computer Physics Communications, 68(1–3):19–42, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190192N.

Heddle:1992:LSG

[HM92] D. P. Heddle and Leonard C. Maximon. LASPE: a subroutine for generating straggling distributions for positrons and electrons. Computer Physics Communications, 70(1):77–91, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290093E.

Houdayer:1999:LLE

[HM99] J. Houdayer and O. C. Martin. Low lying excitations in a frustrated and disordered system. Computer Physics Communications, 121–122:659, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700649.

Hossain:1992:CHO

[HMG92] Murshed Hossain, William H. Matthaeus, and Sanjoy Ghosh. On computing high order Galerkin products. Computer Physics Communications, 69(1):1–6, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290123G.

Hasnaoui:1999:CFC

[HMH99] A. Hasnaoui, A. Menai, and A. Hoummada. Correlation factor and concentration profiles for interstitial-substitutional mechanism with impurity-vacancy interaction in the FCC lattice by means of Monte Carlo method. Computer Physics Communications, 121–122:650, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700558.

Houssa:1999:CSD

[HMR99] Mohammed Houssa, Simon C. McGrother, and Luis F. Rull. Computer simulations of dipolar liquid crystal phases. Computer Physics Communications, 121–122:259– 261, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003252.

Hernandez:1997:JPE

[HN97a] Juan José Hernández and Sergio Navas. JASP: a program to estimate discovery and exclusion limits in prospective studies of searches. Computer Physics Communications, 100(1–2):119–127, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001336.

Hiyama:1997:GET

[HN97b] Miyabi Hiyama and Hiroki Nakamura. Gaussian expansions of the two-center Coulomb functions. Computer Physics Communications, 103(2–3):197–208, July 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000477.

Hiyama:1997:TCC

[HN97c] Miyabi Hiyama and Hiroki Nakamura. Two-center Coulomb functions. Computer Physics Communications, 103(2–3):209– 216, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000453.

Hinzke:1999:MCS

[HN99] D. Hinzke and U. Nowak. Monte Carlo simulation of magnetization switching in a Heisenberg model for small ferromagnetic particles. Computer Physics Communications, 121–122:334–337, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003483.

Hnizdo:1994:VPP

[Hni94] V. Hnizdo. Vacuum-polarization potentials of extended nuclear charges. Computer Physics Communications, 83(1): 95–106, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900361.

Heermann:1990:HMD

[HNR90] Dieter W. Heermann, Peter Nielaba, and Mauro Rovere. Hybrid molecular dynamics. Computer Physics Communications, 60(3):311–318, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900305.

Hrach:1999:CSF

[HNS99] R. Hrach, D. Novotný, and M. Sobotka. Characterisation of semicontinuous film morphology. Computer Physics Communications, 121–122:663, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700686.

Hinrichsen:1999:CIC

[HÓ99] Haye Hinrichsen and Géza Ódor. Correlated initial condition simulations in directed percolation. Computer Physics Communications, 121–122:392–394, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003628.

Hockney:1992:BRB

[Hoc92] Roger W. Hockney. Book review: Parallel algorithms in computational sciences: Dieter W. Heermann and Anthony N. Burkitt, Springer Series in Information Sciences Vol. 24, Springer Verlag, Berlin, 1991. 183 + xiii pages. Hardcover price DM 60.00. ISBN 3-540-53418-0. Computer Physics Communications, 69(2–3):487–488, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901863.

Hoek:1990:LTC

[Hoe90] Jaap Hoek. Lattice topological charges and the H-gauge. Computer Physics Communications, 61(3):304–308, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900464.

Hofer:1999:SAC

[Hof99a] Eduard Hofer. Sensitivity analysis in the context of uncertainty analysis for computationally intensive models. Computer Physics Communications, 117(1–2):21–34, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001532.

Hoffmann:1999:SRD

[Hof99b] Karl Heinz Hoffmann. Slow relaxation dynamics — from spin glasses to stochastic optimization. Computer Physics Communications, 121–122:30–33, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002726.

Hollmann:1996:MVP

[Hol96] Helia Hollmann. A Maple V package for field theoretic Poisson brackets using object-oriented techniques. Computer Physics Communications, 96(2–3):209–216, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000768.

Hopkins:1997:BRB

[Hop97] Tim Hopkins. Book review: Numerical algorithms with Fortran by G. Engeln–Műllges and F. Uhlig (Springer, Berlin, 1996), ISBN 3-540-60529-0: Numerical algorithms with C by G. Engeln–Műllges and F. Uhlig (Springer, Berlin, 1996), ISBN 3-540-60530-4. Computer Physics Communications, 103(1):100–101, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000246.

Horacek:1990:NLP

[Hor90] Helmut Horacek. Natural language processing. Computer Physics Communications, 61(1–2):76–92, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090107C.

Horbatsch:1991:SIE

[Hor91] Marko Horbatsch. A semiclassical independent electron model for ion-atom collisions. Computer Physics Communications, 63(1–3):115–125, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190243E.

Hai-Ping:1993:IMC

[HP93] Fang Hai-Ping. Improved Monte Carlo treatment of multiparticle phase-space at high energy. Computer Physics Communications, 77(3):374–376, November 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390183D.

Haynes:1997:LSW

[HP97a] P. D. Haynes and M. C. Payne. Localised spherical-wave basis set for O(N) total-energy pseudopotential calculations. Computer Physics Communications, 102(1–3):17–27, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000283.

Hirvi:1997:DTP

[HP97b] K. P. Hirvi and J. P. Pekola. Determination of thermometric parameters from the conductance curve of the normal metal based tunnel junction array. Computer Physics Communications, 106(1–2):69–75, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000714.

Hughes:1992:CMR

[HPS92] M. H. Hughes, M. W. Phillips, and R. G. Storer. A computer model for resistive MHD analysis. Computer Physics Communications, 72(1):76–95, October 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290007L.

Hahn:1999:AOL

[HPV99] T. Hahn and M. Pérez-Victoria. Automated one-loop calculations in four and D dimensions. Computer Physics Communications, 118(2–3):153–165, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001738.

Hammonds:1991:CSA

[HR91] Kenton D. Hammonds and Jean-Paul Ryckaert. On the convergence of the SHAKE algorithm. Computer Physics Communications, 62(2–3):336–351, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190105T.

Homeier:1995:ISO

[HRK95] Herbert H. H. Homeier, Sebastian Rast, and Hartmut Krienke. Iterative solution of the Ornstein–Zernike equation with various closures using vector extrapolation. Computer Physics Communications, 92(2–3):188–202, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001160.

Hagino:1999:PCC

[HRK99] K. Hagino, N. Rowley, and A. T. Kruppa. A program for coupled-channel calculations with all order couplings for heavy-ion fusion reactions. Computer Physics Communications, 123(1–3):143–152, December 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900243X.

Henk:1993:SPC

[HS93a] J. Henk and W. Schattke. A subroutine package for computing Green’s functions of relaxed surfaces by the renormalization method. Computer Physics Communications, 77(1):69–83, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390038E.

Homeier:1993:PEN

[HS93b] H. H. H. Homeier and E. O. Steinborn. Programs for the evaluation of nuclear attraction integrals with B functions. Computer Physics Communications, 77(1):135–151, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390042B.

Hawley:1995:MNT

[HS95] John F. Hawley and James M. Stone. MOCCT: a numerical technique for astrophysical MHD. Computer Physics Communications, 89(1–3):127–148, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500190Q.

Homeier:1996:CGQ

[HS96] Herbert H. H. Homeier and E. Otto Steinborn. Comment on: a Gaussian quadrature for the optimal evaluation of integrals involving lorentzians over a semi-infinite interval. Computer Physics Communications, 99(1):77–80, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001154.

Hofler:1999:SPS

[HS99]  Kai Höfler and Stefan Schwarzer.  Simulating parti-

cle suspensions by extended point-force distributions in a Navier–Stokes scheme. Computer Physics Communications, 121–122:657, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700625.

Holmstrom:1995:NMT

[HSM95] Lasse Holmström, Stephan R. Sain, and Hannu E. Miettinen. A new multivariate technique for top quark search. Computer Physics Communications, 88(2–3):195–210, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500040M.

Heikkinen:1993:MCS

[HSP93] J. A. Heikkinen, S. K. Sipilä, and T. J. H. Pättikangas. Monte Carlo simulation of runaway electrons in a toroidal geometry. Computer Physics Communications, 76(2):215–230, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390133W.

Heggarty:1998:CAT

[HSSB98] J. W. Heggarty, M. P. Scott, N. S. Scott, and P. G. Burke. Computational aspects of the two-dimensional propagation of R-matrices on MPPs. Computer Physics Communications, 114(1–3):195–209, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001374.

Hoyle:1991:SEV

[HSW91] J. M. Hoyle, F. T. Smith, and J. D. A. Walker. On sublayer eruption and vortex formation. Computer Physics Communications, 65(1–3):151–157, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190167J.

Hofler:1999:SHS

[HSW99] Kai Höfler, Stefan Schwarzer, and Bernd Wachmann. Simulation of hindered settling in bidisperse suspensions of rigid spheres. Computer Physics Communications, 121–122:268– 269, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003288.

Henderson:1993:DPM

[HT93] James R. Henderson and Jonathan Tennyson. DVR1D: programs for mixed pointwise/basis set calculation of rovibrational spectra. Computer Physics Communications, 75 (3):365–378, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390049I.

Higuchi:1999:SDL

[HTI99] M. Higuchi, H. Takano, and M. Itoh. Simulation of the dynamic light scattering on hydrodynamically interacting particles — memory effect in collective motion. Computer Physics Communications, 121–122:655, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700601.

Hahn:1999:CGS

[HTK99] Oliver Hahn, Wolfgang Tschöp, and Kurt Kremer. Coarse grained simulation of polycarbonates. Computer Physics Communications, 121–122:647, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700522.

Haas:1998:RMN

[HTZ^{+}98] S. Haas, D. A. Thornley, M. Zhu, R. W. Dobinson, R. Heeley, N. A. H. Madsen, and B. Martin. Results from the Macramé 1024 node switching network. Computer Physics Communications, 110(1–3):206–210, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001781.

Harp:1998:CPE

[HUCS98] G. R. Harp, Y. Ueda, X. Chen, and D. K. Saldin. Computation of photo-electron and Auger-electron diffraction II. Multiple scattering cluster calculation PAD2. Computer Physics Communications, 112(1):80–90, July 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000319.

Humpert:1990:CSN

[Hum90a] B. Humpert. A comparative study of neural networks architectures. Computer Physics Communications, 58(3):223– 256, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590900615.

Humpert:1990:UNN

[Hum90b] B. Humpert. On the use of neural networks in high-energy physics experiments. Computer Physics Communications, 56(3):299–311, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090016T.

Humpert:1990:SPA

[Hum90c] B. Humpert. Solving problems with automated reasoning, expert systems and neural networks. Computer Physics Communications, 61(1–2):58–75, November 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090106B.

Hutson:1994:CCM

[Hut94] Jeremy M. Hutson. Coupled channel methods for solving the bound-state Schrödinger equation. Computer Physics Communications, 84(1–3):1–18, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902003.

Hauck:1998:SCS

[HvSA98a] A. Hauck, L. von Smekal, and R. Alkofer. Solving a coupled set of truncated QCD Dyson–Schwinger equations. Computer Physics Communications, 112(2–3):166–182, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000459.

Hauck:1998:SGD

[HvSA98b] A. Hauck, L. von Smekal, and R. Alkofer. Solving the gluon Dyson–Schwinger equation in the Mandelstam approximation. Computer Physics Communications, 112(2–3):149–165, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000460.

Homeier:1995:REL

[HW95] H. H. H. Homeier and E. J. Weniger. On remainder estimates for levin-type sequence transformations. Computer Physics Communications, 92(1):1–10, November 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500103M.

Homeier:1992:PEO

[HWS92] H. H. H. Homeier, E. J. Weniger, and E. O. Steinborn. Programs for the evaluation of overlap integrals with B functions. Computer Physics Communications, 72(2–3):269–287, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290154Q.

Horacek:1993:CSM

[HZQ93] J. Horácek, L. Zejda, and N. M. Queen. Comparative study of methods for the construction of Padé approximants of type III. Computer Physics Communications, 74(2):187–192, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390088T.

Im:1998:CSM

[IBR98] Wonpil Im, Dmitrii Beglov, and Benoît Roux. Continuum solvation model: Computation of electrostatic forces from numerical solutions to the Poisson–Boltzmann equation. Computer Physics Communications, 111(1–3):59–75, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000162.

Iitaka:1995:SSM

[ICS95] Toshiaki Iitaka, Nicolae Carjan, and Dan Strottman. Stability of the symmetric multistep methods for the timedependent Schrödinger equation. Computer Physics Communications, 90(2–3):251–259, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500085T.

Ixaru:1999:SPS

[IDV99] L. Gr. Ixaru, H. De Meyer, and G. Vanden Berghe. SLCPM12 — a program for solving regular Sturm–Liouville problems. Computer Physics Communications, 118(2–3):259–277, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001817.

Ixaru:1997:EFP

[IDVV97] L. Gr. Ixaru, H. De Meyer, G. Vanden Berghe, and M. Van Daele. EXPFIT4 — a FORTRAN program for the numerical solution of systems of nonlinear secondorder initial-value problems. Computer Physics Communications, 100(1–2):71–80, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001464.

Izus:1999:CSF

[IDW99] G. G. Izús, R. R. Deza, and H. S. Wio. Critical slowing-down in the FitzHugh–Nagumo model: a nonequilibrium potential approach. Computer Physics Communications, 121–122:406–407, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003689.

Ingelman:1997:LMC

[IER97] G. Ingelman, A. Edin, and J. Rathsman. LEPTO 6.5- A Monte Carlo generator for deep inelastic lepton-nucleon scattering. Computer Physics Communications, 101(1–2):108– 134, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001579.

Ito:1996:SPC

[IEY^{+}96] Tomoyoshi Ito, Hesham Eldeib, Kenji Yoshida, Shinya Takahashi, Takashi Yabe, and Tomoaki Kunugi. Specialpurpose computer for holography HORN-2. Computer Physics Communications, 93(1):13–20, January 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001255.

Indjin:1992:SMV

[IIM92] D. Indjin, Z. Ikonić, and V. Milanović. On shooting method variations for the 1-D Schrödinger equation and their accuracy. Computer Physics Communications, 72(2–3):149–153, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290146P.

Irback:1992:MST

[IJV92] A. Irbäck, J. Jurkiewicz, and S. Varsted. Measuring the string tension in random surface models with extrinsic curvature. Computer Physics Communications, 70(1):59–68, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290091C.

Ilyin:1996:ARP

[IK96] V. A. Ilyin and A. P. Kryukov. ATENSOR — REDUCE program for tensor simplification. Computer Physics Communications, 96(1):36–52, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000604.

Itoh:1999:MDS

[IK99] H. Itoh and K. Kawamura. Molecular dynamics studies of lattice vibrations of proton-ordered ice XI. Computer Physics Communications, 121–122:666, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700716.

Ido:1999:AFM

[IKH^{+}99] S. Ido, M. Kawashima, R. Hirose, H. Shoji, and M. Kashiwagi. Application of the free-mesh finite element method to electromagnetic analysis. Computer Physics Communications, 121–122:664, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700698.

Ido:1999:CSG

[IKTY99] S. Ido, M. Kashiwagi, M. Takahashi, and T. Yoshida. Computational studies on generation and control in magnetron sputtering plasma. Computer Physics Communications, 121–122:665, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700704.

Iung:1991:ETE

[IL91] Christophe Iung and Claude Leforestier. Exact time evolution methods for large bound systems. Computer Physics Communications, 63(1–3):135–153, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190245G.

Ivanov:1994:ESS

[IMA94] K. N. Ivanov, M. A. Manolova, and T. G. Apostolov. An effective solution scheme of a three-dimensional reactor core model in hexagonal geometry. Computer Physics Communications, 82(1):1–16, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901260.

Ito:1990:SPB

[IMES90] Tomoyoshi Ito, Junichiro Makino, Toshikazu Ebisuzaki, and Daiichiro Sugimoto. A special-purpose N-body machine GRAPE-1. Computer Physics Communications, 60(2):187– 194, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090003J.

Ivantchev:1999:MCS

[IMG99] S. Ivantchev, J. M. Pérez Mato, and Alberto García. Monte Carlo simulation of dielectric response near a phase transition: displacive vs. order-disorder character. Computer Physics Communications, 121–122:668, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550670073X.

Itoh:1995:OTB

[IOM95] Satoshi Itoh, Pablo Ordejón, and Richard M. Martin. Order-N tight-binding molecular dynamics on parallel computers. Computer Physics Communications, 88(2–3):173–185, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500031A.

Iqbal:1995:CSR

[Iqb95a] M. Iqbal. On comparison of spline regularization with exponential sampling method for Laplace transform inversion. Computer Physics Communications, 88(1):43–50, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500026C.

Iqbal:1995:RNI

[Iqb95b] M. Iqbal. On regularized numerical inversion of Mellin transform. Computer Physics Communications, 90(2–3):293–300, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500074P.

Irisa:1996:EAA

[Iri96] Masayuki Irisa. An elegant algorithm of the analytical calculation for the volume of fused spheres with different radii. Computer Physics Communications, 98(3):317–338, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000823.

Ingelman:1997:AMC

[IRS97] G. Ingelman, J. Rathsman, and G. A. Schuler. AROMA 2.2-A Monte Carlo generator for heavy flavour events in ep collisions. Computer Physics Communications, 101(1–2):135–142, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001452.

Ixaru:1995:PPM

[IRV95] L. Gr. Ixaru, M. Rizea, and T. Vertse. Piecewise perturbation methods for calculating eigensolutions of a complex optical potential. Computer Physics Communications, 85(2):217–230, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400135O.

Iskander:1991:CTB

[Isk91] Magdy F. Iskander. Computational techniques in bioelectromagnetics. Computer Physics Communications, 68(1–3): 224–254, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190202V.

Iglesias:1991:MAI

[ISS91] E. Iglesias, T. L. Sordo, and J. A. Sordo. Molecular associations from ab initio pair potentials. Computer Physics Communications, 67(2):268–284, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190022D.

Itoh:1999:CEI

[IST99] M. Itoh, K. Sakiyama, and H. Takano. Configuration and energetics of ion-induced embryos in nucleation. Computer Physics Communications, 121–122:667, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506700728.

Indjin:1995:NSS

[ITMI95] D. Indjin, G. Todorović, V. Milanović, and Z. Ikonić. On numerical solution of the Schrödinger equation: the shooting method revisited. Computer Physics Communications, 90(1):87–94, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500071M.

Ito:1994:SRP

[Ito94] Masaaki Ito. SYMCD — a REDUCE package for finding symmetries and conserved densities of systems of nonlinear evolution equations. Computer Physics Communications, 79(3):547–554, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901937.

Idomura:1997:CBP

[ITW97] Y. Idomura, S. Tokuda, and M. Wakatani. Chaotic behaviour in PIC simulation and its relation to computational errors. Computer Physics Communications, 102(1–3):68–80, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000106.

Ixaru:1997:FSE

[IVDV97] L. Gr. Ixaru, G. Vanden Berghe, H. De Meyer, and M. Van Dacle. Four-step exponential-fitted methods for nonlinear physical problems. Computer Physics Communications, 100(1–2):56–70, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001476.

Ishikawa:1992:MHS

[IWWY92] T. Ishikawa, P. Y. Wang, K. Wakui, and T. Yabe. A method for the high-speed generation of random numbers with arbitrary distributions. Computer Physics Communications, 70(3):501–509, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290112C.

Ixaru:1997:OOF

[Ixa97] L. Gr. Ixaru. Operations on oscillatory functions. Computer Physics Communications, 105(1):1–19, September 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000672.

Ida:1995:ICC

[IY95] M. Ida and T. Yabe. Implicit CIP (cubic-interpolated propagation) method in one dimension. Computer Physics Communications, 92(1):21–26, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559592245C.

Ito:1994:SPC

[IYOY94] Tomoyoshi Ito, Takashi Yabe, Masashi Okazaki, and Masanori Yanagi. Special-purpose computer HORN-1 for reconstruction of virtual image in three dimensions. Computer Physics Communications, 82(2–3):104–110, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901597.

Izvekov:1999:PND

[Izv99] S. V. Izvekov. Polarons by nonlocal dynamical coherent potential method. Computer Physics Communications, 121–122:669, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700741.

Jackson:1991:TDM

[Jac91] Bret Jackson. Time-dependent mean-field approach to molecule-surface scattering at finite temperatures. Computer Physics Communications, 63(1–3):154–170, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190246H.

Jackson:1994:QSC

[Jac94] Bret Jackson. Quantum and semiclassical calculations of gas-surface energy transfer and sticking. Computer Physics Communications, 80(1–3):119–144, March 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901007.

Jacob:1999:ECD

[Jac99] Christian Jacob. Evolution and coevolution of developmental programs. Computer Physics Communications, 121–122: 46–50, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002775.

Jakubi:1998:GPE

[Jak98] Alejandro S. Jakubi. Generalized power expansions in cosmology. Computer Physics Communications, 115(2–3):284– 299, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001313.

James:1990:RPN

[Jam90] F. James. A review of pseudorandom number generators. Computer Physics Communications, 60(3):329–344, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090032V.

James:1994:RFI

[Jam94] F. James. RANLUX: a Fortran implementation of the high-quality pseudorandom number generator of Lüscher. Computer Physics Communications, 79(1):111–114, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559490233X. See erratum

[Jam96b].

James:1996:E

[Jam96a] F. James. Erratum. Computer Physics Communications, 97 (3):357, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000653.

James:1996:ERF

[Jam96b] F. James. Erratum: RANLUX: A Fortran implementation of the high-quality pseudorandom number generator of Lüscher [Comput. Phys. Commun. 79 (1994) 111–114]. Computer Physics Communications, 97(3):357, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000653. See [Jam94].

James:1996:NE

[Jam96c] Frederick James. A note from the editor. Computer Physics Communications, 97(1–2):xii, August 2, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596900934.

Janicki:1990:CPF

[Jan90] Christian Janicki. A computer program for the free-free and bound-free Gaunt factors of rydberg systems. Computer Physics Communications, 60(3):281–296, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090027X.

Jansen:1995:MCS

[Jan95] A. P. J. Jansen. Monte Carlo simulations of chemical reactions on a surface with time-dependent reaction-rate constants. Computer Physics Communications, 86(1–2):1–12, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400155U.

Jansen:1999:AVD

[Jan99] Michiel J. W. Jansen. Analysis of variance designs for model output. Computer Physics Communications, 117(1–2):35–43, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001544.

Jaquet:1990:AFE

[Jaq90] Ralph Jaquet. Application of the finite element method to eigenvalue problems I. one dimensional calculations using optimized elements. Computer Physics Communications, 58 (3):257–269, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465590900626.

Jabbarzadeh:1997:PSS

[JAT97] A. Jabbarzadeh, J. D. Atkinson, and R. I. Tanner. Parallel simulation of shear flow of polymers between structured walls by molecular dynamics simulation on PVM. Computer Physics Communications, 107(1–3):123–136, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559700088X.

Jaun:1995:GWR

[JAVV95] A. Jaun, K. Appert, J. Vaclavik, and L. Villard. Global waves in resistive and hot tokamak plasmas. Computer Physics Communications, 92(2–3):153–187, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001056.

Janke:1999:MOS

[JBB99] Wolfhard Janke, Bernd A. Berg, and Alain Billoire. Multioverlap simulations of free-energy barriers in the 3D Edwards– Anderson Ising spin glass. Computer Physics Communications, 121–122:176–179, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003069.

Johnson:1999:LAE

[JBH99] C. W. Johnson, G. F. Bertsch, and W. D. Hazelton. Lanczos algorithm and energy-weighted sum rules for linear response. Computer Physics Communications, 120(2–3):155–161, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002489.

Jacobsen:1997:ASD

[JBSZ97] Heiko Jacobsen, Attila Bérces, David P. Swerhone, and Tom Ziegler. Analytic second derivatives of molecular energies: a density functional implementation. Computer Physics Communications, 100(3):263–276, March 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001191.

Jamieson:1991:CNC

[JD91a] M. J. Jamieson and M. L. Du. Comments on the Numerov– Cooley and Hartree–Cooley methods for finding eigenvalues. Computer Physics Communications, 64(3):360–362, June 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591901314.

Jamieson:1991:LDM

[JD91b] M. J. Jamieson and M. L. Du. Log-derivative methods based on linear two-step integration formulae. Computer Physics Communications, 66(2–3):150–156, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190063Q.

Jamieson:1992:EEM

[JF92] M. J. Jamieson and R. S. Friedman. Error estimation in molecular vibrational eigenenergies calculated by the Hartree and Numerov methods. Computer Physics Communications, 70(1):53–58, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290090L.

Jonsson:1997:SPR

[JF97] P. Jönsson and C. Froese Fischer. SMS92: a program for relativistic isotope shift calculations. Computer Physics Communications, 100(1–2):81–92, February 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559600118X.

Jamieson:1999:EEM

[JF99]  Michael J. Jamieson and Ronald S. Friedman.  Er-

ror estimation in molecular vibrational energies calculated by Johnson’s log-derivative method. Computer Physics Communications, 118(2–3):145–152, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002039.

Joslin:1993:CQM

[JGP93] C. G. Joslin, C. G. Gray, and J. D. Poll. Calculation of the quantum-mechanical propagator at large real times. Computer Physics Communications, 75(1–2):1–9, April 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390160E.

Jolicard:1991:STD

[JH91a] Georges Jolicard and Jeannette Humbert. Stationary and time-dependent wave-operator formulation in molecular dynamics. Computer Physics Communications, 63(1–3):216– 227, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190251F.

Jumper:1991:STD

[JH91b] E. J. Jumper and R. J. Hugo. Simple theories of dynamic stall that are helpful in interpreting computational results. Computer Physics Communications, 65(1–3):158–163, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190168K.

James:1997:MSS

[JHK97] F. James, J. Hoogland, and R. Kleiss. Multidimensional sampling for simulation and integration: measures, discrepancies, and quasi-random numbers. Computer Physics Communications, 99(2–3):180–220, January 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001087.

Jemmer:1997:EBF

[J̈K97a] Patrick J̈emmer and Peter J. Knowles. Erratum to Generation of functional derivatives in Kohn–Sham density-functional theory [Comput. Phys. Commun. 100 (1997) 93–98]. Computer Physics Communications, 103(1):95–96, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000465. See [J̈K97b].

Jemmer:1997:GFD

[J̈K97b] Patrick J̈emmer and Peter J. Knowles. Generation of functional derivatives in Kohn–Sham density-functional theory. Computer Physics Communications, 100(1–2):93–98, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001610. See [J̈K97a].

Jain:1999:EGC

[JK99] Sanjay Jain and Sandeep Krishna. Emergence and growth of complex networks in adaptive systems. Computer Physics Communications, 121–122:116–121, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002933.

Jay:1999:ESC

[JKSC99] Laurent O. Jay, Hanchul Kim, Yousef Saad, and James R. Chelikowsky. Electronic structure calculations for planewave codes without diagonalization. Computer Physics Communications, 118(1):21–30, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001921.

Jadach:1991:TLM

[JKW91] Stanisław Jadach, Johann H. Kühn, and Zbigniew Was. TAUOLA — a library of Monte Carlo programs to simulate decays of polarized τ leptons. Computer Physics Communications, 64(2):275–299, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190038M.

Jamin:1993:TVM

[JL93] Matthias Jamin and Markus E. Lautenbacher. TRACER version 1.1. A Mathematica package for γ-algebra in arbitrary dimensions. Computer Physics Communications, 74(2): 265–288, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390097V.

Jansen:1997:ISI

[JL97] Karl Jansen and Chuan Liu. Implementation of Symanzik’s improvement program for simulations of dynamical Wilson fermions in lattice QCD. Computer Physics Communications, 99(2–3):221–234, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001282.

Jensen:1991:MCS

[JLFS91] Geir U. Jensen, Bjørnar Lund, Tor A. Fjeldly, and Michael Shur. Monte Carlo simulation of semiconductor devices. Computer Physics Communications, 67(1):1–61, August 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190220F.

Jimenez-Morales:1999:ICA

[JMK99] F. Jiménez-Morales and H. Karma. Intermittency in cellular automata with high connectivity. Computer Physics Communications, 121–122:671, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700765.

Johnson:1993:CAT

[JMR^{+}93] J. L. Johnson, D. A. Monticello, A. H. Reiman, A. Salas, A. L. Fraguas, and S. P. Hirshman. Comparison of ATF and TJ-II stellarator equilibria as computed by the 3-D VMEC and PIES codes. Computer Physics Communications, 77(1):1–10, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593900317.

Jentschura:1999:CAC

[JMSW99] Ulrich D. Jentschura, Peter J. Mohr, Gerhard Soff, and Ernst Joachim Weniger. Convergence acceleration via combined nonlinear-condensation transformations. Computer Physics Communications, 116(1):28–54, January 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001118.

Jarvis:1994:LBP

[JN94] R. D. Jarvis and Paul Nelson. LOCFES-B: a program for solving the one-dimensional particle transport equation with user selected CLOF methods (new version). Computer Physics Communications, 82(2–3):265–286, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901732.

Jones:1990:BRB

[Jon90] I. P. Jones. Book review: Numerical Combustion: A. Dervieux and B. Larrouturou, eds., Lecture Notes in Physics 351, Springer-Verlag, Berlin, 1989. DM88. Computer Physics Communications, 61(3):436, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900567.

Joo:1997:ORC

[Joó97] I. Joó. Order-α radiative correction calculations for oriented allowed nuclear β decays. Computer Physics Communications, 105(2–3):99–107, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000751.

Jerie:1998:CAD

[JOP98] M. Jerie, J. E. R. O’Connor, and G. E. Prince. Computer algebra determination of symmetries in general relativity. Computer Physics Communications, 115(2–3):363–380, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001246.

Jorquera:1995:SAS

[Jor95] Héctor Jorquera. Simple algorithm for solving linear integrodifferential equations with variable limits. Computer Physics Communications, 86(1–2):91–96, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400156V.

Jadach:1992:MCPb

[JP92] Stanisław Jadach and Wiesław Placzek. The Monte Carlo program LESKO-F for deep inelastic e^{±}p → e^{±}X scattering at HERA including QED Bremsstrahlung from the lepton line. Computer Physics Communications, 72(2–3):221–237, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290150W.

Jonsson:1996:HPR

[JPF96] P. Jönsson, F. A. Parpia, and C. Froese Fischer. HFS92: a program for relativistic atomic hyperfine structure calculations. Computer Physics Communications, 96(2–3):301–310, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000574.

Jadach:1997:UMC

[JPRW^{+}97] S. Jadach, W. Placzek, E. Richter-Was, B. F. L. Ward, and Z. Was. Upgrade of the Monte Carlo program BH-LUMI for Bhabha scattering at low angles to version 4.04. Computer Physics Communications, 102(1–3):229–251, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001567.

Jadach:1999:MCP

[JPS^{+}99] S. Jadach, W. Placzek, M. Skrzypek, B. F. L. Ward, and Z. Wass. Monte Carlo program KoralW 1.42 for all fourfermion final states in e^{+}e^{−} collisions. Computer Physics Communications, 119(2–3):272–311, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002192.

Jadach:1992:MCPa

[JRWWW92] S. Jadach, E. Richter-Was, B. F. L. Ward, and Z. Was. Monte Carlo program BHLUMI 2.01 for Bhabha scattering at low angles with Yennie–Frautschi–Suura exponentiation. Computer Physics Communications, 70(2):305–344, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465592901966.

Jost:1997:EEC

[JTAW97a] G. Jost, T. M. Tran, K. Appert, and S. Wüthrich. Effects of electron-cyclotron instabilities on gyrotron beam quality. Computer Physics Communications, 100(1–2):47–55, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001439.

Jost:1997:GPT

[JTAW97b] G. Jost, T. M. Tran, K. Appert, and S. Wüthrich. G2DEM: a parallel two-dimensional electromagnetic PIC code for the study of electron-cyclotron instabilities of relativistic electron beams in cylindrical cavities. Computer Physics Communications, 104(1–3):188–200, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700057X.

Jung:1995:HDS

[Jun95] Hannes Jung. Hard diffractive scattering in high energy ep collisions and the Monte Carlo Generator RAPGAP. Computer Physics Communications, 86(1–2):147–161, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400150Z.

Janke:1999:PRL

[JV99] W. Janke and R. Villanova. 3D Poissonian random lattices with Ising spin systems. Computer Physics Communications, 121–122:670, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700753.

Jadach:1990:YSO

[JW90] Stanisław Jadach and B. F. L. Ward. YFS2 — the second-order Monte Carlo program for fermion pair production at LEP/SLC, with the initial state radiation of two hard and multiple soft photons. Computer Physics Communications, 56(3):351–384, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900202.

Jadach:1991:KVU

[JW91] Stanisław Jadach and Zbigniew Was. KORALB version 2.1. An upgrade with the TAUOLA library of τ decays. Computer Physics Communications, 64(2):267–274, May 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190037L.

Jadach:1995:KUV

[JW95] Stanisław Jadach and Zbigniew Was. KORALB — an upgrade to version 2.4. Computer Physics Communications, 85(3):453– 462, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400091F.

Jadach:1993:DLT

[JWDK93] S. Jadach, Z. Was, R. Decker, and J. H. Kühn. The τ decay library TAUOLA, version 2.4. Computer Physics Communications, 76(3):361–380, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390061G.

Jonsson:1993:PCM

[JWF93] Per Jönsson, Claes-Göran Wahlström, and Charlotte Froese Fischer. A program for computing magnetic dipole and electric quadrupole hyperfine constants from MCHF wavefunctions. Computer Physics Communications, 74(3):399–414, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593900225.

Jeabek:1992:DLT

[JWJK92] M. Jeabek, Z. Was, S. Jadach, and J. H. Kühn. The τ decay library TAUOLA, update with exact O(α) QED corrections in τ → μ(e)vv decay modes. Computer Physics Communications, 70(1):69–76, May 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290092D.

Judge:1996:CAA

[JWM^{+}96] R. H. Judge, E. D. Womeldorf, R. A. Morris, D. E. Shimp, D. J. Clouthier, D. L. Joo, and D. C. Moule. Computer-assisted analysis of singlet-triplet rotational spectra: application to case (A), case (B) and case (AB) coupling cases in polyatomic molecules. Computer Physics Communications, 93(2–3):241–264, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000879.

Jadach:1991:MCP

[JWW91] Stanisław Jadach, B. F. L. Ward, and Zbigniew Was. The Monte Carlo program KORALZ, version 3.8, for the lepton or quark pair production at LEP/SLC energies. Computer Physics Communications, 66(2–3):276–292, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190077X.

Jadach:1994:MCP

[JWW94] S. Jadach, B. F. L. Ward, and Z. Was. The Monte Carlo program KORALZ version 4.0 for lepton or quark pair production at LEP/SLC energies. Computer Physics Communications, 79(3):503–522, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901902.

Wang:1995:NFP

[jWzHxL95] Jia jun Wang, Qi zhi Han, and Yu xin Liu. A new Fortran program for CFPs of an identical fermion system. Computer Physics Communications, 85(1):99–109, January 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400123J.

Jaffe:1997:EBE

[JY97a] David E. Jaffe and Saul Youssef. Erratum to Bayesian estimate of the effect of B^{00} mixing measurements on the CKM matrix elements [Computer Physics Communications 101 (1997) 206]. Computer Physics Communications, 105(2–3):284, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001033. See [JY97b].

Jaffet:1997:BEE

[JY97b] David E. Jaffet and Saul Youssef. Bayesian estimate of the effect of B ^{0} B ^{0} mixing measurements on the CKM matrix elements. Computer Physics Communications, 101(3):206–212, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001713. See erra-

tum [JY97a].

Koning:1995:KPC

[KA95] A. J. Koning and J. M. Akkermans. KAPSIES: a program for the calculation of multi-step direct reaction cross sections. Computer Physics Communications, 85(1):110–126, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400116J.

Kadlecsik:1996:RCP

[Kad96] József Kadlecsik. Ricci calculus package in REDUCE. Computer Physics Communications, 93(2–3):265–282, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001379.

Kaeding:1995:PPG

[Kae95] Thomas A. Kaeding. Pascal program for generating tables of SU(3) Clebsch–Gordan coefficients. Computer Physics Communications, 85(1):82–88, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400115I.

Kagawa:1992:PCF

[Kag92] Takashi Kagawa. A program to calculate fractional parentage coefficients for jj-coupling states with equivalent particles. Computer Physics Communications, 72(2–3):165–174, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290148R.

Kahane:1991:CRD

[Kah91] Sylvian Kahane. Calculating real Delbrück amplitudes on parallel processors. Computer Physics Communications, 67 (2):233–244, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190019H.

Kaluza:1993:PBG

[Kal93] Matjaz Kaluza. Program for Birkhoff–Gustavson normal form for N degrees of freedom — BIRKHOFF 1.2. Computer Physics Communications, 74(3):441–449, March 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900258.

Kaluza:1994:ALM

[Kal94] Matjaz Kaluza. Analytical Lanczos method: quantum eigenstates of anharmonic oscillators in one or more dimensions. Computer Physics Communications, 79(3):425–446, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901864.

Kalkreuter:1996:SCW

[Kal96] Thomas Kalkreuter. Study of Cullum’s and Willoughby’s Lanczos method for Wilson fermions. Computer Physics Communications, 95(1):1–16, May 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000045.

Kaneko:1995:FGG

[Kan95] Toshiaki Kaneko. A feynman-graph generator for any order of coupling constants. Computer Physics Communications, 92(2–3):127–152, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001226.

Karawatzki:1991:PAD

[Kar91] R. Karawatzki. A practical algorithm for the determination of the positions of superstable cycles. Computer Physics Communications, 64(1):64–66, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190050U.

Karimaki:1992:FCF

[Kar92] Veikko Karimäki. Fast code to fit circular arcs. Computer Physics Communications, 69(1):133–141, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290134K.

Kasahara:1991:TRC

[Kas91a] K. Kasahara. A technique for reducing the computational time of electron-photon cascade-shower simulation. Computer Physics Communications, 64(2):235–240, May 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190034I.

Kasahara:1991:TAA

[Kas91b] K. Kasahara. Tools for the analysis of air shower data. Computer Physics Communications, 64(1):98–108, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190052M.

Kawabata:1995:NVM

[Kaw95] Setsuya Kawabata. A new version of the multi-dimensional integration and event generation package BASES/SPRING. Computer Physics Communications, 88(2–3):309–326, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500028E.

Kleiss:1994:BMC

[KB94] R. Kleiss and H. Burkhardt. BBBREM — Monte Carlo simulation of radiative Bhabha scattering in the very forward direction. Computer Physics Communications, 81(3):372–380, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559490085X.

Kamenski:1998:CKW

[KB98] I. V. Kamenski and G. G. Borg. A 1D cylindrical kinetic wave code for helicon plasma sources. Computer Physics Communications, 113(1):10–32, September 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000770.

Kitowski:1999:PVM

[KB99] J. Kitowski and K. Boryczko. Parallel visualization for molecular dynamics simulations. Computer Physics Communications, 121–122:678, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700832.

Kublbeck:1990:FAC

[KBD90] J. Küblbeck, M. Böhm, and A. Denner. Feyn arts — computer-algebraic generation of Feynman graphs and amplitudes. Computer Physics Communications, 60(2):165–180, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090001H.

Krech:1998:FSD

[KBL98] M. Krech, Alex Bunker, and D. P. Landau. Fast spin dynamics algorithms for classical spin systems. Computer Physics Communications, 111(1–3):1–13, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000095.

Kalyanaraman:1998:TDW

[KC98] C. Kalyanaraman and David C. Clary. Time-dependent wavepacket calculations on polyatomic reactive scattering. Computer Physics Communications, 108(2–3):191–199, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001148.

Khanfir:1996:RSE

[KCC96] R. Khanfir, G. Chanteur, and J. P. Croisille. Resolution of the system of equations of ideal magnetohydrodynamics by a finite volume kinetic-type method. Computer Physics Communications, 98(3):301–316, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000951.

Knight:1992:FCT

[KCH92] P. J. Knight, M. Cox, and T. C. Hender. A filament code for tokamak axisymmetric stability. Computer Physics Communications, 71(1–2):99–109, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290075A.

Kim:1999:NCS

[KCK99] Jai Sam Kim, Yoon Sang Chae, and Jung Dae Kim. Numerical computation of solar neutrino flux attenuated by the MSW mechanism. Computer Physics Communications, 120(1):41– 56, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002180.

Kaabar:1995:DVM

[KD95] W. Kaabar and R. Devonshire. DIFCARS: a versatile model of CARS signal generation. Computer Physics Communications, 86(1–2):162–174, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400125L.

Kozlowski:1998:ADI

[KDH^{+}98] Thomas Kozlowski, Ed Desmond, John Haggerty, Hyon-Joo Kehayias, Martin Purschke, and Chris Witzig. Analysis, design, and implementation of PHENIX on-line computing systems software using Shlaer–Mellor Object-Oriented Analysis and Recursive Design. Computer Physics Communications, 110(1–3):164–169, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001720.

Kalia:1993:MDS

[KdLNV93] Rajiv K. Kalia, Simon de Leeuw, Aiichiro Nakano, and Priya Vashishta. Molecular-dynamics simulations of Coulombic systems on distributed-memory MIMD machines. Computer Physics Communications, 74(3):316–326, March 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900155.

Keravnou:1990:WDE

[Ker90a] E. T. Keravnou. What is a deep expert system? an analysis of first-generation limitations and a review of second-generation architectures. Computer Physics Communications, 61(1–2):3–12, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590901027.

Kerpedjiev:1990:TWF

[Ker90b] Stephan M. Kerpedjiev. Transformation of weather forecasts from textual to cartographic form. Computer Physics Communications, 61(1–2):246–256, November 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090124J.

Kuchta:1999:MCS

[KF99] B. Kuchta and L. Firlej. Monte Carlo simulations of FCC and HCP molecular plastic crystals. Computer Physics Communications, 121–122:681, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700868.

Kramer:1995:HFT

[KH95] K. M. Kramer and W. N. G. Hitchon. A highly flexible tool for large scale computational problems in applied physics. Computer Physics Communications, 85(2):167–175, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400132L.

Kramer:1996:SMH

[KH96] K. M. Kramer and W. N. G. Hitchon. Strategies for mesh-handling and model specification within a highly flexible simulation framework. Computer Physics Communications, 93(2–3):179–211, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001301.

Kiviniemi:1997:CCM

[KH97] T. P. Kiviniemi and J. A. Heikkinen. Choice of constants of motion coordinates in numerical solving of the threedimensional Fokker–Planck equation for tokamaks. Computer Physics Communications, 107(1–3):149–154, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001069.

Kashiwagi:1999:NAG

[KI99] M. Kashiwagi and S. Ido. Numerical analyses of generation and features of ion source plasma. Computer Physics Communications, 121–122:673, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700789.

Katayama:1998:BCM

[KIMS98] Nobu Katayama, Ryosuke Itoh, Atsushi Manabe, and Takashi Sasaki. Belle computing model. Computer Physics Communications, 110(1–3):22–25, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001483.

Kinzel:1999:SPN

[Kin99] Wolfgang Kinzel. Statistical physics of neural networks. Computer Physics Communications, 121–122:86–93, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002878.

Kishk:1991:FIS

[Kis91] Ahmed A. Kishk. Formulations of impedance surfaces coated with dielectric materials. Computer Physics Communications, 68(1–3):145–156, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190197S.

Kandil:1991:CCQ

[KK91] Osama A. Kandil and Hamdy A. Kandil. Computation of compressible quasi-axisymmetric slender vortex flow and breakdown. Computer Physics Communications, 65(1–3): 164–172, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190169L.

Kroger:1995:PCA

[KK95] Sophie Kröger and Martin Kröger. A program to compute the angular coefficients of the relativistic one-electron hyperfine structure parameters. Computer Physics Communications, 90(2–3):381–387, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500078T. See erratum [KK97].

Kisel:1996:ENB

[KK96a] I. Kisel and V. Kovalenko. Elastic net for broken multiple scattered tracks. Computer Physics Communications, 98(1–2):45– 51, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000884.

Kroger:1996:NAO

[KK96b] Martin Kröger and Bernd Kröger. A novel algorithm to optimize classification trees. Computer Physics Communications, 95(1):58–72, May 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000021.

Kroger:1997:EBP

[KK97] Sophie Kröger and Martin Kröger. Erratum to A program to compute the angular coefficients of the relativistic one-electron hyperfine structure parameters [Comput. Phys. Commun. 90 (1995) 381–387]. Computer Physics Communications, 103(1):97–99, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000349. See [KK95].

Karolyi:1999:GML

[KK99] Antal Károlyi and János Kertész. Granular medium lattice gas model: the algorithm. Computer Physics Communications, 121–122:290–293, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003367.

Kim:1992:NPN

[KKH^{+}92] B. T. Kim, M. C. Kyum, S. W. Hong, M. H. Park, and T. Udagawa. NLOM — a program for nonlocal optical model calculations. Computer Physics Communications, 71(1–2):150–158, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290080I.

Kandhai:1998:LBH

[KKH^{+}98] D. Kandhai, A. Koponen, A. G. Hoekstra, M. Kataja, J. Timonen, and P. M. A. Sloot. Lattice-Boltzmann hydrodynamics on parallel systems. Computer Physics Communications, 111(1–3):14–26, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000253.

Kanda:1991:MAC

[KKK91] S. Kanda, S. Kim, and K. Kondo. Moment analysis of charged fragment distributions and separation of quark and gluon jets. Computer Physics Communications, 67(2):223–232, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190018G.

Kobayashi:1995:FPN

[KKK95] R. Kobayashi, M. Konuma, and S. Kumano. FORTRAN program for a numerical solution of the nonsinglet Altarelli–Parisi equation. Computer Physics Communications, 86(3):264–278, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400159Y.

Kirianov:1999:DMF

[KKKK99] D. Kirianov, E. Kirianova, N. Kozlov, and V. Kuznetsov. Dynamical model of forest. Computer Physics Communications, 121–122:676, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700819.

Kneller:1995:NPP

[KKKS95] Gerald R. Kneller, Volker Keiner, Meinhard Kneller, and Matthias Schiller. nMOLDYN: a program package for a neutron scattering oriented analysis of Molecular Dynamics simulations. Computer Physics Communications, 91(1–3):191–214, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500048K.

Kertesz:1999:LSS

[KKKT99] J. Kertesz, L. Kullmann, K. Kimmo, and J. Tolyi. Limits of scaling in the statistics of returns in stock market indices. Computer Physics Communications, 121–122:675, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700807.

Karolyi:1999:FSD

[KKM^{+}99] A. Karolyi, J. Kertesz, H. Makse, S. Havlin, and H. E. Stanley. Filling a silo: density profiles and friction induced segregation. Computer Physics Communications, 121–122: 672, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700777.

Kiryanov:1999:CFD

[KKMS99] Yu. F. Kir’yanov, M. L. Kudryavtseva, M. V. Maslov, and I. V. Shestakova. On the conservative finite difference scheme for 2D nonlinear Schrödinger equation. Computer Physics Communications, 121–122:677, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700820.

Kreuzberger:1990:SSC

[KKS90] T. Kreuzberger, W. Kummer, and M. Schweda. Susycal — symbolic computations in supersymmetric theories. Computer Physics Communications, 58(1–2):89–104, February/ March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090138Q.

Kohler:1999:FCS

[KKY99] H. S. Köhler, N. H. Kwong, and Hashim A. Yousif. A Fortran code for solving the Kadanoff–Baym equations for a homogeneous fermion system. Computer Physics Communications, 123(1–3):123–142, December 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900260X.

Kumano:1992:FPN

[KL92] S. Kumano and J. T. Londergan. A FORTRAN program for numerical solution of the Altarelli–Parisi equations by the Laguerre method. Computer Physics Communications, 69(2–3):373–396, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290176Y.

Kurpick:1993:CIP

[KL93] Peter Kürpick and Hans Jürgen Lüdde. Calculation of inclusive probabilities from single-particle amplitudes. Computer Physics Communications, 75(1–2):127–134, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390170H.

Kharraziha:1999:LV

[KL99a] Hamid Kharraziha and Leif Lönnblad. LDCMC version 1.0. Computer Physics Communications, 123(1–3):153–163, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002301.

Kolafa:1999:TLF

[KL99b] J. Kolafa and S. Labík. The thermodynamic limit in fluid simulations and the bridge function. Computer Physics Communications, 121–122:680, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700856.

Kleiss:1992:ACC

[Kle92]  Ronald Kleiss.  Average-case complexity distributions:

a generalization of the woźniakowski lemma for multidimensional numerical integration. Computer Physics Communications, 71(1–2):39–46, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290070F.

Klioner:1998:NSI

[Kli98] Sergei A. Klioner. New system for indicial computation and its applications in gravitational physics. Computer Physics Communications, 115(2–3):231–244, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001611.

Kruger:1991:SSP

[KLS91] Peter Krüger, Mathias Lüke, and Angelika Szameit. SIMLYS — a software package for trajectory analysis of molecular dynamics simulations. Computer Physics Communications, 62(2–3):371–380, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190108W.

Kobus:1996:NHF

[KLS96] Jacek Kobus, Leif Laaksonen, and Dage Sundholm. A numerical Hartree–Fock program for diatomic molecules. Computer Physics Communications, 98(3):346–358, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000987.

Kato:1991:NMC

[KM91] Kiyoshi Kato and Tomo Munehisa. NLLjet: a Monte Carlo code for e^{+}e^{−} QCD jets including next-to-leading order terms. Computer Physics Communications, 64(1):67–97, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190051L.

Knoll:1995:NKM

[KM95] D. A. Knoll and P. R. McHugh. Newton–Krylov methods applied to a system of convection-diffusion-reaction equations. Computer Physics Communications, 88(2–3):141–160, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500062K.

Katsanevas:1998:SMC

[KM98]  Stavros Katsanevas and Peter Morawitz.  SUSYGEN

2.2- A Monte Carlo event generator for MSSM sparticle production at e^{+}e^{−} colliders. Computer Physics Communications, 112(2–3):227–269, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000435.

Koch:1998:SMZ

[KML98] Justus H. Koch, Hubertus R. Mall, and Stefan Lenz. Stochastic methods for zero energy quantum scattering. Computer Physics Communications, 108(2–3):115–146, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001306.

Katoh:1999:SCH

[KMT^{+}99] H. Katoh, K. Matsuo, K. Takayama, K. Iinuma, and Y. Sudo. Simulation code for high power microwave propagation through the atmosphere. Computer Physics Communications, 121–122:674, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700790.

Kang:1991:HAH

[KMz91] Feng Kang and Qin Meng-zhao. Hamiltonian algorithms for Hamiltonian systems and a comparative numerical study. Computer Physics Communications, 65(1–3):173–187, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190170P.

Kutteh:1995:EDI

[KN95a] Ramzi Kutteh and John B. Nicholas. Efficient dipole iteration in polarizable charged systems using the cell multipole method and application to polarizable water. Computer Physics Communications, 86(3):227–235, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400019X.

Kutteh:1995:ICM

[KN95b] Ramzi Kutteh and John B. Nicholas. Implementing the cell multipole method for dipolar and charged dipolar systems.

Computer Physics Communications, 86(3):236–254, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594000203.

Kniehl:1990:QCV

[Kni90] Bernd A. Kniehl. QCD corrections to vector boson self-energies in the standard model. Computer Physics Communications, 58(3):293–303, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900659.

Kral:1995:SSM

[KNK95] Daniel Král, Pavel Neogrády, and Vladimir Kellö. Simple sparse matrix multiplication algorithm. Computer Physics Communications, 85(2):213–216, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400120Q.

Kotler:1991:MTD

[KNN91] Zvi Kotler, Eyal Neria, and Abraham Nitzan. Multiconfiguration time-dependent self-consistent field approximations in the numerical solution of quantum dynamical problems. Computer Physics Communications, 63(1–3):243–258, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190253H.

Knowles:1990:LAC

[Kno90] I. G. Knowles. A linear algorithm for calculating spin correlations in hadronic collisions. Computer Physics Communications, 58(3):271–284, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900637.

Knobloch:1998:AC

[Kno98] Jürgen Knobloch. ATLAS computing. Computer Physics Communications, 110(1–3):26–31, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001495.

Kobus:1994:VAM

[Kob94] Jacek Kobus. Vectorizable algorithm for the (multicolour) successive overrelaxation method. Computer Physics Communications, 78(3):247–255, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900035.

Koike:1992:EFA

[Koi92] F. Koike. Explicit formulae of angular momentum coupling coefficients. Computer Physics Communications, 72 (2–3):154–164, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290147Q.

Kooy:1993:CMO

[Koo93] H. J. Kooy, Jr. The construction of matrices of orthogonal operators for d ^{N} configurations. Computer Physics Communications, 74(1):142–148, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390111O.

Korek:1999:ODS

[Kor99] M. Korek. A one directional shooting method for the computation of diatomic centrifugal distortion constants. Computer Physics Communications, 119(2–3):169–178, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001805.

Kot:1992:ASM

[Kot92] M. Kot. Averaging system for measurement of semiconductor lasers and LEDs optical response. Computer Physics Communications, 73(1–3):217–219, December 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290043X.

Kotrla:1996:NST

[Kot96] M. Kotrla. Numerical simulations in the theory of crystal growth. Computer Physics Communications, 97(1–2):82–100, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000239.

Kleiss:1994:WOM

[KP94] Ronald Kleiss and Roberto Pittau. Weight optimization in multichannel Monte Carlo. Computer Physics Communications, 83(2–3):141–146, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900434.

Kopf:1997:MTS

[KPD97] A. Kopf, W. Paul, and B. Dünweg. Multiple time step integrators and momentum conservation. Computer Physics Communications, 101(1–2):1–8, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001658.

Kulander:1991:EPT

[KR91] K. C. Kulander and T. N. Rescigno. Effective potentials for time-dependent calculations of multiphoton processes in atoms. Computer Physics Communications, 63(1–3):523–528, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190273N.

Kirsch:1995:SPM

[KR95a] Stefan Kirsch and Tord Riemann. SMATASY — a program for the model independent description of the Z resonance. Computer Physics Communications, 88(1):89–107, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595000169.

Krall:1995:TIP

[KR95b] Nicholas A. Krall and Stephen E. Rosenthal. A technique for including 3D plasma turbulence in a two-dimensional plasma simulation. Computer Physics Communications, 87(1–2):95– 116, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400181Z.

Kutner:1999:BEC

[KR99] Ryszard Kutner and Marcin Regulski. Bose–Einstein condensation shown by Monte Carlo simulation. Computer Physics Communications, 121–122:586–590, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004129.

Kreckel:1997:PAM

[Kre97] Richard Kreckel. Parallelization of adaptive MC integrators. Computer Physics Communications, 106(3):258–266, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000994.

Krebs:1999:RDS

[Kre99] Stefan Krebs. A review on the derivation of the spin-Restricted Hartree–Fock (RHF) Self-Consistent Field (SCF) equations for open-shell systems. Description of different methods to handle the off-diagonal Lagrangian multipliers coupling closed and open shells. Computer Physics Communications, 116(2–3):137–277, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001003.

Kronert:1990:IIS

[Krö90] Günther Krönert. Importance of the ISO standard 8613 for document interchange: Basic standard, functional standards, initial experience, benefits. Computer Physics Communications, 61(1–2):209–218, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090118K.

Kroes:1992:RRE

[Kro92] G. J. Kroes. The royal road to an energy-conserving predictorcorrector method. Computer Physics Communications, 70(1):

41–52, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290089H.

Kroger:1996:OCT

[Krö96] Martin Kröger. Optimization of classification trees: strategy and algorithm improvement. Computer Physics Communications, 99(1):81–93, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001233.

Kroger:1999:EHA

[Krö99] Martin Kröger. Efficient hybrid algorithm for the dynamic creation of wormlike chains in solutions, brushes, melts and glasses. Computer Physics Communications, 118(2–3):278– 298, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559800160X.

Kotschenreuther:1995:CIV

[KRT95] Mike Kotschenreuther, G. Rewoldt, and W. M. Tang. Comparison of initial value and eigenvalue codes for kinetic toroidal plasma instabilities. Computer Physics Communications, 88(2–3):128–140, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500035E.

Kruger:1996:IFB

[Kru96] A. Kruger. Implementation of a fast box-counting algorithm. Computer Physics Communications, 98(1–2):224–234, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559600080X.

Kruse:1998:COE

[Kru98] Andrés Kruse. CMS online event filter software. Computer Physics Communications, 110(1–3):103–106, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001616.

Kravanja:1998:ZMS

[KRVZ98] P. Kravanja, O. Ragos, M. N. Vrahatis, and F. A. Zafiropoulos. ZEBEC: a mathematical software package for computing simple zeros of Bessel functions of real order and complex argument. Computer Physics Communications, 113(2–3):220–238, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000642.

Kniehl:1992:ZPC

[KS92a] Bernd A. Kniehl and Robin G. Stuart. Z0POLE — a program to calculate the electroweak and QCD radiative corrections to e^{+}e^{−} → ff near the Z^{0} resonance. Computer Physics Communications, 72(2–3):175–220, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290149S.

Kruger:1992:SV

[KS92b] Peter Krüger and Angelika Szameit. SIMLYS version 2.0. Computer Physics Communications, 72(2–3):265–268, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290153P.

Kokubo:1993:HEB

[KS93] Sachiko Kokubo and Akio Sugamoto. High-energy behavior of the membrane scattering amplitudes. Computer Physics Communications, 75(3):311–338, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390047G.

Kalkreuter:1996:ACG

[KS96] Thomas Kalkreuter and Hubert Simma. An accelerated conjugate gradient algorithm to compute low-lying eigenvalues — a study for the Dirac operator in SU(2) lattice QCD. Computer Physics Communications, 93(1):33–47, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001263.

Koutras:1998:ADW

[KS98] A. Koutras and Jim E. F. Skea. An algorithm for determining whether a space–time is homothetic. Computer Physics Communications, 115(2–3):350–362, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001325.

Kottapalli:1991:ACW

[KSA^{+}91] K. Kottapalli, T. K. Sarkar, R. Adve, Y. Hua, E. K. Miller, and G. J. Burke. Accurate computation of wideband response of electromagnetic systems utilizing narrowband information. Computer Physics Communications, 68(1–3):126–144, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190196R.

Kwong:1991:SQE

[KSG91] N. H. Kwong, K. J. Schaudt, and J. D. Garcia. Solution of quantum evolution equations by finite-difference. Computer Physics Communications, 63(1–3):171–178, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190247I.

Kwiatkowski:1992:HEG

[KSM92] A. Kwiatkowski, H. Spiesberger, and H.-J. Möhring. HER-ACLES: an event generator for ep interactions at HERA energies including radiative processes. Computer Physics Communications, 69(1):155–172, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290136M.

Kulyk:1999:XRD

[KSSB99] Y. Kulyk, A. Staovoitov, A. Sapelkin, and S. Bayliss. Xray diffraction of diamond for improved high pressure EX-AFS. Computer Physics Communications, 121–122:683, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700881.

Kalogeropoulos:1997:ISO

[KSZ97] T. E. Kalogeropoulos, Y. G. Saridakis, and M. S. Zakynthinaki. Improved stochastic optimization algorithms for adaptive optics. Computer Physics Communications, 99 (2–3):255–269, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001014.

Kasahara:1991:SPF

[KT91] K. Kasahara and S. Torii. A subroutine package for fast simulation of air showers and response of surface detectors. Computer Physics Communications, 64(1):109–120, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190053N.

Kawata:1996:GGO

[KTM96] S. Kawata, N. Takura, and Y. Manabe. Grid generation with orthogonality and uniformity of line-spacing changing ratio. Computer Physics Communications, 94(1):19–24, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001387.

Kim:1998:MCO

[KTT98] Jai Sam Kim, J. C. Tolédano, and P. Tolédano. Monte Carlo optimization applied to symmetry breaking. Computer Physics Communications, 109(2–3):207–226, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001392.

Kudrtyavtsev:1999:NMC

[Kud99] A. Yu. Kudrtyavtsev. New method to calculate the intersection volume of 3D convex bodies aimed for 3D computer simulation of the fluid dynamics. Computer Physics Communications, 121–122:682, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670087X.

Kulhanek:1997:TMN

[Kul97] J. Kulhánek. A transformation method for numerical solution of the Schrödinger equation. Computer Physics Communications, 105(1):37–41, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000088.

Kupenova:1999:SMD

[Kup99] T. Kupenova. Smooth multi-dimensional approximation. Computer Physics Communications, 121–122:684, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700893.

Kutteh:1999:NAM

[Kut99] Ramzi Kutteh. New approaches for molecular dynamics simulations with nonholonomic constraints. Computer Physics Communications, 119(2–3):159–168, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598002045.

Krauss-Varban:1990:SAF

[KV90] D. Krauss-Varban. Some aspects of finite element calculations applied to ULF magnetospheric cavity waves. Computer Physics Communications, 59(1):85–94, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090158W.

Knecht:1999:GNS

[KV99] K. Knecht and H. Verschelde. A general numerical scheme for Feynman diagrams. Computer Physics Communications, 121–122:679, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700844.

Kermode:1990:FPI

[KW90] J. P. Kermode and D. Weaire. 2D-FROTH: a program for the investigation of 2-dimensional froths. Computer Physics Communications, 60(1):75–109, August 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090080K.

Kaeding:1996:PGT

[KW96] Thomas A. Kaeding and H. Thomas Williams. Program for generating tables of SU(3) ⊃ SU(2) ⊗ U(1) coupling coefficients. Computer Physics Communications, 98(3):398–414, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000859.

Ktitarev:1999:CAG

[KW99] Dmitri V. Ktitarev and Dietrich E. Wolf. A cellular automaton for grains in a rotating drum. Computer Physics Communications, 121–122:303–305, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003392.

Kohler:1996:FCA

[KWS^{+}96] Bernd Kohler, Steffen Wilke, Matthias Scheffler, Robert Kouba, and Claudia Ambrosch-Draxl. Force calculation and atomic-structure optimization for the full-potential linearized augmented plane-wave code WIEN. Computer Physics Communications, 94(1):31–48, March 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001395.

Kazeminezhad:1995:PMU

[KZS95] F. Kazeminezhad, S. Zalesak, and D. Spicer. A particle model on an unstructured mesh. Computer Physics Communications, 90(2–3):267–292, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500091S.

Llobet:1990:SP

[LABV90] X. Llobet, K. Appert, A. Bondeson, and J. Vaclavik. On spectral pollution. Computer Physics Communications, 59(2):199– 216, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590901706.

Lagana:1992:IPC

[Lag92]  Antonio Laganà.  The impact of parallel computing

on reactive scattering calculations. Computer Physics Communications, 70(2):223–241, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290190A.

Lagana:1999:ICD

[Lag99] Antonio Laganà. Innovative computing and detailed properties of elementary reactions using time independent approaches. Computer Physics Communications, 116(1):1–16, January 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001842.

Laina:1998:XFD

[Lai98] A. Laina. Xfey, a Feynman diagram editor. Computer Physics Communications, 111(1–3):217–242, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000277.

Lamberti:1996:ISSa

[Lam96a] C. Lamberti. Interface simulation of strained and non-abrupt III-V quantum wells. part 1: band profile calculation. Computer Physics Communications, 93(1):53–81, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001182.

Lamberti:1996:ISSb

[Lam96b] C. Lamberti. Interface simulation of strained and nonabrupt III-V quantum wells. part 2: energy level calculation versus experimental data. Computer Physics Communications, 93(1):82–119, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001190.

Lammel:1998:CMC

[Lam98] Stephan Lammel. Computing models of CDF and DØ in run II. Computer Physics Communications, 110(1–3):32–37, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001501.

Langdon:1992:EGL

[Lan92] A. Bruce Langdon. On enforcing Gauss’ law in electromagnetic particle-in-cell codes. Computer Physics Communications, 70(3):447–450, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592901058.

Lankford:1998:TCD

[Lan98] A. J. Lankford. Trends and challenges in data acquisition and control systems summary of CHEP97 Session B. Computer Physics Communications, 110(1–3):69–75, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001550.

Landau:1999:CPC

[Lan99a] Rubin H. Landau. Computational physics: a course and a Web-enhanced book. Computer Physics Communications, 121–122:685, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670090X.

Landau:1999:WEC

[Lan99b] Rubin H. Landau. Web, education, and computational physics; good, bad, and ugly. Computer Physics Communications, 121–122:550–556, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004051.

Lee:1991:PFH

[LB91] Jae Koo Lee and C. K. Birdsall. Particle-fluid hybrid simulations for weak beam-plasma interactions. Computer Physics Communications, 64(2):214–220, May 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190032G.

Lapenta:1995:CNP

[LB95] Giovanni Lapenta and J. U. Brackbill. Control of the number of particles in fluid and MHD particle in cell methods. Computer Physics Communications, 87(1–2):139–154, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400180A.

Liu:1999:IPC

[LB99] D. H. Liu and A. Bondeson. Improved poloidal convergence of the MARS code for MHD stability analysis. Computer Physics Communications, 116(1):55–64, January 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001453.

Leviatan:1991:AES

[LBB91] Yehuda Leviatan, Amir Boag, and Alona Boag. Analysis of electromagnetic scattering using a current model method. Computer Physics Communications, 68(1–3):331– 345, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591902072.

Liem:1991:MDS

[LBC91] S. Y. Liem, D. Brown, and Julian H. R. Clarke. Molecular dynamics simulations on distributed memory machines. Computer Physics Communications, 67(2):261–267, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190021C.

Lamberti:1993:MSS

[LBC^{+}93] C. Lamberti, S. Bordiga, G. Cerrato, C. Morterra, D. Scarano, G. Spoto, and A. Zecchina. MINUIT subroutine for spectra deconvolution. Computer Physics Communications, 74 (1):119–141, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390110X.

Luczak:1999:MBD

[LBDL99] F. Luczak, F. Brosens, J. T. Devreese, and L. F. Lemmens. Many-body diffusion algorithm for interacting harmonic fermions. Computer Physics Communications, 121–122:449– 451, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559900380X.

Lutjens:1992:AME

[LBR92] H. Lütjens, A. Bondeson, and A. Roy. Axisymmetric MHD equilibrium solver with bicubic Hermite elements. Computer Physics Communications, 69(2–3):287–298, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290167W.

Lutjens:1996:CCT

[LBS96] H. Lütjens, A. Bondeson, and O. Sauter. The CHEASE code for toroidal MHD equilibria. Computer Physics Communications, 97(3):219–260, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559600046X.

Lai:1990:ECJ

[LC90a] Shan-Tao Lai and Ying-Nan Chiu. Exact computation of the 3-j and 6-j symbols. Computer Physics Communications, 61(3):350–360, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900497.

Lai:1990:NBT

[LC90b] Shan-Tao Lai and Ying-Nan Chiu. A note on Biedenharn’s triple correlation-angle functions. Computer Physics Communications, 61(3):331–349, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590900486.

Lai:1992:ECJ

[LC92] Shan-Tao Lai and Ying-Nan Chiu. Exact computation of the 9-j symbols. Computer Physics Communications, 70(3):544– 556, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290115F.

Lister:1993:GOD

[LC93a] G. G. Lister and S. E. Coe. GLOMAC : a one dimensional numerical model for steady state low pressure mercury-noble gas discharges. Computer Physics Communications, 75(1–2):160– 184, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390173A.

Lubian:1993:PCE

[LC93b] J. Lubian and R. Cabezas. A program for the calculation of energy spectra and reduced nuclear matrix elements in the frame of the davydov-chaban model. Computer Physics Communications, 76(1):98–112, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390124U.

Liu:1996:IPC

[LC96a] Zhiqiang Liu and Sam J. Cipolla. ISICS: a program for calculating K-, L- and M-shell cross sections from ECPSSR theory using a personal computer. Computer Physics Communications, 97(3):315–330, September 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000689.

Louvel:1996:PDH

[LC96b] S. Louvel and J.-F. Chamayou. Packing and depacking histograms with statistical processing. Computer Physics Communications, 93(2–3):289–302, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001298.

Ludde:1991:MOP

[LD91] H. J. Lüdde and R. M. Dreizler. Microscopic optical potentials for time-dependent quantum systems. Computer Physics Communications, 63(1–3):345–350, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190261I.

Lemon:1995:FMD

[LDOO95] A. P. Lemon, P. Dauber-Osguthorpe, and D. J. Osguthorpe. FOCUS: a molecular dynamics analysis program: New features for the characterisation of lipid bilayers and solvated systems. Computer Physics Communications, 91(1–3):97– 109, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500044G.

Losada:1999:CPO

[LEBB99] J. C. Losada, J. M. Estebaranz, R. M. Benito, and F. Borondo. Calculation of periodic orbits in classically chaotic systems by frequency analysis. Computer Physics Communications, 121–122:691, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700959.

Lee:1990:ATF

[Lee90a] Deok Kyo Lee. Application of theta functions for numerical evaluation of complete elliptic integrals of the first and second kinds. Computer Physics Communications, 60(3):319–327, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090031U.

Lee:1990:TDM

[Lee90b] L. C. Lee. Time-dependent magnetic reconnection: Twoand three-dimensional MHD simulations. Computer Physics Communications, 59(1):163–174, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090166X.

Lee:1992:CCP

[Lee92] D. K. Lee. Calculation of coefficients in a power-series expansion of the nome q = exp[−πK(k′)K(k)]. Computer Physics Communications, 70(2):292–296, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901944.

Leherte:1999:MAE

[Leh99] L. Leherte. Multiresolution analysis of electron density maps. Computer Physics Communications, 121–122:688, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700923.

Lemoine:1996:NAM

[Lem96] Didier Lemoine. Nonreactive atom/molecule-surface scattering within the finite basis wave packet method. Computer Physics Communications, 97(3):331–344, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000641.

Lemoine:1997:OCS

[Lem97] Didier Lemoine. Optimal cylindrical and spherical Bessel transforms satisfying bound state boundary conditions. Computer Physics Communications, 99(2–3):297–306, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001415.

Lenoski:1998:MDO

[Len98] Daniel Lenoski. Multiprocessor design options and the Silicon Graphics S^{2} MP architecture. Computer Physics Communications, 110(1–3):59–68, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001549.

Leszczynski:1999:NMR

[Les99] Henryk Leszczyński. Numerical methods for reaction-diffusion equations with functionals. Computer Physics Communications, 121–122:689, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700935.

Levine:1990:LGR

[Lev90] Michael J. S. Levine. A L^{A}TEX graphics routine for drawing Feynman diagrams. Computer Physics Communications, 58(1–2):181–198, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090144P.

Larsen:1992:EEF

[LF92] Finn Larsen and Henrik Flyvbjerg. Efficient evaluation of Feynman diagrams on lattices. Computer Physics Communications, 69(1):59–64, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290128L.

Luchsinger:1993:VRM

[LG93a] Rolf Luchsinger and Christoph Grab. Vertex reconstruction by means of the method of Kalman filtering. Computer Physics Communications, 76(3):263–280, August 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390055H.

Lynas-Gray:1993:VFS

[LG93b] A. E. Lynas-Gray. VOIGTL — a fast subroutine for voigt function evaluation on vector processors. Computer Physics Communications, 75(1–2):135–142, April 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901718.

Lehoucq:1998:VEF

[LGZL98] R. B. Lehoucq, S. K. Gray, D.-H. Zhang, and J. C. Light. Vibrational eigenstates of four-atom molecules: a parallel strategy employing the implicitly restarted Lanczos method. Computer Physics Communications, 109(1):15–25, March 1998.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000022.

Lomba:1992:HPC

[LH92] Enrique Lomba and Johan S. Høye. HNCR — a program to calculate the structure and thermodynamics of binary mixtures of charged hard spheres. Computer Physics Communications, 69(2–3):420–428, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901807.

Law:1997:NPI

[LH97] Mark M. Law and Jeremy M. Hutson. I-NoLLS: a program for interactive nonlinear least-squares fitting of the parameters of physical models. Computer Physics Communications, 102(1–3):252–268, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000131.

Linke:1995:SVL

[LHA95] Andreas Linke, Dieter W. Heermann, and Peter Altevogt. Simulating very large Ising systems for short timescales. Computer Physics Communications, 90(1):66–72, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559592816Y.

Larson:1995:CFP

[LHL95] David J. Larson, Dennis W. Hewett, and A. Bruce Langdon. Correction factors for PIC accumulation on radial grids. Computer Physics Communications, 90(2–3):260–266, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500104N.

Lu:1998:PCP

[LHL98] Dinghui Lu, Guangliang He, and Rubin H. Landau. PiN: Computation of pion-nucleon scattering and bound states with the color dielectric quark model. Computer Physics Communications, 108(1):75–89, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001227.

Levitt:1995:PEF

[LHSD95] Michael Levitt, Miriam Hirshberg, Ruth Sharon, and Valerie Daggett. Potential energy function and parameters for simulations of the molecular dynamics of proteins and nucleic acids in solution. Computer Physics Communications, 91(1–3):215–231, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500049L.

Li:1995:NSK

[Li95] Ping-Wah Li. On the numerical study of the KdV equation by the Semi-Implicit and Leap-frog Method. Computer Physics Communications, 88(2–3):121–127, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500060S.

Liang:1995:PHF

[Lia95] Shoudan Liang. A perfect hashing function for exact diagonalization of many-body systems of identical particles. Computer Physics Communications, 92(1):11–15, November 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500108R.

Lill:1993:CCL

[Lil93] J. V. Lill. The construction of crystalline lattice vectors in specified orientations for atomistic simulations. Computer Physics Communications, 78(1–2):130–140, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465593901497.

Lill:1994:IMD

[Lil94] J. V. Lill. The integration of molecular dynamics simulations with imposed temperature and stress. Computer Physics Communications, 79(2):219–248, April 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900701.

Lima:1991:NPC

[Lim91a] P. M. Lima. A new program for calculating matrix elements in atomic structure. Computer Physics Communications, 66(1):99–114, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591900119.

Lima:1991:PDR

[Lim91b] P. M. Lima. A program for deriving recoupling coefficients formulae. Computer Physics Communications, 66(1):89–98, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190010I.

Lim:1996:ZVA

[Lim96] Kieran F. Lim. ZhuKè : a vibrationally-adiabatic impulsive dissociation model. Computer Physics Communications, 97(3):345–356, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000598.

Lin:1990:NCF

[Lin90] Yuanzhang Lin. Numerical calculation of force-free magnetic field for solar active regions and its application to prediction of solar flares. Computer Physics Communications, 59(1):139– 143, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090163U.

Ling:1991:TDM

[Lin91] R. T. Ling. A time-dependent method for the numerical solution of wave equations in electromagnetic scattering problems. Computer Physics Communications, 68(1–3):213–223, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190201U.

Liolios:1997:ANM

[Lio97] Th. E. Liolios. Algebraic and numerical manipulation of the even-power-series central potentials by means of the hypervirial theorems technique. Computer Physics Communications, 105(2–3):254–262, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000532.

Lewis:1996:UHP

[LK96] H. Ralph Lewis and Peter J. Kostelec. The use of Hamilton’s principle to derive time-advance algorithms for ordinary differential equations. Computer Physics Communications, 96(2–3):129–151, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000392.

Lamarche:1990:EVI

[LL90] François Lamarche and Claude Leroy. Evaluation of the volume of intersection of a sphere with a cylinder by elliptic integrals. Computer Physics Communications, 59(2):359–369, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590901843.

Lutjens:1996:CBF

[LL96] H. Lütjens and J. F. Luciani. A class of basis functions for non-ideal magnetohydrodynamic computations. Computer Physics Communications, 95(1):47–57, May 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000069.

Ledue:1999:TDT

[LL99a] D. Ledue and D. P. Landau. Temperature-driven transition of the eight-state Potts model on the quasiperiodic octagonal tiling. Computer Physics Communications, 121–122:686–687, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700911.

Lobaskin:1999:CSH

[LL99b] V. Lobaskin and P. Linse. Computer simulation of highly asymmetric electrolytes with charge asymmetry 60 : 1 and 60 : 2. Computer Physics Communications, 121–122:690, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506700947.

Lyubartsev:1999:RPI

[LL99c] Alexander P. Lyubartsev and Aatto Laaksonen. Reconstruction of pair interaction potentials from radial distribution functions. Computer Physics Communications, 121–122:57–59, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002805.

Lagaris:1997:ANN

[LLF97] I. E. Lagaris, A. Likas, and D. I. Fotiadis. Artificial neural network methods in quantum mechanics. Computer Physics Communications, 104(1–3):1–14, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000544.

Liu:1991:MMF

[LLM91] C. Liu, Z. Liu, and S. McCormick. Multigrid methods for flow transition in a planar channel. Computer Physics Communications, 65(1–3):188–200, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190171G.

Leyh:1998:NMT

[LLW98] H. Leyh, D. Loffhagen, and R. Winkler. A new multiterm solution technique for the electron Boltzmann equation of weakly ionized steady-state plasmas. Computer Physics Communications, 113(1):33–48, September 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000629.

Lassig:1991:HSS

[LM91a] Michael Lässig and Giuseppe Mussardo. Hilbert space and structure constants of descendant fields in two-dimensional conformal theories. Computer Physics Communications, 66 (1):71–88, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190009A.

Lin:1991:STD

[LM91b] F. J. Lin and J. T. Muckerman. Solution of the timedependent Schrödinger equation employing a basis of explicit discrete-coordinate eigenfunctions: spherical and azimuthal symmetry, adiabaticity, and multiphoton excitation of a rotating Morse oscillator. Computer Physics Communications, 63(1–3):538–568, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190275P.

Lenz:1995:MCA

[LM95] Stefan Lenz and Hubertus Mall. A Monte Carlo approach to zero energy quantum scattering. Computer Physics Communications, 90(2–3):215–234, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595000992.

Labudde:1997:MCS

[LMF97] Dirk Labudde, Reinhard Mahnke, and Vilnis Frischfeld. Monte Carlo simulation of thermodynamic systems with cluster formation under different boundary conditions. Computer Physics Communications, 106(3):181–198, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001045.

LaPenna:1997:MDM

[LMM^{+}97] G. La Penna, V. Minicozzi, S. Morante, G. C. Rossi, and G. Salina. Molecular dynamics with the massively parallel APE computers. Computer Physics Communications, 106(1–2):53–68, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000866.

Lopez:1993:APC

[LMS^{+}93] R. López, M. I. Menéndez, D. Suárez, T. L. Sordo, and J. A. Sordo. ANACAL: a program to carry out a configurational analysis of the wave function of reactive systems. Computer Physics Communications, 76(2):235–249, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390135Y.

Li:1994:PAO

[LMS94] Ming Li, K. J. M. Moriarty, and S. Srinivas. Parallelism analysis and optimization in SPEFY, a programming environment. Computer Physics Communications, 79(1):1–12, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594902259.

Lyons:1994:DMB

[LMW94] J. Lyons, D. Moncrieff, and S. Wilson. Diagrammatic manybody perturbation expansion for atoms and molecules: X. automatic generation and analysis of fifth-order Hugenholtz energy diagrams. Computer Physics Communications, 84 (1–3):91–101, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594902054.

Long:1996:MDS

[LMW96] Lyle N. Long, Michael M. Micci, and Brian C. Wong. Molecular dynamics simulations of droplet evaporation. Computer Physics Communications, 96(2–3):167–172, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000501.

Lonnblad:1992:MEG

[LN92] Leif Lönnblad and Anders Nilsson. The MC++ event generator toolkit — version 0. Computer Physics Communications, 71(1–2):1–14, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900679.

Larsson:1993:AER

[LN93] Lars-Göran Larsson and Erik Nilsson. An Ada environment for relativistic cross-section calculations. Computer Physics Communications, 74(1):41–57, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390104K.

Lobanov:1996:DCF

[Lob96] Yu. Yu. Lobanov. Deterministic computation of functional integrals. Computer Physics Communications, 99(1):59–72, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001130.

Lobanov:1999:CFI

[Lob99] Yu. Yu. Lobanov. Computation of functional integrals in problems of quantum and statistical physics. Computer Physics Communications, 121–122:60–63, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002817.

Lonnblad:1992:AVP

[Lön92] Leif Lönnblad. Ariadne version 4- A program for simulation of QDC cascades implementing the colour dipole model. Computer Physics Communications, 71(1–2):15–31, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290068A.

Lonnblad:1994:CPD

[Lön94] Leif Lönnblad. CLHEP — a project for designing a C++ class library for high energy physics. Computer Physics Communications, 84(1–3):307–316, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902178.

Lonnblad:1999:DSP

[Lön99] Leif Lönnblad. Development strategies for Pythia version 7. A new HEP event generator. Computer Physics Communications, 118(2–3):213–228, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598002008.

Lu:1990:TDP

[LP90] Donghao R. Lu and Kinam Park. A three-dimensional protein graphic program. Computer Physics Communications, 60(2):257–263, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090012P.

Lagaris:1993:CEP

[LP93] I. E. Lagaris and D. G. Papageorgiou. CONVUS — an efficient package for calculating three-dimensional convolutiontype integrals. Computer Physics Communications, 76(1):80– 86, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390122S.

Lorenzini:1999:TNM

[LP99] R. Lorenzini and L. Passoni. Test of numerical methods for the integration of kinetic equations in tropospheric chemistry. Computer Physics Communications, 117(3):241–249, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001039.

Lai:1994:ODS

[LPC94] Shan-Tao Lai, Pancracio Palting, and Ying-Nan Chiu. Onedimensional Schrödinger equation in the harmonic oscillator basis with various potentials. Computer Physics Communications, 82(2–3):221–232, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901708.

Lonnblad:1991:SON

[LPPR91] Leif Lönnblad, Carsten Peterson, Hong Pi, and Thorsteinn Rögnvaldsson. Self-organizing networks for extracting jet features. Computer Physics Communications, 67(2):193–209, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190016E.

Lonnblad:1992:PRH

[LPR92] Leif Lönnblad, Carsten Peterson, and Thorsteinn Rögnvalsson. Pattern recognition in high energy physics with artificial neural networks — JETNET 2.0. Computer Physics Communications, 70(1):167–182, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290099K.

Luana:1993:QMC

[LPR^{+}93] Víctor Luaña, Angel Martín Pendás, JoséManuel Recio, Evelio Francisco, and Margarita Bermejo. Quantum mechanical cluster calculations of ionic materials: the ab initio perturbed ion (version 7) program. Computer Physics Communications, 77(1):107–134, September 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390041A.

Landau:1999:MCS

[LPS99] D. P. Landau, S. Pal, and Y. Shim. Monte Carlo simulations of film growth. Computer Physics Communications, 121–122:341–346, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003501.

Lopez-Pineiro:1992:MRE

[LPSM92] A. Lopez-Piñeiro, M. L. Sanchez, and B. Moreno. MORS-MATEL: a rapid and efficient code to calculate vibrationrotational matrix elements for r-dependent operators of two Morse oscillators. Computer Physics Communications, 70 (2):355–361, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592901988.

Linn:1992:DIG

[LR92a] Stephan L. Linn and Phillip Rulon. A distributed implementation of the Geant Monte Carlo program. Computer Physics Communications, 72(2–3):144–148, November 1992.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290145O.

Lopez:1992:SSS

[LR92b] Jorge Alberto López and Jørgen Randrup. SOS: sequential or simultaneous nuclear multifragmentation. Computer Physics Communications, 70(1):92–96, May 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290094F.

Lowekamp:1993:EAN

[LS93] Bruce B. Lowekamp and John C. Schug. An efficient algorithm for nucleation studies of particles using the octree data structure. Computer Physics Communications, 76(3):281–293, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390056I.

Liu:1992:NFP

[LSZ92] Yu Xin Liu, Hong Zhou Sun, and En Guang Zhao. A new FORTRAN program for the CFP’s of a system with identical bosons. Computer Physics Communications, 70(1):154– 166, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290098J.

Luchini:1994:ECI

[Luc94a] Paolo Luchini. End-correction integration formulae with optimized terminal sampling points. Computer Physics Communications, 83(2–3):236–244, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900515.

Luchini:1994:FAN

[Luc94b] Paolo Luchini. Fourier analysis of numerical integration formulae. Computer Physics Communications, 83(2–3):227–235, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900507.

Lucic:1995:DAS

[Luč95] Vladan Lučić. Dill: an algorithm and a symbolic software package for doing classical supersymmetry calculations. Computer Physics Communications, 92(1):90–110, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500081P.

Ludwig:1991:GMT

[Lud91] A. C. Ludwig. The generalized multipole technique. Computer Physics Communications, 68(1–3):306–314, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190205Y.

Luo:1996:MDF

[Luo96] Xiang-Qian Luo. Molecular dynamics for full QCD simulations with an improved action. Computer Physics Communications, 94(2–3):119–127, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500128X.

Luscher:1994:PHQ

[Lüs94] Martin Lüscher. A portable high-quality random number generator for lattice field theory simulations. Computer Physics Communications, 79(1):100–110, February 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902321.

Lanza:1994:BFF

[LVC^{+}94] A. Lanza, P. Vitulo, C. Cattaneo, M. Caresana, and F. Panetsos. Behaviour of feed-forward neural networks in invariant track finding. Computer Physics Communications, 79(3):364– 372, May 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594901821.

Lambin:1990:CSE

[LVL90] Ph. Lambin, J.-P. Vigneron, and A. A. Lucas. Computation of the surface electron-energy-loss spectrum in specular geometry for an arbitrary plane-stratified medium. Computer Physics Communications, 60(3):351–364, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090034X.

Lavery:1995:JJM

[LZS95] R. Lavery, K. Zakrzewska, and H. Sklenar. JUMNA (junction minimisation of nucleic acids). Computer Physics Communications, 91(1–3):135–158, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500046I.

Mirin:1994:CSM

[MAB^{+}94] A. A. Mirin, J. J. Ambrosiano, J. H. Bolstad, A. J. Bourgeois, J. C. Brown, B. Chan, W. P. Dannevik, P. B. Duffy, P. G. Eltgroth, C. Matarazzo, and M. F. Wehner. Climate system modeling using a domain and task decomposition message-passing approach. Computer Physics Communications, 84(1–3):278–296, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902151.

Macaraeg:1991:ISM

[Mac91] Michéle G. Macaraeg. Investigation of supersonic modes and three-dimensionality in bounded, free shear flows. Computer Physics Communications, 65(1–3):201–208, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190172H.

MacLeod:1992:CSS

[Mac92a] Allan J. MacLeod. Chebyshev series solution of the Thomas– Fermi equation. Computer Physics Communications, 67(3): 389–391, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900473.

MacLeod:1992:NCT

[Mac92b] Allan J. MacLeod. The numerical computation of transport integrals. Computer Physics Communications, 69(2– 3):229–234, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290162R.

Mottez:1998:NGC

[MAH98] F. Mottez, J. C. Adam, and A. Heron. A new guiding centre PIC scheme for electromagnetic highly magnetized plasma simulation. Computer Physics Communications, 113 (2–3):109–130, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000769.

Makri:1991:FPI

[Mak91] Nancy Makri. Feynman path integration in quantum dynamics. Computer Physics Communications, 63(1–3):389–414, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190265M.

Makino:1993:SPR

[Mak93] Jun Makino. On the structure of parallelized random number sources. Computer Physics Communications, 78(1–2): 105–112, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593901464.

Malegat:1990:DPC

[Mal90] L. Malegat. DCS — a program for calculating differential cross sections for the electronic excitation of diatomic molecules at fixed nuclei. Computer Physics Communications, 60(3):391–404, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090036Z.

Malmgren:1997:IND

[Mal97] T. G. M. Malmgren. An iterative nonlinear discriminant analysis program: IDA 1.0. Computer Physics Communications, 106(3):230–236, November 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001008.

Mann:1991:RSP

[Man91] Patrick J. Mann. A relativistic smoothed particle hydrodynamics method tested with the shock tube. Computer Physics Communications, 67(2):245–260, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190020L.

Mann:1993:FEM

[Man93] Patrick J. Mann. A finite element method in space and time for relativistic spherical collapse. Computer Physics Communications, 75(1–2):10–30, April 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901615.

Manson:1994:MAS

[Man94] Joseph R. Manson. Multiphonon atom-surface scattering. Computer Physics Communications, 80(1–3):145–167, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901015.

Mantegna:1999:IHS

[Man99] R. N. Mantegna. Information and hierarchical structure in financial markets. Computer Physics Communications, 121–122:153–156, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599003021.

Martinec:1991:PCL

[Mar91] Zdenek Martinec. Program to calculate the least-squares estimates of the spherical harmonic expansion coefficients of an equally angular-gridded scalar field. Computer Physics Communications, 64(1):140–148, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190057R.

Marksteiner:1996:HPC

[Mar96] Peter Marksteiner. High-performance computing — an overview. Computer Physics Communications, 97(1–2):16–35, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000185.

Marin:1997:EDH

[Mar97] Mauricio Marín. Event-driven hard-particle molecular dynamics using bulk-synchronous parallelism. Computer Physics Communications, 102(1–3):81–96, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000118.

Miller:1991:UMB

[MB91] E. K. Miller and G. J. Burke. Using model-based parameter estimation to increase the physical interpretability and numerical efficiency of computational electromagnetics. Computer Physics Communications, 68(1–3):43–75, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190193O.

Muller:1994:ASG

[MB94] Marcus Müller and Kurt Binder. An algorithm for the semi-grand-canonical simulation of asymmetric polymer mixtures. Computer Physics Communications, 84(1–3):173–185, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594902100.

Marin:1999:DMC

[MBC99a] J. M. Marín, J. Boronat, and J. Casulleras. A diffusion Monte Carlo characterization of the free surface of superfluid ^{4}He. Computer Physics Communications, 121–122: 694, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700984.

Munoz:1999:NPR

[MBC^{+}99b] M. A. Muñoz, G. Bianconi, C. Castellano, A. Gabrielli, M. Marsili, and L. Pietronero. Non perturbative renormalization group approach to surface growth. Computer Physics Communications, 121–122:358–362, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003537.

Mertig:1991:FCC

[MBD91] R. Mertig, M. Böhm, and A. Denner. Feyn Calc — computeralgebraic calculation of Feynman amplitudes. Computer Physics Communications, 64(3):345–359, June 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190130D.

Morgado:1999:GEL

[MBMM99] W. A. M. Morgado, M. Bahiana, S. Martins, and M. S. O. Massunaga. Gravity effects on lamellar pattern formation in confined systems. Computer Physics Communications, 121–122:327–329, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900346X.

Madura:1995:EDM

[MBW^{+}95] Jeffry D. Madura, James M. Briggs, Rebecca C. Wade, Malcolm E. Davis, Brock A. Luty, Andrew Ilin, Jan Antosiewicz, Michael K. Gilson, Babak Bagheri, L. Ridgway Scott, and J. Andrew McCammon. Electrostatics and diffusion of molecules in solution: simulations with the University of Houston Brownian Dynamics Program. Computer Physics Communications, 91(1–3):57–95, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500043F.

Maksymowicz:1999:IMO

[MBZM99a] A. Z. Maksymowicz, M. Bubak, K. Zajac, and M. Magdoń. Impact of mutation and overhunting on population extinction in the Penna model. Computer Physics Communications, 121–122:693, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506700972.

Maksymowicz:1999:SAE

[MBZM99b] A. Z. Maksymowicz, M. Bubak, K. Zajac, and M. Magdoń. Simulation of aging with an extended Penna model. Computer Physics Communications, 121–122:113–115, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002921.

Mak:1993:TDH

[MC93] S. S. Mak and W. K. Chung. A two-dimensional hybrid simulation of ionic fluids. Computer Physics Communications, 78(1–2):121–129, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901486.

Marin:1995:EAP

[MC95] Mauricio Marín and Patricio Cordero. An empirical assessment of priority queues in event-driven molecular dynamics simulation. Computer Physics Communications, 92(2–3):214–224, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001202.

Martines:1997:PII

[MC97] E. Martines and R. Cavazzana. Parallel implementation of the iterative Monte Carlo method for the simulation of the presheath of an object immersed in a plasma. Computer Physics Communications, 103(1):43–50, June 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000258.

Martorell:1999:CRE

[MC99] J. Ma. Martorell and A. Carnicer. Computational requirements to evaluate diffraction integrals near the aperture. Computer Physics Communications, 121–122:697, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506701011.

Malonda:1999:ECE

[MCCC99] A. Grau Malonda, A. Grau Carles, P. Grau Carles, and G. Galiano Casas. EMI2, the counting efficiency for electron capture by a KL _{1} L _{2} L _{3} M model. Computer Physics Communications, 123(1–3):114–122, December 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002635.

McDonnell:1990:DMS

[McD90] J. McDonnell. Domain modelling in second-generation expert systems. Computer Physics Communications, 61(1–2):13–21, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590901038.

MacLaren:1991:PLS

[MCMD91] J. M. MacLaren, D. P. Clougherty, M. E. McHenry, and M. M. Donovan. Parameterised local spin density exchangecorrelation energies and potentials for electronic structure calculations I. zero temperature formalism. Computer Physics Communications, 66(2–3):383–391, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190084X.

Mizokami:1999:SSP

[MCO99] S. Mizokami, Y. Chen, and H. Ohashi. Simulation of singlephase flow using hydrodynamic lattice gases. Computer Physics Communications, 121–122:702, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701060.

MacLaren:1990:LKK

[MCV^{+}90] J. M. MacLaren, S. Crampin, D. D. Vvedensky, R. C. Albers, and J. B. Pendry. Layer Korringa–Kohn–Rostoker electronic structure code for bulk and interface geometries. Computer Physics Communications, 60(3):365–389, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090035Y.

Mayer:1998:EGS

[MCV98] A. Mayer, A. Castiaux, and J.-P. Vigneron. Electronic Green scattering with n-fold symmetry axis from block circulant matrices. Computer Physics Communications, 109(1):81–89, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001410.

Marchand:1996:CQO

[MD96] R. Marchand and M. Dumberry. CARRE: a quasi-orthogonal mesh generator for 2D edge plasma modelling. Computer Physics Communications, 96(2–3):232–246, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000525.

McCoy:1997:PEA

[MD97] R. Alan McCoy and Yuefan Deng. Parallel embedded-atom method simulations with delayed electron density approximations. Computer Physics Communications, 100(1–2):41–46, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001361.

Martinez-Davalos:1999:CBS

[MDMR99] A. Martínez-Dávalos and A. Menchaca-Rocha. A comparison of bulk- and surface-dynamic models applied to liquiddrop fusion. Computer Physics Communications, 121–122: 695, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700996.

Mobius:1999:TPM

[MDSF^{+}99] A. Möbius, A. Díaz-Sánchez, B. Freisleben, M. Schreiber, A. Fachat, K. H. Hoffmann, P. Merz, and A. Neklioudov. Two physically motivated algorithms for combinatorial optimization: thermal cycling and iterative partial transcription. Computer Physics Communications, 121–122:34–36, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002738.

Morriss:1991:CAC

[ME91] Gary P. Morriss and Denis J. Evans. A constraint algorithm for the computer simulation of complex molecular liquids. Computer Physics Communications, 62(2–3):267–278, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591900997.

McGrath:1990:MCR

[MEA^{+}90] R. L. McGrath, A. Elmaani, J. M. Alexander, P. A. DeYoung, T. Ethvignot, M. S. Gordon, and E. Renshaw. A Monte Carlo reaction simulation for small-angle correlations between light charged particles. Computer Physics Communications, 59(3):507–519, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090092F.

Meglicki:1994:VAS

[Meg94] Zdzisław Meglicki. Verification and accuracy of smoothed particle magnetohydrodynamics. Computer Physics Communications, 81(1–2):91–104, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594901139.

Mendoza:1999:IP

[Men99] C. Mendoza. The IRON project. Computer Physics Communications, 121–122:74–77, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002842.

Merazzi:1990:ASE

[Mer90] Silvio Merazzi. The ASTRID system: an example of an integrated numerical analysis system. Computer Physics Communications, 61(1–2):133–140, November 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090111D.

Mess:1990:AC

[Mes90a] K.-H. Mess. Accelerator control. Computer Physics Communications, 61(1–2):115–132, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090110M.

Mohamed:1990:BMS

[MES90b] A. Raouf Mohamed and M. A. El-Sharkawy. 16-bit microprocessor in second-order state-space digital filter design and implementation. Computer Physics Communications, 59(2):283– 288, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465590901773.

Metcalf:1996:CPM

[Met96] Michael Metcalf. Comment on a paper by maley et al. Computer Physics Communications, 99(1):1, December 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001488.

Micha:1994:CET

[MF94] David A. Micha and Eric Q. Feng. The calculation of electron transfer probabilities in slow ion-metal surface collisions. Computer Physics Communications, 80(1–3):242–258, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559490104X.

Mansky:1995:AGA

[MF95a] E. J. Mansky and M. R. Flannery. Automatic generation of analytical matrix elements for electron-atom scattering. Computer Physics Communications, 88(2–3):278–292, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400163V.

Mansky:1995:MET

[MF95b] E. J. Mansky and M. R. Flannery. The multichannel eikonal theory program for electron-atom scattering. Computer Physics Communications, 88(2–3):249–277, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400162U.

Meyer:1998:SPS

[MFG^{+}98] P. Meyer, M. Fromm, J. E. Groetz, F. Torrealba, and A. Chambaudet. Simulation of processes in a SSNTD exposed by monoenergetic neutrons. Computer Physics Communications, 109(1):6–14, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001380.

Monjoie:1990:ULU

[MG90] F. S. Monjoie and H. P. Garnir. The university of Liège user interface for laboratory experiment control. Computer Physics Communications, 61(1–2):267–273, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090127M.

Monjoie:1993:FSE

[MG93] F. S. Monjoie and H. P. Garnir. Fit of a sum of exponential functions to experimental data points. Computer Physics Communications, 74(1):1–8, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390101H.

Marushchenko:1997:BAF

[MGMR97] N. Marushchenko, U. Gasparino, H. Maaßberg, and M. Rome’. Bounce-averaged Fokker–Planck code for the description of ECRH in a periodic magnetic field. Computer Physics Communications, 103(2–3):145–156, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000386.

Miyamoto:1996:MCP

[MH96] Akiya Miyamoto and Hisaki Hayashii. A Monte Carlo program to generate mini-jet events in the two-photon process. Computer Physics Communications, 96(1):87–104, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000422.

Munoz:1999:EBH

[MH99] Jose D. Muñoz and Hans J. Herrmann. Extending the broad histogram method for continuous systems. Computer Physics Communications, 121–122:13–15, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002684.

Mikulla:1999:LSM

[MHHL99] R. Mikulla, J. E. Hammerberg, B. L. Holian, and P. S. Lomdahl. Large scale molecular dynamics simulation of dry friction. Computer Physics Communications, 121–122: 701, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701059.

Madsen:1999:PEA

[MHN99] L. B. Madsen, J. P. Hansen, and H. M. Nilsen. Procedures for the evaluation of atomic transition matrix elements in the interaction with laser light. Computer Physics Communications, 120(2–3):231–237, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002507.

Mickens:1991:NEF

[Mic91] Ronald E. Mickens. Novel explicit finite-difference schemes for time-dependent Schrödinger equations. Computer Physics Communications, 63(1–3):203–208, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190249K.

Michalewicz:1994:MPC

[Mic94] Marek T. Michalewicz. Massively parallel calculations of the electronic structure of non-periodic micro-crystallites of transition metal oxides. Computer Physics Communications, 79 (1):13–23, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594902267.

Midy:1992:STE

[Mid92] P. Midy. Scaling transformations and extrapolation algorithms for vector sequences. Computer Physics Communications, 70(2):285–291, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592901933.

Mino:1991:VAC

[Min91] Hidetoshi Mino. A vectorized algorithm for cluster formation in the Swendsen–Wang dynamics. Computer Physics Communications, 66(1):25–30, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591900045.

Mitsutake:1996:MCE

[MIO96] Ayori Mitsutake, Toshiaki Iitaka, and Yuko Okamoto. A method for calculating the eigenvalues of large Hermitian matrices by second-order recursion formulae. Computer Physics Communications, 96(2–3):217–231, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000677.

Mitas:1996:ESQ

[Mit96] Lubos Mitas. Electronic structure by quantum Monte Carlo: atoms, molecules and solids. Computer Physics Communications, 96(2–3):107–117, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559600063X.

Minter:1990:BCS

[MJ90] Anthony Minter and David A. Jenkins. Bremsstrahlung cross section for a point, spinless target. Computer Physics Communications, 59(3):499–505, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090091E.

Moura:1991:DML

[MK91a] L. Moura and R. Kitney. A direct method for least-squares circle fitting. Computer Physics Communications, 64(1):57– 63, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190049Q.

Mowrey:1991:IMC

[MK91b] R. C. Mowrey and D. J. Kouri. Inelastic moleculecorrugated surface scattering using the close-coupling wavepacket method. Computer Physics Communications, 63(1–3): 100–114, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190242D.

Monaghan:1995:SSM

[MK95] J. J. Monaghan and A. Kocharyan. SPH simulation of multiphase flow. Computer Physics Communications, 87(1–2):225– 235, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400174Z.

Miyama:1996:NSE

[MK96] M. Miyama and S. Kumano. Numerical solution of Q ^{2} evolution equations in a brute-force method. Computer Physics Communications, 94(2–3):185–215, April 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000136.

Metten:1999:MDS

[MKL99] B. Metten, T. Kurz, and W. Lauterborn. Molecular dynamics simulation versus continuum mechanics for a sonoluminescing bubble. Computer Physics Communications, 121–122:699, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506701035.

Munehisa:1997:GTP

[MKPGS97] T. Munehisa, K. Kato, D. Perret-Gallix, and Y. Shimizu. GGPS: a two-photon event generator based on the parton shower method. Computer Physics Communications, 106 (1–2):139–156, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597000854.

Maley:1996:FSA

[MKS^{+}96] D. Maley, P. L. Kilpatrick, E. W. Schreiner, N. S. Scott, and G. H. F. Diercksen. The formal specification of abstract data types and their implementation in Fortran 90: implementation issues concerning the use of pointers. Computer Physics Communications, 98(1–2):167–180, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000938.

Mancini:1991:CAS

[MKWH91] R. C. Mancini, D. P. Kilcrease, L. A. Woltz, and C. F. Hooper, Jr. Calculational aspects of the Stark line broadening of multielectron ions in plasmas. Computer Physics Communications, 63(1–3):314–322, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190258M.

Martoff:1992:CPE

[ML92] C. J. Martoff and P. D. Lewin. COSMO — a program to estimate spallation radioactivity produced in a pure substance by exposure to cosmic radiation on the earth. Computer Physics Communications, 72(1):96–103, October 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290008M.

Marchand:1993:AIA

[MLB93] R. Marchand, S. Lafortune, and X. Bonnin. Average ion approximation for modelling impurity transport in tokamaks. Computer Physics Communications, 76(2):203–214, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390132V.

Mastrangelo:1996:SMD

[MLM96] V. Mastrangelo, D. Lippmann, and I. Mehilli. Stochastic modelling of diffusion equations. implementation and performance of the Mixage3D code on a multinode parallel machine based on the transputer T9000. Computer Physics Communications, 95(1):23–38, May 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000112.

Makino:1991:GSR

[MM91] Jun Makino and Osamu Miyamura. Generation of shift register random numbers on vector processors. Computer Physics Communications, 64(3):363–368, June 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901325.

Malon:1998:OCT

[MM98] David M. Malon and Edward N. May. An ODMG-compatible testbed architecture for scalable management and analysis of physics data. Computer Physics Communications, 110 (1–3):120–124, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597001641.

Marshall:1995:ICS

[MMC95] David Marshall, Mark McCartney, and Derrick S. F. Crothers. ION2 — calculation of single ionisation total cross section using the continuum distorted-wave eikonal initial-state approximation. Computer Physics Communications, 86(3):279–288, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400154T.

Martinec:1993:SPN

[MMCH93] Zdenek Martinec, Ctirad Matyska, Ondrej Cadek, and Pavel Hrdina. The stokes problem with 3D Newtonian rheology in a spherical shell. Computer Physics Communications, 76(1):63– 79, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390121R.

Mastrangelo:1993:SMD

[MMGS93] M. Mastrangelo, V. Mastrangelo, D. Gassilloud, and F. Simon. Stochastic modelling of diffusion equations on a parallel machine. Computer Physics Communications, 76(2):159– 183, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390129Z.

Macak:1999:MCS

[MMH99] Karol Macàk, Peter Macàk, and Ulf Helmersson. Monte Carlo simulations of the transport of sputtered particles. Computer Physics Communications, 120(2–3):238–254, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002453.

Margio:1994:BCR

[MMPT94] C. Margio, K. J. M. Moriarty, B. Plache, and T. Trappenberg. Bank conflict resolution. Computer Physics Communications, 83(2–3):125–129, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900418.

McNamara:1993:ADC

[MMS93] B. McNamara, K. J. M. Moriarty, and S. Sanielevici. Application-driven considerations for the design of a multipleprocessor teraflop/s computer system. Computer Physics Communications, 75(3):283–297, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390045E.

Moriarty:1993:DCP

[MMSM93] K. J. M. Moriarty, G. Murdeshwar, S. Sanielevici, and B. McNamara. The design of a CASE package: Scotia programming environment and Facility. Computer Physics Communications, 77(3):325–341, November 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390180K.

Mirin:1994:FTD

[MMTK94] A. A. Mirin, M. G. McCoy, G. P. Tomaschke, and J. Killeen. FPPAC94: a two-dimensional multispecies nonlinear Fokker–Planck package for UNIX systems. Computer Physics Communications, 81(3):403–408, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900876.

McKay:1999:EPU

[MMU99] Michael D. McKay, John D. Morrison, and Stephen C. Upton. Evaluating prediction uncertainty in simulation models. Computer Physics Communications, 117(1–2):44–51, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001556.

Medgyesi-Mitschang:1991:HMC

[MMW91] L. N. Medgyesi-Mitschang and Dau-Sing Wang. Hybrid methods in computational electromagnetics: a review. Computer Physics Communications, 68(1–3):76–94, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190194P.

Maksymowicz:1996:CSD

[MMW96] Andrzej Z. Maksymowicz, Maria Magdoń, and Jeremy S. S. Whiting. Computer simulation of dynamics of surface growth. Computer Physics Communications, 97(1–2):101–105, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000240.

Morales:1990:TIL

[MN90] Juan J. Morales and María J. Nuevo. A technique for improving the link-cell method. Computer Physics Communications, 60(2):195–199, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090004K.

Morales:1992:CLC

[MN92] Juan J. Morales and Maria J. Nuevo. Comparison of link-cell and neighbourhood tables on a range of computers. Computer Physics Communications, 69(2–3):223–228, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290161Q.

Mohankumar:1997:AEP

[MN97a] N. Mohankumar and A. Natarajan. The accurate evaluation of a particular Fermi–Dirac integral. Computer Physics Communications, 101(1–2):47–53, April 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559600166X.

Mohankumar:1997:MNE

[MN97b] N. Mohankumar and A. Natarajan. A method for the numerical evaluation of a Hadamard finite part integral. Computer Physics Communications, 105(2–3):187–196, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000805.

Montagna:1995:WMC

[MNP95] Guido Montagna, Oreste Nicrosini, and Fulvio Piccinini. WWGENPV — a Monte Carlo event generator for fourfermion production in e^{+}e^{−} → W^{+}W^{−} → 4f. Computer Physics Communications, 90(1):141–150, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500077S.

Montagna:1996:NMC

[MNP96] Guido Montagna, Oreste Nicrosini, and Fulvio Piccinini. NUNUGPV — a Monte Carlo event generator for e^{+}e^{−} → vvνγ(γ) events at LEP. Computer Physics Communications, 98(1–2):206–214, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000811.

Montagna:1996:TPC

[MNPP96] Guido Montagna, Oreste Nicrosini, Giampiero Passarino, and Fulvio Piccinini. TOPAZO 2.0- A program for computing de-convoluted and realistic observables around the Z ^{0} peak. Computer Physics Communications, 93(1):120–126, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001271.

Montagna:1999:TNV

[MNPP99] Guido Montagna, Oreste Nicrosini, Fulvio Piccinini, and Giampiero Passerino. TOPAZ0 4.0 — a new version of a computer program for evaluation of deconvoluted and realistic observables at LEP 1 and LEP 2. Computer Physics Communications, 117(3):278–289, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000800.

Munoz:1999:BPT

[MNS99] A. Muñoz, I. Nieves, and E. Suarez. BANDAS: a program for teaching band theory (LCAO) in solid state physics. Computer Physics Communications, 121–122:706, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701102.

Marsaglia:1990:RNG

[MNZ90] George Marsaglia, B. Narasimhan, and Arif Zaman. A random number generator for PC’s. Computer Physics Communications, 60(3):345–349, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090033W.

Morrison:1993:PIA

[MÖ93] H. G. Morrison and U. Öpik. A program for ion-atom collisions involving one electron. Computer Physics Communications, 77(3):403–428, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390186G.

McKenzie:1991:STD

[MOC91] J. S. McKenzie, M. R. O’Brien, and M. Cox. Solution of three-dimensional Fokker–Planck equations for tokamak plasmas using an operator splitting technique. Computer Physics Communications, 66(2–3):194–206, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190068V.

Monkenbusch:1991:SRE

[Mon91] Michael Monkenbusch. A set of routines for efficient and accurate computation of lattice sums of 1r ^{n}-potentials. Computer Physics Communications, 67(2):343–355, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190027I.

Montvay:1998:QOP

[Mon98] I. Montvay. Quadratically optimized polynomials for fermion simulations. Computer Physics Communications, 109(2– 3):144–160, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000071.

Monk:1999:BRD

[Mon99] P. Monk. Book review: D. Funaro, Spectral Elements for Transport-Dominated Equations (1997) Springer, Cambridge 3-540-62649-2. Computer Physics Communications, 118(1):92–93, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001490.

Mooney:1993:STF

[Moo93] J. W. Mooney. Solution of a Thomas–Fermi problem using linear approximants. Computer Physics Communications, 76(1): 51–57, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390119W.

Mooney:1994:SET

[Moo94] J. W. Mooney. Solution of emden-type problems using accurate, efficient discretisation schemes. Computer Physics Communications, 83(2–3):245–254, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900523.

Meyer:1999:EST

[MOPR99] M. Meyer, G. Onida, M. Palummo, and L. Reining. Electronic structure of tin oxides. Computer Physics Communications, 121–122:700, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701047.

Morrison:1992:PIC

[Mor92] T. P. Morrison. POP — an interactive charged particle transport system design tool. Computer Physics Communications, 69(2–3):477–485, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290184Z.

Morita:1998:CK

[Mor98] Y. Morita. Computing at KEK. Computer Physics Communications, 110(1–3):1–5, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597001446.

Molisch:1993:RTPb

[MOSM93a] Andreas F. Molisch, Bernhard P. Oehry, Walter Schupita, and Gottfried Magerl. RAD-TRAP 2, a program for the solution of the Holstein equation of radiation trapping. Computer Physics Communications, 77(2):255–262, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900092.

Molisch:1993:RTPa

[MOSM93b] Andreas F. Molisch, Bernhard P. Oehry, Walter Schupita, and Gottfried Magerl. RAD-TRAP, a program for the computation of the eigenvalues and eigenfunctions of the holstein radiation-trapping equation. Computer Physics Communications, 74(1):81–90, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390108O.

Molisch:1996:MPC

[MOSM96] Andreas F. Molisch, Bernhard P. Oehry, Walter Schupita, and Gottfried Magerl. McTrap, a program for the computation of radiation trapping in 3-level atoms including bleaching effects. Computer Physics Communications, 93(1):127–135, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595000933.

Morgan:1990:ETD

[MP90a] W. L. Morgan and B. M. Penetrante. ELENDIF: a timedependent Boltzmann solver for partially ionized plasmas. Computer Physics Communications, 58(1–2):127–152, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090141M.

Muller-Plathe:1990:PMD

[MP90b] Florian Müller-Plathe. Parallelising a molecular dynamics algorithm on a multi-processor workstation. Computer Physics Communications, 61(3):285–293, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090043Z.

Muller-Plathe:1993:YMS

[MP93] Florian Müller-Plathe. YASP: a molecular simulation package. Computer Physics Communications, 78(1–2):77–94, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593901442.

Mata:1995:GPA

[MP95] G. J. Mata and E. Pestana. A general purpose algorithm for the integration of functions with power-law singularities of arbitrary exponent. Computer Physics Communications, 92(2–3):225–228, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595000695.

Mombach:1999:CMB

[MP99] José Carlos M. Mombach and Lisiane Preissler. Computer model for biological cell sorting. Computer Physics Communications, 121–122:703, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701072.

Muller-Plathe:1991:MCA

[MPB91] Florian Müller-Plathe and David Brown. Multi-colour algorithms in molecular simulation: vectorisation and parallelisation of internal forces and constraints. Computer Physics Communications, 64(1):7–14, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190045M.

Martonak:1996:MCS

[MPB96] R. Martonák, W. Paul, and K. Binder. Monte Carlo simulation of crystalline polyethylene. Computer Physics Communications, 99(1):2–8, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001142.

Montagna:1993:PPC

[MPN93a] Guido Montagna, Fulvio Piccinini, and Oreste Nicrosini. PHIPHI — a program for computing radiative Bhabha scattering cross sections at DAΦNE. Computer Physics Communications, 78(1–2):155–171, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901512.

Montagna:1993:TPC

[MPN^{+}93b] Guido Montagna, Fulvio Piccinini, Oreste Nicrosini, Giampiero Passarino, and Roberto Pittau. TOPAZ0 — a program for computing observables and for fitting cross sections and forward-backward asymmetries around the Z ^{0} peak. Computer Physics Communications, 76(3):328–360, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390060P.

McDonough:1997:IMD

[MPRS97] A. McDonough, A. Panjkov, S. Russo, and I. K. Snook. Implementation of a molecular dynamics simulation on a transputer array in parallel C. Computer Physics Communications, 103(2–3):157–169, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000362.

Miguel:1999:GOC

[MPS99] M.-Carmen Miguel and R. Pastor-Satorras. Growth of oriented chains in dipolar colloids. Computer Physics Communications, 121–122:262–264, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003264.

Muller-Plathe:1994:PBP

[MPSvG94] Florian Müller-Plathe, Walter Scott, and Wilfred F. van Gunsteren. PARALLACS: a benchmark for parallel molecular dynamics. Computer Physics Communications, 84(1–3): 102–114, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594902062.

Marquina:1998:TDD

[MPT98] M. Marquina, R. Ramos Pollán, and A. Taddei. A taxonomically distributed data retrieval model. Computer Physics Communications, 110(1–3):198–205, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700177X.

Mignani:1995:MBB

[MR95] S. Mignani and R. Rosa. The moving block bootstrap to assess the accuracy of statistical estimates in Ising model simulations. Computer Physics Communications, 92(2–3):203–213, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001147.

Martin:1996:NSC

[MR96] V. Martin and L. M. Robledo. Numerical and symbolic calculation of the multipole matrix elements in the axial and triaxial harmonic oscillator basis. Computer Physics Communications, 99(1):113–127, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001051.

Mattson:1999:NNC

[MR99a] William Mattson and Betsy M. Rice. Near-neighbor calculations using a modified cell-linked list method. Computer Physics Communications, 119(2–3):135–148, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598002033.

Mitroy:1999:OPP

[MR99b] J. Mitroy and G. G. Ryzhikh. Orthogonalising pseudopotentials in electronic structure calculations. Computer Physics Communications, 123(1–3):103–113, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004452.

Mansfield:1998:NRC

[MRC98] Elizabeth L. Mansfield, Gregory J. Reid, and Peter A. Clarkson. Nonclassical reductions of a 3 + 1-cubic nonlinear Schrödinger system. Computer Physics Communications, 115(2–3):460–488, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001362.

Mrenna:1997:SSE

[Mre97] S. Mrenna. SPYTHIA, a supersymmetric extension of PYTHIA 5.7. Computer Physics Communications, 101(3): 232–240, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597000039.

Maggio:1994:ASA

[MRPC94] A. Maggio, F. Reale, G. Peres, and A. Ciaravella. The analysis system for astrophysical plasmas (ASAP) of the Osservatorio Astronomico di Palermo. Computer Physics Communications, 81(1–2):105–119, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901147.

McCurdy:1991:EWR

[MS91a] C. William McCurdy and Carrie K. Stroud. Eliminating wavepacket reflection from grid boundaries using complex coordinate contours. Computer Physics Communications, 63(1–3):323–330, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190259N.

Moheb:1991:NFM

[MS91b] H. Moheb and L. Shafai. Numerical field modelling of objects with edges and corners using a conformal mapping. Computer Physics Communications, 68(1–3):95–125, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190195Q.

Micu:1995:SPC

[MS95a] Alexandru M. Micu and Jeremy C. Smith. SERENA: a program for calculating X-ray diffuse scattering intensities from molecular dynamics trajectories. Computer Physics Communications, 91(1–3):331–338, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500057M.

Muller:1995:SSG

[MS95b] Ewald Müller and Matthias Steinmetz. Simulating selfgravitating hydrodynamic flows. Computer Physics Communications, 89(1–3):45–58, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594001855.

Maillard:1996:TPG

[MS96] J. Maillard and J. Silva. Track parallelisation in GEANT detector simulations. Computer Physics Communications, 95(1): 17–22, May 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001077.

Mertig:1998:TMP

[MS98] R. Mertig and R. Scharf. TARCER — a Mathematica program for the reduction of two-loop propagator integrals. Computer Physics Communications, 111(1–3):265–273, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000423.

Ma:1993:BCM

[MSD93] S. Ma, R. D. Sydora, and J. M. Dawson. Binary collision model in gyrokinetic simulation plasmas. Computer Physics Communications, 77(2):190–206, October 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390003U.

Meyer:1996:NVP

[MSFR96] J. Meyer, W.-D. Sepp, B. Fricke, and A. Rosén. A new version of the program TSYM generating relativistic molecular symmetry orbitals for finite double point groups. Computer Physics Communications, 96(2–3):263–287, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000586.

Mohammadian:1991:CES

[MSH91] Alireza H. Mohammadian, Vijaya Shankar, and William F. Hall. Computation of electromagnetic scattering and radiation using a time-domain finite-volume discretization procedure. Computer Physics Communications, 68(1–3):175–196, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190199U.

Montavez:1999:MCS

[MSJ99] J. P. Montávez, E. Sánchez, and J. I. Jiménez. A Monte Carlo simulation of the longwave radiation balance in urban structures. Computer Physics Communications, 121–122:704, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701084.

Maley:1998:CAT

[MSK98] David Maley, Ivor Spence, and Peter Kilpatrick. Config: a GRACE tool for constructing configuration trees. Computer Physics Communications, 114(1–3):271–294, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000927.

Mitnik:1998:EII

[MSP^{+}98] D. M. Mitnik, J. A. Shaw, M. S. Pindzola, D. C. Griffin, and N. R. Badnell. Electron-impact ionization of Fe^{14+} and other atomic ions in the Mg isoelectronic sequence. Computer Physics Communications, 114(1–3):368–377, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000666.

Mann:1996:MTW

[MSR96] M. Mann, E. Shamonina, and K. H. Ringhofer. Modelling of two wave mixing experiments in sillenite crystals. Computer Physics Communications, 96(1):61–86, July 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000343.

Michalewicz:1992:EMM

[MSTH92] Marek T. Michalewicz, Herbert B. Shore, N. Tit, and J. W. Halley. Equation of motion method for the electronic structure of disordered transition metal oxides. Computer Physics Communications, 71(3):222–234, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290011M.

Moriarty:1993:EMC

[MSTK93] K. J. M. Moriarty, S. Sanielevici, T. Trappenberg, and D. W. Kuba. An efficiently microtasked Cray Y-MP C90 version of the Kuba–Moriarty SU(3) gauge theory simulation program. Computer Physics Communications, 76(1):87–97, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390123T.

Mankiewicz:1992:PMC

[MSV92] L. Mankiewicz, A. Schäfer, and M. Veltri. PEPSI — a Monte Carlo generator for polarized leptoproduction. Computer Physics Communications, 71(3):305–318, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290016R.

Moncrieff:1992:DMBb

[MSW92a] D. Moncrieff, V. R. Saunders, and S. Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules: IX. on the use of dynamic load balancing on a multi-processor computer. Computer Physics Communications, 72(2–3):115–128, November 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290143M.

Moncrieff:1992:DMBa

[MSW92b] D. Moncrieff, V. R. Saunders, and S. Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules VIII. ccMBPT-4_{t}. Computer Physics Communications, 70 (2):345–354, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592901977.

Mountain:1991:EMS

[MT91] Raymond D. Mountain and D. Thirumalai. Ergodic measures for the simulation of dialectric properties of water. Computer Physics Communications, 62(2–3):352–359, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190106U.

Mittal:1992:NSS

[MT92a] S. Mittal and T. E. Tezduyar. Notes on the stabilized space– time finite-element formulation of unsteady incompressible flows. Computer Physics Communications, 73(1–3):93–112, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290031S.

Morales:1992:CNT

[MT92b] Juan J. Morales and Søren Toxvaerd. The cell-neighbor table method in molecular dynamics simulations. Computer Physics Communications, 71(1–2):71–76, August 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900738.

Morgan:1998:UMR

[MTG98] Lesley A. Morgan, Jonathan Tennyson, and Charles J. Gillan. The UK molecular R-matrix codes. Computer Physics Communications, 114(1–3):120–128, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000563.

Mori:1999:APC

[MTH^{+}99] W. B. Mori, K.-C. Tzeng, R. G. Hemker, F. S. Tsung, B. J. Duda, and T. Katsouleas. The application of particle-in-cell algorithms to modeling plasma-based accelerator experiments. Computer Physics Communications, 121–122:705, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506701096.

Makino:1992:GSR

[MTM92] Jun Makino, Tetsuya Takaishi, and Osamu Miyamura. Generation of shift register random numbers on distributed memory multiprocessors. Computer Physics Communications, 70 (3):495–500, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290111B.

Marx:1999:QDA

[MTM99] Dominik Marx, Mark E. Tuckerman, and Glenn J. Martyna. Quantum dynamics via adiabatic ab initio centroid molecular dynamics. Computer Physics Communications, 118(2–3):166– 184, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002088.

Mussa:1998:RVC

[MTNA98] Hamse Y. Mussa, Jonathan Tennyson, C. J. Noble, and R. J. Allan. Rotation-vibration calculations using massively parallel computers. Computer Physics Communications, 108(1):29– 37, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001252.

Moriarty:1994:DMP

[MTR94] Kevin J. M. Moriarty, Thomas Trappenberg, and Claudio Rebbi. Development of massively parallel applications. Computer Physics Communications, 81(1–2):153–162, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901171.

Mardahl:1997:CCE

[MV97] P. J. Mardahl and J. P. Verboncoeur. Charge conservation in electromagnetic PIC codes; spectral comparison of Boris/DADI and Langdon–Marder methods. Computer Physics Communications, 106(3):219–229, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000945.

Marchesini:1992:HMC

[MWA^{+}92] G. Marchesini, B. R. Webber, G. Abbiendi, I. G. Knowles, M. H. Seymour, and L. Stanco. HERWIG 5.1 — a Monte Carlo event generator for simulating hadron emission reactions with interfering gluons. Computer Physics Communications, 67(3):465–508, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900554.

Maidantchik:1998:WSD

[MXR98] C. Maidantchik, G. B. Xexéo, and A. R. C. Rocha. A WWW software development environment to support cooperative and spread working groups. Computer Physics Communications, 110(1–3):181–191, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001756.

Malevsky:1992:LSN

[MY92] A. V. Malevsky and D. A. Yuen. Large-scale numerical simulations of turbulent non-Newtonian thermal convection using method of characteristics. Computer Physics Communications, 73(1–3):61–71, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290029X.

Nadrchal:1990:MMC

[Nad90a] J. Nadrchal. Modern methods of computer applications. Computer Physics Communications, 61(1–2):1–2, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590901016.

Nadrchal:1990:P

[Nad90b] Jaroslav Nadrchal. Preface. Computer Physics Communications, 61(1–2):xi, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090099M.

Nadrchal:1992:P

[Nad92] J. Nadrchal. Preface. Computer Physics Communications, 73 (1–3):xi, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290024S.

Nadrchal:1996:P

[Nad96] Jaroslav Nadrchal. Preface. Computer Physics Communications, 97(1–2):xi, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596900910.

Nakano:1991:NAL

[Nak91] Hisamatsu Nakano. A numerical approach to line antennas printed on dielectric materials. Computer Physics Communications, 68(1–3):441–450, November 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591902124.

Nakano:1997:FCA

[Nak97a] Aiichiro Nakano. Fuzzy clustering approach to hierarchical molecular dynamics simulation of multiscale materials phenomena. Computer Physics Communications, 105(2–3): 139–150, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597000441.

Nakano:1997:PMP

[Nak97b] Aiichiro Nakano. Parallel multilevel preconditioned conjugategradient approach to variable-charge molecular dynamics. Computer Physics Communications, 104(1–3):59–69, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000416.

Narayanan:1992:CPK

[Nar92] K. S. S. Narayanan. Comment on a paper by K.J.F. gaemers: (comput. phys. commun. 22 (1981) 115). Computer Physics Communications, 69(1):73–75, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290130Q.

Narayan:1996:CMG

[Nar96]  K. Lakshmi Narayan.  Computer modelling of grain

microstructure in three dimensions. Computer Physics Communications, 93(1):136–140, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595000348.

Neuhauser:1991:ATD

[NBJK91] Daniel Neuhauser, Michael Baer, Richard S. Judson, and Donald J. Kouri. The application of time-dependent wavepacket methods to reactive scattering. Computer Physics Communications, 63(1–3):460–481, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190270U.

Ng:1996:SRP

[NC96] C. N. M. Ng and M. J. Carvalho. The sp(3, R ) representations of an A-particle nucleus. Computer Physics Communications, 96(2–3):288–300, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000458.

Nguyen:1996:PCP

[NCC^{+}96] H. V. Nguyen, J. M. Campbell, G. P. Couchell, S. Li, D. J. Pullen, W. A. Schier, E. H. Seabury, and S. V. Tipnis. Programs in C for parameterizing measured 5′′ × 5′′ NaI gamma response functions and unfolding of continuous gamma spectra. Computer Physics Communications, 93(2–3):303–321, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595000968.

Nelson:1993:LPA

[NE93] Paul Nelson and David S. Ek. LOCFES — a program for automatic determination of the order of spatial approximations in one-dimensional particle transport. Computer Physics Communications, 74(1):91–118, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390109P.

Neubert:1999:ECC

[Neu99] B. Neubert. Efficient calculation of complex multidimensional phase diagrams of discrete-lattice models. Computer Physics Communications, 120(2–3):162–176, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002222.

Nelson:1995:MVC

[NHK^{+}95] Mark Nelson, William Humphrey, Richard Kufrin, Attila Gursoy, Andrew Dalke, Laxmikant Kale, Robert Skeel, and Klaus Schulten. MDScope — a visual computing environment for structural biology. Computer Physics Communications, 91(1–3):111–133, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500045H.

Nievergelt:1994:CCS

[Nie94] Yves Nievergelt. Computing circles and spheres of arithmitic least squares. Computer Physics Communications, 81(3):343– 350, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900825.

Nieto:1995:ESM

[Nie95] Agustin Nieto. Evaluating sums over the Matsubara frequencies. Computer Physics Communications, 92(1):54–64, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500061J.

Newman:1991:IMM

[NK91] E. H. Newman and K. Kingsley. An introduction to the method of moments. Computer Physics Communications, 68(1–3):1–18, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190191M.

Natarajan:1993:NEG

[NK93] A. Natarajan and N. Mohan Kumar. On the numerical evaluation of the generalised Fermi–Dirac integrals. Computer Physics Communications, 76(1):48–50, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390118V.

Najima:1996:CTL

[NK96] Ryuichi Najima and Yoshimasa Kurihara. Complete tree level evaluation for the reaction e^{+}e^{−} → v_{e}vν_{e}W^{+}W^{−} by an automatical amplitude generator. Computer Physics Communications, 99(1):73–76, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596000872.

Nadal:1999:BSS

[NK99] J. P. Nadal and E. Korutcheva. Blind source separation of sources with different magnitudes. Computer Physics Communications, 121–122:707, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701114.

Novotny:1999:SFP

[NKR99] M. A. Novotny, M. Kolesik, and P. A. Rikvold. Slow forcing in the projective dynamics method. Computer Physics Communications, 121–122:330–333, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003471.

Nakano:1994:MMD

[NKV94] Aiichiro Nakano, Rajiv K. Kalia, and Priya Vashishta. Multiresolution molecular dynamics algorithm for realistic materials modeling on parallel computers. Computer Physics Communications, 83(2–3):197–214, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900485.

Neto:1991:NSV

[NL91] J. J. Soares Neto and Jan Linderberg. A numerical study of various finite-element method schemes applied to quantum mechanical calculations. Computer Physics Communications, 66(1):55–65, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591900078.

Natarajan:1997:NNE

[NM97] A. Natarajan and N. Mohankumar. A note on the numerical evaluation of the H function. Computer Physics Communications, 107(1–3):54–60, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000957.

Novak:1993:CPN

[NN93] Milos H. Novak and Edwin S. Nowak. The CAV program for numerical evaluation of laminar natural convection heat transfer in vertical rectangular cavities. Computer Physics Communications, 78(1–2):95–104, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465593901453.

Nguyen:1998:SCU

[NND98] Nam-Anh Nguyen and T. T. Nguyen-Dang. Symbolic calculations of unitary transformations in quantum dynamics. Computer Physics Communications, 115(2–3):183–199, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001295.

Nesbitt:1998:IAW

[NOC98] B. S. Nesbitt, S. F. C. O’Rourke, and D. S. F. Crothers. Ion-atom-wave — calculation of single ionization cross sections in ion-atom collisions. Computer Physics Communications, 114(1–3):385–400, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000721.

Nordbeck:1994:SPC

[Nor94] Ralf-Peter Nordbeck. SCATTAMPDIAMOL — a program to calculate scattering amplitudes for electron scattering from state selected rotating diatomic molecules. Computer Physics Communications, 82(2–3):233–246, September 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901716.

Nemnyugin:1996:MCC

[NP96] S. A. Nemnyugin and A. M. Petrov. Monte Carlo calculation of muonic molecules. Computer Physics Communications, 97(1–2):175–184, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000306.

Nomura:1990:PSA

[NRG90] Toshio Nomura, Michael Ramek, and Bruno Gruber. Programs for symmetry adaption coefficients for semisimple symmetry chains: the completely symmetric representations. Computer Physics Communications, 61(3):410–432, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900545.

Nordlund:1990:DSS

[NS90] Å. Nordlund and R. F. Stein. 3-D simulations of solar and stellar convection and magnetoconvection. Computer Physics Communications, 59(1):119–125, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090161S.

Nguyen:1995:NCC

[NSE95] Mau Chung Nguyen, P. Santini, and Paul Erdős. Numerical calculation of correlation functions by boundary condition averaging: test on the one-dimensional Hubbard model. Computer Physics Communications, 92(1):16–20, November 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500109S.

Nunn:1990:NSV

[Nun90] D. Nunn. The numerical simulation of VLF nonlinear wave-particle interactions in collision-free plasmas using the Vlasov hybrid simulation technique. Computer Physics Communications, 60(1):1–25, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090074B.

Ngo:1996:FCS

[NVC96] Dat Ngo, Gorden Videen, and Petr Chýlek. A FOR-TRAN code for the scattering of EM waves by a sphere with a nonconcentric spherical inclusion. Computer Physics Communications, 99(1):94–112, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001099.

Nijboer:1997:CMF

[NvdHPG97] R. J. Nijboer, B. v. d Holst, S. Poedts, and J. P. Goedbloed. Calculating magnetohydrodynamic flow spectra. Computer Physics Communications, 106(1–2):39–52, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000829.

Nieto-Vesperinas:1993:FRE

[NVFNP93] M. Nieto-Vesperinas, F. J. Fuentes, R. Navarro, and M. J. Perez-Ilzarbe. A FORTRAN routine to estimate a function of two variables from its autocorrelation. Computer Physics Communications, 78(1–2):211–217, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901567.

Nakano:1993:PMT

[NVK93] Aiichiro Nakano, Priya Vashishta, and Rajiv K. Kalia. Parallel multiple-time-step molecular dynamics with three-body interaction. Computer Physics Communications, 77(3):303–312, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390178F.

Nakano:1994:MPA

[NVK94] Aiichiro Nakano, Priya Vashishta, and Rajiv K. Kalia. Massively parallel algorithms for computational nanoelectronics based on quantum molecular dynamics. Computer Physics Communications, 83(2–3):181–196, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900477.

Nakamura:1999:CIP

[NY99] Takashi Nakamura and Takashi Yabe. Cubic interpolated propagation scheme for solving the hyper-dimensional Vlasov–Poisson equation in phase space. Computer Physics Communications, 120(2–3):122–154, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002477.

Ortiz:1995:DLE

[OA95] Félix Ortiz and JoséM. Los Arcos. DYEFIC: LSC efficiency computation at the dynodic output of PMTs. Computer Physics Communications, 86(1–2):123–128, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400157W.

Ortiz:1996:MLS

[OA96] Félix Ortiz and José M. Los Arcos. MCBETH: liquid scintillation counting spectra computation at the dynodic output of the photomultipliers. Computer Physics Communications, 93(2–3):283–288, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001115.

Otten:1992:DHS

[OBRR92] J. Otten, A. Bledowski, K. H. Ringhofer, and R. A. Rupp. Dynamical holographic storage in photorefractive crystals. Computer Physics Communications, 69(1):187–200, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290139P.

ORourke:1998:LCL

[OC98] S. F. C. O’Rourke and D. S. F. Crothers. LMD — calculation of longitudinal momentum distributions in the single ionization of helium by ion impact. Computer Physics Communications, 109(2–3):184–192, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000198.

Opper:1999:RD

[OD99] Manfred Opper and Sigurd Diederich. Replicator dynamics. Computer Physics Communications, 121–122:141–144, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003008.

Odorico:1990:CMC

[Odo90] R. Odorico. Cojets 5.12: a Monte Carlo simulation program for p–p and p–p collisions. Computer Physics Communications, 59(3):527–544, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090094H.

Odorico:1992:CMC

[Odo92] R. Odorico. COJETS 6.23: a Monte Carlo simulation program for p–p, p–p collisions and e^{+}e^{−} annihilation. Computer Physics Communications, 72(2–3):238–248, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290151N.

Odorico:1996:NPN

[Odo96] R. Odorico. Neural 2.00- A program for neural net and statistical pattern recognition. Computer Physics Communications, 96(2–3):314–330, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000100.

Oeffner:1999:IPG

[OE99] R. D. Oeffner and S. R. Elliott. Interatomic potential for germanium dioxide fitted to an ab-initio energy surface. Computer Physics Communications, 121–122:708, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506701126.

Oleari:1992:USC

[OF92] Claudio Oleari and Guido Formaleoni. (θ, δ) uniformscale chromaticity diagram. Computer Physics Communications, 69(1):112–132, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290133J.

Osswald:1991:SDS

[OGG91] G. A. Osswald, K. N. Ghia, and U. Ghia. Simulation of dynamic stall phenomenon using unsteady Navier–Stokes equations. Computer Physics Communications, 65(1–3):209–218, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190173I.

Ohira:1999:EDD

[Ohi99] Toru Ohira. Encryption with delayed dynamics. Computer Physics Communications, 121–122:54–56, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002799.

Ohl:1992:HLS

[Ohl92] Thorsten Ohl. hepawk — a language for scanning high energy physics events. Computer Physics Communications, 70 (1):120–146, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290096H.

Ohlsson:1993:EEE

[Ohl93] Mattias Ohlsson. Extensions and explorations of the elastic arms algorithm. Computer Physics Communications, 77(1):19–32, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593900339.

Ohl:1995:DFD

[Ohl95] Thorsten Ohl. Drawing Feynman diagrams with L^{A}TEX and . Computer Physics Communications, 90(2–3): 340–354, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559590137S.

Ohl:1997:CVB

[Ohl97] Thorsten Ohl. circe version 1.02: beam spectra for simulating linear collider physics. Computer Physics Communications, 101(3):269–288, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001671.

Ohl:1999:VRA

[Ohl99] T. Ohl. Vegas revisited: Adaptive Monte Carlo integration beyond factorization. Computer Physics Communications, 120(1):13–19, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559900209X.

Oleksy:1996:CAM

[Ole96] Cz. Oleksy. A convergence acceleration method of Fourier series. Computer Physics Communications, 96(1):17–26, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000446.

Olness:1998:BRW

[Oln98] F. Olness. Book review: W. Kinzel, G. Reents, Physics by Computer: Programming of Physical Problems Using Mathematica and C (1998) Springer,New York 3-540-62743-X. Computer Physics Communications, 112(2–3):270–272, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000484.

Olson:1996:ETC

[Ols96] Kevin M. Olson. Efficient tree codes on SIMD computer architectures. Computer Physics Communications, 98(3):267–287, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000963.

Okamoto:1993:NAC

[OM93] Yuko Okamoto and Humphrey J. Maris. A novel algorithm for calculation of the extreme eigenvalues of large Hermitian matrices. Computer Physics Communications, 76(2):191–202, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390131U.

Oss:1993:VCP

[OMC93] Stefano Oss, Nicola Manini, and Renato Lemus Casillas. VIBR3AT: a computer program for triatomic molecular spectroscopy in an algebraic approach. Computer Physics Communications, 74(2):164–186, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390087S.

Omelyan:1997:EST

[Ome97] Igor P. Omelyan. Ewald summation technique for interaction site models of polar fluids. Computer Physics Communications, 107(1–3):113–122, December 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001094.

Omelyan:1998:NIE

[Ome98] Igor P. Omelyan. Numerical integration of the equations of motion for rigid polyatomics: The matrix method. Computer Physics Communications, 109(2–3):171–183, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000241.

Ovesen:1996:CTM

[OPP96] Jens H. Ovesen, Henrik G. Petersen, and John W. Perram. Comparison of two methods for solving linear equations occurring in molecular dynamics applications. Computer Physics Communications, 94(1):1–18, March 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001352.

Ososkov:1999:RTS

[OPT99] G. Ososkov, V. Palichik, and E. Tikhonenko. Robust technique with sub-optimal weight function for track fitting in CMS muon strip chamber. Computer Physics Communications, 121–122:709, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701138.

Ohlsson:1992:TFD

[OPY92] Mattias Ohlsson, Carsten Peterson, and Alan L. Yuille. Track finding with deformable templates — the elastic arms approach. Computer Physics Communications, 71(1–2):77–98, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465592900749.

Otto:1990:RMC

[Ott90] A. Otto. 3D resistive MHD computations of magnetospheric physics. Computer Physics Communications, 59(1):185–195, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090168Z.

Ogata:1999:SCI

[OY99] Y. Ogata and T. Yabe. Shock capturing with improved numerical viscosity in primitive Euler representation. Computer Physics Communications, 119(2–3):179–193, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599001885.

Pacios:1991:PTB

[Pac91] L. Fernandez Pacios. A program to test basis sets for quantum calculations with the option to include effective core potentials. Computer Physics Communications, 67(2):309–324, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190025G.

Pinches:1998:HSC

[PAC^{+}98] S. D. Pinches, L. C. Appel, J. Candy, S. E. Sharapov, H. L. Berk, D. Borba, B. N. Breizman, T. C. Hender, K. I. Hopcraft, G. T. A. Huysmans, and W. Kerner. The HAGIS selfconsistent nonlinear wave-particle interaction model. Computer Physics Communications, 111(1–3):133–149, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000344.

Paetzold:1991:VDL

[Pae91] O. Paetzold. Vectorization of diffusion of lattice gas without double occupancy. Computer Physics Communications, 64(1):1–6, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190044L.

Page:1994:ESC

[Pag94] Michael Page. Electronic structure computations and theoretical chemical kinetics: developments at the interface. Computer Physics Communications, 84(1–3):115–130, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902070.

Painter:1993:DPA

[Pai93] Scott Painter. Data-parallel algorithms for Monte Carlo simulation of neoclassical transport in magnetically confined plasmas. Computer Physics Communications, 77(3):342–356, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390181B.

Pan:1993:MCE

[Pan93] Jicai Pan. A Monte Carlo event generator for high energy hadron-hadron collisions — ECCO 2.0. Computer Physics Communications, 78(1–2):190–198, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901545.

Panayotaros:1999:TWW

[Pan99]  Panayotis Panayotaros.  Traveling water waves on a

sphere. Computer Physics Communications, 121–122:386– 387, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003604.

Papp:1992:CBR

[Pap92a] Zoltán Papp. Calculating bound and resonant states in local and nonlocal coulomb-like potentials. Computer Physics Communications, 70(2):426–434, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290205D.

Papp:1992:CSS

[Pap92b] Zoltán Papp. Calculating scattering states in local and nonlocal coulomb-like potentials. Computer Physics Communications, 70(2):435–439, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290206E.

Papadopoulos:1997:EEG

[Pap97] Costas G. Papadopoulos. ERATO: Event generator for fourfermion production at LEP2 energies and beyond. Computer Physics Communications, 101(1–2):183–195, April 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559600149X.

Papadopoulos:1999:SLP

[Pap99]  Costas G. Papadopoulos.  Single leptoquark produc-

tion at high-energy e^{+}e^{−} colliders. Computer Physics Communications, 118(1):81–91, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599001939.

Partl:1990:HMT

[Par90] Hubert Partl. How to make TEX and L^{A}TEX international. Computer Physics Communications, 61(1–2):190–200, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090116I.

Parker:1999:BRM

[Par99] D. J. Parker. Book review: M. Bertero, P. Boccacci, Introduction to Inverse Problems in Imaging (1999) Inst. of Physics Publ. 0-750-30435-9. Computer Physics Communications, 120(2–3):269–270, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559900257X.

Passarino:1996:WDA

[Pas96] Giampiero Passarino. WTO — a deterministic approach to 4-fermion physics. Computer Physics Communications, 97 (3):261–303, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000410.

Plothow-Besch:1993:PLA

[PB93] H. Plothow-Besch. PDFLIB: a library of all available parton density functions of the nucleon, the pion and the photon and the corresponding α_{S} calculations. Computer Physics Communications, 75(3):396–416, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390051D.

Pisani:1996:UHF

[PB96] C. Pisani and U. Birkenheuer. Unrestricted Hartree–Fock treatment of paramagnetic defect centers in non-magnetic crystals VI. Computer Physics Communications, 96(2–3): 152–166, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000549.

Park:1994:NDS

[PBDZ94] S. C. Park, C. Bahri, J. P. Draayer, and S.-Q. Zheng. Numerical database system based on a weighted search tree. Computer Physics Communications, 82(2–3):247–264, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901724.

Partridge:1994:IES

[PBL94] Harry Partridge, Charles W. Bauschlicher, Jr., and Stephen R. Langhoff. Ab initio electronic structure studies in molecular spectroscopy and chemical thermodynamics. Computer Physics Communications, 78(3):223–237, January 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900019.

Perger:1993:NEG

[PBN93] Warren F. Perger, Atul Bhalla, and Mark Nardin. A numerical evaluator for the generalized hypergeometric series. Computer Physics Communications, 77(2):249–254, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390008Z.

Pais:1990:CMC

[PC90] V. A. Pais and A. Caruso. A collisional method to correct the mesh distortion in two-dimensional Lagrangian hydrocodes. Computer Physics Communications, 58(1–2):55–62, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090134M.

Pisani:1994:HFPe

[PC94a] C. Pisani and F. Corà. Hartree–Fock perturbed-cluster treatment of defects in crystals V. the overlap correction in relaxation studies. Computer Physics Communications, 82(2–3):187–192, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901678.

Pisani:1994:HFPc

[PC94b] C. Pisani and F. Corà. Hartree–Fock perturbed-cluster treatment of local defects in crystals III. the use of local symmetry. Computer Physics Communications, 82(2–3):168–178, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594901651.

Pisani:1994:HFPd

[PC94c] C. Pisani and F. Corà. Hartree–Fock perturbed-cluster treatment of local defects in crystals IV. the cluster electro-chemical potential. Computer Physics Communications, 82(2–3):179–186, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559490166X.

Palassini:1999:FSS

[PC99] M. Palassini and S. Caracciolo. Finite size scaling in 3D Ising spin glasses. Computer Physics Communications, 121–122:180–182, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003070.

Pisani:1994:HFPb

[PCC94] C. Pisani, S. Casassa, and F. Corà. Hartree–Fock perturbedcluster treatment of local defects in crystals II. the energy dependent coupling matrices. Computer Physics Communications, 82(2–3):157–167, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901643.

Pearlman:1995:APC

[PCC^{+}95] David A. Pearlman, David A. Case, James W. Caldwell, Wilson S. Ross, Thomas E. Cheatham III, Steve DeBolt, David Ferguson, George Seibel, and Peter Kollman. AM-BER, a package of computer programs for applying molecular mechanics, normal mode analysis, molecular dynamics and free energy calculations to simulate the structural and energetic properties of molecules. Computer Physics Communications, 91(1–3):1–41, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500041D.

Pisani:1994:HFPa

[PCNO94] C. Pisani, F. Corà, R. Nada, and R. Orlando. Hartree– Fock perturbed-cluster treatment of local defects in crystals: I. the EMBED program: general features. Computer Physics Communications, 82(2–3):139–156, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901635.

Pastres:1999:GSA

[PCSD99] R. Pastres, K. Chan, C. Solidoro, and C. Dejak. Global sensitivity analysis of a shallow-water 3D eutrophication model. Computer Physics Communications, 117(1–2):62–74, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001647.

Parker:1990:DNS

[PD90] R. D. Parker and R. L. Dewar. 2-D, nonlinear spectral simulation of reconnective transitions on a periodic, planar current sheet with (1) smooth and (2) corrugated conducting wall boundary conditions with flow. Computer Physics Communications, 59(1):1–12, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090151P.

Parkes:1996:ATF

[PD96] E. J. Parkes and B. R. Duffy. An automated tanhfunction method for finding solitary wave solutions to non-linear evolution equations. Computer Physics Communications, 98(3):288–300, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559600104X.

Potapenko:1999:NSH

[PdA99] I. F. Potapenko and C. A. de Azevedo. Numerical simulation of heating problems for a weakly collisional plasma. Computer Physics Communications, 121–122:274–277, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003306.

Papageorgiou:1998:MMO

[PDL98a] D. G. Papageorgiou, I. N. Demetropoulos, and I. E. Lagaris. Merlin-3.0: a multidimensional optimization environment. Computer Physics Communications, 109(2–3):227–249, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000058.

Papageorgiou:1998:MCL

[PDL98b] D. G. Papageorgiou, I. N. Demetropoulos, and I. E. Lagaris. The merlin control language for strategic optimization. Computer Physics Communications, 109(2–3):250–275, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559800006X.

Park:1992:FSM

[PDZ92] S. C. Park, J. P. Draayer, and S.-Q. Zheng. Fast sparse matrix multiplication. Computer Physics Communications, 70(3):557–568, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290116G.

Peskov:1993:MMC

[Pes93] N. V. Peskov. Microcomputer Monte Carlo simulations of III-V semiconductor growth during molecular beam epitaxy. Computer Physics Communications, 77(1):64–68, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390037D.

Petersen:1992:IPD

[Pet92] Johnny Petersen. Introduction to programming on distributed memory multiprocessors. Computer Physics Communications, 73(1–3):72–92, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900303.

Puertas:1999:BDS

[PFBdlN99] A. M. Puertas, A. Fernández-Barbero, and F. J. de las Nieves. Brownian dynamics simulation of diffusive mesoscopic particle aggregation. Computer Physics Communications, 121–122:353–357, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003525.

Parpia:1996:GPL

[PFG96] F. A. Parpia, C. Froese Fischer, and I. P. Grant. GRASP92: a package for large-scale relativistic atomic structure calculations. Computer Physics Communications, 94(2–3):249–271, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001360.

Porta:1999:EII

[PFVC99] Marcel Porta, Carlos Frontera, Eduard Vives, and Teresa Castán. Effect of interatomic interactions on antiphase domain growth in two-dimensional ternary alloys. Computer Physics Communications, 121–122:719, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701229.

Pfalzner:1994:HTC

[PG94] Susanne Pfalzner and Paul Gibbon. A 3D hierarchical tree code for dense plasma simulation. Computer Physics Communications, 79(1):24–38, February 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594902275.

Pollock:1996:CPF

[PG96] E. L. Pollock and Jim Glosli. Comments on P^{3}M, FMM, and the Ewald method for large periodic Coulombic systems. Computer Physics Communications, 95(2–3):93–110, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000434.

Palmer:1991:CCS

[PGJ91] D. R. Palmer, T. M. Georges, and R. M. Jones. Classical chaos and the sensitivity of the acoustic field to small-scale ocean structure. Computer Physics Communications, 65(1–3):219–223, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190174J.

Poedts:1990:NSS

[PGK90a] Stefaan Poedts, Marcel Goossens, and Wolfgang Kerner. Numerical simulation of the stationary state of periodically driven coronal loops. Computer Physics Communications, 59(1):75–84, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090157V.

Poedts:1990:TER

[PGK90b] Stefaan Poedts, Marcel Goossens, and Wolfgang Kerner. Temporal evolution of resonant absorption in solar coronal loops. Computer Physics Communications, 59(1):95–103, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090159X.

Patterson:1999:MBT

[PHN99] C. H. Patterson, C. D. Hogan, and M. Nicastro. Many-body theory applied to optical properties of silicon surfaces. Computer Physics Communications, 121–122:711, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506701151.

Perger:1993:CWS

[PHT93] W. F. Perger, Z. Halabuka, and D. Trautmann. Continuum wavefunction solver for GRASP. Computer Physics Communications, 76(2):250–262, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390136Z.

Pi:1992:EGI

[Pi92] Hong Pi. An event generator for interactions between hadrons and nuclei — FRITIOF version 7.0. Computer Physics Communications, 71(1–2):173–192, August 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290082A.

Perger:1995:RCS

[PI95] Warren F. Perger and Muhammad Idrees. Relativistic calculation of specific mass shifts for Ar^{+}, Ni, Kr^{+}, and Ce^{+} using a multi-configuration Dirac–Fock approach. Computer Physics Communications, 85(3):389–397, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400109F.

Palagyi:1999:MCS

[PI99] G. Palágyi and F. Iglói. Monte Carlo simulations of twodimensional random Potts models. Computer Physics Communications, 121–122:710, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670114X.

Pittau:1997:SMM

[Pit97] R. Pittau. A simple method for multi-leg loop calculations. Computer Physics Communications, 104(1–3):23–36, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000659.

Pittau:1998:SMM

[Pit98] R. Pittau. A simple method for multi-leg loop calculations 2: a general algorithm. Computer Physics Communications, 111(1–3):48–52, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559800040X.

Perez-Jorda:1998:PSS

[PJ98] JoséM. Pérez-Jordá. In-place self-sorting fast Fourier transform algorithm with local memory references. Computer Physics Communications, 108(1):1–8, January 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001161.

Perez-Jorda:1999:FAT

[PJ99] José M. Pérez-Jordá. Fast algorithms through divideand combine-and-conquer strategies. Computer Physics Communications, 117(3):239–240, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001106.

Perez-Jorda:1993:SEM

[PJSF93] JoséM. Pérez-Jordá and Emilio San-Fabián. A simple, efficient and more reliable scheme for automatic numerical integration. Computer Physics Communications, 77(1):46–56, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390035B.

Perez-Jorda:1992:SRE

[PJSFM92] JoséM. Pérez-Jordá, Emilio San-Fabián, and Federico Moscardó. A simple, reliable and efficient scheme for automatic numerical integration. Computer Physics Communications, 70(2):271–284, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592901922.

Perger:1991:RCW

[PK91] W. F. Perger and Vasan Karighattam. Relativistic continuum wavefunction solver. Computer Physics Communications, 66(2–3):392–402, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190085Y.

Putz:1998:OTP

[PK98] M. Pütz and A. Kolb. Optimization techniques for parallel molecular dynamics using domain decomposition. Computer Physics Communications, 113(2–3):145–167, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000745.

Pederiva:1999:FMC

[PK99] Francesco Pederiva and M. H. Kalos. Fermion Monte Carlo. Computer Physics Communications, 121–122:440– 445, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003781.

Papageorgiou:1990:MDP

[PL90] D. G. Papageorgiou and I. E. Lagaris. MERLIN-2.1 double precision. Computer Physics Communications, 58(1–2): 119–125, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090140V.

Popov:1999:TDM

[PL99] A. M. Popov and Y. Q. Liu. Three-dimensional magnetohydrodynamic NFTC code for simulation of nonlinear instabilities in toroidal plasmas. Computer Physics Communications, 121–122:718, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701217.

Pernot:1991:QTD

[PLC91] P. Pernot, W. A. Lester, Jr., and Ch. Cerjan. Quantum time-dependent treatment of molecular collisions: scattering of He by H_{2} (B^{1}Σ_{u}^{+} ). Computer Physics Communications, 63(1–3):259–264, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190254I.

Pletzer:1996:SFE

[Ple96] A. Pletzer. Symplectic finite element scheme: application to a driven problem with a regular singularity. Computer Physics Communications, 96(1):1–9, July 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000471.

Pryce:1991:NMP

[PM91] John D. Pryce and Marco Marletta. A new multi-purpose software package for Schrödinger and Sturm–Liouville computations. Computer Physics Communications, 62(1):42–52, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901196.

Perales:1999:DMR

[PMB99] A. Perales, M. Moscoso, and L. L. Bonilla. Drift model of resonant tunneling in superlattices including a time delay simulating electron scattering mechanisms. Computer Physics Communications, 121–122:713–714, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701175.

Preische:1993:PPP

[PMJ93] S. Preische, J. Manickam, and J. L. Johnson. A postprocessor for the PEST code. Computer Physics Communications, 76(3):318–327, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390059L.

Plummer:1998:CMI

[PMM98] M. Plummer, J. F. McCann, and L. Bojer Madsen. The calculation of multiphoton ionization rates of the hydrogen molecular ion. Computer Physics Communications, 114(1–3):94–119, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000812.

Pacher:1990:TPM

[PMUV90] Pál Pacher, Géza Márk, László Udvardi, and Imre Varga. Teaching physics by means of computer modeling. Computer Physics Communications, 61(1–2):260–266, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090126L.

Pabon-Ortiz:1992:LPN

[POA92] Carlos U. Pabon-Ortiz and M. Artoni. Laguerre polynomials: novel properties and numerical generation scheme for analysis of a discrete probability distribution. Computer Physics Communications, 71(3):215–221, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290010V.

Podolak:1999:FNN

[Pod99] Igor T. Podolak. Feedforward neural network’s sensitivity to input data representation. Computer Physics Communications, 117(1–2):181–188, March 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001726.

Popelier:1996:MPA

[Pop96] Paul L. A. Popelier. MORPHY, a program for an automated “atoms in molecules” analysis. Computer Physics Communications, 93(2–3):212–240, February 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001131.

Popowicz:1997:S

[Pop97] Ziemowit Popowicz. SUSY 2. Computer Physics Communications, 100(3):277–296, March 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001518.

Popelier:1998:MIA

[Pop98] P. L. A. Popelier. A method to integrate an atom in a molecule without explicit representation of the interatomic surface. Computer Physics Communications, 108(2–3):180–190, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001215.

Portilho:1990:MMP

[Por90] Oyanarte Portilho. MP — a multiple precision package. Computer Physics Communications, 59(2):345–358, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590901832.

Portugal:1998:AST

[Por98] R. Portugal. An algorithm to simplify tensor expressions. Computer Physics Communications, 115(2–3):215–230, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001179.

Poschl:1998:BSF

[Pös98] W. Pöschl. B-spline finite elements and their efficiency in solving relativistic mean field equations. Computer Physics Communications, 112(1):42–66, July 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000034.

Potvliege:1998:SPT

[Pot98] R. M. Potvliege. STRFLO: a program for time-independent calculations of multiphoton processes in one-electron atomic systems I. quasienergy spectra and angular distributions. Computer Physics Communications, 114(1–3):42–93, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000733.

Potter:1999:JCO

[Pöt99] B. Pötter. JetViP 1.1: Calculating one- and two-jet cross sections with virtual photons in NLO QCD. Computer Physics Communications, 119(1):45–66, June 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001903.

Pozela:1991:MCS

[Poz91] Juras Pozela. Monte Carlo simulation of charge-carrier behavior in electric fields. Computer Physics Communications, 67(1):105–118, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591902249.

Popova:1992:LEG

[PP92] L. Popova and P. Petrova. LIPAOL: an event generator for hadron interaction simulations. Computer Physics Communications, 72(1):39–60, October 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290005J.

Pi:1994:DPF

[PP94] Hong Pi and Carsten Peterson. Delta 2.0- A program for finding dependencies in tables of data. Computer Physics Communications, 83(2–3):293–306, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900566.

Pazirandeh:1999:FNS

[PP99] A. Pazirandeh and B. Piroozfar. Fission neutron scattering simulation in ice using Monte Carlo method. Computer Physics Communications, 121–122:712, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701163.

Pfalzner:1993:TFM

[PR93] S. Pfalzner and S. J. Rose. Thomas–Fermi model for dense plasmas with strong electric fields. Computer Physics Communications, 75(1–2):98–104, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390167B.

Penna:1999:GOC

[PRdM99] T. J. P. Penna, A. Racco, and M. Argollo de Menezes. Getting older with computers. Computer Physics Communications, 121–122:108–112, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900291X.

Perlt:1990:CQG

[PRH90] H. Perlt, J. Ranft, and J. Heinrich. Calculation of QED graphs with the spinor technique. Computer Physics Communications, 56(3):385–390, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090021R.

Peterson:1994:JVA

[PRL94] Carsten Peterson, Thorsteinn Rögnvaldsson, and Leif Lönnblad. JETNET 3.0 — a versatile artificial neural network package. Computer Physics Communications, 81(1–2):185–220, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901201.

Pulkkinen:1999:SOE

[PRRL99] Ulla Pulkkinen, Tapio T. Rantala, Tuomo S. Rantala, and Vilho Lantto. Simulation of oxygen exchange of SnO_{2} surface. Computer Physics Communications, 121–122:720, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701230.

Pettitt:1995:P

[PS95] B. Montgomery Pettitt and Jeremy C. Smith. Preface. Computer Physics Communications, 91(1–3):xii, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559590011X.

Portugal:1997:AMG

[PS97] R. Portugal and S. L. Sautú. Applications of Maple to general relativity. Computer Physics Communications, 105 (2–3):233–253, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597000787.

Peysson:1998:AFP

[PS98a] Y. Peysson and M. Shoucri. An approximate factorization procedure for solving nine-point elliptic difference equations application for a fast 2-D relativistic Fokker–Planck solver. Computer Physics Communications, 109(1):55–80, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001434.

Pickering:1998:MPM

[PS98b] S. Pickering and I. K. Snook. A massively parallel molecular dynamics algorithm for the maspar supercomputer. Computer Physics Communications, 108(2–3):200–210, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001136.

Pastor-Satorras:1999:BAC

[PSR99] R. Pastor-Satorras and J. M. Rubí. Ballistic adsorption of colloidal magnetic particles. Computer Physics Communications, 121–122:265–267, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003276.

Puzynin:1999:HOA

[PSV99] I. V. Puzynin, A. V. Selin, and S. I. Vinitsky. A high-order accuracy method for numerical solving of the time-dependent Schrödinger equation. Computer Physics Communications, 123(1–3):1–6, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002246.

Phillips:1992:ASO

[PT92] G. M. Phillips and P. J. Taylor. Algorithms for spline and other approximations to functions and data. Computer Physics Communications, 73(1–3):1–21, December 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290025T.

Parker:1995:PIP

[PT95] Jonathan Parker and K. T. Taylor. Parallelized implicit propagators for the finite-difference Schrödinger equation. Computer Physics Communications, 88(2–3):217–228, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500036F.

Perez-Tudela:1999:AEC

[PTMUJdFM99] J. Pérez-Tudela, M. Montes-Usategui, I. Juvells, and J. R. de F. Moneo. Analysis and evaluation of converging diffractometers for optical correlators. Computer Physics Communications, 121–122:715, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701187.

Pech:1997:DCI

[PTU97] J. Pech, A. Tarancón, and C. L. Ullod. A dedicated computer for ising-like spin glass models. Computer Physics Communications, 106(1–2):10–20, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000490.

Pullan:1997:GOA

[Pul97] W. J. Pullan. Genetic operators for the atomic cluster problem. Computer Physics Communications, 107(1–3):137–148, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000921.

Purrington:1990:PVC

[Pur90] R. D. Purrington. Particle-vibration coupling model for odd-A nuclei. Computer Physics Communications, 58(1–2): 211–215, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559090146R.

Puri:1999:KPS

[Pur99] Sanjay Puri. Kinetics of phase separation near surfaces. Computer Physics Communications, 121–122:312– 316, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003422.

Pesola:1999:ISD

[PvBMN99] M. Pesola, J. von Boehm, T. Mattila, and R. M. Nieminen. Ab-initio spin-density pseudopotential calculation of the local vibrations of the oxygen complexes in silicon. Computer Physics Communications, 121–122:716, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701199.

Poschl:1996:AFE

[PVR96] W. Pöschl, D. Vretenar, and P. Ring. Application of the finite element method in self-consistent relativistic mean field calculations. Computer Physics Communications, 99(1):128–148, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001002.

Poschl:1997:RHB

[PVR97] W. Pos̈chl, D. Vretenar, and P. Ring. Relativistic Hartree– Bogoliubov theory in coordinate space: Finite element solution for a nuclear system with spherical symmetry. Computer Physics Communications, 103(2–3):217–250, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000428.

Poschl:1997:AFE

[PVRR97] W. Pöschl, D. Vretenar, A. Rummel, and P. Ring. Application of finite element methods in relativistic mean-field theory: spherical nucleus. Computer Physics Communications, 101(1–2):75–107, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597845833.

Platkowski:1999:EDV

[PW99] T. Platkowski and W. Waluś. An efficient discrete-velocity method for the Boltzmann equation. Computer Physics Communications, 121–122:717, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701205.

Parpia:1993:PGC

[PWG93] Farid A. Parpia, Wasantha P. Wijesundera, and Ian P. Grant. A program for generating complete active spaces for relativistic atomic structure calculations. Computer Physics Communications, 76(1):127–139, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390126W.

Piomelli:1991:LES

[PZ91] Ugo Piomelli and Thomas A. Zang. Large-eddy simulation of transitional channel flow. Computer Physics Communications, 65(1–3):224–230, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190175K.

Parker:1996:MPS

[PZH^{+}96] Gregory A. Parker, Wei Zhu, Youhong Huang, David K. Hoffman, and Donald J. Kouri. Matrix pseudo-spectroscopy: iterative calculation of matrix eigenvalues and eigenvectors of large matrices using a polynomial expansion of the Dirac delta function. Computer Physics Communications, 96(1):27–35, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001441.

QCDTARO:1995:QCS

[QAd^{+}95] QCD TARO Collaboration, K. Akemi, Ph. deForcrand, M. Fujisaki, T. Hashimoto, H. C. Hege, S. Hioki, O. Miyamura, A. Nakamura, M. Okuda, I. O. Stamatescu, Y. Tago, and T. Takaishi. Quantum chromodynamics simulations on a non-dedicated highly parallel computer. Computer Physics Communications, 90(2–3):201–214, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500082Q.

Quigley:1998:QPA

[QBP98] Lisa Quigley, Keith Berrington, and John Pelan. The QB program: Analysing resonances using R-matrix theory. Computer Physics Communications, 114(1–3):225–235, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000599.

Qin:1999:SVA

[QTR99] H. Qin, W. M. Tang, and G. Rewoldt. Symbolic vector analysis in plasma physics. Computer Physics Communications, 116(1):107–120, January 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001441.

Quartapelle:1995:SST

[QV95] L. Quartapelle and M. Verri. On the spectral solution of the three-dimensional Navier–Stokes equations in spherical and cylindrical regions. Computer Physics Communications, 90(1):1–43, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500072N.

Radjai:1999:MDG

[Rad99a] Farhang Radjai. Multicontact dynamics of granular systems. Computer Physics Communications, 121–122:294– 298, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003379.

Romera:1999:FIU

[RAD99b] E. Romera, J. C. Angulo, and J. S. Dehesa. Fisher information and uncertainty-like relationships in many-body systems. Computer Physics Communications, 121–122:729, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701321.

Raine:1991:BRB

[Rai91] Andrew C. R. Raine. Book review: Computer modelling of fluids, polymers and solids: C. R. A. Catlow, S. C. Parker and M. P. Allen, eds., Kluwer, Dordrecht, 1990. 540 pages. Hardcover price Dfl. 265,-. ISBN 0-7923-0549-3. Computer Physics Communications, 62(2–3):381–382, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190109X.

Randrup:1990:MSM

[Ran90] Jørgen Randrup. Microcanonical sampling of momenta in simulations of many-particle systems. Computer Physics Communications, 59(3):439–446, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090085F.

Randrup:1993:WSN

[Ran93] Jørgen Randrup. WIX: statistical nuclear multifragmentation with collective expansion and Coulomb forces. Computer Physics Communications, 77(2):153–166, October 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390001S.

Rapaport:1991:MMPa

[Rap91a] D. C. Rapaport. Multi-million particle molecular dynamics I. design considerations for vector processing. Computer Physics Communications, 62(2–3):198–216, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900953.

Rapaport:1991:MMPb

[Rap91b] D. C. Rapaport. Multi-million particle molecular dynamics: II. design considerations for distributed processing. Computer Physics Communications, 62(2–3):217–228, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591900964.

Rapaport:1993:MMP

[Rap93] D. C. Rapaport. Multi-million particle molecular dynamics: III. design considerations for data-parallel processing. Computer Physics Communications, 76(3):301–317, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390058K.

Rybicki:1996:SLS

[RARF96] J. Rybicki, W. Alda, A. Rybicka, and S. Feliziani. Structure of lead-silicate glasses via constant-pressure MD simulations. Computer Physics Communications, 97(1–2):191–194, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000367.

Rabitz:1999:EIO

[RASS99] Herschel Rabitz, Ömer F. Alis, Jeffrey Shorter, and Kyurhee Shim. Efficient input-output model representations. Computer Physics Communications, 117(1–2):11–20, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001520.

Raab:1999:TNM

[RB99] A. Raab and G. Beikert. Two new Monte Carlo methods for point flux estimation. Computer Physics Communications, 123(1–3):27–45, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002350.

Roy:1993:RAG

[RBB93] Dhiranjan Roy, Ranjan Bhattacharya, and Siddhartha Bhowmick. Rational approximants generated by the utransform. Computer Physics Communications, 78(1–2): 29–54, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390141X.

Roy:1996:RAU

[RBB96] Dhiranjan Roy, Ranjan Bhattacharya, and Siddhartha Bhowmick. Rational approximants using Levin–Weniger transforms. Computer Physics Communications, 93(2–3): 159–178, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465595001069.

Roy:1998:MLW

[RBB98] Dhiranjan Roy, Ranjan Bhattacharya, and Siddhartha Bhowmick. Multipoint Levin–Weniger approximants and their application to the ground state energies of quantum anharmonic oscillators. Computer Physics Communications, 113(2–3):131–144, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000952.

Romero:1999:OGL

[RBG99] David Romero, Carlos Barrón, and Susana Gómez. The optimal geometry of Lennard-Jones clusters: 148-309. Computer Physics Communications, 123(1–3):87–96, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002593.

Reale:1990:ODH

[RBP^{+}90] F. Reale, F. Brugé, G. Peres, S. L. Fornili, V. Martorana, and S. Serio. One-dimensional hydrodynamic modeling of coronal plasmas on transputer arrays. Computer Physics Communications, 60(2):201–210, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090005L.

Reale:1992:FPI

[RBS92] F. Reale, M. Barbera, and S. Sciortino. Fast parallel implementation of multidimensional data-domain FORTRAN codes on distributed-memory processor arrays. Computer Physics Communications, 72(2–3):129–143, November 1992.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290144N.

Reale:1994:PCU

[RBS94] F. Reale, F. Bocchino, and S. Sciortino. Parallel computing on Unix workstation arrays. Computer Physics Communications, 83(2–3):130–140, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900426.

Riccardo:1998:SSP

[RCL98] V. Riccardo, G. G. M. Coppa, and G. Lapenta. Spices 1-A smart-particle code for kinetic plasma simulation. Computer Physics Communications, 113(2–3):199–219, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000757.

Riba:1999:MAP

[RCVJ99] J. R. Riba, A. Carnicer, S. Vallmitjana, and I. Juvells. Multivariate analysis for pattern classification. Computer Physics Communications, 121–122:723, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701266.

Rohmer:1990:PSI

[RDB^{+}90] Marie-Madeleine Rohmer, Jean Demuynck, Marc Bénard, Roland Wiest, Christian Bachmann, Charles Henriet, and René Ernenwein. A program system for ab initio MO calculations on vector and parallel processing machines: II. SCF closed-shell and open-shell iterations. Computer Physics Communications, 60(1):127–144, August 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090082C.

Ratis:1993:CCH

[RdC93] Yu. L. Ratis and P. Fernández de Córdoba. A code to calculate (high order) Bessel functions based on the continued fractions method. Computer Physics Communications, 76(3):381–388, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390062H.

Rejcek:1997:AFK

[RDF^{+}97] J. M. Rejcek, S. Datta, N. G. Fazleev, J. L. Fry, and A. Korzeniowski. Application of the Feynman–Kac path integral method in finding the ground state of quantum systems. Computer Physics Communications, 105(2–3):108–126, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000611.

Racco:1999:MAT

[RdMP99] A. Racco, M. Argollo de Menezes, and T. J. P. Penna. Mutation accumulation theory and species-dependent mortality curves. Computer Physics Communications, 121–122: 721, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701242.

Roy:1995:EGT

[RDMS95] U. Roy, L. J. Dubé, P. Mandal, and N. C. Sil. Evaluation of a general three-denominator Lewis integral. Computer Physics Communications, 92(2–3):277–289, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001214.

Reale:1995:TCD

[Rea95] F. Reale. Thermal conduction in a 2-D FCT plasma hydrodynamic code. Computer Physics Communications, 86(1–2):13– 24, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500002W.

Rebbi:1993:BRB

[Reb93] Claudio Rebbi. Book review: Computational methods in field theory: Proceedings of the 31. Internationale Universitätswochen für Kern- und Teilchenphysik, Schladming, Austria, February 1992. Edited by H. Gausterer and C. B. Lang, Springer-Verlag, Berlin, 1992. 274 + xii pages. Hardcover price DM 86.00. ISBN 3-540-55997-3. Computer Physics Communications, 78(1–2):220–221, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901589.

Reid:1990:INP

[Rei90] Ian Reid. Interfaces to numerical packages. Computer Physics Communications, 61(1–2):141–149, November 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090112E.

Reid:1993:LCO

[Rei93] John S. Reid. Laue crystallographic orientation with a fixed detector. Computer Physics Communications, 75(1–2):259– 273, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390177E.

Renner:1992:CRP

[Ren92] Friedrich Renner. A constructive REDUCE package based upon the Painlevé analysis of nonlinear evolutions equations in Hamiltonian and / or normal form. Computer Physics Communications, 70(2):409–416, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290203B.

Reed:1994:HMD

[RF94] Mark S. C. Reed and K. M. Flurchick. Hybrid molecular dynamics: an approach to low density simulations. Computer Physics Communications, 81(1–2):56–64, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901104.

Rossi:1995:SAR

[RF95] M. Rossi and A. P. Flitney. Symbolic algebra and renormalization of gauge theories. Computer Physics Communications, 90(2–3):189–200, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500098Z.

Reed:1996:IAD

[RF96] Mark S. C. Reed and K. M. Flurchick. Investigation of artifacts due to periodic boundary conditions. Computer Physics Communications, 95(1):39–46, May 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001409.

Reynders:1995:OOO

[RFH^{+}95] John V. W. Reynders, David W. Forslund, Paul J. Hinker, Marydell Tholburn, David G. Kilman, and William F. Humphrey. OOPS: an object-oriented particle simulation class library for distributed architectures. Computer Physics Communications, 87(1–2):212–224, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400172X.

Richardson:1999:CCA

[RFMM99] Steven L. Richardson, Joseph S. Francisco, Alexander Mebel, and Keiji Morokuma. Can chlorine anion catalyze the reaction of HOC1 with HCI? Computer Physics Communications, 121–122:724, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701278.

Ramek:1992:PSA

[RG92] Michael Ramek and Bruno Gruber. Programs for symmetry adaptation coefficients for semisimple symmetry chains: the general case. Computer Physics Communications, 70(2):371– 388, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290200I.

Ring:1997:CPR

[RGL97] P. Ring, Y. K. Gambhir, and G. A. Lalazissis. Computer program for the relativistic mean field description of the ground state properties of even-even axially deformed nuclei. Computer Physics Communications, 105(1):77–97, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000222.

Ritley:1998:PSE

[RGL^{+}98] K. A. Ritley, V. J. Ghosh, K. G. Lynn, M. McKeown, and D. O. Welch. POS-SPRITE — an extensible calculation of positron and electron implantation in metals. Computer Physics Communications, 109(2–3):93–110, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000083.

Roldan:1999:CSS

[RGLV^{+}99] J. B. Roldán, F. Gámiz, J. A. López-Villanueva, J. E. Carceller, and P. Cartujo. A computational study of the strained-si MOSFET: a possible alternative for the next century electronics industry. Computer Physics Communications, 121–122:547–549, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900404X.

Rombouts:1999:GQM

[RH99] S. M. A. Rombouts and K. Heyde. A general quantum Monte Carlo method for fermions, free of discretization errors. Computer Physics Communications, 121–122:446–448, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003793.

Riffert:1995:NAS

[RHFR95] Harald Riffert, Heinz Herold, Olaf Flebbe, and Hanns Ruder. Numerical aspects of the smoothed particle hydrodynamics method for simulating accretion disks. Computer Physics Communications, 89(1–3):1–16, August 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001822.

Rodriguez-Hernandez:1999:BOC

[RHGDM99] P. Rodríguez-Hernández, M. González-Díaz, and A. Muñoz. Band offset control in ZnSe/BeTe heterojunctions: an ab initio study. Computer Physics Communications, 121–122: 725, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550670128X.

Rhoades:1992:FAC

[Rho92] Clifford E. Rhoades, Jr. A fast algorithm for calculating particle interactions in smooth particle hydrodynamic simulations. Computer Physics Communications, 70(3):478–482, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290109C.

Rhodin:1997:IIM

[Rho97] Andreas Rhodin. IMAS Integrated Modeling and Analysis System for the solution of optimal control problems. Computer Physics Communications, 107(1–3):21–38, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001070.

Rathsman:1997:MMC

[RI97] J. Rathsman and G. Ingelman. MAJOR 1.5- A Monte Carlo generator for heavy Majorana neutrinos in ep collisions. Computer Physics Communications, 101(1–2):143–154, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001324.

Richardson:1991:VQS

[Ric91] John L. Richardson. Visualizing quantum scattering on the CM-2 supercomputer. Computer Physics Communications, 63(1–3):84–94, February 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190240L.

Rieger:1999:DSN

[Rie99] Heiko Rieger. Disordered systems near quantum critical points. Computer Physics Communications, 121–122:505– 509, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003938.

Ristow:1996:PMS

[Ris96] Gerald H. Ristow. Particles moving in spatially bounded, viscous fluids. Computer Physics Communications, 99(1):43–52, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001075.

Rittger:1992:HTB

[Rit92] E. Rittger. Handling three-body potentials in computer simulation. Computer Physics Communications, 67(3):412–434, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465592900509.

Rizzato:1999:BMA

[Riz99] F. B. Rizzato. Breakdown of modulational approximations in nonlinear wave interaction. Computer Physics Communications, 121–122:420–422, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003720.

Rycerz:1990:MDSa

[RJ90] Zbigniew A. Rycerz and Patrick W. M. Jacobs. Molecular dynamics simulation program of order N for condensed matter: I. MDPYRS1: scalar pyramid, short-range interactions. Computer Physics Communications, 60(1):53–74, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090079G.

Rycerz:1991:VPOa

[RJ91a] Zbigniew A. Rycerz and Patrick W. M. Jacobs. Vectorized program of order N for molecular dynamics simulation of condensed matter: I. MDPYRV1: Vector pyramid, short-range interactions. Computer Physics Communications, 62(1):125–144, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591901266.

Rycerz:1991:VPOb

[RJ91b] Zbigniew A. Rycerz and Patrick W. M. Jacobs. Vectorized program of order N for molecular dynamics simulation of condensed matter: II. MDSLAB1: Slab, short-range interactions. Computer Physics Communications, 62(1):145–161, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901277.

Rapaport:1999:SSA

[RJS99] D. C. Rapaport, J. E. Johnson, and J. Skolnick. Supramolecular self-assembly: molecular dynamics modeling of polyhedral shell formation. Computer Physics Communications, 121–122:231–235, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003197.

Reggiani:1991:MCC

[RKV91] Lino Reggiani, Tilmann Kuhn, and Luca Varani. Monte Carlo calculation of electronic noise in semiconductors. Computer Physics Communications, 67(1):135–143, August 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190226B.

Roediger:1998:MSB

[RL98] G. A. Roediger and W. P. Lidinsky. The multi-session bridge. Computer Physics Communications, 110(1–3):149–154, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001690.

Rico:1997:CMC

[RLER97] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Calculation of many-centre two-electron molecular integrals with STO. Computer Physics Communications, 105(2–3):216–224, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559700060X.

Rybicki:1996:CGF

[RLF96] J. Rybicki, R. Laskowski, and S. Feliziani. Crystallisation and glass formation in liquid lead: a molecular dynamics study. Computer Physics Communications, 97(1–2):185–190, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000355.

Reber:1999:QSM

[RLP^{+}99] M. Reber, D. Löding, M. Presber, Chr. Rickwardt, and P. Nielaba. Quantum simulations in materials science: molecular monolayers and crystals. Computer Physics Communications, 121–122:524–527, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003975.

Rico:1991:CTC

[RLPR91] Jaime Fernández Rico, Rafael López, Miguel Paniagua, and Guillermo Ramírez. Calculation of two-center oneelectron molecular integrals with STOs. Computer Physics Communications, 64(2):329–342, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190041I. See erratum [RLPR94].

Rico:1994:EBT

[RLPR94] J. Fernández Rico, R. López, M. Paniagua, and G. Ramírez. Erratum: Calculation of two-center one-electron molecular integrals with STOs (Comput. Phys. Commun. 64 (1991) 329– 342). Computer Physics Communications, 82(2–3):306–307, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594901759. See [RLPR91].

Rasch:1997:NES

[RLWW97] J. Rasch, S. P. Lucey, Colm T. Whelan, and H. R. J. Walters. On the numerical evaluation of six dimensional integrals occurring in scattering problems. Computer Physics Communications, 101(3):197–205, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465596001646.

Rafelski:1990:PFP

[RM90] H. E. Rafelski and B. Müller. PC-Fortran programs for muon reactivation calculations in muon-catalyzed fusion. Computer Physics Communications, 59(3):521–526, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090093G.

Rodriguez-Miguel:1999:TMO

[RMMUJ99] F. J. Rodríguez-Miguel, M. Montes-Usategui, and I. Juvells. 3D target modeling for optical correlators. Computer Physics Communications, 121–122:726, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701291.

Rodriguez-Navarro:1999:CSC

[RNGR99] A. Rodriguez-Navarro and J. M. Garcia-Ruiz. Computer simulation of competitive crystal growth. Computer Physics Communications, 121–122:727, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701308.

Rosa:1999:COD

[RNR^{+}99] R. R. Rosa, C. R. Neto, F. M. Ramos, A. S. Sharma, and J. A. Valdivia. Computational operators for dynamical complex pattern recognition. Computer Physics Communications, 121–122:731, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701345.

Roberts:1997:LDM

[Rob97] A. J. Roberts. Low-dimensional modelling of dynamics via computer algebra. Computer Physics Communications, 100 (3):215–230, March 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001622.

Robertson:1998:CTS

[Rob98a] Les Robertson. Computing trends and strategy at CERN. Computer Physics Communications, 110(1–3):6–11, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001458.

Robinson:1998:JPI

[Rob98b] A. W. Robinson. JavaFit: a platform independent program for interactive nonlinear least-squares fitting using the Levenberg–Marquardt method. Computer Physics Communications, 112(2–3):183–190, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000526.

RodriguezAzara:1992:MPG

[Rod92] Jose Luis Rodriguez Azara. A Maple program for the generation of the Lie-series solution of systems of nonlinear ordinary differential equations. Computer Physics Communications, 67(3):537–542, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900587.

Roux:1995:CPM

[Rou95] Benoît Roux. The calculation of the potential of mean force using computer simulations. Computer Physics Communications, 91(1–3):275–282, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500053I.

Refson:1991:CPC

[RP91] K. Refson and G. S. Pawley. Correlations in the plastic crystal phase of n-butane. Computer Physics Communications, 62(2–3):279–288, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190100Y.

Ramon:1995:PKV

[RP95] Javier Ramón and Jorge Peña. Parallelization of KENO-Va Monte Carlo code. Computer Physics Communications, 88(1): 76–82, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559500025B.

Romiszowski:1999:CSD

[RS99] Piotr Romiszowski and Andrzej Sikorski. Computer simulations of dynamics of protein-like heteropolymers. Computer Physics Communications, 121–122:730, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701333.

Rupp:1992:SMP

[RSW92] Bernhard Rupp, Bryan Smith, and Joe Wong. SEXIE — a microcomputer program for the calculation of coordination shells and geometrics. Computer Physics Communications, 67(3):543–547, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465592900598.

Rouhi:1996:NOS

[RSW96] Ali Rouhi, Roy Schult, and Jon Wright. A new operator splitting method for the numerical solution of partial differential equations II. Computer Physics Communications, 97(3):209–218, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559600077X.

Rieger:1999:GST

[RSW^{+}99] Martin M. Rieger, L. Steinbeck, I. D. White, H. N. Rojas, and R. W. Godby. The GW space–time method for the self-energy of large systems. Computer Physics Communications, 117(3):211–228, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800174X.

Rabadan:1998:RPC

[RT98] Ismanuel Rabadán and Jonathan Tennyson. ROTIONS: a program for the calculation of rotational excitation cross sections in electron-molecular ion collisions. Computer Physics Communications, 114(1–3):129–141, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800068X.

Righetti:1992:CIP

[RTLM92] F. Righetti, H. Telley, Th. M. Leibling, and A. Mocellin. 2D-CELL: image processing software for extraction and analysis of 2-dimensional cellular structures. Computer Physics Communications, 67(3):509–526, January 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592900565.

Raines:1992:CNF

[RU92] P. E. Raines and T. Uzer. Computing normal forms of nonseparable Hamiltonians by symbolic manipulation. Computer Physics Communications, 70(3):569–578, July 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290117H.

Rutz:1998:TBT

[Rut98] Solange F. Rutz. Theorems of Birkhoff type in Finsler spaces. Computer Physics Communications, 115(2–3):300–315, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001209.

Rodriguez-Villafuerte:1999:NMS

[RVB99] M. Rodríguez-Villafuerte and M. E. Brandan. Numerical model of the supralinear response of heavy chargedparticle TL dosemeters. Computer Physics Communications, 121–122:728, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550670131X.

Ramos:1999:BME

[RVCF99] F. M. Ramos, H. F. Campos Velho, J. C. Carvalho, and N. J. Ferreira. Beyond maximum entropy: novel approaches to entropic regularization. Computer Physics Communications, 121–122:722, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701254.

Rouhi:1995:NOS

[RW95] Ali Rouhi and Jon Wright. A new operator splitting method for the numerical solution of partial differential equations. Computer Physics Communications, 85(1):18–28, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400119M.

Rasch:1997:NEC

[RW97] J. Rasch and Colm T. Whelan. On the numerical evaluation of a class of integrals occurring in scattering problems. Computer Physics Communications, 101(1–2):31–46, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559600152X.

Rasch:1998:DIS

[RWL^{+}98] J. Rasch, Colm T. Whelan, S. P. Lucey, C. Dal Cappello, and H. R. J. Walters. 6-dimensional integrals and supercomputers. Computer Physics Communications, 114(1–3):378–384, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000691.

Rybowicz:1998:BRB

[Ryb98] M. Rybowicz. Book review: Introduction to scientific programming. Computational problem solving using Maple and C by Joseph L. Zachary (Springer, Berlin, 1996) ISBN 0-387-94630-6. Computer Physics Communications, 109(1):90–92, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000046.

Rycerz:1990:AMD

[Ryc90a] Zbigniew A. Rycerz. Acceleration of molecular dynamics simulation of order N with neighbour list. Computer Physics Communications, 60(3):297–303, October 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090028Y.

Rycerz:1990:MDSb

[Ryc90b] Zbigniew A. Rycerz. Molecular dynamics simulation program of order N for condensed-matter systems: II. MD-SPNL: pyramid with neighbour list, short-range interactions. Computer Physics Communications, 61(3):361–373, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090050B.

Revenga:1999:BCD

[RZE99] M. Revenga, I. Zúñiga, and P. Español. Boundary conditions in dissipative particle dynamics. Computer Physics Communications, 121–122:309–311, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003410.

Stimpfl-Abele:1991:RDC

[SA91] Georg Stimpfl-Abele. Recognition of decays of charged tracks with neural network techniques. Computer Physics Communications, 67(2):183–192, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190015D.

Shida:1992:REL

[SA92] Koichiro Shida and Yuichiro Anzai. Reduction of the eventlist for molecular dynamic simulation. Computer Physics Communications, 69(2–3):317–329, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901704.

Stimpfl-Abele:1995:VID

[SA95] Georg Stimpfl-Abele. Validation of input data for trained neural-nets. Computer Physics Communications, 85(2): 176–188, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559400142O.

Sanchez:1999:TIN

[SA99a] J. R. Sanchez and C. M. Arizmendi. Transfer of information in noise induced transport. Computer Physics Communications, 121–122:374–375, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003574.

Seppala:1999:DWR

[SA99b] E. T. Seppälä and M. J. Alava. Domain walls in the 3D random field Ising model. Computer Physics Communications, 121–122:732, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701357.

Stariolo:1999:EDS

[SA99c] Daniel A. Stariolo and Jeferson Arenzon. Off-equilibrium dynamics of the site frustrated percolation model. Computer Physics Communications, 121–122:739, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701424.

Sagar:1991:GQC

[Sag91a] Robin P. Sagar. A Gaussian quadrature for the calculation of generalized Fermi–Dirac integrals. Computer Physics Communications, 66(2–3):271–275, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190076W.

Stimpfl-Abele:1991:FTF

[SAG91b] Georg Stimpfl-Abele and Lluís Garrido. Fast track finding with neural networks. Computer Physics Communications, 64(1):46–56, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190048P.

Srinivasan:1997:DDD

[SAJ^{+}97a] S. G. Srinivasan, I. Ashok, Hannes Jônsson, Gretchen Kalonji, and John Zahorjan. Dynamic-domain-decomposition parallel molecular dynamics. Computer Physics Communications, 102(1–3):44–58, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000167.

Srinivasan:1997:PSR

[SAJ^{+}97b] S. G. Srinivasan, I. Ashok, Hannes Jônsson, Gretchen Kalonji, and John Zahorjan. Parallel short-range molecular dynamics using the adhara runtime system. Computer Physics Communications, 102(1–3):28–43, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000155.

Salin:1991:MMT

[Sal91] Antoine Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. IV. Computer Physics Communications, 62(1):58–64, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190121Z.

Salgado:1994:GRC

[Sal94] M. Salgado. 3 + 1 General Relativity by computer. Computer Physics Communications, 79(2):309–328, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900752.

Salam:1997:OOE

[Sal97] G. P. Salam. OEDIPUS: Onium Evolution, Dipole Interaction and Perturbative Unitarisation Simulation. Computer Physics Communications, 105(1):62–76, September 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000660.

Schweinzer:1995:TPC

[SAP^{+}95] J. Schweinzer, F. Aumayr, P. Platzer, M. Schneider, D. Wutte, and H. P. Winter. Two programs for calculations of collisional atomic data for lithium beam plasma spectroscopy. Computer Physics Communications, 88(1):83–88, July 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500032B.

Sattin:1997:LFP

[Sat97a] Fabio Sattin. Lyap: a Fortran 90 program to compute the Lyapunov exponents of a dynamical system from a time series. Computer Physics Communications, 107(1–3):253–257, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001240.

Sattin:1997:RCE

[Sat97b] Fabio Sattin. A routine to compute the energy and wave function forone-electron two-nuclei molecular systems. Computer Physics Communications, 105(2–3):225–232, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000581.

Stimpfl-Abele:1993:HSN

[SAY93] Georg Stimpfl-Abele and Pablo Yepes. Higgs search and neural-net analysis. Computer Physics Communications, 78(1–2):1–16, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593901383.

Stefanou:1990:ENM

[SAZ90] N. Stefanou, H. Akai, and R. Zeller. An efficient numerical method to calculate shape truncation functions for wigner-seitz atomic polyhedra. Computer Physics Communications, 60(2):231–238, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090009P.

Schack:1997:CLU

[SB97] Rüdiger Schack and Todd A. Brun. A C++ library using quantum trajectories to solve quantum master equations. Computer Physics Communications, 102(1–3):210–228, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000192.

Sethian:1991:PIR

[SBGM91] J. A. Sethian, Jean-Philippe Brunet, Adam Greenberg, and Jill P. Mesirov. A parallel implementation of the random vortex method. Computer Physics Communications, 65(1– 3):231–237, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190176L.

Schroer:1998:DTD

[SBK98] A. Schröer, G. T. Birk, and A. Kopp. DENISIS — a threedimensional partially ionized dusty magnetoplasma code. Computer Physics Communications, 112(1):7–22, July 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000447.

Sasik:1991:NSN

[ŠČ91] R. Šášik and R. Černý. Numerical solution of the nonisothermal moving boundary problem in heat conduction. Computer Physics Communications, 64(2):241–251, May 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190035J.

Santos:1999:CSC

[SC99a] Maria Augusta Santos and Marta Chaves. Cluster statistics for a critical nonequilibrium model. Computer Physics Communications, 121–122:408–410, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003690.

Segredo:1999:LRI

[SC99b] Nélido González Segredo and Peter V. Coveney. Long-range interactions in a lattice gas automaton model for immiscible fluids. Computer Physics Communications, 121–122: 645, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700509.

Schafer:1991:EAT

[Sch91a] K. J. Schafer. The energy analysis of time-dependent numerical wave functions. Computer Physics Communications, 63(1–3):427–434, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190267O.

Schwenke:1991:CRV

[Sch91b] David W. Schwenke. Compact representation of vibrational wave functions for diatomic molecules. Computer Physics Communications, 62(1):1–7, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901152.

Schwenke:1992:CRV

[Sch92] David W. Schwenke. On the computation of ro-vibrational energy levels of triatomic molecules. Computer Physics Communications, 70(1):1–14, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290086E.

Scholz:1993:APM

[Sch93] T. T. Scholz. Averaging of pseudoresonant T -matrix elements. Computer Physics Communications, 74(2):256–264, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390096U.

Schuler:1998:TPP

[Sch98] Gerhard A. Schuler. Two-photon physics with GALUGA 2.0. Computer Physics Communications, 108(2–3):279–303, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001276.

Sciutto:1991:HPM

[Sci91] S. J. Sciutto. HUBCOM: a program for Monte Carlo simulations on generalized Hubbard models. Computer Physics Communications, 62(1):90–106, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901244.

Sciutto:1993:PSP

[Sci93a] S. J. Sciutto. POWEV: a subroutine package to evaluate eigenvalues and eigenvectors of large sparse matrices. Computer Physics Communications, 77(1):95–106, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390040J.

Sciutto:1993:SSP

[Sci93b] S. J. Sciutto. SPARSEM: a subroutine package to operate with large sparse matrices. Computer Physics Communications, 77(1):84–94, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390039F.

Sciutto:1994:CST

[Sci94] S. J. Sciutto. Comparative study of two algorithms for the calculation of the extreme eigenvalues of large matrices. Computer Physics Communications, 79(2):215–218, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900698.

Scrinzi:1995:DFE

[Scr95] Armin Scrinzi. A 3-dimensional finite elements procedure for quantum mechanical applications. Computer Physics Communications, 86(1–2):67–80, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594001604.

Seixas:1996:NSL

[SCS^{+}96] J. M. Seixas, L. P. Caloba, M. N. Souza, A. L. Braga, and A. P. Rodrigues. Neural second-level trigger system based on calorimetry. Computer Physics Communications, 95(2– 3):143–157, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000124.

Saltelli:1999:P

[SCS99] Andrea Saltelli, Karen Chan, and Marian Scott. Preface. Computer Physics Communications, 117(1–2):xi–xiv, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001507.

Scalerandi:1996:TEG

[SDB96] M. Scalerandi, P. P. Delsanto, and S. Biancotto. Time evolution of growth phenomena in the KPZ model. Computer Physics Communications, 97(1–2):195–198, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000379.

Segura:1997:CEM

[SdCR97] J. Segura, P. Fernández de Córdoba, and Yu. L. Ratis. A code to evaluate modified Bessel functions based on the continued fraction method. Computer Physics Communications, 105(2–3):263–272, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000696.

Secada:1999:NEH

[Sec99] José D. Secada. Numerical evaluation of the Hankel transform. Computer Physics Communications, 116(2–3):278–294, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001088.

Seiler:1991:SRP

[Sei91] Werner M. Seiler. SUPERCALC — a REDUCE package for commutator calculations. Computer Physics Communications, 66(2–3):363–376, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190082V.

Seiler:1994:PDO

[Sei94] Werner M. Seiler. Pseudo differential operators and integrable systems in AXIOM. Computer Physics Communications, 79(2):329–340, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900760.

Sekera:1992:VPH

[Sek92] Z. Sekera. Vectorization and parallelization on high performance computers. Computer Physics Communications, 73 (1–3):113–138, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290032T.

Selke:1996:MCS

[Sel96] W. Selke. Monte Carlo simulations of Ising models. Computer Physics Communications, 97(1–2):106–110, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000495.

Semenov:1998:LPA

[Sem98] A. Semenov. LanHEP — a package for automatic generation of Feynman rules from the Lagrangian. Computer Physics Communications, 115(2–3):124–139, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559800143X.

Seymour:1995:MEC

[Sey95] Michael H. Seymour. Matrix-element corrections to parton shower algorithms. Computer Physics Communications, 90 (1):95–101, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500064M.

Serrano:1999:LTT

[SEZ99a] M. Serrano, P. Español, and I. Zúñiga. Long time tails and hydrodynamics in dissipative particle dynamics. Computer Physics Communications, 121–122:734, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701370.

Serrano:1999:CED

[SEZ99b] M. Serrano, P. Español, and Ignacio Zúñiga. Collective effects in dissipative particle dynamics. Computer Physics Communications, 121–122:306–308, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003409.

Sturesson:1993:LPG

[SF93] Lennart Sturesson and Charlotte Froese Fischer. LSGEN — a program to generate configuration-state lists of LScoupled basis functions. Computer Physics Communications, 74(3):432–440, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465593900247.

Smith:1994:PMSa

[SF94a] W. Smith and T. R. Forester. Parallel macromolecular simulations and the replicated data strategy: I. the computation of atomic forces. Computer Physics Communications, 79(1):52–62, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902291.

Smith:1994:PMSb

[SF94b] W. Smith and T. R. Forester. Parallel macromolecular simulations and the replicated data strategy: II. the RD-SHAKE algorithm. Computer Physics Communications, 79(1):63–77, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594902305.

Stathopoulos:1994:DPF

[SF94c] Andreas Stathopoulos and Charlotte Froese Fischer. A Davidson program for finding a few selected extreme eigenpairs of a large, sparse, real, symmetric matrix. Computer Physics Communications, 79(2):268–290, April 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900736.

Stedman:1998:TQM

[SF98] M. L. Stedman and W. M. C. Foulkes. Talus — a quantum Monte Carlo modelling suite. Computer Physics Communications, 113(2–3):180–198, October 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000873.

Saunders:1994:EPL

[SFFDR94] V. R. Saunders, C. Freyria-Fava, R. Dovesi, and C. Roetti. On the electrostatic potential in linear periodic polymers. Computer Physics Communications, 84(1–3):156–172, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902097.

Schneider:1996:SBE

[SFM^{+}96] Johannes Schneider, Christine Froschhammer, Ingo Morgenstern, Thomas Husslein, and Johannes Maria Singer. Searching for backbones — an efficient parallel algorithm for the traveling salesman problem. Computer Physics Communications, 96(2–3):173–188, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000628.

Salvat:1995:ANS

[SFVW95] F. Salvat, J. M. Fernández-Varea, and W. Williamson, Jr. Accurate numerical solution of the radial Schrödinger and Dirac wave equations. Computer Physics Communications, 90(1):151–168, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500039I.

Sobh:1991:PCG

[SG91] Nahil Atef Sobh and Karl Gustafson. Preconditioned conjugate gradient and finite element methods for massively data-parallel architectures. Computer Physics Communications, 65(1–3):253–267, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190179O.

Sanna:1998:DCS

[SG98a] N. Sanna and F. A. Gianturco. Differential cross sections for electron/positron scattering from polyatomic molecules. Computer Physics Communications, 114(1–3):142–167, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000915.

Segura:1998:PCF

[SG98b] J. Segura and A. Gil. Parabolic cylinder functions of integer and half-integer orders for nonnegative arguments. Computer Physics Communications, 115(1):69–86, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000976.

Segura:1999:EGT

[SG99] J. Segura and A. Gil. ELF and GNOME: Two tiny codes to evaluate the real zeros of the Bessel functions of the first kind for real orders. Computer Physics Communications, 117(3):250–262, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001933.

Scott:1993:PMD

[SGB^{+}93] W. Scott, A. Gunzinger, B. Bäumle, P. Kohler, U. A. Müller, H-R. Vonder Mühll, A. Eichenberger, W. Guggenbühl, N. Ironmonger, F. Müller-Plathe, and W. F. van Gunsteren. Parallel molecular dynamics on a multi signal processor system. Computer Physics Communications, 75(1–2):65–86, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593901659.

Sarsa:1999:VMC

[SGB99] Antonio Sarsa, Francisco J. Gálvez, and Enrique Buendía. Variational Monte Carlo calculation of two body properties in atoms: Importance sampling considerations. Computer Physics Communications, 121–122:493–495, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003896.

Schweizer:1999:DVM

[SGF99a] W. Schweizer and R. González-Férez. Discrete variable method for non-integrable quantum systems. Computer Physics Communications, 121–122:480–482, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003860.

Sternberg:1999:NNO

[SGF99b] Michael Sternberg, Giulia Galli, and Thomas Frauenheim. NOON — a non-orthogonal localised orbital order-N method.

Computer Physics Communications, 118(2–3):200–212, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002027.

Schreiber:1999:ELQ

[SGRZ99] Michael Schreiber, Uwe Grimm, Rudolf A. Römer, and Jianxin Zhong. Energy levels of quasiperiodic Hamiltonians, spectral unfolding, and random matrix theory. Computer Physics Communications, 121–122:499–501, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003914.

Stuart:1990:ARO

[SGT90] Robin G. Stuart and A. Góngora-T. Algebraic reduction of one-loop Feynman diagrams to scalar integrals. II. Computer Physics Communications, 56(3):337–350, January 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090019W.

Storey:1991:FCE

[SH91] P. J. Storey and D. G. Hummer. Fast computer evaluation of radiative properties of hydrogenic systems. Computer Physics Communications, 66(1):129–141, July 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190013B.

Sun:1997:FCP

[SH97] Yang Sun and Kenji Hara. Fortran code of the Projected Shell Model: feasible shell model calculations for heavy nuclei. Computer Physics Communications, 104(1–3):245–258, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597000647.

Shafai:1991:P

[Sha91] L. Shafai. Preface. Computer Physics Communications, 68(1–3):xiii, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190190V.

Sharan:1992:SCW

[Sha92] Pankaj Sharan. Symbolic computation of Wigner–Kirkwood expansion to O(h̄8). Computer Physics Communications, 69 (2–3):235–242, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290163S.

Shchur:1999:QRN

[Shc99] Lev N. Shchur. On the quality of random number generators with taps. Computer Physics Communications, 121–122:83–85, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002866.

Steckler:1995:PNV

[SHL^{+}95] Rozeanne Steckler, Wei-Ping Hu, Yi-Ping Liu, Gillian C. Lynch, Bruce C. Garrett, Alan D. Isaacson, Vasilios S. Melissas, Da hong Lu, Thanh N. Truong, Sachchida N. Rai, Gene C. Hancock, J. G. Lauderdale, Tomi Joseph, and Donald G. Truhlar. POLYRATE 6.5: a new version of a computer program for the calculation of chemical reaction rates for polyatomics. Computer Physics Communications, 88(2–3):341–343, August 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500037G.

Sunderland:1998:PMP

[SHNS98] A. G. Sunderland, J. W. Heggarty, C. J. Noble, and N. S. Scott. Parallelization of R-matrix propagation methods on distributed memory computers. Computer Physics Communications, 114(1–3):183–194, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000939.

Shukhman:1994:GQR

[Shu94] B. Shukhman. Generation of quasi-random (LP_{τ}) vectors for parallel computation. Computer Physics Communications, 78(3):279–286, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559490006X.

Signer:1997:MPJ

[Sig97] Adrian Signer. MENLO PARC, a program for e^{+}e^{−} → 4 jets at next-to-leading order. Computer Physics Communications, 106(1–2):125–138, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000933.

Sikorski:1999:PND

[Sik99] Andrzej Sikorski. Properties of de novo designed helical proteins. A Monte Carlo study of degrado dimers. Computer Physics Communications, 121–122:736, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701394.

Simos:1992:EFM

[Sim92] T. E. Simos. Exponential fitted methods for the numerical integration of the Schrödinger equation. Computer Physics Communications, 71(1–2):32–38, August 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290069B.

Simos:1993:HOM

[Sim93] T. E. Simos. High-order methods with minimal phaselag for the numerical integration of the special secondorder initial value problem and their application to the one-dimensional Schrödinger equation. Computer Physics Communications, 74(1):63–66, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559390106M.

Simos:1997:EOM

[Sim97] T. E. Simos. Eighth order methods for accurate computations for the Schrödinger equation. Computer Physics Communications, 105(2–3):127–138, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559700074X.

Simos:1998:EFR

[Sim98] T. E. Simos. An exponentially-fitted Runge–Kutta method for the numerical integration of initial-value problems with periodic or oscillating solutions. Computer Physics Communications, 115(1):1–8, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000885.

Simos:1999:EEO

[Sim99] T. E. Simos. Explicit eighth order methods for the numerical integration of initial-value problems with periodic or oscillating solutions. Computer Physics Communications, 119(1):32– 44, June 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559800188X.

Singer:1990:MMC

[Sin90] Sherwin J. Singer. Multiparticle Monte Carlo moves: Algorithm for solids with free-energy determination. Computer Physics Communications, 59(3):463–470, July 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090088I.

Sinclair:1994:EFG

[Sin94] R. Sinclair. The evaluation of Feynman graphs for a (2 + 1)-dimensional nonrelativistic electron gas. Computer Physics Communications, 81(1–2):173–184, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901198.

Sivaloganathan:1991:ULM

[Siv91] S. Sivaloganathan. The use of local mode analysis in the design and comparison of multigrid methods. Computer Physics Communications, 65(1–3):246–252, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190178N.

Suedan:1991:TDS

[SJ91] G. A. Suedan and E. V. Jull. Three-dimensional scalar beam diffraction by a half plane. Computer Physics Communications, 68(1–3):346–352, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465591902083.

Scott:1994:GPS

[SJB^{+}94] N. S. Scott, J. Johnston, V. M. Burke, C. J. Noble, and D. W. Busby. GRACE: the problem specification stage. II. Computer Physics Communications, 84(1–3):317–334, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594902186.

Sjostrand:1994:HEP

[Sjö94] Torbjörn Sjöstrand. High-energy-physics event generation with PYTHIA 5.7 and JETSET 7.4. Computer Physics Communications, 82(1):74–89, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901325.

Skrzypek:1996:MCP

[SJPW96] M. Skrzypek, S. Jadach, W. Placzek, and Z. Was. Monte Carlo program KORALW 1.02 for W -pair production at LEP2NLC energies with Yennie–Frautschi–Suura exponentiation. Computer Physics Communications, 94(2–3):216–248, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595001336.

Sklenar:1990:CAL

[SK90] Ivan Sklenár and Václav Kríz. Computer aided lexicography. Computer Physics Communications, 61(1–2):257–259, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090125K.

Skala:1990:NVP

[Ská90] Lubomír Skála. A new version of the program for the generation of symmetry-adapted functions for molecular calculations. Computer Physics Communications, 58(3):343–344, April 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090068C.

Schafer:1991:DBS

[SKG91] K. J. Schafer, N. H. Kwong, and J. D. Garcia. Development of B-spline techniques for dynamic ion-metal interaction theories. Computer Physics Communications, 63(1–3):306–313, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190257L.

Scott:1994:FSA

[SKM94] N. S. Scott, P. L. Kilpatrick, and D. Maley. The formal specification of abstract data types and their implementation in Fortran 90. Computer Physics Communications, 84(1–3):201–225, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902127.

Solvason:1995:RCE

[SKPP95] Dorthe Sølvason, Jirí Kolafa, Henrik G. Petersen, and John W. Perram. A rigorous comparison of the Ewald method and the fast multipole method in two dimensions. Computer Physics Communications, 87(3):307–318, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500027D.

Sun:1991:TDW

[SKST91] Yan Sun, Donald J. Kouri, David W. Schwenke, and Donald G. Truhlar. Time-dependent wavepacket algorithm for inelastic molecule-molecule scattering. Computer Physics Communications, 63(1–3):51–62, February 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190237F.

Shi:1999:GSI

[SKWH99] Zhuoer Shi, Donald J. Kouri, G. W. Wei, and David K. Hoffman. Generalized symmetric interpolating wavelets. Computer Physics Communications, 119(2–3):194–218, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559900185X.

Salanon:1994:HSD

[SL94a] B. Salanon and J. Lapujoulade. Helium scattering from disordered surfaces. Computer Physics Communications, 80 (1–3):32–52, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900973.

Stelzer:1994:AGT

[SL94b]  T. Stelzer and W. F. Long.  Automatic generation

of tree level helicity amplitudes. Computer Physics Communications, 81(3):357–371, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900841.

Stade:1995:GDF

[SL95] Eric Stade and E. G. Layton. Generalized discrete Fourier transforms: the discrete fourier-riccati–bessel transform. Computer Physics Communications, 85(3):336–370, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400124K.

Smithe:1997:ESA

[SL97] David Smithe and Larry Ludeking. An eigenmode solution algorithm based on high-order power iteration with fractally ordered shifts. Computer Physics Communications, 106(1–2):95–104, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000817.

Sobol:1999:UVR

[SL99] I. M. Sobol’ and Yu. L. Levitan. On the use of variance reducing multipliers in Monte Carlo computations of a global sensitivity index. Computer Physics Communications, 117(1–2):52–61, March 1, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598001568.

Somasundaram:1999:PGT

[SLB99] T. Somasundaram and R. M. Lynden-Bell. Permeability of gases through foams. Computer Physics Communications, 121–122:738, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701412.

Schmalian:1996:SCS

[SLGB96] J. Schmalian, M. Langer, S. Grabowski, and K. H. Bennemann. Self-consistent summation of many-particle diagrams on the real frequency axis and its application to the FLEX approximation. Computer Physics Communications, 93(2–3):141–151, February 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595001344.

Salvat:1991:ANS

[SM91a] Francesc Salvat and Ricardo Mayol. Accurate numerical solution of the Schrödinger and Dirac wave equations for central fields. Computer Physics Communications, 62(1):65–79, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901222.

Srivastava:1991:CTP

[SM91b] Deepak Srivastava and David A. Micha. Complex-time path-integration with time-dependent Hamiltonians for extended molecular systems. Computer Physics Communications, 63(1–3):331–344, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190260R.

Salvat:1993:ESE

[SM93] Francesc Salvat and Ricardo Mayol. Elastic scattering of electrons and positrons by atoms. Schrödinger and Dirac partial wave analysis. Computer Physics Communications, 74(3):358–374, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593900199.

Sim:1997:FPF

[SM97] Eunji Sim and Nancy Makri. Filtered propagator functional for iterative dynamics of quantum dissipative systems. Computer Physics Communications, 99(2–3):335–354, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001300.

Scott:1993:GMA

[SMB^{+}93] N. S. Scott, A. McMinn, P. G. Burke, V. M. Burke, and C. J. Noble. Graphical R-matrix atomic collision environment (G R ACE): the problem specification stage. Computer Physics Communications, 78(1–2):67–76, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390143Z.

Schnack:1990:MSC

[SMBV90] D. D. Schnack, Z. Mikić, D. C. Barnes, and G. Van Hoven. Magnetohydrodynamic simulation of coronal magnetic fields. Computer Physics Communications, 59(1):21–37, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090153R.

Socorro:1998:CAG

[SMH98] José Socorro, Alfredo Macías, and Friedrich W. Hehl. Computer algebra in gravity: Reduce–Excalc programs for (non- ) Riemannian space–times. I. Computer Physics Communications, 115(2–3):264–283, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001337.

Smith:1991:MDH

[Smi91] W. Smith. Molecular dynamics on hypercube parallel computers. Computer Physics Communications, 62(2–3):229–248, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591900975.

Smith:1992:EMB

[Smi92a] Richard W. Smith. Energy minimization in binary alloy models via genetic algorithms. Computer Physics Communications, 71(1–2):134–146, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290078D.

Smith:1992:RDM

[Smi92b] W. Smith. A replicated data molecular dynamics strategy for the parallel Ewald sum. Computer Physics Communications, 67(3):392–406, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465592900484.

Schreiber:1999:ESA

[SMR99] Michael Schreiber, Frank Milde, and Rudolf A. Römer. Electronic states in the Anderson model of localization: benchmarking eigenvalue algorithms. Computer Physics Communications, 121–122:517–523, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003963.

Schmid:1990:VMC

[SMV90] U. Schmid, N. C. Myers, and J. A. Van Vechten. VID-SIM — a Monte Carlo program for the simulation of atomic diffusion in diamond and zinc-blende structures. Computer Physics Communications, 58(3):329–341, April 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090067B.

Solomon:1999:BRS

[Sol99] Sorin Solomon. Behaviorly realistic simulations of stock market traders with a soul. Computer Physics Communications, 121–122:161–167, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003045.

Sordo:1998:QPA

[Sor98] J. A. Sordo. QUALITY: a program to assess basis set quality. Computer Physics Communications, 113(1):85–104, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598000630.

Sourlas:1999:URS

[Sou99] Nicolas Sourlas. Universality in random systems: the case of the 3D random field Ising model. Computer Physics Communications, 121–122:183–187, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003082.

Sanchez:1993:HCC

[SP93a] M. L. Sánchez and A. López Piñeiro. HYDMATEL: a code to calculate matrix elements for hydrogen-like atoms. Computer Physics Communications, 75(1–2):185–190, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390174B.

Seesselberg:1993:IAE

[SP93b] M. Seeßelberg and F. Petruccione. An improved algorithm for the estimation of the mean first passage time of ordinary stochastic differential equations. Computer Physics Communications, 74(2):247–255, February 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390095T.

Seesselberg:1993:NIS

[SP93c] Markus Seeßelberg and Francesco Petruccione. Numerical integration of stochastic partial differential equations. Computer Physics Communications, 74(3):303–315, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593900144.

Simonson:1995:DPP

[SP95a] Thomas Simonson and David Perahia. Dielectric properties of proteins from simulations: tools and techniques. Computer Physics Communications, 91(1–3):291–303, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500054J.

Smith:1995:EEE

[SP95b] Paul E. Smith and B. Montgomery Pettitt. Efficient Ewald electrostatic calculations for large systems. Computer Physics Communications, 91(1–3):339–344, September 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500058N.

Seppala:1999:ODR

[SPA99] E. T. Seppälä, V. Petäjä, and M. J. Alava. Order, disorder and roughening in the 2D random field Ising model. Computer Physics Communications, 121–122:733, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701369.

Santosuosso:1998:IRT

[SPP^{+}98] Kevin Santosuosso, Denis Pollney, Nicos Pelavas, Peter Musgrave, and Kayll Lake. Invariants of the Riemann tensor for class B warped product space–times. Computer Physics Communications, 115(2–3):381–394, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001349.

Smyth:1998:NIT

[SPT98] Edward S. Smyth, Jonathan S. Parker, and K. T. Taylor. Numerical integration of the time-dependent Schrödinger equation for laser-driven helium. Computer Physics Communications, 114(1–3):1–14, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000836.

Shi:1991:OCB

[SR91] Shenghua Shi and Herschel Rabitz. Optimal control of bond selectivity in unimolecular reactions. Computer Physics Communications, 63(1–3):71–83, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190239H.

Schultz:1998:EES

[SR98] D. R. Schultz and C. O. Reinhold. Elastic — elastic scattering of electrons from ions and atoms. Computer Physics Communications, 114(1–3):342–355, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000848.

Stotler:1994:AAS

[SRB94] D. P. Stotler, W. T. Reiersen, and G. Bateman. ASPECT: an advanced specified-profile evaluation code for tokamaks. Computer Physics Communications, 81(1–2):261–291, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901244.

Stampfer:1994:TFE

[SRF94] D. Stampfer, M. Regler, and R. Frühwirth. Track fitting with energy loss. Computer Physics Communications, 79(2):157– 164, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900647.

Sonzogni:1992:CDE

[SRFN92] A. A. Sonzogni, A. S. M. A. Romo, W. R. Frosch, and S. J. Nassiff. A code to determine the energy distribution, the incident energy and the flux of a beam of light ions into a stack of foils. Computer Physics Communications, 69(2–3):429–438, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290181W.

Speith:1995:PTF

[SRR95] R. Speith, H. Riffert, and H. Ruder. The photon transfer function for accretion disks around a Kerr black hole. Computer Physics Communications, 88(2–3):109–120, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500067P.

Sararu:1990:SFL

[SS90a] A. Sararu and M. Sararu. Stability of finite linear combinations of vectors under changes of their coefficients. an application to approximation problems. Computer Physics Communications, 60(1):47–52, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090078F.

Sherborne:1990:RGC

[SS90b] Brad S. Sherborne and G. E. Stedman. Recursive generation of Cartesian angular momentum coupling trees for SO(3). Computer Physics Communications, 59(2):417–428, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590901887.

Sebak:1991:GSE

[SS91a] A. Sebak and L. Shafai. Generalized solutions for electromagnetic scattering by elliptical structures. Computer Physics Communications, 68(1–3):315–330, November 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190206Z.

Spina:1991:PSH

[SS91b] Alejandro Spina and Rex. T. Skodje. The phase-space hydrodynamic model for the quantum standard map. Computer Physics Communications, 63(1–3):279–305, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190256K.

Shevchenko:1993:ANH

[SS93a] Ivan I. Shevchenko and Andrej G. Sokolsky. Algorithms for normalization of Hamiltonian systems by means of computer algebra. Computer Physics Communications, 77(1):11–18, September 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593900328.

Shoucri:1993:FPC

[SS93b] M. Shoucri and I. Shkarofsky. A Fokker–Planck code for the numerical solution of the plasma heating and the current drive problems with synergetic effects. Computer Physics Communications, 78(1–2):199–210, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901556.

Shoucri:1994:FFP

[SS94a] M. Shoucri and I. Shkarofsky. A fast 2D Fokker–Planck solver with synergetic effects. Computer Physics Communications, 82(2–3):287–305, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901740.

Stumpf:1994:SCE

[SS94b] Roland Stumpf and Matthias Scheffler. Simultaneous calculation of the equilibrium atomic structure and its electronic ground state using density-functional theory. Computer Physics Communications, 79(3):447–465, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901872.

Simons:1991:TLM

[SSBA91] N. R. S. Simons, A. A. Sebak, E. Bridges, and Y. M. M. Antar. Transmission-line matrix (TLM) method for scattering problems. Computer Physics Communications, 68(1–3):197–212, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591902005.

Shkarofsky:1992:NSF

[SSF92] I. P. Shkarofsky, M. M. Shoucri, and V. Fuchs. Numerical solution of the Fokker–Planck equation with a dc electric field. Computer Physics Communications, 71(3):269–284, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290014P.

Sinha:1994:PAP

[SSH94] Amitabh B. Sinha, Klaus Schulten, and Helmut Heller. Performance analysis of a parallel molecular dynamics program. Computer Physics Communications, 78(3):265–278, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900051.

Selberherr:1991:NST

[SSHT91] S. Selberherr, M. Stiftinger, O. Heinreichsberger, and K. P. Traar. On the numerical solution of the three-dimensional semiconductor device equations on vector-concurrent computers. Computer Physics Communications, 67(1):145–156, August 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190227C.

Shida:1997:NET

[SSK97] Koichiro Shida, Ryotaro Suzuki, and Toshio Kawai. Numerical error of total energy. dependence on timestep. Computer Physics Communications, 102(1–3):59–67, May 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000076.

Sanders:1992:SCC

[SSPD92] David E. Sanders, Mark S. Stave, Leslie S. Perkins, and Andrew E. DePristo. SCT89: a computer code for atomic and molecular scattering from clean and adsorbate covered surfaces. Computer Physics Communications, 70(3):579–608, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290118I.

Sagar:1991:GQO

[SSS91] Robin P. Sagar, Vedene H. Smith, Jr., and Alfredo M. Simas. A Gaussian quadrature for the optimal evaluation of integrals involving lorentzians over a semi-infinite interval. Computer Physics Communications, 62(1):16–24, January 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465591901174.

Solomon:1994:DAM

[SSW94] S. Solomon, D. Stanhill, and K. Wolowelsky. Dynamical algebraic multi-grid in simulations of free fields on random triangulated surfaces. Computer Physics Communications, 83 (1):23–29, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594900329.

Sullivan:1992:FDN

[ST92] E. C. Sullivan and A. Temkin. Further developments in the noniterative method of solving PDE’s in electron scattering. Computer Physics Communications, 71(3):319–342, September 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290017S.

Sa:1995:EGF

[ST95] Ben-Hao Sa and An Tai. An event generator for the firecracker model and the rescattering in high energy pA and AA collisions LUCIAE version 2.0. Computer Physics Communications, 90(1):121–140, September 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500066O.

Stibbe:1998:TPD

[ST98] Darian T. Stibbe and Jonathan Tennyson. TIMEDEL: a program for the detection and parameterization of resonances using the time-delay matrix. Computer Physics Communications, 114(1–3):236–242, November 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000708.

Salazar:1999:HSA

[ST99] Rafael Salazar and Raúl Toral. Hybrid simulated annealing using Tsallis statistics. Computer Physics Communications, 121–122:40–42, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002751.

Stapleton:1996:NSR

[Sta96] David P. Stapleton. Numerical simulation of a rigid rotating body by obrechkoff integration. Computer Physics Communications, 98(1–2):153–166, October 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000707.

Stangerup:1999:ISC

[Sta99] Paul Stangerup. Implementation of sensitivity calculations in a general-purpose simulation program. Computer Physics Communications, 117(1–2):130–141, March 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001696.

Stepanek:1997:PDR

[Ste97] Jiri Stepanek. A program to determine the radiation spectra due to a single atomic-subshell ionisation by a particle or due to deexcitation or decay of radionuclides. Computer Physics Communications, 106(3):237–257, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000684.

Stiegler:1994:SMC

[Sti94] U. Stiegler. SCOT, a Monte Carlo generator to study the Lorentz structure of e^{+}e^{−} → Z → τ^{+}τ^{−}. Computer Physics Communications, 81(1–2):221–236, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559490121X.

Sato:1999:NCS

[STI99] Yuzuru Sato, Makoto Taiji, and Takashi Ikegami. NPcompleteness of k SAT and multifractals. Computer Physics Communications, 121–122:51–53, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002787.

Scholer:1990:RFC

[STJ90] M. Scholer, T. Terasawa, and F. Jamitzky. Reconnection and fluctuations in compressible MHD: a comparison of different numerical methods. Computer Physics Communications, 59(1):175–184, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090167Y.

Stoll:1998:FCC

[Sto98] E. Stoll. A fast cluster counting algorithm for percolation on and off lattices. Computer Physics Communications, 109(1): 1–5, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001422.

Storer:1999:RMS

[Sto99] Robin Storer. Resistive magnetohydrodynamic spectra and tokamak plasma response. Computer Physics Communications, 121–122:740, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701436.

Schmitz:1991:TPC

[STR91a] Joachim Schmitz, Hans-Rainer Trebin, and Ulrich Rössler. TRS: a program to calculate Landau levels and direct dipole transitions in uniaxially stressed semiconductors. Computer Physics Communications, 66(2–3):308–318, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190079Z.

Straton:1991:CPG

[Str91b] Jack C. Straton. CMPT — a program giving the analytical result of Nordsieck’s integral including powers of the radius from −1 to +6. Computer Physics Communications, 66(1): 66–70, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591900089.

Schmitz:1992:TNV

[STR92] Joachim Schmitz, Hans-Rainer Trebin, and Ulrich Rössler. TRSS: a new version of program TRS for a different geometry. Computer Physics Communications, 69(2–3):369– 372, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559290175X.

Strilka:1993:PCN

[Str93a] Richard Strilka. A parallel code for the numerical simulation of flux tubes in superconductors and cosmic strings. Computer Physics Communications, 78(1–2):141–154, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390150B.

Strnadl:1993:SCP

[Str93b] C. F. Strnadl. Symbolic computation of the pair-distribution function for hard-sphere systems in the whole R-range. Computer Physics Communications, 75(1–2):47–54, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465593901637.

Stuart:1995:ARF

[Stu95a] Robin G. Stuart. Algebraic reduction of Feynman diagrams to scalar integrals: a Mathematica implementation of LERG-I. Computer Physics Communications, 85(2):267–277, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400141N. See erratum

[Stu95b].

Stuart:1995:EBR

[Stu95b] Robin G. Stuart. Erratum: Algebraic reduction of Feynman diagrams to scalar integrals: a Mathematica implementation of LERG-I [Comput. Phys. Commun. 85 (1995) 267–277]. Computer Physics Communications, 88(2–3):347, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500063L. See [Stu95a].

Styer:1990:SLP

[Sty90] Daniel F. Styer. Subroutine library for partial differential approximants. Computer Physics Communications, 61(3): 374–386, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465590900512.

Sugar:1991:NSQ

[Sug91] Robert L. Sugar. Numerical simulation of quantum chromodynamics. Computer Physics Communications, 65(1–3): 268–280, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190180S.

Sugiyama:1999:OVM

[Sug99] Yuki Sugiyama. Optimal velocity model for traffic flow. Computer Physics Communications, 121–122:399–401, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003665.

Sulsky:1996:EBP

[Sul96] Deborah Sulsky. Erratum: Application of a particle-in-cell method to solid mechanics [Comput. Phys. Commun. 87 (1995) 236–252]. Computer Physics Communications, 96 (1):105–106, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000616.

See [SZS95].

Sunko:1997:CSN

[Sun97] D. K. Sunko. Code SPINDIS: Nuclear shell-model level densities at high energies with a simplified pairing interaction.

Computer Physics Communications, 101(1–2):171–182, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559600135X.

Shimizu:1999:FPB

[SUYH99] H. Shimizu, K. Uehara, K. Yamamoto, and Y. Hiwatari. Folding property of di-block polyampholytes. Computer Physics Communications, 121–122:735, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701382.

Sauter:1994:GAP

[SV94] O. Sauter and J. Vaclavik. Gyrokinetic approach to the propagation of electromagnetic waves in nonuniform bounded plasma slabs. Computer Physics Communications, 84(1–3): 226–242, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465594902135.

Severijs:1992:DCA

[SVBD92] C. A. Severijs, G. Verbist, H. H. Brongersma, and J. T. Devreese. Differential correction algorithm for a function of two continuous variables: application to the collision integrals. Computer Physics Communications, 70(3):459–466, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290107A.

Sibani:1999:LME

[SvdPS99] Paolo Sibani, Ruud van der Pas, and J. Christian Schön. The lid method for exhaustive exploration of metastable states of complex systems. Computer Physics Communications, 116(1):17–27, January 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001763.

Searles:1992:FPP

[SvNF92] D. J. Searles and E. I. von Nagy-Felsobuki. A fitting program for potential energy surfaces of bent triatomic molecules. Computer Physics Communications, 67(3):527–536, January 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465592900576.

Stroganova:1999:EVD

[SVV99] S. V. Stroganova, M. I. Vasilevskiy, and O. V. Vikhrova. The effect of vibrational degrees of freedom on the phase transition in the 2D Ising model. Computer Physics Communications, 121–122:741, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701448.

Shang:1991:OAF

[SW91a] E. C. Shang and Y. Y. Wang. Ocean acoustic field simulations for monitoring large-scale ocean structures. Computer Physics Communications, 65(1–3):238–245, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190177M.

Srikameswaran:1991:ISP

[SW91b] K. Srikameswaran and R. Wallace. 3D isovalue surface plots. Computer Physics Communications, 62(1):80–89, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591901233.

Sanders:1998:CMF

[SW98] Jan A. Sanders and Jing Ping Wang. Combining Maple and Form to decide on integrability questions. Computer Physics Communications, 115(2–3):447–459, December 2, 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001222.

Solla:1999:OOL

[SW99] Sara A. Solla and Ole Winther. Optimal online learning: a Bayesian approach. Computer Physics Communications, 121–122:94–97, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900288X.

Swarte:1991:PSN

[Swa91] J. H. Swarte. Precision of solutions in nonlinear leastsquares model fitting. Computer Physics Communications, 66(2–3):143–149, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190062P.

Swendsen:1991:AMM

[Swe91] RobertH. Swendsen. Acceleration methods for Monte Carlo computer simulations. Computer Physics Communications, 65(1–3):281–288, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190181J.

Schwartz:1999:EPU

[SWG99] Ira B. Schwartz, Yvette K. Wood, and Ioannis T. Georgiou. Extreme parametric uncertainty and instant chaos in coupled structural dynamics. Computer Physics Communications, 121–122:425–428, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003744.

Singleton:1991:PMC

[SWT91] Gregory L. Singleton, Chwan-Hwa Wu, and Jyun-Hwei Tsai. A parallel Monte Carlo simulation for gaseous electronics on a dynamically reconfigurable multiprocessor system. Computer Physics Communications, 66(2–3):181–193, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190067U.

Shida:1995:REL

[SY95] Koichiro Shida and Shin’ichi Yamada. Reduced eventlist on an array for many-body simulation. Computer Physics Communications, 86(3):289–296, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400161T.

Stefanou:1998:HPC

[SYM98] N. Stefanou, V. Yannopapas, and A. Modinos. Heterostructures of photonic crystals: frequency bands and transmission coefficients. Computer Physics Communications, 113(1): 49–77, September 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000605.

Strandlie:1999:PTI

[SZ99] A. Strandlie and J. Zerubia. Particle tracking with iterated Kalman filters and smoothers: the PMHT algorithm. Computer Physics Communications, 123(1–3):77–86, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002581.

Smirnova:1999:CMD

[SZG99] Julia A. Smirnova, Leonid V. Zhigilei, and Barbara J. Garrison. A combined molecular dynamics and finite element method technique applied to laser induced pressure wave propagation. Computer Physics Communications, 118(1):11–16, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001751.

Szmytkowski:1995:RMC

[Szm95] Radosław Szmytkowski. The relativistic multi-channel variable phase method for solving asymptotic equations in electron-atom and electron-ion scattering. Computer Physics Communications, 90(2–3):244–250, October 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559500095W.

Szolnoki:1999:NSS

[Szo99] Attila Szolnoki. Novel stationary state in a two-temperature model with competing dynamics. Computer Physics Communications, 121–122:742, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550670145X.

Sulsky:1995:APC

[SZS95] Deborah Sulsky, Shi-Jian Zhou, and Howard L. Schreyer. Application of a particle-in-cell method to solid mechanics. Computer Physics Communications, 87(1–2):236–252, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594001707. See erratum

[Sul96].

Takada:1992:PAC

[Tak92] K. Takada. Programs for algebraic calculation of angular momentum coupling. Computer Physics Communications, 69(1):142–154, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290135L.

Teo:1990:SCC

[TAMSW90] K. H. Teo, W. Allegretto, J. N. McMullin, and H. G. Schmidt-Weinmar. Self-consistent calculation of the densityof-states mass of holes in 2-D silicon structures. Computer Physics Communications, 59(2):277–282, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465590901762.

Tanaka:1990:NCH

[Tan90] Hidekazu Tanaka. Numerical calculation of helicity amplitudes for processes involving massive fermions. Computer Physics Communications, 58(1–2):153–168, February/March 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090142N.

Tanaka:1995:MEP

[Tan95] Motohiko Tanaka. The macro-EM particle simulation method and a study of collisionless magnetic reconnection. Computer Physics Communications, 87(1–2):117–138, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594001682.

Tarutin:1999:PDC

[Tar99] O. B. Tarutin. Power of DSP and computers changes the concept of acoustic measurement in solids under high pressure. Computer Physics Communications, 121–122:744, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701473.

Taubmann:1992:PCF

[Tau92] G. Taubmann. Parabolic cylinder functions U (n, x ) for natural n and positive x. Computer Physics Communications, 69(2–3):415–419, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901793.

Taubmann:1995:ICL

[Tau95] Gerhard Taubmann. Improvement of the convergence of the linear variation method due to a variation of the basis. Computer Physics Communications, 90(2–3):235–243, October 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500101K.

Taylor:1999:BRC

[Tay99] Peter R. Taylor. Book review: Y. C. Pao, Engineering analysis: Interactive methods and programs with FOR-TRAN, QuickBasic, MATLAB, and Mathematica (1999) CRC Press, Bristol 0-8493-2016-X. Computer Physics Communications, 120(2–3):271–272, August 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002374.

Thirumalai:1991:MST

[TB91] D. Thirumalai and B. J. Berne. Methods for simulating time correlation functions in quantum systems. Computer Physics Communications, 63(1–3):415–426, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190266N.

Toukmaji:1996:EST

[TB96] Abdulnour Y. Toukmaji and John A. Board, Jr. Ewald summation techniques in perspective: a survey. Computer Physics Communications, 95(2–3):73–92, June 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000161.

Taylor:1998:SCL

[TBA^{+}98] M. B. Taylor, G. D. Barrera, N. L. Allan, T. H. K. Barron, and W. C. Mackrodt. Shell: a code for lattice dynamics and structure optimisation of ionic crystals. Computer Physics Communications, 109(2–3):135–143, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000186.

Toomre:1990:TDC

[TBCH90] Juri Toomre, Nicholas Brummell, Fausto Cattaneo, and Neal E. Hurlburt. Three-dimensional compressible convection at low prandtl numbers. Computer Physics Communications, 59(1):105–117, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465590901603.

Tuzun:1998:ACI

[TBS98] Robert E. Tuzun, Paul Burkhardt, and Don Secrest. Accurate computation of individual and tables of 3-j and 6-j symbols. Computer Physics Communications, 112(2–3):112–148, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000654.

Toral:1993:GGD

[TC93] Raúl Toral and Amitabha Chakrabarti. Generation of Gaussian distributed random numbers by using a numerical inversion method. Computer Physics Communications, 74(3):327– 334, March 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593900166.

Teboul:1997:CEM

[TC97] Victor Teboul and Stéphane Chaussedent. Cutoff effect in molecular dynamics simulations of interaction induced light scattering spectra. Computer Physics Communications, 105(2–3):151–158, October 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000799.

Taraskin:1999:LFV

[TCE99a] S. N. Taraskin, G. Cormier, and S. R. Elliott. Lowfrequency vibrational modes in disordered structures. Computer Physics Communications, 121–122:743, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506701461.

Taraskin:1999:PVE

[TCE99b] S. N. Taraskin, G. Cormier, and S. R. Elliott. Propagation of vibrational excitations in disordered atomic structures by a normal-mode analysis. Computer Physics Communications, 121–122:540–541, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004014.

Todd:1999:NAU

[TD99] B. D. Todd and Peter J. Daivis. A new algorithm for unrestricted duration nonequilibrium molecular dynamics simulations of planar elongational flow. Computer Physics Communications, 117(3):191–199, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001787.

Trunec:1996:SDS

[TDV96] D. Trunec, J. Drímal, and M. Vicar. A subprogram for direct solution of Poisson equation in cylindrically symmetric geometry. Computer Physics Communications, 98(3):339–345, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000999.

Thiers:1993:OOA

[TEJJ93] A. H. M. Thiers, M. J. Ephraïm, T. Janssen, and A. Janner. An object oriented approach towards a crystal symmetry environment. Computer Physics Communications, 77(2):167–189, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390002T.

Tennyson:1997:PFE

[Ten97] Jonathan Tennyson. Phase factors in electron-molecule collision calculations. Computer Physics Communications, 100 (1–2):26–30, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001440.

Thorleifsson:1998:IAS

[TF98] Gudmar Thorleifsson and Marco Falcioni. Improved algorithms for simulating crystalline membranes. Computer Physics Communications, 109(2–3):161–170, April 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000149.

Tsuda:1995:NMM

[TFY95] N. Tsuda, A. Fujitsu, and T. Yukawa. Note on the Metropolis Monte Carlo method on random lattices. Computer Physics Communications, 87(3):372–374, June 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400129P.

Tian:1996:BTW

[TG96] Bo Tian and Yi-Tian Gao. Beyond travelling waves: a new algorithm for solving nonlinear evolution equations. Computer Physics Communications, 95(2–3):139–142, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000148.

Tawfik:1999:MLT

[TG99] A. M. Tawfik and E. Ganssauge. MIRACLE Lab: Track recognition and event track reconstruction in nuclear emulsion chambers. Computer Physics Communications, 118(1):49–60, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001969.

Tran:1996:DPS

[TGAW96] T. M. Tran, R. Gruber, K. Appert, and S. Wüthrich. A direct parallel sparse matrix solver. Computer Physics Communications, 96(2–3):118–128, August 1, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000070.

Todesco:1997:NCN

[TGG97] E. Todesco, M. Gemmi, and M. Giovannozzi. NERO: a code for the nonlinear evaluation of resonances in one-turn mappings. Computer Physics Communications, 106(1–2):169–180, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559700091X.

Tafelmayer:1995:SFC

[TH95] R. Tafelmayer and K. H. Hoffmann. Scaling features in complex optimization problems. Computer Physics Communications, 86(1–2):81–90, April 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500004Y.

Theuns:1994:PPE

[The94] T. Theuns. Parallel P3M with exact calculation of short range forces. Computer Physics Communications, 78(3):238–246, January 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900027.

Tennyson:1995:DFP

[THF95] Jonathan Tennyson, James R. Henderson, and Nicholas G. Fulton. DVR3D: for the fully pointwise calculation of ro-vibrational spectra of triatomic molecules. Computer Physics Communications, 86(1–2):175–198, April 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400139S.

Truong:1993:MPD

[ThLL^{+}93] Thanh N. Truong, Da hong Lu, Gillian C. Lynch, Yi-Ping Liu, Vasilios S. Melissas, James J. P. Stewart, Rozeanne Steckler, Bruce C. Garrett, Alan D. Isaacson, Angels Gonzalez-Lafont, Sachchida N. Rai, Gene C. Hancock, Tomi Joseph, and Donald G. Truhlar. MORATE: a program for direct dynamics calculations of chemical reaction rates by semiempirical molecular orbital theory. Computer Physics Communications, 75(1–2):143–159, April 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465593901729.

Tiszauer:1991:CME

[TK91] D. H. Tiszauer and K. C. Kulander. Computation methods employed in the self-consistent time-dependent, Hartree– Fock calculation for a reactive molecular collision. Computer Physics Communications, 63(1–3):351–364, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190262J.

Tanaka:1991:NCF

[TKS91] Hidekazu Tanaka, Toshiaki Kaneko, and Yoshimitsu Shimizu. Numerical calculation of Feynman amplitudes for electroweak theories and an application to e^{+}e^{−} → W^{+}W^{−}γ. Computer Physics Communications, 64(1):149–166, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190058S.

Tsou:1994:MCS

[TLK94] R. H. Tsou, Simon C. Lin, and L. L. Kiang. Monte Carlo simulation for Compton suppression spectrometer. Computer Physics Communications, 83(1):30–44, October 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900337.

Torres:1999:ESF

[TMG99] J. J. Torres, J. Marro, and P. L. Garrido. On the effect of synaptic fluctuations during retrieval processes in neural network models. Computer Physics Communications, 121–122:98–102, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002891.

Tennyson:1993:TPC

[TML93] Jonathan Tennyson, Steven Miller, and C. Ruth Le Sueur. Triatom: programs for the calculation of ro-vibrational spectra of triatomic molecules. Computer Physics Communications, 75(3):339–364, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390048H.

Tabor-Morris:1994:SUC

[TMR94] Anne E. Tabor-Morris and Bernhard Rupp. SEXIE 3.0 — an updated computer program for the calculation of coordination shells and geometries. Computer Physics Communications, 82(1):23–29, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901287.

Trebin:1999:MDS

[TMS^{+}99] Hans-Rainer Trebin, Ralph Mikulla, Jörg Stadler, Gunther Schaaf, and Peter Gumbsch. Molecular dynamics simulations of crack propagation in quasicrystals. Computer Physics Communications, 121–122:536–539, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004002.

Tobochnik:1999:TSW

[Tob99] Jan Tobochnik. Teaching students to write computer simulations in Java. Computer Physics Communications, 121–122:562–568, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004075.

Takagishi:1995:RIC

[TON95] Kunio Takagishi, Mitsushi Ohkura, and Shinobu Nakazaki. Radial integrals in the Coulomb–Born approximation. Computer Physics Communications, 85(2):293–305, February 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400126M.

Topper:1996:BRB

[Top96] Robert Q. Topper. Book review: Quantum mechanics simulation: By John R. Hiller, Ian D. Johnston and Daniel F. Styer, The consortium for upper-level physics software. John Wiley & Sons, Chichester, 1995. 221 pages and 1 MS-DOS formatted diskette. UK£24.95 (paperback). ISBN 0-471-54884-7. Computer Physics Communications, 94(2–3):272–273, April 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596900235.

Torrens:1991:PPC

[TOSM91] Francisco Torrens, Enrique Ortí, and José Sánchez-Marín. Pair potential calculation of molecular associations: a vectorized version. Computer Physics Communications, 66(2–3): 341–362, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190081U.

Toxvaerd:1999:CRP

[Tox99] S. Toxvaerd. Chemical reactions and phase separations in condensed fluids simulated by molecular dynamics. Computer Physics Communications, 121–122:251–255, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003239.

Tokarev:1999:SNS

[TP99] M. V. Tokarev and E. V. Potrebenikova. Study of new scaling of direct photon production in pp collisions at high energies using MC simulation. Computer Physics Communications, 117(3):229–238, March 11, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001404.

Theuns:1993:CSR

[TR93] T. Theuns and M. E. Rathsack. Calculating short range forces on a massively parallel computer: SPH on the Connection Machine. Computer Physics Communications, 76(2):141–158, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390128Y.

Torrens:1998:ANV

[TRSM98] Francisco Torrens, Mercedes Rubio, and José Sánchez-Marín. AMYR 2: a new version of a computer program for pair potential calculation of molecular associations. Computer Physics Communications, 115(1):87–89, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000988.

Truhlar:1993:BRBa

[Tru93a] Donald G. Truhlar. Book review: Directory of chemistry software 1992: Edited by Wendy Warr, Pater Willett, and Geoff Downs, Cherwell Scientific Publishing and the American Chemical Society, Oxford, UK, 1992. 204 + iii pages. ISBN 0-9518236-0-4. Computer Physics Communications, 74(1):149, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390112P.

Truhlar:1993:BRBb

[Tru93b] Donald G. Truhlar. Book review: The Monte Carlo Method in Condensed Matter Physics: K. Binder, editor, Topics in Applied Physics Series, Volume 71, Springer-Verlag, Berlin, 1992. xvi + 392 pages. 83 figures. Hardcover price DM 99.00. ISBN 3-540-54369-4. Computer Physics Communications, 74(1):150, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390113Q.

Truhlar:1994:LST

[Tru94] Donald G. Truhlar. Long-standing themes in computational chemical dynamics. Computer Physics Communications, 84(1–3):78–90, November 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594902046.

Truhlar:1995:F

[Tru95] Donald G. Truhlar. Foreword. Computer Physics Communications, 91(1–3):xi, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465595900098.

Tuzun:1992:NEK

[TS92] Robert E. Tuzun and Don Secrest. Numerical evaluation of kratzer oscillator matrix elements. Computer Physics Communications, 70(2):362–370, June 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901999.

Teleman:1991:ESM

[TSJ91] Olle Teleman, Bo Svensson, and Bo Jönsson. Efficiency in statistical mechanical simulations of biomolecules — computer programs for molecular and continuum modelling. Computer Physics Communications, 62(2–3):307–326, March 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190103R.

Thomas:1997:NTC

[TSK^{+}97] E. Thomas, Jr., G. E. Sasser, S. F. Knowlton, J. D. Hanson, and R. F. Gandy. A new technique for calculating toroidal harmonic coefficients for curl-free magnetic fields in helical toroidal devices. Computer Physics Communications, 100(1–2):31–40, February 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001397.

Tsivline:1999:BTF

[TSL^{+}99] D. Tsivline, V. S. Stepanyuk, N. Levanov, W. Hergert, and A. A. Katsnelson. Bonding trends in free and supported metal clusters. Computer Physics Communications, 121–122:747, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506701503.

Tsang:1999:CIA

[TT99] I. J. Tsang and I. R. Tsang. Cluster identification algorithm for lattice animals. Computer Physics Communications, 121–122:746, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701497.

Turtia:1999:MOW

[Tur99] S. B. Turtia. Multicriteria optimization: the way to decrease the sensitivity of the solution from the input parameters. Computer Physics Communications, 121–122:748, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701515.

Torda:1991:RNS

[TvG91] Andrew E. Torda and Wilfred F. van Gunsteren. The refinement of NMR structures by molecular dynamics simulation. Computer Physics Communications, 62(2–3):289–296, March 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190101P.

Torres:1999:SCS

[TVH^{+}99] J. J. Torres, P. Varona, R. Huerta, H. D. I. Abarbanel, and M. I. Rabinovich. Spatial clusters of synchronization in neural networks of chaotic spiking-bursting neurons. Computer Physics Communications, 121–122:745, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701485.

Ting:1992:VWP

[TYJ92] Julian J.-L. Ting, J. M. Yuan, and T.-F. Jiang. Vectorizable wave propagation FORTRAN code for calculations of multiphoton dissociation. Computer Physics Communications, 70(2):417–425, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290204C.

Towler:1996:DFT

[TZC96] Michael D. Towler, Ales Zupan, and Mauro Causà. Density functional theory in periodic systems using local Gaussian basis sets. Computer Physics Communications, 98(1–2):181–205, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465596000781.

Utsumi:1997:SAC

[UKA97] Takayuki Utsumi, Tomoaki Kunugi, and Takayuki Aoki. Stability and accuracy of the Cubic Interpolated Propagation scheme. Computer Physics Communications, 101(1–2):9–20, April 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001531.

Ukawa:1998:CPP

[Uka98] Akira Ukawa. The CP-PACS parallel computer. Computer Physics Communications, 110(1–3):220–224, May 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597001811.

Utsumi:1998:NMS

[UKK98] Takayuki Utsumi, Tomoaki Kunugi, and James Koga. A numerical method for solving the one-dimensional Vlasov– Poisson equation in phase space. Computer Physics Communications, 108(2–3):159–179, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597001197.

Ueda:1994:SNH

[UOMO94] Hiroko Ueda, Yoshiharu Omura, Hiroshi Matsumoto, and Takashi Okuzawa. A study of the numerical heating in electrostatic particle simulations. Computer Physics Communications, 79(2):249–259, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559490071X.

Unonius:1990:USV

[UP90] L. Unonius and P. Paatero. Use of singular value decomposition for analyzing repetitive measurements. Computer Physics Communications, 59(2):225–243, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090172W.

Umar:1991:NMN

[US91] A. S. Umar and M. R. Strayer. Numerical methods for nuclear mean-field dynamics. Computer Physics Communications, 63(1–3):179–202, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190248J.

Uhrlandt:1999:MSN

[USW99] D. Uhrlandt, M. Schmidt, and R. Winkler. A method to solve the nonlinear kinetic equation of the “nonlocal approach” including Coulomb interaction of electrons. Computer Physics Communications, 118(2–3):185–199, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001982.

Ueda:1994:NSI

[UTI94] Tamotsu Ueda, Kouji Tominaga, and Yasushi Ikegami. New spline interpolation for physics and some applications including nucleon-nucleon scattering phase-shifts. Computer Physics Communications, 81(1–2):163–172, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559490118X.

VanScoy:1996:DSP

[Van96] Frances L. Van Scoy. Developing software for parallel computing systems. Computer Physics Communications, 97(1– 2):36–44, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465596000197.

Vargas:1993:NSS

[Var93] William E. Vargas. A numerical scheme to solve the Korteweg–de Vries equation. Computer Physics Communications, 74(1):58–62, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390105L.

Varga:1997:CMF

[Var97] Kálmán Varga. A combined Mathematica–Fortran program package for analytical calculation of the matrix elements of the microscopic cluster model. Computer Physics Communications, 104(1–3):259–274, August 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000313.

Vos:1991:CEN

[VB91] J. B. Vos and C. M. Bergman. Chemical equilibrium and nonequilibrium inviscid flow simulations using an explicit scheme. Computer Physics Communications, 65(1–3):289–298, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190182K.

Villasenor:1992:RCC

[VB92a] John Villasenor and Oscar Buneman. Rigorous charge conservation for local electromagnetic field solvers. Computer Physics Communications, 69(2–3):306–316, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290169Y.

Vu:1992:CIF

[VB92b] H. X. Vu and J. U. Brackbill. CELEST1D: an implicit, fully kinetic model for low-frequency, electromagnetic plasma simulation. Computer Physics Communications, 69(2–3):253–276, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290165U.

Vamosi:1996:TPM

[VB96] J. Vámosi and S. Biri. TrapCAD — a program to model magnetic traps of charged particles. Computer Physics Communications, 98(1–2):215–223, October 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000537.

Velgakis:1997:IA

[VBO97] M. J. Velgakis, G. A. Baker, Jr., and J. Oitmaa. Integral approximants. Computer Physics Communications, 99 (2–3):307–322, January 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465596001117.

Vasilevskiy:1996:VPI

[VBS96] M. I. Vasilevskiy, O. V. Baranova, and S. V. Stroganova. Vibrational properties of a 2D Ising alloy. Computer Physics Communications, 97(1–2):199–204, August 2, 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000380.

Voinov:1994:IEC

[VD94] V. G. Voinov and V. E. Dorof̂eyev. Integral equation of convolution type: a new algorithm for analytical deconvolver computing. Computer Physics Communications, 81(3):351– 356, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900833.

vandenBos:1993:CEA

[vdB93] A. van den Bos. Critical errors associated with parameter resolvability. Computer Physics Communications, 76(2):184– 190, July 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593901305.

Verbruggen:1992:ASE

[VdN92] M. H. W. Verbruggen and J. M. M. de Nijs. Analysis of spectroscopic ellipsometric measurements. Computer Physics Communications, 69(1):201–214, February 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290140T.

VandenBerghe:1999:EFE

[VDVV99] G. Vanden Berghe, H. De Meyer, M. Van Daele, and T. Van Hecke. Exponentially-fitted explicit Runge–Kutta methods. Computer Physics Communications, 123(1–3):7–15, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599003653.

Vermaseren:1994:A

[Ver94] J. A. M. Vermaseren. Axodraw. Computer Physics Communications, 83(1):45–58, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900345.

Verges:1999:CIK

[Ver99] J. A. Vergés. Computational implementation of the Kubo formula for the static conductance: Application to two-dimensional quantum dots. Computer Physics Communications, 118(1):71–80, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002064.

Veseli:1998:MIH

[Ves98] Sinisa Veseli. Multidimensional integration in a heterogeneous network environment. Computer Physics Communications, 108(1):9–19, January 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465597001203.

Vojta:1999:HFB

[VES99] Thomas Vojta, Frank Epperlein, and Michael Schreiber. Hartree–Fock based diagonalization: An efficient method for simulating disordered interacting electrons. Computer Physics Communications, 121–122:489–492, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003884.

Veit:1994:OCE

[VFO94] Th. Veit, J. Friedrich, and E. A. J. M. Offermann. Optimization of complex experiments with respect to a maximum gain of information. Computer Physics Communications, 82(2–3):129–138, September 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901627.

Vargas:1999:AMD

[VG99] C. A. Vargas and V. Márquez García. Analysis of models of deposition on surfaces. Computer Physics Communications, 121–122:749, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701527.

vanGunsteren:1995:CSP

[vGHM^{+}95] W. F. van Gunsteren, P. H. Hünenberger, A. E. Mark, P. E. Smith, and I. G. Tironi. Computer simulation of protein motion. Computer Physics Communications, 91(1–3):305–319, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500055K.

VanDerLinden:1990:CFE

[VGK90] R. A. M. Van Der Linden, M. Goossens, and W. Kerner. A combined finite element/Fourier series method for the numerical study of the stability of line-tied magnetic plasmas. Computer Physics Communications, 59(1):61–73, May 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090156U.

Velli:1990:SWE

[VGM90] Marco Velli, Roland Grappin, and André Mangeney. Solar wind expansion effects on the evolution of hydromagnetic turbulence in the interplanetary medium. Computer Physics Communications, 59(1):153–162, May 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090165W.

vanGisbergen:1999:ITD

[vGSB99] S. J. A. van Gisbergen, J. G. Snijders, and E. J. Baerends. Implementation of time-dependent density functional response equations. Computer Physics Communications, 118(2–3):119– 138, May 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599001873.

vanHameren:1999:CAA

[vHK99] André van Hameren and Ronald Kleiss. Computer-aided analysis of Riemann sheet structures. Computer Physics Communications, 116(2–3):311–318, February 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000794.

vanHameren:1997:GLD

[vHKH97] André van Hameren, Ronald Kleiss, and Jiri Hoogland. Gaussian limits for discrepancies I. Asymptotic results. Computer Physics Communications, 107(1–3):1–20, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001057.

Vattulainen:1995:CSS

[VKSAN95] I. Vattulainen, K. Kankaala, J. Saarinen, and T. Ala-Nissila. A comparative study of some pseudorandom number generators. Computer Physics Communications, 86(3):209–226, May 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595000158.

Verboncoeur:1995:OOE

[VLG95] J. P. Verboncoeur, A. B. Langdon, and N. T. Gladd. An object-oriented electromagnetic PIC code. Computer Physics Communications, 87(1–2):199–211, May 2, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400173Y.

vanMilligen:1991:FPF

[vMC91] B. Ph. van Milligen and N. J. Lopes Cardozo. Function parametrization: a fast inverse mapping method. Computer Physics Communications, 66(2–3):243–258, September/ October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190073T.

vanMilligen:1994:EVM

[vMF94] B. Ph. van Milligen and A. Lopez Fraguas. Expansion of vacuum magnetic fields in toroidal harmonics. Computer Physics Communications, 81(1–2):74–90, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901120.

Verwoerd:1998:ADM

[VN98] W. S. Verwoerd and V. Nolting. Angle decomposition of matrices. Computer Physics Communications, 108(2–3):218–239, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001331.

vanOldenborgh:1991:FPE

[vO91] G. J. van Oldenborgh. FF — a package to evaluate oneloop Feynman diagrams. Computer Physics Communications, 66(1):1–15, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591900023.

vanOldenborgh:1994:WGF

[vOFB94] Geert Jan van Oldenborgh, Paula J. Franzini, and Arianna Borrelli. WWF: a generator for e^{+}e^{−} → 4 fermions +γ. Computer Physics Communications, 83(1):14–22, October 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900310.

Valdes-Perez:1996:JMS

[VP96] Raúl E. Valdés-Pérez. On the justification of multiple selection rules of conservation in particle physics phenomenology. Computer Physics Communications, 94(1):25–30, March 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001425.

Valdes-Perez:1994:SIP

[VPE94] Raúl E. Valdés-Pérez and Michael Erdmann. Systematic induction and parsimony of phenomenological conservation laws. Computer Physics Communications, 83(2–3):171–180, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900469.

Vrahatis:1995:RPP

[VRS^{+}95] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions. Computer Physics Communications, 92(2–3):252–266, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465595001159. See erratum [VRS^{+}99].

Vrahatis:1999:EBP

[VRS^{+}99] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. Erratum to: RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions [Comput. Phys. Commun. 92 (1995) 252–266]. Computer Physics Communications, 117(3):290, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559800109X. See [VRS^{+}95].

Vahedi:1995:MCC

[VS95] V. Vahedi and M. Surendra. A Monte Carlo collision model for the particle-in-cell method: applications to argon and oxygen discharges. Computer Physics Communications, 87(1–2):179– 198, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559400171W.

Varga:1997:SFB

[VS97] K. Varga and Y. Suzuki. Solution of few-body problems with the stochastic variational method I. central forces with zero orbital momentum. Computer Physics Communications, 106(1–2):157–168, October 2, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000593.

Vojta:1999:DSR

[VS99] Thomas Vojta and Michael Schreiber. Damage spreading in random field systems. Computer Physics Communications, 121–122:750, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506701539.

Vackar:1991:IFF

[VSS91] J. Vackár, A. Simunek, and O. Sipr. Integration of free-free radiative-transition matrix elements. Computer Physics Communications, 66(2–3):259–265, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190074U.

Vladuca:1994:PPS

[VST94] G. Vladuca, M. Sin, and A. Tudora. PROBFIS — a program for subbarrier prompt and isomeric fission probabilities calculations for even-even nuclei. Computer Physics Communications, 83(2–3):266–274, December 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559490054X.

Vakhidov:1999:CAM

[VV99] Akmal A. Vakhidov and Nickolay N. Vasiliev. Computer algebra methods for developing an artificial satellite motion theory within elliptic functions. Computer Physics Communications, 118(1):17–20, April 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001775.

Visscher:1994:RQC

[VVA^{+}94] L. Visscher, O. Visser, P. J. C. Aerts, H. Merenga, and W. C. Nieuwpoort. Relativistic quantum chemistry: the MOLFDIR program package. Computer Physics Communications, 81(1–2):120–144, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901155.

vanVlimmeren:1996:CND

[vVF96] B. A. C. van Vlimmeren and J. G. E. M. Fraaije. Calculation of noise distribution in mesoscopic dynamics models for phase separation of multicomponent complex fluids. Computer Physics Communications, 99(1):21–28, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465596001257.

Vorontsov-Velyaminov:1999:BFP

[VVGI99] P. N. Vorontsov-Velyaminov, R. I. Gorbunov, and S. D. Ivanov. Bead–Fourier path integral Monte Carlo method applied to systems of identical particles. Computer Physics Communications, 121–122:64–66, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002829.

Vorontsov-Velyaminov:1999:EEM

[VVIBA99] P. N. Vorontsov-Velyaminov, S. D. Ivanov, A. V. Broukhno, and A. V. Akinshina. Expanded ensemble Monte Carlo of entropy and other properties for free and restricted lattice polymers. Computer Physics Communications, 121–122: 751, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506701540.

VanBelle:1995:ELF

[VW95] D. Van Belle and S. J. Wodak. Extended Lagrangian formalism applied to temperature control and electronic polarization effects in molecular dynamics simulations. Computer Physics Communications, 91(1–3):253–262, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500051G.

Weigel:1994:BSE

[WA94] H. Weigel and R. Alkofer. The Bethe–Salpeter equation for mesons as quark — anti-quark bound states in a soliton background. Computer Physics Communications, 82(1):57–73, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901317.

Widmann:1997:CES

[WA97] Albert H. Widmann and David B. Adolf. A comparison of Ewald summation techniques for planar surfaces. Computer Physics Communications, 107(1–3):167–186, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000908.

Wagner:1992:MCP

[Wag92] Robert G. Wagner. A Monte Carlo program for generating QED radiative decay W and Z events. Computer Physics Communications, 70(1):15–31, May 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559290087F.

Wang:1993:ACF

[Wan93] Jian-Xiong Wang. Automatic calculation of Feynman loopdiagrams: I. generation of a simplified form of the amplitude. Computer Physics Communications, 77(2):263–285, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390010A.

Wang:1999:TMM

[Wan99a] Jian-Sheng Wang. Transition matrix Monte Carlo method. Computer Physics Communications, 121–122:22–25, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002702.

Wang:1999:QMC

[Wan99b] Xiaoqun Wang. Quasi-Monte Carlo integration of characteristic functions and the rejection sampling method. Computer Physics Communications, 123(1–3):16–26, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002532.

Wolle:1999:RMC

[WBB99] B. Wolle, G. Beikert, and T. Baloui. Refined Monte Carlo modelling of fusion neutron emission in magnetically confined plasmas. Computer Physics Communications, 123(1– 3):46–55, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002623.

Wenzien:1995:GFC

[WBS95] Bernd Wenzien, Jörg Bormet, and Matthias Scheffler. Green function for crystal surfaces I. Computer Physics Communications, 88(2–3):230–248, August 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559400127N.

Weniger:1990:RAM

[WC90] Ernst Joachim Weniger and Jiři Cížek. Rational approximations for the modified Bessel function of the second kind. Computer Physics Communications, 59(3):471–493, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090089J.

Wu:1992:AIA

[WCJ92] S. Y. Wu, J. A. Cocks, and C. S. Jayanthi. An accelerated inversion algorithm using the resolvent matrix method. Computer Physics Communications, 71(1–2):125–133, August 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559290077C.

Whitlock:1999:QMC

[WCK99] P. A. Whitlock, G. V. Chester, and B. Krishnamachari. Quantum Monte Carlo simulation of the second layer helium film on graphite. Computer Physics Communications, 121–122:460–465, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003823.

Wang:1990:VAE

[WCL90] S. C. Wang, C. K. Chan, and S. P. Li. A vectorized algorithm on the ETA-10Q for MD simulation of particles in a box interacting by long-ranged forces. Computer Physics Communications, 60(2):181–186, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090002I.

Wiest:1991:PSI

[WDB^{+}91] Roland Wiest, Jean Demuynck, Marc Bénard, Marie-Madeleine Rohmer, and René Ernenwein. A program system for ab initio MO calculations on vector and parallel processing machines: III. integral reordering and four-index transformation. Computer Physics Communications, 62(1):107–124, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465591901255.

Weese:1992:RFM

[Wee92] Jürgen Weese. A reliable and fast method for the solution of fredhol integral equations of the first kind based on Tikhonov regularization. Computer Physics Communications, 69(1):99–111, February 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290132I.

Weese:1993:RMN

[Wee93] Jürgen Weese. A regularization method for nonlinear ill-posed problems. Computer Physics Communications, 77(3):429–440, November 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390187H.

Wegmann:1991:BRB

[Weg91] Rudolf Wegmann. Book review: Numerische Strömungsmechanik, inkompressible Strömungen mit komplexen Berandungen: B. E. Schönung, Springer-Verlag, Berlin, 1990. 308 + xiii pages, 101 figures. Broschiert DM64.00. ISBN 3-540-53137-8. Computer Physics Communications, 67(2):360–361, December 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190030O.

Weinketz:1993:TDM

[Wei93] Sieghard Weinketz. A two-dimensional monotonical logical grid algorithm. Computer Physics Communications, 74(2): 228–232, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559390092Q.

Wei:1999:UAE

[Wei99] Liqiang Wei. Unified approach for exact calculation of angular momentum coupling and recoupling coefficients. Computer Physics Communications, 120(2–3):222–230, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002325. See erratum [Wei11].

Wei:2011:DEC

[Wei11] Liqiang Wei. Direct and exact compute and table of entire 3j, 6j, and 9j symbols: Erratum to CPC 120 (1999) 222– 230. Computer Physics Communications, 182(5):1199, May 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465511000282. See [Wei99].

Weniger:1991:DIS

[Wen91] Ernst Joachim Weniger. On the derivation of iterated sequence transformations for the acceleration of convergence and the summation of divergent series. Computer Physics Communications, 64(1):19–45, April 1, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190047O.

Wako:1995:FPS

[WENG95] Hiroshi Wako, Shigeru Endo, Kuniaki Nagayama, and Nobuhiro Go. FEDER/2: program for static and dynamic conformational energy analysis of macro-molecules in dihedral angle space. Computer Physics Communications, 91(1–3):233–251, September 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559500050P.

West:1993:FTP

[Wes93] Todd H. West. FeynmanParameter and trace — programs for expressing Feynman amplitudes as integrals over Feynman parameters. Computer Physics Communications, 77(2):286–298, October 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390011Z.

Wu:1990:TAT

[WF90] Ernest Yue Wu and Robert J. Friauf. Techniques for achieving thermal equilibrium in molecular dynamics calculations for solids. Computer Physics Communications, 59(2):259– 266, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090174Y.

Wu:1993:ESF

[WF93] Ruqian Wu and A. J. Freeman. An efficient step-forward way to solve the Schrödinger eigenvalue equation in self-consistent calculations. Computer Physics Communications, 76(1):58– 62, June 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465593901202.

Wolf:1998:DSC

[WGH^{+}98] A. Wolf, C. Gwon, K. Honscheid, J. Lorenc, R. Wanke, E. Lipeles, A. Shapiro, A. Weinstein, F. Würthwein, A. Bean, D. Coppage, C. Darling, B. Forrest, T. Noor, C. Strohman, V. Fadejev, J. Staeck, and I. Volobouev. The DAQ system for CLEO III. Computer Physics Communications, 110(1–3):91– 94, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001598.

Whaley:1994:TIT

[Wha94] K. Birgitta Whaley. Time-independent theory of quantum scattering from disordered surfaces. Computer Physics Communications, 80(1–3):17–31, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900965.

Whittle:1993:BRB

[Whi93] John Whittle. Book review: Advances in the Free-Lagrange Method — Proceedings, Jackson Lake Lodge, WY, USA, 3-7 June 1990: H. E. Trease, M. J. Fritts and W. P. Crowley, eds., Lecture Notes in Physics, Vol. 395, Springer-Verlag, Berlin, 1991. 327 + xi pages. Hardcover price DM 69.00. ISBN 3-540-54960-9. Computer Physics Communications, 74(1):151, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559390114R.

White:1998:MIS

[Whi98] Vicky White. MISCOMP — an information system tool-kit. Computer Physics Communications, 110(1–3):177–180, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001744.

Wieder:1995:SPE

[Wie95] T. Wieder. SBGBBG, a program to evaluate the macroscopic strain/stress tensor of a polycrystalline sample from X-ray reflection positions. Computer Physics Communications, 85(3):398–414, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559400128O.

Wieder:1996:WCP

[Wie96] Thomas Wieder. WVM: a computer program for the determination of lattice parameters and strains in thin films. Computer Physics Communications, 96(1):53–60, July 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000550.

Wilson:1990:UBS

[Wil90] Stephen Wilson. Universal basis sets and Cholesky decomposition of the two-electron integral matrix. Computer Physics Communications, 58(1–2):71–81, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090136O.

Winters:1991:BSC

[Win91] K. H. Winters. Bifurcation and stability: a computational approach. Computer Physics Communications, 65(1–3):299–309, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559190183L.

Wong:1991:NST

[WJ91] Yau Shu Wong and Hong Jiang. Numerical simulations for transonic aerodynamic flows. Computer Physics Communications, 65(1–3):310–319, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190184M.

Wiggs:1994:PIC

[WJ94] James Wiggs and Hannes Jónsson. A parallel implementation of the Car–Parrinello method by orbital decomposition. Computer Physics Communications, 81(1–2):1–18, June 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594901082.

Wiggs:1995:HDP

[WJ95] James Wiggs and Hannes Jónsson. A hybrid decomposition parallel implementation of the Car–Parrinello method. Computer Physics Communications, 87(3):319–340, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400153S.

Wittmann:1990:VVB

[WK90] Hans-Peter Wittmann and Kurt Kremer. Vectorized version of the bond fluctuation method for lattice polymers.

Computer Physics Communications, 61(3):309–330, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900475.

Wei:1998:WDAa

[WKH98] G. W. Wei, D. J. Kouri, and D. K. Hoffman. Wavelets and distributed approximating functionals. Computer Physics Communications, 112(1):1–6, July 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000514.

Wang:1995:EPP

[WLD95] J. Wang, P. Liewer, and V. Decyk. 3D electromagnetic plasma particle simulations on a MIMD parallel computer. Computer Physics Communications, 87(1–2):35–53, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400167Z.

Wahnstrom:1999:QMC

[WM99] Göran Wahnström and Thomas R. Mattsson. Quantum Monte Carlo simulation of atomic motion. Computer Physics Communications, 121–122:477–479, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003859.

Wahle:1999:MST

[WNS99] J. Wahle, L. Neubert, and M. Schreckenberg. Modeling and simulation of traffic flow. Computer Physics Communications, 121–122:402–405, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003677.

Wolbers:1998:SDC

[Wol98a] Stephen Wolbers. Strategic directions of computing at Fermilab. Computer Physics Communications, 110(1–3):18–21, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597001471.

Wolf:1998:SEK

[Wol98b] Thomas Wolf. Structural equations for killing tensors of arbitrary rank. Computer Physics Communications, 115(2– 3):316–329, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001234.

Wonneberger:1990:SDP

[Won90] Reinhard Wonneberger. Structured document processing: the L^{A}TEX approach. Computer Physics Communications, 61(1–2):177–189, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559090115H.

Westerhof:1996:MBC

[WP96] E. Westerhof and A. G. Peeters. A model for bootstrap current calculations with bounce averaged Fokker–Planck codes. Computer Physics Communications, 95(2–3):131–138, June 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001433.

Ward:1998:PCP

[WP98] A. J. Ward and J. B. Pendry. A program for calculating photonic band structures and Green’s functions using a nonorthogonal FDTD method. Computer Physics Communications, 112(1):23–41, July 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000496.

Wolff:1999:TPM

[WR99] D. Wolff and W. G. Rudd. Tabulated potentials in molecular dynamics simulations. Computer Physics Communications, 120(1):20–32, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002179.

Webster:1991:NPC

[WRF91] Frank Webster, P. J. Rossky, and R. A. Friesner. Nonadiabatic processes in condensed matter: semi-classical theory and implementation. Computer Physics Communications, 63(1–3):494–522, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190272M.

Warren:1995:PPP

[WS95a] Michael S. Warren and John K. Salmon. A portable parallel particle program. Computer Physics Communications, 87(1– 2):266–290, May 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://qso.lanl.gov/

papers/cpc/v9.ps;   http://www.sciencedirect.com/science/
article/pii/0010465594001774.

Widmann:1995:PMC

[WS95b] Albert H. Widmann and Ulrich W. Suter. Parallelization of a Monte Carlo algorithm for the simulation of polymer melts. Computer Physics Communications, 92(2–3):229–251, December 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465595000920.

Wojciechowski:1999:CSE

[WT99] K. W. Wojciechowski and K. V. Tretiakov. Computer simulation of elastic properties of solids under pressure. Computer Physics Communications, 121–122:528–530, September/ October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599003987.

Wuchterl:1995:TDC

[Wuc95] Günther Wuchterl. Time-dependent convection on selfadaptive grids. Computer Physics Communications, 89(1– 3):119–126, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559500189M.

Wang:1990:CVM

[WW90] Chia-Jiu Wang and Chwan-Hwa Wu. Concurrent and vectorized Monte Carlo simulation of the evolution of an assembly of particles increasing in number. Computer Physics Communications, 58(1–2):63–70, February/March 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090135N.

Wang:1993:SHC

[WW93] J. B. Wang and J. F. Williams. Study of the helium 3^{1}D → 2^{1}P → ^{1}S cascade with algebraic computing. Computer Physics Communications, 75(3):275–282, May 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390044D.

Weissman:1999:CCC

[WY99] Z. Weissman and R. Z. Yahel. Communication courseware for college students. Computer Physics Communications, 121–122:578–582, September/October 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004105.

Wybourne:1994:CWC

[Wyb94] B. G. Wybourne. Comment on “Wbase: a C package to reduce tensor products of Lie algebra representations”. Computer Physics Communications, 83(2–3):332–333, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594900582. See [Can94b, Can94a].

Wei:1998:WDAb

[WZA^{+}98] G. W. Wei, D. S. Zhang, S. C. Althorpe, D. J. Kouri, and D. K. Hoffman. Wavelet-distributed approximating functional method for solving the Navier–Stokes equation. Computer Physics Communications, 115(1):18–24, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001131.

Wei:1998:DAF

[WZKH98a] G. W. Wei, D. S. Zhang, D. J. Kouri, and D. K. Hoffman. Distributed approximating functional approach to Burgers’ equation in one and two space dimensions. Computer Physics Communications, 111(1–3):93–109, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000411.

Wei:1998:RRA

[WZKH98b] G. W. Wei, D. S. Zhang, D. J. Kouri, and D. K. Hoffman. A robust and reliable approach to nonlinear dynamical problems. Computer Physics Communications, 111(1–3):87–92, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598000204.

Liu:1995:NVC

[xLzHjW95] Yu xin Liu, Qi zhi Han, and Jia jun Wang. A new version of CFPSIB: fractional parentage of an identical boson system. Computer Physics Communications, 85(1):89–98, January 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559400122I.

Liu:1994:CMC

[xLzSlLzH94] Yu xin Liu, Hong zhou Sun, Gui lu Long, and Qi zhi Han. A checking method on calculating code for the eigenvalue problem of many particle systems. Computer Physics Communications, 81(1–2):145–152, June 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901163.

Xiao:1999:OSS

[XYE99] F. Xiao, T. Yabe, and T. Ebisuzaki. An oscillation suppressing semi-Lagrangian solver for advection equation. Computer Physics Communications, 116(1):121–135, January 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000940.

Xiao:1996:COP

[XYI96] F. Xiao, T. Yabe, and T. Ito. Constructing oscillation preventing scheme for advection equation by rational function. Computer Physics Communications, 93(1):1–12, January 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465595001247.

Xiao:1997:ASS

[XYIT97] F. Xiao, T. Yabe, T. Ito, and M. Tajima. An algorithm for simulating solid objects suspended in stratified flow. Computer Physics Communications, 102(1–3):147–160, May 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000234.

Xiao:1996:CMD

[XYNI96] F. Xiao, T. Yabe, G. Nizam, and T. Ito. Constructing a multi-dimensional oscillation preventing scheme for the advection equation by a rational function. Computer Physics Communications, 94(2–3):103–118, April 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465596000082.

Xie:1991:PAS

[XZ91] Ganquan Xie and Qisu Zou. A parallel algorithm for solving the 3D inverse scattering problem. Computer Physics Communications, 65(1–3):320–326, April 2, 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190185N.

Yabe:1991:USHa

[YA91]  T. Yabe and T. Aoki.  A universal solver for hy-

perbolic equations by cubic-polynomial interpolation I. one-dimensional solver. Computer Physics Communications, 66(2–3):219–232, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190071R.

Yanez:1999:CSA

[YAD99] R. J. Yáñez, J. C. Angulo, and J. S. Dehesa. Computational studies of atomic information entropies. Computer Physics Communications, 121–122:752, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701552.

Yatsunenko:1999:RDC

[Yat99] Yu. A. Yatsunenko. Reconstruction of the drift characteristic by the integral equation. Computer Physics Communications, 118(1):1–10, April 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598001052.

Yousif:1992:FCS

[YB92] Hashim A. Yousif and Edward Boutros. A FORTRAN code for the scattering of EM plane waves by an infinitely long cylinder at oblique incidence. Computer Physics Communications, 69(2–3):406–414, March/April 1992. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592901782.

Yavor:1999:HOS

[YB99] M. I. Yavor and A. S. Berdnikov. High order slices — a way to make an accurate approximation of twodimensional electrostatic and magnetostatic fields near the axis of symmetry. Computer Physics Communications, 121–122:753, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506701564.

Yi:1991:SAS

[YBS91] Jae-Yel Yi, J. Bernholc, and Peter Salamon. Simulated annealing strategies for molecular dynamics. Computer Physics Communications, 66(2–3):177–180, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559190066T.

Yeung:1991:PLB

[Yeu91] W. Yeung. Parallelising the LMTO band-structure program on transputer arrays. Computer Physics Communications, 66(1):47–54, July 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465591900067.

Yabe:1991:USHb

[YIW^{+}91] T. Yabe, T. Ishikawa, P. Y. Wang, T. Aoki, Y. Kadota, and F. Ikeda. A universal solver for hyperbolic equations by cubic-polynomial interpolation II. two- and three-dimensional solvers. Computer Physics Communications, 66(2–3):233– 242, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190072S.

Yongming:1992:MIR

[YJ92] Lou Yongming and Börje Johansson. A multivariant interpolation routine for a random distribution of data points. Computer Physics Communications, 70(2):389–400, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465592902019.

Yousif:1990:FCS

[YK90] Hashim A. Yousif and Sigurd Köhler. A Fortran code for the scattering of EM plane waves by two cylinders at normal incidence. Computer Physics Communications, 59(2):371–385, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590901854.

Yorke:1995:UMN

[YK95] Harold W. Yorke and Michael Kaisig. Use of multiply nested grids for the solution of flux-limited radiation diffusion and hydrodynamics. Computer Physics Communications, 89(1–3):29–44, August 2, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594001844.

Yano:1996:HSM

[YKM^{+}96] T. Yano, K. Kitani, H. Miyatake, M. Otsuka, S. Tomiyoshi, S. Matsushima, T. Wada, and Y. Ezawa. A highspeed method for eigenvalue problems. IV. sturm–liouvilletype differential equations. Computer Physics Communications, 96(2–3):247–262, August 1, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465596000513.

Yano:1994:HSM

[YKO^{+}94] T. Yano, K. Kitani, M. Otsuka, S. Tomiyoshi, S. Matsushima, T. Wada, and Y. Ezawa. A high-speed method for eigenvalue problems III case of unsymmetrical potentials in Milne’s method. Computer Physics Communications, 81(3):409–424, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900884.

Yasar:1992:RMA

[YM92] Osman Yasar and Gregory A. Moses. R-MHD: an adaptivegrid radiation-magnetohydrodynamics computer code. Computer Physics Communications, 69(2–3):439–458, March/ April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290182X.

Yousif:1997:BFF

[YM97] Hashim A. Yousif and Richard Melka. Bessel function of the first kind with complex argument. Computer Physics Communications, 106(3):199–206, November 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465597000878.

Youchison:1991:ITA

[YN91] D. L. Youchison and M. D. Nahemow. Ion Trajectory Analysis Program (ITAP). Computer Physics Communications, 64(1):167–182, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190059T.

Young:1994:SRA

[YN94] I. G. Young and P. H. Norrington. Solution of the relativistic asymptotic equations in electron-ion scattering. Computer Physics Communications, 83(2–3):215–226, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465594900493.

Young:1999:DBM

[You99] Peter Young. Data-based mechanistic modelling, generalised sensitivity and dominant mode analysis. Computer Physics Communications, 117(1–2):113–129, March 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001684.

Yoon:1999:USI

[YY99] Seong Y. Yoon and T. Yabe. The unified simulation for incompressible and compressible flow by the predictorcorrector scheme based on the CIP method. Computer Physics Communications, 119(2–3):149–158, June 2, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599001927.

Yano:1992:HSM

[YYK^{+}92] T. Yano, T. Yokota, K. Kawabata, M. Otsuka, S. Matsushima, Y. Ezawa, and S. Tomiyoshi. A high-speed method for eigenvalue problems. II calculation of the eigenfunction in Milne’s method. Computer Physics Communications, 72(1):61–75, October 1992. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559290006K.

Yokota:1990:HSM

[YYOM90] Toshiaki Yokota, Tadashi Yano, Masashi Otsuka, and Shigeo Matsushima. A high-speed method for solving eigenvalue problems: use of the spline function in Milne’s method. Computer Physics Communications, 61(3):387–394, December 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465590900523.

Zio:1999:SUA

[ZA99] Enrico Zio and George E. Apostolakis. Sensitivity and uncertainty analysis within a methodology for evaluating environmental restoration technologies. Computer Physics Communications, 117(1–2):1–10, March 1, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001519.

Zagrodzinski:1999:STP

[Zag99] J. A. Zagrodziński. Soliton type phenomena in discrete systems. Computer Physics Communications, 121–122:437– 439, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046559900377X.

Zanarini:1990:KUI

[Zan90] Pietro Zanarini. KUIP: the user interface of PAW. Computer Physics Communications, 61(1–2):103–114, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090109E.

Zatsarinny:1996:GPC

[Zat96] Oleg Zatsarinny. A general program for computing matrix elements in atomic structure with nonorthogonal orbitals. Computer Physics Communications, 98(1–2):235–254, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465596000793.

Zauli:1994:EVA

[Zau94] C. Zauli. An extended version of “ASYROT”. Computer Physics Communications, 79(3):555–560, May 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901945.

Zuckert:1994:SCS

[ZAW94] U. Zückert, R. Alkofer, and H. Weigel. Self-consistent solution to a complex fermion determinant with space dependent fields. Computer Physics Communications, 82(1):42–56, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594901309.

Zeiri:1994:TSS

[Zei94] Y. Zeiri. Theoretical studies of surface diffusion and of photo-induced surface processes. Computer Physics Communications, 80(1–3):200–241, March 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594901031.

Zeiri:1997:SLE

[Zei97] Yehuda Zeiri. Study of the lowest energy structure of atomic clusters using a genetic algorithm. Computer Physics Communications, 103(1):28–42, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465597000064.

Zilberberg:1992:CHT

[ZFZ92] I. L. Zilberberg, M. J. Filatov, and G. M. Zhidomirov. Coulomb hybrid two-electron integrals for transition metal atoms: comparison of theoretical integrals with those evaluated from spectroscopic data. Computer Physics Communications, 73(1–3):205–208, December 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559290041V.

Zhang:1991:MQW

[Zha91] John Z. H. Zhang. Multichannel quantum wave packet propagation in the interaction picture: application to gassurface scattering. Computer Physics Communications, 63 (1–3):28–37, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559190235D.

Zhang:1994:QDS

[Zha94]  Lijun Zhang.  Quantum dynamics simulation of few

body Coulomb systems. Computer Physics Communications, 83(2–3):156–170, December 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465594900450.

Zhang:1996:ECH

[Zha96] Fei Zhang. Energy corrections in Hamiltonian dynamics simulations. Computer Physics Communications, 99(1):53–58, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001129.

Zhang:1998:EBP

[Zha98a] Bin Zhang. Erratum to ZPC 1.0.1: a parton cascade for ultrarelativistic heavy ion collisions [Comput. Phys. Commun. 109 (1998) 193–206]. Computer Physics Communications, 111(1–3):276, June 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465598000356.

See [Zha98b].

Zhang:1998:ZPC

[Zha98b] Bin Zhang. ZPC 1.0.1: a parton cascade for ultrarelativistic heavy ion collisions. Computer Physics Communications, 109(2–3):193–206, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000101. See erratum [Zha98a].

Zhang:1998:MER

[Zha98c] G. P. Zhang. Modified explicitly restarted Lanczos algorithm. Computer Physics Communications, 109(1):27–33, March 1998. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465597001367.

Zheng:1999:MCS

[Zhe99] B. Zheng. Monte Carlo simulations of short-time critical dynamics. Computer Physics Communications, 121–122:338– 340, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003495.

Ziegler:1998:NAM

[Zie98] Udo Ziegler. NIRVANA^{+}: An adaptive mesh refinement code for gas dynamics and MHD. Computer Physics Communications, 109(2–3):111–134, April 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465598000228.

Ziegler:1999:TDC

[Zie99] Udo Ziegler. A three-dimensional Cartesian adaptive mesh code for compressible magnetohydrodynamics. Computer Physics Communications, 116(1):65–77, January 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598001398.

Zharekeshev:1999:ALD

[ZK99] I. Kh. Zharekeshev and B. Kramer. Advanced Lanczos diagonalization for models of quantum disordered systems. Computer Physics Communications, 121–122:502– 504, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003926.

Zhang:1999:DAF

[ZKHG99] De S. Zhang, Donald J. Kouri, David K. Hoffman, and Gemunu H. Gunaratne. Distributed approximating functional treatment of noisy signals. Computer Physics Communications, 120(1):1–12, July 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599002155.

Zeng:1999:APA

[ZL99] Chen Zeng and P. L. Leath. Application of polynomial algorithms to a random elastic medium. Computer Physics Communications, 121–122:531–535, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003999.

Zlokazov:1990:AHA

[Zlo90] V. B. Zlokazov. Analysis of hidden anharmonic periodicities. Computer Physics Communications, 59(2):217–223, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559090171V.

Zlokazov:1995:APA

[Zlo95] V. B. Zlokazov. AUTOX — a program for autoindexing reflections from multiphase polycrystals. Computer Physics Communications, 85(3):415–422, March 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559400151Q.

Zhi-meng:1997:HUS

[ZmSjFyJ97] Teng Zhi-meng, Cai Shi-jie, Zhang Fu-yan, and Zheng Jiang. A hybrid upwind scheme to reduce the crosswind diffusion inherent in the SG scheme for semiconductor device simulations. Computer Physics Communications, 103(2–3):187– 196, July 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465596001373.

Zhu:1993:NMT

[ZN93] Chaoyuan Zhu and Hiroki Nakamura. Numerical method for the two-state linear curve crossing: nonadiabatic tunneling case. Computer Physics Communications, 74(1):9–17, January 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390102I.

Zheng:1991:NFA

[ZS91] Wenxin Zheng and Staffan Ström. The null-field approach to electromagnetic resonance of composite objects. Computer Physics Communications, 68(1–3):157–174, November 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190198T.

Zou:1991:PCA

[ZX91] Qisu Zou and Ganquan Xie. A parallel computational algorithm for an inverse problem of low frequency. Computer Physics Communications, 65(1–3):327–332, April 2, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559190186O.

Ziegler:1997:NGR

[ZY97] Udo Ziegler and Harold W. Yorke. A nested grid refinement technique for magnetohydrodynamical flows. Computer Physics Communications, 101(1–2):54–74, April 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465596001634.

Zybin:1999:EAM

[Zyb99] Sergey V. Zybin. Ensemble-averaged molecular dynamics simulation of shock wave in Lennard-Jones crystals. Computer Physics Communications, 121–122:755, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506701588.

Zhimeng:1994:GSG

[ZYQ94] Teng Zhimeng, He Ye, and Tong Qinyi. Generalized Scharfetter–Gummel scheme reducing the crosswind effect for the current continuity equation including energy balance. Computer Physics Communications, 79(2):190–200, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465594900671.

Zimmerer:1991:TMT

[ZZGS91] P. Zimmerer, M. Zimmermann, N. Grün, and W. Scheid. Trajectory method for the time-dependent Schrödinger and Thomas–Fermi equations. Computer Physics Communications, 63(1–3):21–27, February 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046559190234C.